### v1

# Discrete conformal structures on surfaces with boundary (II)—Rigidity and Existence

Abstract.
This is a continuation of [16] studying the discrete conformal structures on surfaces with boundary, in which we gave a classification of the discrete conformal structures on surfaces with boundary. In this paper, we prove the rigidity and existence of these discrete conformal structures on surfaces with boundary.
Key words and phrases:
Discrete conformal structures; Rigidity; Existence; Surfaces with boundary
MSC (2020): 52C25,52C26
1. Introduction
1.1. Basic definition and notations
Let $\widetilde{\mathcal{T}}={(V,\widetilde{E},\widetilde{F})}$ be a triangulation of a closed surface $\widetilde{\Sigma}$, where $V,\widetilde{E},\widetilde{F}$ represent the sets of vertices, edges and faces respectively and $V$ is a finite subset of $\widetilde{\Sigma}$ with $|V|=N$. Denote $N(V)$ as a small open regular disjoint neighborhood of the union of all vertices $V$. Then $\Sigma=\widetilde{\Sigma}-N(V)$ is a compact surface with $N$ boundary components. The intersection $\mathcal{T}=\widetilde{\mathcal{T}}\cap\Sigma$ is called an ideal triangulation of the surface $\Sigma$. The intersections $E=\widetilde{E}\cap\Sigma$ and $F=\widetilde{F}\cap\Sigma$ are the sets of the ideal edges and ideal faces of $\Sigma$ in the ideal triangulation $\mathcal{T}$ respectively. The intersection of an ideal face and the boundary $\partial\Sigma$ is called a boundary arc. For simplicity, we use $B=\{1,2,...,N\}$ to denote the boundary components and label them as $i\in B$. We use $E_{+}$ to denote the set of oriented edges. Denote $\{ij\}\in E$ as the ideal edge between two adjacent boundary components $i,j\in B$. Label the ordered ideal edge as $(i,j)$. Denote $\{ijk\}\in F$ as the ideal face adjacent to the boundary components $i,j,k\in B$. In fact, the ideal face $\{ijk\}$ is a hexagon, which corresponds to the triangle $v_{i}v_{j}v_{k}\in\widetilde{F}$ in $\widetilde{\mathcal{T}}$. The sets of real valued functions on $B,E$ and $E_{+}$ are denoted by $\mathbb{R}^{N}$, $\mathbb{R}^{E}$ and $\mathbb{R}^{E_{+}}$ respectively.
An edge length function associated to $\mathcal{T}$ is a vector $l\in\mathbb{R}^{E}_{>0}$ assigning each ideal edge $\{ij\}\in E$ a positive number $l_{ij}=l_{ji}$. For any three positive numbers, there exists a unique right-angled hyperbolic hexagon (up to isometry) with the lengths of three non-pairwise adjacent edges given by these three positive numbers [7]. Hence, for any ideal face $\{ijk\}\in F$, there exists a unique right-angled hyperbolic hexagon whose three non-pairwise adjacent edges have lengths $l_{ij},l_{ik},l_{jk}$. Gluing all such right-angled hyperbolic hexagons isomorphically along the ideal edges in pairs, one can construct a hyperbolic surface with totally geodesic boundary from the ideal triangulation $\mathcal{T}$. Conversely, any ideally triangulated hyperbolic surface $(\Sigma,\mathcal{T})$ with totally geodesic boundary produces a function $l\in\mathbb{R}^{E}_{>0}$ with $l_{ij}$ given by the length of the shortest geodesic connecting the boundary components $i,j\in B$. The edge length function $l\in\mathbb{R}^{E}_{>0}$ is called a discrete hyperbolic metric on $(\Sigma,\mathcal{T})$. The length $K_{i}$ of the boundary component $i\in B$ is called the generalized combinatorial curvature of the discrete hyperbolic metric $l$ at $i$. Namely, the generalized combinatorial curvature $K_{i}$ is defined to be
$K_{i}=\sum_{\{ijk\}\in F}\theta^{jk}_{i},$
where the summation is taken over all the right-angled hyperbolic hexagons adjacent to $i$ and the generalized angle $\theta^{jk}_{i}$ is the length of the boundary arc of the right-angled hyperbolic hexagon $\{ijk\}$ at $i$.
Motivated by Glickenstein’s work in [2] and Glickenstein-Thomas’ work in [3] for triangulated closed surfaces, we introduced the following two definitions of partial edge length and discrete conformal structures on an ideally triangulated surface with boundary.
Definition 1.1 ([16], Definition 1.1).
Let $(\Sigma,\mathcal{T})$ be an ideally triangulated surface with boundary. An assignment of partial edge lengths is a map $d\in\mathbb{R}^{E_{+}}$ such that $l_{ij}=d_{ij}+d_{ji}>0$ for every edge $\{ij\}\in E$ and
(1) $\sinh d_{ij}\sinh d_{jk}\sinh d_{ki}=\sinh d_{ji}\sinh d_{kj}\sinh d_{ik}$
for every ideal face $\{ijk\}\in F$.
Definition 1.2 ([16], Definition 1.2).
Let $(\Sigma,\mathcal{T})$ be an ideally triangulated surface with boundary. A discrete conformal structure $d=d(f)$ on $(\Sigma,\mathcal{T})$ is a smooth map, sending a function $f\in\mathbb{R}^{N}$ defined on boundary components $B$ to a partial edge length function $d\in\mathbb{R}^{E_{+}}$, such that
(2) $\frac{\partial l_{ij}}{\partial f_{i}}=\coth d_{ij}$
for each $(i,j)\in E_{+}$ and
(3) $\frac{\partial d_{ij}}{\partial f_{k}}=0$
if $k\neq i$ and $k\neq j$. The function $f\in\mathbb{R}^{N}$ is called a discrete conformal factor.
Remark 1.3.
In general, the partial edge lengths $d\in\mathbb{R}^{E_{+}}$ do not satisfy the symmetry condition, i.e., $d_{ij}\neq d_{ji}$. Denote $E_{ij}$ as the hyperbolic geodesic in the hyperbolic plane $\mathbb{H}^{2}$ extending the ideal edge $\{ij\}\in E$. The partial edge length $d_{ij}$ represents the signed distance from a point $c_{ij}$ in $E_{ij}$ to $i\in B$ and $d_{ji}$ represents the signed distance from $c_{ij}$ to $j\in B$. This point $c_{ij}$ is called the edge center of the ideal edge $\{ij\}$. Note that the signed distance $d_{ij}$ of $c_{ij}$ to $i$ is positive if $c_{ij}$ is on the same side of $i$ as $j$ along the hyperbolic geodesic $E_{ij}$, and negative otherwise. Please refer to Figure 1.
Figure 1. Poincaré dual to the ideal edge $\{ij\}\in E$ with the position determined by $d_{ij}$ and $d_{ji}$ in the Klein model (dotted). The red lines represent the boundary arcs of the ideal face $\{ijk\}$. The black triangle $v_{i}v_{j}v_{k}$ is the corresponding hyper-ideal hyperbolic triangle.
Remark 1.4.
We briefly explain the motivation of Definition 1.1 for the partial edge length on surfaces with boundary. One can refer to [16] for more details. The hyperbolic geodesic orthogonal to the hyperbolic geodesic $E_{ij}$ through $c_{ij}$ is called the edge perpendicular, denoted by $P_{ij}$. We can identify the perpendicular $P_{ij}$ with a unique 2-dimensional time-like subspace $\widetilde{P}_{ij}$ of the Lorentzian space $\mathbb{R}^{3}$ such that $\widetilde{P}_{ij}\cap\mathbb{H}^{2}=P_{ij}$. Since the 2-dimensional time-like subspaces $\widetilde{P}_{ij}$ and $\widetilde{P}_{jk}$ associated to $P_{ij}$ and $P_{jk}$ intersect in a 1-dimensional subspace of $\mathbb{R}^{3}$, then the intersection $\widetilde{P}_{ij}\cap\widetilde{P}_{jk}$ corresponds to a point $q$ in the Klein model of hyperbolic plane $\mathbb{H}^{2}$. We view this point $q$ as the intersection point of the perpendiculars $P_{ij}$ and $P_{jk}$. Note that the point $q$ may be not in $\mathbb{H}^{2}$, i.e., the perpendiculars $P_{ij},P_{jk}$ may not intersect within $\mathbb{H}^{2}$. Lemma 2.4 in [16] shows that the perpendiculars $P_{ij},P_{jk},P_{ki}$ of $\{ijk\}$ intersect at a common point $c_{ijk}$ (called face center) if and only if the condition (1) holds. Hence, the condition (1) in Definition 1.1 ensures the existence of a dual geometric structure on the Poincaré dual of an ideally triangulated surface with boundary.
Remark 1.5.
The geometric motivation for Definition 1.2 is that the vertex $v_{k}$ and face center $c_{ijk}$ of a hyper-ideal hyperbolic triangle should be in the 2-dimensional vector subspace of $\mathbb{R}^{3}$ including the vertex $\overline{v}_{k}$ of the new hyper-ideal hyperbolic triangle under the $f_{k}$-conformal variation. Please refer to [16, Remark 1.3] for more details.
The main result in [16] is the following theorem, which gives the explicit forms of the discrete conformal structures in Definition 1.2.
Theorem 1.6 ([16], Theorem 1.4).
Let $(\Sigma,\mathcal{T})$ be an ideally triangulated surface with boundary and let $d=d(f)$ be a discrete conformal structure on $(\Sigma,\mathcal{T})$. There exist constant vectors $\alpha\in\mathbb{R}^{N}$, $\eta\in\mathbb{R}_{>0}^{E}$ and $\zeta\in\mathbb{R}^{E}$ satisfying $\eta_{ij}=\eta_{ji}$ and $\zeta_{ij}=\zeta_{ji}$, such that for any right-angled hyperbolic hexagon $\{ijk\}\in F$,
(A1):
if $\frac{\sinh d_{ij}}{\sinh d_{ji}}=\sqrt{\frac{1+\alpha_{i}e^{2f_{i}}}{1+\alpha_{j}e^{2f_{j}}}}>0$, then the discrete conformal structure $d=d(f)$ is
(4) $\coth d_{ij}=-\frac{\alpha_{i}e^{2f_{i}}}{\sinh l_{ij}}\sqrt{\frac{1+\alpha_{j}e^{2f_{j}}}{1+\alpha_{i}e^{2f_{i}}}}+\frac{\eta_{ij}e^{f_{i}+f_{j}}}{\sinh l_{ij}}$
with
(5) $\cosh l_{ij}=-\sqrt{(1+\alpha_{i}e^{2f_{i}})(1+\alpha_{j}e^{2f_{j}})}+\eta_{ij}e^{f_{i}+f_{j}};$
(A2):
if $\frac{\sinh d_{ij}}{\sinh d_{ji}}=e^{\frac{1}{2}C_{ij}(f_{i}-f_{j})}>0$, then the discrete conformal structure $d=d(f)$ is
(6) $\coth d_{ij}=\frac{\sinh(f_{j}-f_{i}-C_{ij})}{\sinh l_{ij}}+\frac{\eta_{ij}e^{f_{i}+f_{j}}}{\sinh l_{ij}}$
with
(7) $\cosh l_{ij}=-\cosh(f_{j}-f_{i}-C_{ij})+\eta_{ij}e^{f_{i}+f_{j}},$
where $C\in\mathbb{R}^{E_{+}}$ is a constant vector satisfying $C_{ij}+C_{jk}+C_{ki}=0$ for any $\{ijk\}\in F$ and $C_{rs}+C_{sr}=0$ for any $\{r,s\}\subseteq\{i,j,k\}$;
(B1):
if $\frac{\sinh d_{ij}}{\sinh d_{ji}}=-\sqrt{\frac{1+\alpha_{i}e^{2f_{i}}}{1+\alpha_{j}e^{2f_{j}}}}<0$, then the discrete conformal structure $d=d(f)$ is
(8) $\coth d_{ij}=\frac{\alpha_{i}e^{2f_{i}}}{\sinh l_{ij}}\sqrt{\frac{1+\alpha_{j}e^{2f_{j}}}{1+\alpha_{i}e^{2f_{i}}}}+\frac{\zeta_{ij}e^{f_{i}+f_{j}}}{\sinh l_{ij}}$
with
(9) $\cosh l_{ij}=\sqrt{(1+\alpha_{i}e^{2f_{i}})(1+\alpha_{j}e^{2f_{j}})}+\zeta_{ij}e^{f_{i}+f_{j}};$
(B2):
if $\frac{\sinh d_{ij}}{\sinh d_{ji}}=-e^{\frac{1}{2}C_{ij}(f_{i}-f_{j})}<0$, then the discrete conformal structure $d=d(f)$ is
$\coth d_{ij}=-\frac{\sinh(f_{j}-f_{i}-C_{ij})}{\sinh l_{ij}}+\frac{\zeta_{ij}e^{f_{i}+f_{j}}}{\sinh l_{ij}}$
with
(10) $\cosh l_{ij}=\cosh(f_{j}-f_{i}-C_{ij})+\zeta_{ij}e^{f_{i}+f_{j}},$
where $C\in\mathbb{R}^{E_{+}}$ is a constant vector satisfying $C_{ij}+C_{jk}+C_{ki}=0$ for any $\{ijk\}\in F$ and $C_{rs}+C_{sr}=0$ for any $\{r,s\}\subseteq\{i,j,k\}$.
Remark 1.7.
As mentioned in [16, Remark 1.5], we can always assume $\alpha:B\rightarrow\{-1,0,1\}$ in the discrete conformal structures (A1) and (B1) and $C\equiv 0$ in the discrete conformal structures (A2) and (B2) for simplicity.
Remark 1.8.
According to [16, Remark 1.7], the discrete conformal structures (A1) and (A2) can exist alone on an ideally triangulated surface with boundary, but they can not exist simultaneously on the same ideally triangulated surface with boundary. Moreover, the discrete conformal structures (B1) and (B2) can not exist alone on an ideally triangulated surface with boundary. Therefore, we consider the mixed type of discrete conformal structures, i.e., the discrete hyperbolic metrics $l_{ij},l_{jk},l_{ki}$ of a right-angled hyperbolic hexagon $\{ijk\}\in F$ are induced by (7) and (10) or (5) and (9), such that the condition (1) holds. We call the former the mixed discrete conformal structure I and the latter the mixed discrete conformal structure II. For example, set $\{r,s,t\}=\{i,j,k\}$, if $d_{rs}<0$, then $d_{rt}<0$ by Lemma 2.17, and $d_{sr}>0,\ d_{tr}>0$ by the conditions $l_{rs}=d_{rs}+d_{sr}>0$ and $l_{rt}=d_{rt}+d_{tr}>0$ in Definition 1.1. We need $d_{st}>0,\ d_{ts}>0$ to make the condition (1) hold. Thus $l_{rs},l_{rt}$ are defined by (10) and $l_{st}$ is defined by (7) or $l_{rs},l_{rt}$ are defined by (9) and $l_{st}$ is defined by (5).
1.2. Rigidity and existence of discrete conformal structures
An important problem in discrete conformal geometry is to obtain the relationships between the discrete conformal structures and their combinatorial curvatures. The main result in this paper is the following theorem, which gives the global rigidity and existence of discrete conformal structures on ideally triangulated surfaces with boundary.
For convenience of description, we use the new parameters $\xi^{(1)}$ and $\xi^{(2)}$ to replace parameters $\eta$ and $\zeta$ in the mixed discrete conformal structures $\mathrm{I}$ and $\mathrm{II}$, respectively.
Theorem 1.9.
Let $(\Sigma,\mathcal{T})$ be an ideally triangulated surface with boundary and let $d=d(f)$ be a discrete conformal structure on $(\Sigma,\mathcal{T})$.
(i):
For the discrete conformal structure (A1), let $\alpha:B\rightarrow\{-1,0,1\}$ and $\eta\in\mathbb{R}_{>0}^{E}$ be the weights on $(\Sigma,\mathcal{T})$ satisfying $\eta_{ij}>\alpha_{i}\alpha_{j}$ if $\alpha_{i}=\alpha_{j}$ for any two adjacent boundary components $i,j\in B$;
(ii):
For the discrete conformal structure (A2), let $\eta\in\mathbb{R}_{>0}^{E}$ be the weight on $(\Sigma,\mathcal{T})$;
(iii):
For the mixed discrete conformal structure $\mathrm{I}$, i.e., the discrete hyperbolic metrics $l_{ij},l_{jk},l_{ki}$ of a right-angled hyperbolic hexagon $\{ijk\}$ are induced by (7) and (10), let $\xi^{(1)}\in\mathbb{R}_{>0}^{E}$ be the weights on $(\Sigma,\mathcal{T})$;
(iv):
For the mixed discrete conformal structure $\mathrm{II}$, i.e., the discrete hyperbolic metrics $l_{ij},l_{jk},l_{ki}$ of a right-angled hyperbolic hexagon $\{ijk\}$ are induced by (5) and (9), let $\alpha:B\rightarrow\{-1,0,1\}$ and $\xi^{(2)}:E\rightarrow(1,+\infty)$ be the weights on $(\Sigma,\mathcal{T})$;
Then the discrete conformal factor $f$ is determined by its generalized combinatorial curvature $K\in\mathbb{R}^{N}_{>0}$. In particular, the map from $f$ to $K$ is a smooth embedding. Furthermore, the image of $K$ is >$\mathbb{R}^{N}_{>0}$ for the cases $\mathrm{(ii)}$ and $\mathrm{(iii)}$, for the case $\mathrm{(i)}$ if $\alpha:B\rightarrow\{0,1\}$, and for the case $\mathrm{(iv)}$ if $\alpha:B\rightarrow\{0,1\}$ and satisfies a restrictive condition.
Remark 1.10.
If $\alpha\equiv 0$, the case $\mathrm{(i)}$ in Theorem 1.9 generalizes Guo’s results [4] and Li-Xu-Zhou’s results [6]. If $\alpha\equiv 1$, the case $\mathrm{(i)}$ in Theorem 1.9 was proved by Guo-Luo [5]. Moreover, the case $\mathrm{(ii)}$ in Theorem 1.9 was also proved by Guo-Luo [5]. We include them in Theorem 1.9 for completeness. For the case $\mathrm{(iii)}$ in Theorem 1.9, the weight $\xi^{(1)}\in\mathbb{R}_{>0}^{E}$ can be extended to a wider range, please see Theorem 5.1. For the case $\mathrm{(iv)}$ in Theorem 1.9, the weight $\xi^{(2)}:E\rightarrow(1,+\infty)$ can also be extended to a wider range, and we will use a new notation $\Xi$ to denote the set of the range of weight because of its complexity, please see Subsection 6.1. Moreover, for the image of $K$ in the mixed discrete conformal structure $\mathrm{II}$, we not only require $\alpha:B\rightarrow\{0,1\}$, but also $\alpha$ satisfies a restrictive condition. Here the restrictive condition means to exclude the case that $\alpha_{j}=\alpha_{k}=1$ for a right-angled hyperbolic hexagon $\{ijk\}\in F$ with the edge lengths $l_{ij}$ and $l_{ik}$ are given by (9) and the edge length $l_{jk}$ is given by (5), please see Subsection 6.3.
1.3. Basic ideas of the proof of Theorem 1.9
The proof for the existence of discrete conformal structures uses the continuity method. We will show that the image of the curvature map is both open and closed subset of >$\mathbb{R}^{N}_{>0}$, and thus is the whole space >$\mathbb{R}^{N}_{>0}$. The proof for the rigidity of discrete conformal structures involves a variational principle, which was pioneered by Colin de Verdière’s work [1] for tangential circle packings on triangulated surfaces. The proof of the rigidity consists of the following three steps.
The first step is to give a characterization of the admissible space of discrete conformal factor $f$ for a right-angled hyperbolic hexagon. By a change of variables $u_{i}=u_{i}(f_{i})$ described in the proof of Lemma 2.22, the admissible space $\Omega_{ijk}$ of $f$ is transferred to the admissible space $\mathcal{U}_{ijk}$ of $u$, which is proved to be convex and simply connected. The second step is to prove the Jacobian of the generalized angles $\theta$ with respect to $u$ for a right-angled hyperbolic hexagon is symmetric and negative definite. As a consequence, the Jacobian of the generalized combinatorial curvature $K$ with respect to $u$ is symmetric and negative definite on the admissible space $\mathcal{U}=\cap\ \mathcal{U}_{ijk}$. The first step and second step enable us to define a strictly concave function on the admissible space $\mathcal{U}$. The third step is to use the following well-known result from analysis:
Lemma 1.11.
If $f:\Omega\rightarrow\mathbb{R}$ is a $C^{1}$-smooth strictly concave function on an open convex set $\Omega\subset\mathbb{R}^{n}$, then its gradient $\nabla f:\Omega\rightarrow\mathbb{R}^{n}$ is injective. Furthermore, $\nabla f$ is a smooth embedding.
The key step is how to get a change of variables $u_{i}=u_{i}(f_{i})$. For this, we need to calculate the variation formulas of the generalized angles, which is as follows.
Theorem 1.12.
Let $(\Sigma,\mathcal{T})$ be an ideally triangulated surface with boundary and let $d=d(f)$ be a discrete conformal structure on $(\Sigma,\mathcal{T})$.
(i):
For any right-angled hyperbolic hexagon $\{ijk\}\in F$,
$\frac{\partial\theta^{jk}_{i}}{\partial f_{j}}=\frac{-1}{\sinh d_{ji}}\frac{\tanh^{\beta}h_{ij}}{\sinh l_{ij}},$
where $\beta=1$ if $c_{ijk}$ is time-like and $\beta=-1$ if $c_{ijk}$ is space-like. If $c_{ijk}$ is light-like, we interpret the formula as $\tanh^{\beta}h_{ij}=\tanh^{\beta}h_{jk}=\tanh^{\beta}h_{ki}=1$.
(ii):
Furthermore, we have the following Glickenstein-Thomas formula for $\frac{\partial\theta^{jk}_{i}}{\partial f_{i}}$, $\frac{\partial\theta^{ik}_{j}}{\partial f_{i}}$ and $\frac{\partial\theta^{ij}_{k}}{\partial f_{i}}$, i.e.,
$\frac{\partial\theta^{jk}_{i}}{\partial f_{i}}=\cosh l_{ij}\frac{\partial\theta^{ik}_{j}}{\partial f_{i}}+\cosh l_{ik}\frac{\partial\theta^{ij}_{k}}{\partial f_{i}}.$
(iii):
Moreover, there is a change of variables $u_{i}=u_{i}(f_{i})$ such that
(11) $\frac{\partial\theta^{jk}_{i}}{\partial u_{j}}=\frac{\partial\theta^{ik}_{j}}{\partial u_{i}}.$
The function $u$ is also called a discrete conformal factor.
Here $h_{ij}$ is the signed distance of the face center $c_{ijk}$ to the hyperbolic geodesic $E_{ij}$, which is positive if $c_{ijk}$ is on the same side of the geodesic $E_{ij}$ as the right-angled hyperbolic hexagon $\{ijk\}$ and negative otherwise (or zero if $c_{ijk}$ is on the geodesic $E_{ij}$).
1.4. Organization of the paper
In Section 2, we prove Theorem 1.12 by some useful lemmas. In Section 3, we prove Theorem 1.9 (i). In Section 4, we prove Theorem 1.9 (ii). In Section 5, we prove Theorem 1.9 (iii). In Section 6, we prove Theorem 1.9 (iv).
2. Discrete conformal variations of generalized angles
In this section, we first present some basic facts in the 2-dimensional hyperbolic geometry, which are discussed in [7, Chapter 3 and 6]. Then we prove Theorem 1.12. Different from the proof of Theorem 7 in Glickenstein-Thomas [3], we are unable to get the variational formulas of generalized angle by similar geometric properties directly from the figure under the $f_{k}$-conformal variation, and we obtain them by complex calculations.
2.1. Some basic knowledge
For any vectors $x,y$ in $\mathbb{R}^{3}$, the Lorentzian inner product * of $x$ and $y$ is given by $x*y=xJy^{T}$, where $J=\mathrm{diag}\{1,1,-1\}$. If $x*y=0$, then we say $x,y\in\mathbb{R}^{3}$ are Lorentz orthogonal. The Lorentzian cross product $\otimes$ of $x$ and $y$ is defined as $x\otimes y:=J(x\times y)$, where $\times$ is the Euclidean cross product.
The Lorentzian norm of $x\in\mathbb{R}^{3}$ is defined to be the complex number $||x||=\sqrt{x*x}$. Here $||x||$ is either positive, zero or positive imaginary. If $||x||$ is positive imaginary, we denote by $|||x|||$ its absolute value. Furthermore, $x\in\mathbb{R}^{3}$ is called space-like if $x*x>0$, light-like if $x*x=0$ and time-like if $x*x<0$. The hyperboloid model of $\mathbb{H}^{2}$ is defined to be
$\mathbb{H}^{2}=\{x=(x_{1},x_{2},x_{3})\in\mathbb{R}^{3}|\ x*x=-1\ \text{and}\ x_{3}>0\},$
which can be embedded in $(\mathbb{R}^{3},*)$. A geodesic (or hyperbolic line) in $\mathbb{H}^{2}$ is the non-empty intersection of $\mathbb{H}^{2}$ with a 2-dimensional vector subspace of $\mathbb{R}^{3}$.
The Klein model of $\mathbb{H}^{2}$ is the center projection of $\mathbb{H}^{2}$ into the plane $x_{3}=1$ of $\mathbb{R}^{3}$. In the Klein model, the interior, boundary and external points of the unit disk represent the time-like, light-like and space-like vectors respectively, and they are called hyperbolic points, ideal points and hyper-ideal points respectively. A geodesic in $\mathbb{H}^{2}$ is a chord on the unit disk. In the Klein model, a triangle $xyz$ with its three vertices $x,y,z$ being hyperbolic is called a hyperbolic triangle. If at least one of $x,y,z$ is ideal or hyper-ideal, then we say the triangle $xyz$ is a generalized hyperbolic triangle. Specially, if $x,y,z$ are all hyper-ideal, we call this triangle as a hyper-ideal hyperbolic triangle for simplicity.
The following proposition follows from the properties of Lorentzian cross product, which is used extensively in this paper.
Proposition 2.1.
Let $xyz$ be a (hyperbolic or generalized hyperbolic) triangle with $x\in\mathbb{H}^{2}$ and a right angle at $x$. Then
$-(z*y)=(z*x)(x*y).$
For two vectors $x,y\in\mathbb{H}^{2}$, we denoted by $d_{\mathbb{H}}(x,y)$ the hyperbolic distance between $x$ and $y$, which is the arc length of the geodesic between $x$ and $y$. We use $\mathrm{Span}(x)$ or $\mathrm{Span}(x,y)$ to denote the vector subspace spanned $x$ or $x,y$. If $x\in\mathbb{R}^{3}$ is a space-like vector, then the Lorentzian complement $x^{\bot}:=\{w\in\mathbb{R}^{3}|x*w=0\}$ of $\mathrm{Span}(x)$ is a 2-dimension vector subspace of $\mathbb{R}^{3}$ and $x^{\bot}\cap\mathbb{H}^{2}$ is non-empty. This implies that $x^{\bot}\cap\mathbb{H}^{2}$ is a geodesic in $\mathbb{H}^{2}$. And the hyperbolic distance between $x$ and $y\in\mathbb{H}^{2}$ is defined to be $d_{\mathbb{H}}(y,x^{\bot})=\inf\{d_{\mathbb{H}}(y,z)|z\in x^{\bot}\cap\mathbb{H}^{2}\}$. Let $x,y\in\mathbb{R}^{3}$ be space-like vectors. If $\mathrm{Span}(x,y)\cap\mathbb{H}^{2}$ is non-empty, then the hyperbolic distance between $x$ and $y$ is defined by $d_{\mathbb{H}}(x^{\bot},y^{\bot})=\inf\{d_{\mathbb{H}}(z,w)|z\in x^{\bot}\cap\mathbb{H}^{2},w\in y^{\bot}\cap\mathbb{H}^{2}\}$. If $\mathrm{Span}(x,y)\cap\mathbb{H}^{2}$ is empty, then $x^{\bot}$ and $y^{\bot}$ intersect in an angle $\angle(x^{\bot},y^{\bot})$ in $\mathbb{H}^{2}$.
Proposition 2.2 ([7], Chapter 3).
Suppose $x,y\in\mathbb{H}^{2}$ are time-like vectors and $z,w\in\mathbb{R}^{3}$ are space-like vectors. Then
(i):
$x*y=-|||x|||\cdot|||y|||\cosh d_{\mathbb{H}}(x,y)$;
(ii):
$|x*z|=|||x|||\cdot||z||\sinh d_{\mathbb{H}}(x,z^{\bot})$, and $x*z<0$ if and only if $x$ and $z$ are on opposite sides of the hyperplane $z^{\bot}$;
(iii):
if $\mathrm{Span}(z,w)\cap\mathbb{H}^{2}$ is non-empty, then $|z*w|=||z||\cdot||w||\cosh d_{\mathbb{H}}(z^{\bot},w^{\bot})$, and $z*w<0$ if and only if $z$ and $w$ are oppositely oriented tangent vectors of $Q$, where $Q$ is the hyperbolic line Lorentz orthogonal to $z^{\bot}$ and $w^{\bot}$;
(iv):
if $\mathrm{Span}(z,w)\cap\mathbb{H}^{2}$ is empty, then $z*w=||z||\cdot||w||\cos\angle(z^{\bot},w^{\bot})$.
The following proposition is useful for identifying perpendicular geodesics in the Klein model.
Proposition 2.3.
([8], Proposition 4) Let $x\in\mathbb{R}^{3}$ be a space-like vector and let $\gamma=x^{\bot}\cap\mathbb{H}^{2}$ be the geodesic corresponding to $x^{\bot}$. Suppose $\omega$ is another geodesic that intersects $\gamma$, and $P$ is the plane such that $\omega=P\cap\mathbb{H}^{2}$. Then $\gamma$ and $\omega$ meet at a right angle if and only if $x\in P$.
Proposition 2.4.
([7], Theorem 3.2.11) Let $P$ be a geodesic in $\mathbb{H}^{2}$ and let $y\in\mathbb{H}^{2}$ be a time-like vector. Then there exists a unique geodesic of $\mathbb{H}^{2}$ through $y$ and Lorentz orthogonal to $P$.
2.2. A geometric explanation
Consider a hyper-ideal hyperbolic triangle $v_{i}v_{j}v_{k}$ in the Kelin model. Set $\{r,s,t\}=\{i,j,k\}$. Set $E_{rs}=\mathrm{Span}(v_{r},v_{s})\cap\mathbb{H}^{2}$, then $E_{rs}$ is a geodesic in $\mathbb{H}^{2}$. Set $v_{r}^{\bot}\cap\mathbb{H}^{2}=E^{\prime}_{st}$, then $E^{\prime}_{st}$ is also a geodesic in $\mathbb{H}^{2}$ and Lorentz orthogonal to the geodesics $E_{rs},E_{rt}$. The geodesic arc in $E^{\prime}_{st}$ bounded by $E_{rs}$ and $E_{rt}$ is called a boundary arc, denoted by $r\in B$. The geodesic arc between two adjacent boundary arcs $r$ and $s$ is denoted by $\{rs\}$, which is a part of $E_{rs}$ and called an ideal edge. Moreover, the six geodesic arcs form a right-angled hyperbolic hexagon $\{ijk\}$. Conversely, for any geodesic $E_{rs}$ in $\mathbb{H}^{2}$, there exists a unique space-like vector $v^{\prime}_{t}$ such that $(v^{\prime}_{t})^{\bot}\cap\mathbb{H}^{2}=E_{rs}$. Then the three points $v^{\prime}_{i},v^{\prime}_{j}$ and $v^{\prime}_{k}$ are hyper-ideal and form a hyper-ideal hyperbolic triangle $v^{\prime}_{i}v^{\prime}_{j}v^{\prime}_{k}$, which is called the polar triangle of the hyper-ideal hyperbolic triangle $v_{i}v_{j}v_{k}$. Please refer to Figure 2.
Figure 2. A hyper-ideal hyperbolic triangle $v_{i}v_{j}v_{k}$ and its polar triangle $v^{\prime}_{i}v^{\prime}_{j}v^{\prime}_{k}$ in the Kelin model (left), and a right-angled hyperbolic hexagon $\{ijk\}\in F$ in the Poincaré model (right).
By the construction of the polar triangle $v^{\prime}_{i}v^{\prime}_{j}v^{\prime}_{k}$, the vertices $v^{\prime}_{i},v^{\prime}_{j},v^{\prime}_{k}$ are determined by the vertices $v_{i},v_{j},v_{k}$ and vice versa. Without loss of generality, we assume $||v_{i}||=||v_{j}||=||v_{k}||=1$, then
(12) $v^{\prime}_{i}=\frac{v_{k}\otimes v_{j}}{||v_{k}\otimes v_{j}||},\ v^{\prime}_{j}=\frac{v_{i}\otimes v_{k}}{||v_{i}\otimes v_{k}||},\ v^{\prime}_{k}=\frac{v_{j}\otimes v_{i}}{||v_{j}\otimes v_{i}||}.$
Therefore, given a hyper-ideal hyperbolic triangle $v_{i}v_{j}v_{k}$, there exists a unique polar triangle $v^{\prime}_{i}v^{\prime}_{j}v^{\prime}_{k}$ with its vertices satisfying (12). The length of the ideal edge $\{rs\}\in E$ is denoted by $l_{rs}$ and the length of the boundary arc $t\in B$ is denoted by $\theta^{rs}_{t}$.
Lemma 2.5.
The vectors $c_{ijk},c_{rs},v^{\prime}_{t}$ lie in the same 2-dimensional vector subspace of $\mathbb{R}^{3}$, where the notations of $c_{ijk},c_{rs}$ follows from Remark 1.4.
Proof.
Without loss of generality, we assume $r=i,s=j,t=k$. Since $c_{ij}\in E_{ij}$, by Proposition 2.4, there exists a unique geodesic of $\mathbb{H}^{2}$ through $c_{ij}$ and Lorentz orthogonal to $E_{ij}$, which is $P_{ij}=\mathrm{Span}(c_{ij},c_{ijk})\cap\mathbb{H}^{2}$. Note that $E_{ij}=(v^{\prime}_{k})^{\bot}\cap\mathbb{H}^{2}$. By Proposition 2.3, since $E_{ij}$ and $P_{ij}$ meet at a right angle, then $v^{\prime}_{k}\in\mathrm{Span}(c_{ij},c_{ijk})$. Q.E.D.
Combining Remark 1.4 and Lemma 2.5, we have $c_{ijk}=\mathrm{Span}(v^{\prime}_{i},c_{jk})\cap\mathrm{Span}(v^{\prime}_{j},c_{ik})\cap\mathrm{Span}(v^{\prime}_{k},c_{ij})$. Due to the equivalence of $v_{i}v_{j}v_{k}$ and $v^{\prime}_{i}v^{\prime}_{j}v^{\prime}_{k}$, it is natural to consider a dual model. Note that for any $c_{ijk}\in\mathbb{R}^{3}$, $\mathrm{Span}(v_{r},c_{ijk})\cap v_{r}^{\bot}$ may not intersect within $\mathbb{H}^{2}$. For the convenience of calculations, we focus on the case that
$\mathrm{Span}(v_{r},c_{ijk})\cap v_{r}^{\bot}\cap\mathbb{H}^{2}=\mathrm{Span}(v_{r},c_{ijk})\cap E^{\prime}_{st}\neq\emptyset,\quad\forall r\in\{i,j,k\}.$
This condition limits the positions of $c_{ijk}$ when $c_{ijk}$ is space-like, i.e., $c_{ijk}$ can only be in some certain open domains. Denote $c^{\prime}_{st}=\mathrm{Span}(v_{r},c_{ijk})\cap E^{\prime}_{st}$ as the dual edge center, which is the unique point along the geodesic $E^{\prime}_{st}$ such that $c^{\prime}_{st}$ is of (signed) distance $\theta_{st}$ from $\{rt\}$ and $\theta_{ts}$ from $\{rs\}$. Thus we have
(13) $\theta_{r}^{st}=\theta_{st}+\theta_{ts}.$
Note that the signed distance $\theta_{st}$ is positive if $c^{\prime}_{st}$ is on the same side of $\{rt\}$ as $\{rs\}$ along the hyperbolic geodesic $E^{\prime}_{st}$, and negative otherwise. The above construction shows the following lemma.
Lemma 2.6.
The vectors $c_{ijk},c^{\prime}_{rs},v_{t}$ lie in the same 2-dimensional vector subspace of $\mathbb{R}^{3}$.
By Proposition 2.4, the edge perpendicular $P^{\prime}_{ij}=\mathrm{Span}(v_{k},c_{ijk})\cap\mathbb{H}^{2}$ is through $c^{\prime}_{ij}$ and Lorentz orthogonal to $E^{\prime}_{ij}$. Hence, the three edge perpendiculars $P^{\prime}_{ij},P^{\prime}_{jk},P^{\prime}_{ki}$ of $\{ijk\}$ also intersect in a point $c_{ijk}$. Similar to [16, Lemma 2.4], we further have that the three perpendiculars $P^{\prime}_{ij},P^{\prime}_{jk},P^{\prime}_{ki}$ of $\{ijk\}$ intersect in a point if and only if
(14) $\sinh\theta_{ij}\sinh\theta_{jk}\sinh\theta_{ki}=\sinh\theta_{ji}\sinh\theta_{kj}\sinh\theta_{ik}.$
Hence, Definition 1.1 is equivalent to the following definition with respect to the parameter $\theta$. For precision, we rewrite the boundary component $i\in B$ as $\{j^{\prime}k^{\prime}\}\in B$. We use $B_{+}$ to denote the sets of oriented boundary components and label them with ordered pairs $(j^{\prime},k^{\prime})\in B_{+}$.
Definition 2.7.
Let $(\Sigma,\mathcal{T})$ be an ideally triangulated surface with boundary. An assignment of partial edge lengths is a map $\theta:B_{+}\rightarrow\mathbb{R}$, such that $\theta_{i}^{jk}=\theta_{jk}+\theta_{kj}>0$ for each oriented boundary arc $(j^{\prime},k^{\prime})\in B_{+}$ and formula (14) holds for every right-angled hyperbolic hexagon $\{ijk\}\in F$.
Recall the definition of $h_{rs}$, which is the signed distance of the face center $c_{ijk}$ to the hyperbolic geodesic $E_{rs}$ and is positive if $c_{ijk}$ is on the same side of the geodesic $E_{rs}$ as the hyper-ideal hyperbolic triangle $v_{i}v_{j}v_{k}$ and negative otherwise (or zero if $c_{ijk}$ is on the geodesic $E_{ij}$). Similarly, we denote $q_{rs}$ as the signed distance from $c_{ijk}$ to $c^{\prime}_{rs}$, which is defined to be positive if $c_{ijk}$ is on the same side of the geodesic $E^{\prime}_{rs}$ as the polar triangle $v^{\prime}_{i}v^{\prime}_{j}v^{\prime}_{k}$ and negative otherwise (or zero if $c_{ijk}$ is on the geodesic $E^{\prime}_{rs}$). Please refer to Figure 3.
Figure 3. Data for two hyper-ideal hyperbolic triangles in the Klein model
In the Klein model, the hyper-ideal hyperbolic triangle $v_{i}v_{j}v_{k}$ and the polar triangle $v^{\prime}_{i}v^{\prime}_{j}v^{\prime}_{k}$ divide the unit disk into 13 open domains, denoted by $D_{1},D_{2},...,D_{13}$ respectively. Please refer to Figure 4.
Figure 4. Division for a right-angled hyperbolic hexagon $\{ijk\}$ in the Klein model. If $c_{ijk}$ is space-like, then it can be in red and blue open domains.
By the definition of $h$ and $q$, one can directly obtain the relationships between the position of $c_{ijk}$ and the signs of $h$ and $q$ in these 13 open domains, as shown in Table 1. For example, suppose $c_{ijk}$ is in the domain $D_{1}$. Then $c_{ijk}$ is on the opposite side of $E_{ik}$ as $v_{i}v_{j}v_{k}$ and hence $h_{ik}<0$. And $c_{ijk}$ is on the same side of $E_{ij},E_{jk}$ as $v_{i}v_{j}v_{k}$ and hence $h_{ij}>0,h_{jk}>0$. Similarly, $c_{ijk}$ is also on the opposite side of $E^{\prime}_{ij}$ as $v^{\prime}_{i}v^{\prime}_{j}v^{\prime}_{k}$ and hence $q_{ij}<0$. And $c_{ijk}$ is on the same side of $E^{\prime}_{ik},E^{\prime}_{jk}$ as $v^{\prime}_{i}v^{\prime}_{j}v^{\prime}_{k}$ and hence $q_{jk}>0,q_{ik}>0$. If $c_{ijk}$ is space-like, then $c_{ijk}$ can be only in one of 12 open domains and each domain is bounded by two tangential lines and unit circle, denoted by $D_{\mathrm{I}},D_{\mathrm{II}},D_{\mathrm{III}},D_{\mathrm{IV}},D_{\mathrm{V}},D_{\mathrm{VI}}$ and $D_{\mathrm{i}},D_{\mathrm{ii}},D_{\mathrm{iii}},D_{\mathrm{iv}},D_{\mathrm{v}},D_{\mathrm{vi}}$ respectively. Please refer to Figure 4. If $c_{ijk}$ is not in one of 12 open domains, then either $\mathrm{Span}(c_{ijk},v_{r})\cap\mathbb{H}^{2}=\emptyset$ or $\mathrm{Span}(c_{ijk},v^{\prime}_{r})\cap\mathbb{H}^{2}=\emptyset$, which contradicts with the assumption that the edge centers $c_{st}\in E_{rs}$ and $c^{\prime}_{st}\in E^{\prime}_{rs}$ are in $\mathbb{H}^{2}$ for $\{r,s,t\}=\{i,j,k\}$. Similar to the case that $c_{ijk}$ is time-like, one can also directly obtain the relationships between the position of $c_{ijk}$ and the signs of $h$ and $q$ in these 12 open domains, as shown in Table 1. By moving $c_{ijk}$ to the light cone continuously, one can obtain similar results. Note that $c_{ijk}$ can only be on the unit disk that removes 12 points.
Domains $D_{1}$($D_{\mathrm{i}}$) $D_{2}$($D_{\mathrm{I}}$) $D_{3}$($D_{\mathrm{ii}}$) $D_{4}$($D_{\mathrm{II}}$) $D_{5}$($D_{\mathrm{iii}}$) $D_{6}$($D_{\mathrm{III}}$) $D_{7}$($D_{\mathrm{iv}}$) $D_{8}$($D_{\mathrm{IV}}$) $D_{9}$($D_{\mathrm{v}}$) $D_{10}$($D_{\mathrm{V}}$) $D_{11}$($D_{\mathrm{vi}}$) $D_{12}$($D_{\mathrm{VI}}$) $D_{13}$
$h_{jk}$ + + - - - + + + + + + + +
$h_{ik}$ + + + + + + + + + - - +
$h_{ij}$ + + + + + + - - - + + + +
$q_{jk}$ + + + + + + + + - - - + +
$q_{ik}$ + + + + - - - + + + + + +
$q_{ij}$ + + + + + + + + + +
Table 1. Relationships between the position of $c_{ijk}$ and the signs of $h$ and $q$
In the following, we provide another way to obtain Table 1. Suppose $c_{ijk}$ is time-like. Without loss of generality, we assume $c_{ijk}$ is in the domain $D_{1}$. By Lemma 2.5, the vectors $c_{ijk},c_{ik},v^{\prime}_{j}$ lie in the same 2-dimensional vector subspace of $\mathbb{R}^{3}$. Due to $c_{ijk}$ is not in the polar triangle $v^{\prime}_{i}v^{\prime}_{j}v^{\prime}_{k}$, then $c_{ik}\in E_{ik}$ is not in $\{ik\}$. Furthermore, since $c_{ijk}$ is in $D_{1}$, then $c_{ik}$ is on the opposite side of $k$ as $i$. By the definition of $d$ in Remark 1.3, we have $d_{ki}<0,d_{ik}>0$. Similarly, we have $d_{kj}<0,d_{jk}>0,d_{ij}>0,d_{ji}>0$. By Lemma 2.6, the vectors $c_{ijk},c^{\prime}_{ij},v_{k}$ lie in the same 2-dimensional vector subspace of $\mathbb{R}^{3}$. Due to $c_{ijk}$ is not in the hyper-ideal hyperbolic triangle $v_{i}v_{j}v_{k}$, then $c^{\prime}_{ij}\in E^{\prime}_{ij}$ is not in $k$. Furthermore, since $c_{ijk}$ is in $D_{1}$, then $c^{\prime}_{ij}$ is on the opposite side of $\{ik\}$ as $\{jk\}$ and hence $\theta_{ji}<0,\theta_{ij}>0$. Similarly, we have $\theta_{jk}<0,\theta_{kj}>0,\theta_{ik}>0,\theta_{ki}>0$. By Corollary 2.11, we have $h_{ik}<0,h_{jk}>0,h_{ij}>0$ and $q_{ij}<0,q_{ik}>0,q_{jk}>0$. The case that $c_{ijk}$ is space-like follows from the following two lemmas.
Lemma 2.8.
The signs of $q$ and $h$ in the domains $D_{\mathrm{I}},D_{\mathrm{II}},D_{\mathrm{III}},D_{\mathrm{IV}},D_{\mathrm{V}},D_{\mathrm{VI}}$ are the same as those in the domains $D_{2},D_{4},D_{6},D_{8},D_{10},D_{12}$, respectively.
Proof.
We only prove the signs of $q$ and $h$ in the domain $D_{\mathrm{V}}$ are the same as those in the domain $D_{10}$, and the remaining cases are similar.
If $c_{ijk}$ is in the domain $D_{10}$, then $q_{jk}<0$ and the rest is positive by Table 1. Suppose $c_{ijk}$ is in the domain $D_{\mathrm{V}}$. By Lemma 2.5, the vectors $c_{ijk},c_{ij},v^{\prime}_{k}$ lie in the same 2-dimensional vector subspace of $\mathbb{R}^{3}$. Due to $c_{ijk}$ is not in the polar triangle $v^{\prime}_{i}v^{\prime}_{j}v^{\prime}_{k}$, then $c_{ij}\in E_{ij}$ is not in $\{ij\}$. Furthermore, since $c_{ijk}$ is in $D_{\mathrm{V}}$, then $c_{ij}$ is on the opposite side of $j$ as $i$ and hence $d_{ij}<0,d_{ji}>0$. Similarly, we have $d_{ik}<0,d_{ki}>0,d_{jk}>0,d_{kj}>0$. By Proposition 2.12, we have $h_{ij}>0,h_{jk}>0,h_{ik}>0$. Note that $c_{ijk}$ is in the polar triangle $v^{\prime}_{i}v^{\prime}_{j}v^{\prime}_{k}$. By Lemma 2.6, the vectors $c_{ijk},c^{\prime}_{rs},v_{t}$ lie in the same 2-dimensional vector subspace of $\mathbb{R}^{3}$ for $\{r,s,t\}=\{i,j,k\}$. Then $c^{\prime}_{rs}\in E^{\prime}_{rs}$ is in $t$. Hence, $\theta_{ij}>0,\theta_{ji}>0,\theta_{jk}<0,\theta_{kj}>0,\theta_{ik}>0,\theta_{ki}>0$. By Proposition 2.12, we have $q_{ij}>0,q_{jk}<0,q_{ik}>0$. Q.E.D.
Lemma 2.9.
The signs of $q$ and $h$ in the domains $D_{\mathrm{i}},D_{\mathrm{ii}},D_{\mathrm{iii}},D_{\mathrm{iv}},D_{\mathrm{v}},D_{\mathrm{vi}}$ are the same as those in the domains $D_{1},D_{3},D_{5},D_{7},D_{9},D_{11}$, respectively.
Proof.
We only prove the signs of $q$ and $h$ in the domain $D_{\mathrm{vi}}$ are the same as those in the domain $D_{11}$, and the remaining cases are similar.
If $c_{ijk}$ is in the domain $D_{11}$, then $q_{i}<0,h_{j}<0$ and the rest is positive by Table 1. Suppose $c_{ijk}$ is in the domain $D_{\mathrm{vi}}$. By Lemma 2.5, the vectors $c_{ijk},c_{ij},v^{\prime}_{k}$ lie in the same 2-dimensional vector subspace of $\mathbb{R}^{3}$. Due to $c_{ijk}$ is not in the polar triangle $v^{\prime}_{i}v^{\prime}_{j}v^{\prime}_{k}$, then $c_{ij}\in E_{ij}$ is not in $\{ij\}$. Furthermore, since $c_{ijk}$ is in $D_{\mathrm{vi}}$, then $c_{ij}$ is on the opposite side of $j$ as $i$ and hence $d_{ij}<0,d_{ji}>0$. Similarly, we have $d_{ik}<0,d_{ki}>0,d_{jk}>0,d_{kj}>0$. By Proposition 2.14, we have $h_{ij}<0,h_{jk}>0,h_{ik}>0$. By Lemma 2.6, the vectors $c_{ijk},c^{\prime}_{jk},v_{i}$ lie in the same 2-dimensional vector subspace of $\mathbb{R}^{3}$. Due to $c_{ijk}$ is not in the hyper-ideal hyperbolic triangle $v_{i}v_{j}v_{k}$, then $c^{\prime}_{jk}\in E^{\prime}_{jk}$ is not in $i$. Furthermore, since $c_{ijk}$ is in $D_{\mathrm{vi}}$, then $c^{\prime}_{jk}$ is on the opposite side of $\{ik\}$ as $\{ij\}$ and hence $\theta_{jk}<0,\theta_{kj}>0$. Similarly, we have $\theta_{ji}<0,\theta_{ij}>0,\theta_{ik}>0,\theta_{ki}>0$. By Proposition 2.14, we have $q_{ij}>0,q_{jk}<0,q_{ik}>0$. Q.E.D.
By direct calculations, we can get the following results. As the proofs of the Lemma 2.10, Proposition 2.12 and Proposition 2.14 are too long, we defer them to Appendix A. For simplicity, we set $h_{t}=h_{rs},\ q_{t}=q_{rs}$ for $\{r,s,t\}=\{i,j,k\}$.
Lemma 2.10.
If $c_{ijk}$ is time-like, then
$\sinh q_{r}=\cosh h_{s}\cdot\sinh d_{rt}=\cosh h_{t}\cdot\sinh d_{rs},$
$\sinh h_{r}=\cosh q_{s}\cdot\sinh\theta_{rt}=\cosh q_{t}\cdot\sinh\theta_{rs}.$
As a direct corollary of Lemma 2.10, we have
Corollary 2.11.
If $c_{ijk}$ is time-like, then the signs of $q_{r}$, $d_{rt}$ and $d_{rs}$ are the same and the signs of $h_{r}$, $\theta_{rs}$ and $\theta_{rt}$ are the same.
Proposition 2.12.
If $c_{ijk}\in\mathbb{R}^{3}$ lies in the open domain $D_{\mathrm{V}}$, then
$\cosh q_{i}=-\sinh h_{j}\cdot\sinh d_{ik}=-\sinh h_{k}\cdot\sinh d_{ij},$
$\cosh q_{j}=\sinh h_{i}\cdot\sinh d_{jk}=\sinh h_{k}\cdot\sinh d_{ji},$
$\cosh q_{k}=\sinh h_{i}\cdot\sinh d_{kj}=\sinh h_{j}\cdot\sinh d_{ki},$
$\cosh h_{i}=\sinh q_{j}\cdot\sinh\theta_{ik}=\sinh q_{k}\cdot\sinh\theta_{ij},$
$\cosh h_{j}=-\sinh q_{i}\cdot\sinh\theta_{jk}=\sinh q_{k}\cdot\sinh\theta_{ji},$
$\cosh h_{k}=-\sinh q_{i}\cdot\sinh\theta_{kj}=\sinh q_{j}\cdot\sinh\theta_{ki}.$
If $c_{ijk}\in\mathbb{R}^{3}$ lies in the open domain $D_{\mathrm{II}}$, then we swap $h$ and $q$, and $\theta$ and $d$.
If $c_{ijk}$ lies in the domains $D_{\mathrm{I}},D_{\mathrm{III}},D_{\mathrm{IV}}$ or $D_{\mathrm{VI}}$, we have similar propositions. Using these propositions, we can prove Lemma 2.8 and further obtain Table 1. Moreover, from Table 1, we see that $c_{ijk}$ lies in the open domains $D_{\mathrm{I}},D_{\mathrm{II}},D_{\mathrm{III}},D_{\mathrm{IV}},D_{\mathrm{V}}$ or $D_{\mathrm{VI}}$ if and only if one of $q_{i},q_{j},q_{k},h_{i},h_{j},h_{k}$ is negative and the others are positive. Hence, these propositions can be written as a unified form.
Lemma 2.13.
Let $c_{ijk}\in\mathbb{R}^{3}$ lie in one of the open domains $D_{\mathrm{I}},D_{\mathrm{II}},D_{\mathrm{III}},D_{\mathrm{IV}},D_{\mathrm{V}},D_{\mathrm{VI}}$. If $q_{r}<0$, then
$\cosh q_{r}=-\sinh h_{s}\cdot\sinh d_{rt}=-\sinh h_{t}\cdot\sinh d_{rs},$
$\cosh q_{s}=\sinh h_{r}\cdot\sinh d_{st}=\sinh h_{t}\cdot\sinh d_{sr},$
$\cosh q_{t}=\sinh h_{r}\cdot\sinh d_{ts}=\sinh h_{s}\cdot\sinh d_{tr},$
$\cosh h_{r}=\sinh q_{s}\cdot\sinh\theta_{rt}=\sinh q_{t}\cdot\sinh\theta_{rs},$
$\cosh h_{s}=-\sinh q_{r}\cdot\sinh\theta_{st}=\sinh q_{t}\cdot\sinh\theta_{sr},$
$\cosh h_{t}=-\sinh q_{r}\cdot\sinh\theta_{ts}=\sinh q_{s}\cdot\sinh\theta_{tr}.$
If $h_{r}<0$, then we swap $h,\ \theta$ and $q,\ d$.
Proposition 2.14.
Let $c_{ijk}\in\mathbb{R}^{3}$ lie in the open domain $D_{\mathrm{vi}}$, then
$\cosh q_{i}=\sinh h_{j}\cdot\sinh d_{ik}=-\sinh h_{k}\cdot\sinh d_{ij},$
$\cosh q_{j}=\sinh h_{i}\cdot\sinh d_{jk}=\sinh h_{k}\cdot\sinh d_{ji},$
$\cosh q_{k}=\sinh h_{i}\cdot\sinh d_{kj}=-\sinh h_{j}\cdot\sinh d_{ki},$
$\cosh h_{i}=\sinh q_{j}\cdot\sinh\theta_{ik}=\sinh q_{k}\cdot\sinh\theta_{ij},$
$\cosh h_{j}=\sinh q_{i}\cdot\sinh\theta_{jk}=-\sinh q_{k}\cdot\sinh\theta_{ji},$
$\cosh h_{k}=-\sinh q_{i}\cdot\sinh\theta_{kj}=\sinh q_{j}\cdot\sinh\theta_{ki}.$
Similarly, if $c_{ijk}$ lies in the domains $D_{\mathrm{i}},D_{\mathrm{ii}},D_{\mathrm{iii}},D_{\mathrm{iv}}$ or $D_{\mathrm{v}}$, we have the corresponding propositions. Using these propositions, we can prove Lemma 2.9 and further obtain Table 1. Moreover, from Table 1, we see that $c_{ijk}$ lies in the open domains $D_{\mathrm{i}},D_{\mathrm{ii}},D_{\mathrm{iii}},D_{\mathrm{iv}},D_{\mathrm{v}}$ or $D_{\mathrm{vi}}$ if and only if $q_{r}<0,h_{s}<0$ and the rest are positive. Hence, these propositions can also be written as a unified form.
Lemma 2.15.
Let $c_{ijk}\in\mathbb{R}^{3}$ lie in one of the open domains $D_{\mathrm{i}},D_{\mathrm{ii}},D_{\mathrm{iii}},D_{\mathrm{iv}},D_{\mathrm{v}},D_{\mathrm{vi}}$. If $q_{r}<0,h_{s}<0$, then
$\cosh q_{r}=\sinh h_{s}\cdot\sinh d_{rt}=-\sinh h_{t}\cdot\sinh d_{rs},$
$\cosh q_{s}=\sinh h_{r}\cdot\sinh d_{st}=\sinh h_{t}\cdot\sinh d_{sr},$
$\cosh q_{t}=\sinh h_{r}\cdot\sinh d_{ts}=-\sinh h_{s}\cdot\sinh d_{tr},$
$\cosh h_{r}=\sinh q_{s}\cdot\sinh\theta_{rt}=\sinh q_{t}\cdot\sinh\theta_{rs},$
$\cosh h_{s}=\sinh q_{r}\cdot\sinh\theta_{st}=-\sinh q_{t}\cdot\sinh\theta_{sr},$
$\cosh h_{t}=-\sinh q_{r}\cdot\sinh\theta_{ts}=\sinh q_{s}\cdot\sinh\theta_{tr}.$
As a direct corollary of Lemma 2.13 and Lemma 2.15, we have
Corollary 2.16.
If $c_{ijk}$ is space-like, then the signs of $q_{r}$, $d_{rt}$ and $d_{rs}$ are the same and the signs of $h_{r}$, $\theta_{rs}$ and $\theta_{rt}$ are the same.
Combining Corollary 2.11 and Corollary 2.16 gives the following lemma.
Lemma 2.17.
For any right-angled hyperbolic hexagon $\{ijk\}\in F$, the signs of $q_{r}$, $d_{rt}$ and $d_{rs}$ are the same and the signs of $h_{r}$, $\theta_{rs}$ and $\theta_{rt}$ are the same.
Proof.
If $c_{ijk}$ is time-like, the conclusion follows from Corollary 2.11. If $c_{ijk}$ is space-like, the conclusion follows from Corollary 2.16. If $c_{ijk}$ is light-like, the conclusion follows from continuity by moving the center to the light cone. Q.E.D.
2.3. Variational formulas of generalized angle
We divide Theorem 1.12 into Lemma 2.18, Lemma 2.20 and Lemma 2.22, and prove them respectively.
Lemma 2.18.
For a right-angled hyperbolic hexagon $\{ijk\}\in F$, we have
(15) $\frac{\partial\theta^{jk}_{i}}{\partial f_{j}}=\frac{-1}{\sinh d_{ji}}\frac{\tanh^{\beta}h_{ij}}{\sinh l_{ij}},$
where $\beta=1$ if $c_{ijk}$ is time-like and $\beta=-1$ if $c_{ijk}$ is space-like. If $c_{ijk}$ is light-like, we interpret the formula as $\tanh^{\beta}h_{ij}=\tanh^{\beta}h_{jk}=\tanh^{\beta}h_{ki}=1$.
Proof.
Without loss of generality, we assume $||v_{i}||=||v_{j}||=||v_{k}||=1$ and $||c_{ijk}||^{2}=\pm 1$. For simplicity, we set $l_{r}=l_{st},\ \theta_{r}=\theta^{st}_{r},\ h_{r}=h_{st},\ q_{r}=q_{st}$ and $A=\sinh l_{r}\sinh l_{s}\sinh\theta_{t}$, where $\{r,s,t\}=\{i,j,k\}$. By generalized hyperbolic cosine law, i.e., $\cosh l_{r}=\cosh\theta_{r}\sinh l_{s}\sinh l_{t}-\cosh l_{s}\cosh l_{t}$, we have
(16) $\frac{\partial\theta_{i}}{\partial l_{i}}=\frac{\sinh l_{i}}{A},\ \frac{\partial\theta_{i}}{\partial l_{j}}=-\frac{\sinh l_{i}\cosh\theta_{k}}{A},\ \frac{\partial\theta_{i}}{\partial l_{k}}=-\frac{\sinh l_{i}\cosh\theta_{j}}{A}.$
By (2) and (3), we have
(17) $\frac{\partial\theta_{i}}{\partial f_{j}}=$ $\frac{\partial\theta_{i}}{\partial l_{i}}\frac{\partial l_{i}}{\partial f_{j}}+\frac{\partial\theta_{i}}{\partial l_{j}}\frac{\partial l_{j}}{\partial f_{j}}+\frac{\partial\theta_{i}}{\partial l_{k}}\frac{\partial l_{k}}{\partial f_{j}}$
$=$ $\frac{\sinh l_{i}}{A}\cdot\coth d_{jk}+\frac{-\sinh l_{i}\cosh\theta_{j}}{A}\cdot\coth d_{ji}$
$=$ $\frac{1}{\sinh l_{k}\sinh\theta_{j}}\cdot\left(\coth d_{jk}-\cosh\theta_{j}\cdot\coth d_{ji}\right)$
$=$ $\frac{1}{\sinh l_{k}\sinh\theta_{j}}\cdot\left[\coth d_{jk}-\cosh(\theta_{ki}+\theta_{ik})\cdot\coth d_{ji}\right]$
$=$ $\frac{1}{\sinh l_{k}\sinh\theta_{j}}\cdot(\coth d_{jk}-\cosh\theta_{ki}\cosh\theta_{ik}\coth d_{ji}$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ -\sinh\theta_{ki}\sinh\theta_{ik}\coth d_{ji}),$
where (13) is used in the forth line. Set
(18) $L_{1}=\coth d_{jk}-\cosh\theta_{ki}\cosh\theta_{ik}\coth d_{ji},\ L_{2}=-\sinh\theta_{ki}\sinh\theta_{ik}\coth d_{ji}.$
Then
(19) $\frac{\partial\theta_{i}}{\partial f_{j}}=\frac{1}{\sinh l_{k}\sinh\theta_{j}}\cdot(L_{1}+L_{2}).$
To calculate $L_{1}$ and $L_{2}$, we first give some formulas. Please refer to Figure 3. Applying Proposition 2.1 to the generalized right-angled hyperbolic triangle $c_{ijk}c_{ij}v_{j}$ gives
(20) $v_{j}\ast c_{ijk}=-(c_{ij}\ast c_{ijk})(v_{j}\ast c_{ij})=\sinh d_{ji}\cdot(c_{ij}\ast c_{ijk}),$
where $v_{j}\ast c_{ij}=-\sinh d_{ji}$ follows from Proposition 2.2 (ii). Similarly, in the generalized right-angled hyperbolic triangle $c_{ijk}c_{jk}v_{j}$, we have
(21) $v_{j}\ast c_{ijk}=-(v_{j}\ast c_{jk})(c_{jk}\ast c_{ijk})=\sinh d_{jk}\cdot(c_{jk}\ast c_{ijk}),$
where $v_{j}\ast c_{jk}=-\sinh d_{jk}$ follows from Proposition 2.2 (ii). Combining (20) and (21) gives
(22) $\sinh d_{ji}\cdot(c_{ij}\ast c_{ijk})=\sinh d_{jk}\cdot(c_{jk}\ast c_{ijk}).$
By Proposition 2.2 $\mathrm{(iii)}$, we have $(c_{ijk}\otimes v_{j})\ast(v_{j}\otimes c_{jk})=||v_{j}\otimes c_{ijk}||\cdot||v_{j}\otimes c_{jk}||(-\cosh\theta_{ik}).$ Moreover,
$(c_{ijk}\otimes v_{j})\ast(v_{j}\otimes c_{jk})=(c_{jk}\ast c_{ijk})||v_{j}||^{2}-(v_{j}\ast c_{ijk})(v_{j}\ast c_{jk})=(c_{jk}\ast c_{ijk})\cdot\cosh^{2}d_{jk},$
where (21) is used in the second equality. Since $||v_{j}\otimes c_{jk}||^{2}=(v_{j}\ast c_{jk})^{2}+1=\cosh^{2}d_{jk}$, then
(23) $\cosh\theta_{ik}=-\frac{(c_{ijk}\otimes v_{j})\ast(v_{j}\otimes c_{jk})}{||v_{j}\otimes c_{ijk}||\cdot||v_{j}\otimes c_{jk}||}=-\frac{(c_{jk}\ast c_{ijk})\cdot\cosh d_{jk}}{||v_{j}\otimes c_{ijk}||}.$
Similarly, $(c_{ijk}\otimes v_{j})\ast(v_{j}\otimes c_{ij})=||v_{j}\otimes c_{ijk}||\cdot||v_{j}\otimes c_{ij}||(-\cosh\theta_{ki}).$ Moreover,
$(c_{ijk}\otimes v_{j})\ast(v_{j}\otimes c_{ij})=(c_{ij}\ast c_{ijk})||v_{j}||^{2}-(v_{j}\ast c_{ijk})(v_{j}\ast c_{ij})=(c_{ij}\ast c_{ijk})\cdot\cosh^{2}d_{ji},$
where (20) is used in the second equality. Since $||v_{j}\otimes c_{ij}||^{2}=(v_{j}\ast c_{ij})^{2}+1=\cosh^{2}d_{ji}$, then
(24) $\cosh\theta_{ki}=-\frac{(c_{ijk}\otimes v_{j})\ast(v_{j}\otimes c_{ij})}{||v_{j}\otimes c_{ijk}||\cdot||v_{j}\otimes c_{ij}||}=-\frac{(c_{ij}\ast c_{ijk})\cdot\cosh d_{ji}}{||v_{j}\otimes c_{ijk}||}.$
By the properties of Lorentzian cross product, we have
(25) $||v_{j}\otimes c_{ijk}||^{2}=(v_{j}\otimes c_{ijk})*(v_{j}\otimes c_{ijk})=(v_{j}\ast c_{ijk})^{2}-||c_{ijk}||^{2},$
and
(26) $||c_{ij}\otimes c_{ijk}||^{2}=(c_{ij}\otimes c_{ijk})*(c_{ij}\otimes c_{ijk})=(c_{ij}\ast c_{ijk})^{2}+||c_{ijk}||^{2}.$
Adding (25) and (26) gives
(27) $||v_{j}\otimes c_{ijk}||^{2}+||c_{ij}\otimes c_{ijk}||^{2}=$ $(v_{j}\ast c_{ijk})^{2}+(c_{ij}\ast c_{ijk})^{2}$
$=$ $(c_{ij}\ast c_{ijk})^{2}\cdot\sinh^{2}d_{ji}+(c_{ij}\ast c_{ijk})^{2}$
$=$ $(c_{ij}\ast c_{ijk})^{2}\cosh^{2}d_{ji},$
where (20) is used in the second line. Then
(28) $\sinh^{2}\theta_{ki}=$ $\cosh^{2}\theta_{ki}-1$
$=$ $\frac{1}{||v_{j}\otimes c_{ijk}||^{2}}[(c_{ij}\ast c_{ijk})^{2}\cosh^{2}d_{ji}-||v_{j}\otimes c_{ijk}||^{2}]$
$=$ $\frac{||c_{ij}\otimes c_{ijk}||^{2}}{||v_{j}\otimes c_{ijk}||^{2}},$
where (24) is used in the second line and (27) is used in the last line.
Now we begin to calculate $L_{1}$ and $L_{2}$ in (18).
$L_{2}=$ $-\sinh\theta_{ki}\sinh\theta_{ik}\cdot\coth d_{ji}$
$=$ $-\sinh\theta_{ki}\cdot\sinh\theta_{ik}\cdot\frac{\cosh d_{ji}}{\sinh d_{ji}}\cdot\frac{||v_{j}\otimes c_{ijk}||}{c_{ij}\ast c_{ijk}}\cdot\frac{c_{ij}\ast c_{ijk}}{||v_{j}\otimes c_{ijk}||}$
$=$ $\sinh\theta_{ki}\sinh\theta_{ik}\cdot\frac{||v_{j}\otimes c_{ijk}||}{(c_{ij}\ast c_{ijk})\sinh d_{ji}}\cdot\cosh\theta_{ki},$
where (24) is used in the last line.
$L_{1}=$ $\coth d_{jk}-\cosh\theta_{ki}\cosh\theta_{ik}\cdot\coth d_{ji}$
$=$ $\frac{\cosh d_{jk}\cdot(c_{jk}\ast c_{ijk})}{\sinh d_{ji}\cdot(c_{ij}\ast c_{ijk})}+\frac{(c_{ij}\ast c_{ijk})\cosh d_{ji}}{||v_{j}\otimes c_{ijk}||}\cdot\cosh\theta_{ik}\cdot\frac{\cosh d_{ji}}{\sinh d_{ji}}$
$=$ $\frac{\cosh\theta_{ik}}{\sinh d_{ji}\cdot(c_{ij}\ast c_{ijk})\cdot||v_{j}\otimes c_{ijk}||}\big{(}-||v_{j}\otimes c_{ijk}||^{2}+(c_{ij}\ast c_{ijk})^{2}\cosh^{2}d_{ji}\big{)}$
$=$ $\frac{\cosh\theta_{ik}\cdot||c_{ij}\otimes c_{ijk}||^{2}}{\sinh d_{ji}\cdot(c_{ij}\ast c_{ijk})\cdot||v_{j}\otimes c_{ijk}||}$
$=$ $\frac{\cosh\theta_{ik}\sinh^{2}\theta_{ki}||v_{j}\otimes c_{ijk}||}{\sinh d_{ji}(c_{ij}\ast c_{ijk})},$
where (22) and (24) are used in the second line, (23) is used in the third line, (27) is used in the forth line, and (28) is used in the last line. Therefore, the formula (19) can be written as
$\frac{\partial\theta_{i}}{\partial f_{j}}=$ $\frac{1}{\sinh l_{k}\sinh\theta_{j}}\cdot(L_{1}+L_{2})$
$=$ $\frac{1}{\sinh l_{k}\sinh\theta_{j}}\cdot\frac{\sinh\theta_{ki}\cdot||v_{j}\otimes c_{ijk}||}{(c_{ij}\ast c_{ijk})\sinh d_{ji}}\cdot(\cosh\theta_{ik}\sinh\theta_{ki}+\sinh\theta_{ik}\cosh\theta_{ki})$
$=$ $\frac{1}{\sinh l_{k}\sinh\theta_{j}}\cdot\frac{\sinh\theta_{ki}\cdot||v_{j}\otimes c_{ijk}||}{(c_{ij}\ast c_{ijk})\sinh d_{ji}}\cdot\sinh\theta_{j}$
$=$ $\frac{1}{\sinh l_{k}\sinh d_{ji}}\cdot\frac{\sinh\theta_{ki}\cdot||v_{j}\otimes c_{ijk}||}{c_{ij}\ast c_{ijk}}.$
(i): Suppose $c_{ijk}$ is time-like, then $||c_{ijk}||^{2}=-1$ and $v_{j}*c_{ijk}=-\sinh q_{j}$. By (25), we have $||v_{j}\otimes c_{ijk}||=\cosh q_{j}$. Then $\sinh\theta_{ki}\cdot||v_{j}\otimes c_{ijk}||=\sinh\theta_{ki}\cdot\cosh q_{j}=\sinh h_{k}$ by Lemma 2.10. Since $c_{ij}\ast c_{ijk}=-\cosh h_{k}$, then
$\frac{\partial\theta_{i}}{\partial f_{j}}=\frac{1}{\sinh d_{ji}\sinh l_{k}}\cdot\frac{\sinh h_{k}}{-\cosh h_{k}}=\frac{-1}{\sinh d_{ji}}\frac{\tanh h_{k}}{\sinh l_{k}}.$
(ii): Suppose $c_{ijk}$ is space-like, then $||c_{ijk}||^{2}=1$ and $v_{j}*c_{ijk}=\pm\cosh q_{j}$, thus $||v_{j}\otimes c_{ijk}||=\sinh q_{j}$ if $q_{j}>0$ and $||v_{j}\otimes c_{ijk}||=-\sinh q_{j}$ if $q_{j}<0$ by (25).
If $c_{ijk}$ is in one of the domains $D_{\mathrm{I}},D_{\mathrm{II}},D_{\mathrm{V}},D_{\mathrm{VI}}$, then $q_{j}>0$ and $h_{k}>0$ by Table 1. Then $\sinh\theta_{ki}\cdot||v_{j}\otimes c_{ijk}||=\sinh\theta_{ki}\cdot\sinh q_{j}=\cosh h_{k}$ by Lemma 2.13. And $c_{ij}\ast c_{ijk}=-\sinh h_{k}$.
If $c_{ijk}$ is in the domain $D_{\mathrm{III}}$, then $q_{j}<0$ and $h_{k}>0$ by Table 1. Then $\sinh\theta_{ki}\cdot||v_{j}\otimes c_{ijk}||=\sinh\theta_{ki}\cdot(-\sinh q_{j})=\cosh h_{k}$ by Lemma 2.13. And $c_{ij}\ast c_{ijk}=-\sinh h_{k}$.
If $c_{ijk}$ is in the domain $D_{\mathrm{IV}}$, then $q_{j}>0$ and $h_{k}<0$ by Table 1. Then $\sinh\theta_{ki}\cdot||v_{j}\otimes c_{ijk}||=\sinh\theta_{ki}\cdot\sinh q_{j}=-\cosh h_{k}$ by Lemma 2.13. Note that $c_{ij}\ast c_{ijk}=-\sinh(d_{\mathbb{H}}(c_{ij},(c_{ijk})^{\bot}))=-\sinh(-h_{k})=\sinh h_{k}$.
If $c_{ijk}$ is in one of the domains $D_{\mathrm{i}},D_{\mathrm{ii}},D_{\mathrm{vi}}$, then $q_{j}>0$ and $h_{k}>0$ by Table 1. Then $\sinh\theta_{ki}\cdot||v_{j}\otimes c_{ijk}||=\sinh\theta_{ki}\cdot\sinh q_{j}=\cosh h_{k}$ by Lemma 2.15. And $c_{ij}\ast c_{ijk}=-\sinh h_{k}$.
If $c_{ijk}$ is in the domain $D_{\mathrm{iii}}$, then $q_{j}<0$ and $h_{k}>0$ by Table 1. Then $\sinh\theta_{ki}\cdot||v_{j}\otimes c_{ijk}||=\sinh\theta_{ki}\cdot(-\sinh q_{j})=\cosh h_{k}$ Lemma 2.15. And $c_{ij}\ast c_{ijk}=-\sinh h_{k}$.
If $c_{ijk}$ is in the domain $D_{\mathrm{v}}$, then $q_{j}>0$ and $h_{k}<0$ by Table 1. Then $\sinh\theta_{ki}\cdot||v_{j}\otimes c_{ijk}||=\sinh\theta_{ki}\cdot\sinh q_{j}=-\cosh h_{k}$ by Lemma 2.15. And $c_{ij}\ast c_{ijk}=\sinh h_{k}$.
If $c_{ijk}$ is in the domain $D_{\mathrm{iv}}$, then $q_{j}<0$ and $h_{k}<0$ by Table 1. Then $\sinh\theta_{ki}\cdot||v_{j}\otimes c_{ijk}||=\sinh\theta_{ki}\cdot(-\sinh q_{j})=-\cosh h_{k}$ by Lemma 2.15. And $c_{ij}\ast c_{ijk}=\sinh h_{k}$.
In summary, if $c_{ijk}$ is space-like, then
$\frac{\partial\theta_{i}}{\partial f_{j}}=\frac{-1}{\sinh d_{ji}}\frac{\coth h_{k}}{\sinh l_{k}}.$
(iii): The case that $c_{ijk}$ is light-like follows from continuity by moving the center to the light cone. In fact, if $c_{ijk}$ is light-like, then $||c_{ijk}||^{2}=0$. Combining (28) and (26) gives
$\frac{\sinh^{2}\theta_{ki}\cdot||v_{j}\otimes c_{ijk}||^{2}}{(c_{ij}\ast c_{ijk})^{2}}=\frac{||c_{ij}\otimes c_{ijk}||^{2}}{(c_{ij}\ast c_{ijk})^{2}}=1.$
Since $||c_{ij}\otimes c_{ijk}||=-(c_{ij}\ast c_{ijk})$, then
$\frac{\partial\theta_{i}}{\partial f_{j}}=\frac{-1}{\sinh d_{ji}\sinh l_{k}}.$
This completes the proof. Q.E.D.
Remark 2.19.
Actually, the formulas in the proof of Lemma 2.18 are the sine and cosine laws of generalized right-angled hyperbolic triangles. In order to avoid discussing one by one, we use Lorentzian inner product and Lorentzian cross product to unify calculations.
Lemma 2.20.
For a right-angled hyperbolic hexagon $\{ijk\}\in F$, we have
(29) $\frac{\partial\theta^{jk}_{i}}{\partial f_{i}}=\cosh l_{ij}\frac{\partial\theta^{ik}_{j}}{\partial f_{i}}+\cosh l_{ik}\frac{\partial\theta^{ij}_{k}}{\partial f_{i}}.$
Proof.
Under the same notations as in the proof of Lemma 2.18, a routine calculation gives
$\frac{\partial\theta_{i}}{\partial f_{i}}=$ $\frac{\partial\theta_{i}}{\partial l_{i}}\frac{\partial l_{i}}{\partial f_{i}}+\frac{\partial\theta_{i}}{\partial l_{j}}\frac{\partial l_{j}}{\partial f_{i}}+\frac{\partial\theta_{i}}{\partial l_{k}}\frac{\partial l_{k}}{\partial f_{i}}$
$=$ $\frac{-\sinh l_{i}\cosh\theta_{k}}{A}\cdot\coth d_{ik}+\frac{-\sinh l_{i}\cosh\theta_{j}}{A}\cdot\coth d_{ij}$
$=$ $\frac{-\sinh l_{i}}{A}\cdot\left(\cosh\theta_{k}\cdot\coth d_{ik}+\cosh\theta_{j}\cdot\coth d_{ij}\right).$
Similar to (17), we have
$\frac{\partial\theta_{j}}{\partial f_{i}}=\frac{\sinh l_{j}}{A}(\coth d_{ik}-\cosh\theta_{i}\coth d_{ij}),$
$\frac{\partial\theta_{k}}{\partial f_{i}}=\frac{\sinh l_{k}}{A}(\coth d_{ij}-\cosh\theta_{i}\coth d_{ik}).$
Then
$\cosh l_{k}\frac{\partial\theta_{j}}{\partial f_{i}}+\cosh l_{j}\frac{\partial\theta_{k}}{\partial f_{i}}$
$=$ $\frac{1}{A}(\cosh l_{k}\sinh l_{j}\coth d_{ik}-\cosh l_{k}\sinh l_{j}\cosh\theta_{i}\coth d_{ij}$
$\ \ \ \ +\cosh l_{j}\sinh l_{k}\coth d_{ij}-\cosh l_{j}\sinh l_{k}\cosh\theta_{i}\coth d_{ik})$
$=$ $\frac{1}{A}[\coth d_{ik}\cdot(\cosh l_{k}\sinh l_{j}-\cosh l_{j}\sinh l_{k}\cdot\cosh\theta_{i})$
$\ \ \ \ +\coth d_{ij}\cdot(\cosh l_{j}\sinh l_{k}-\cosh l_{k}\sinh l_{j}\cosh\theta_{i})]$
$=$ $\frac{1}{A}[\coth d_{ik}\cdot(\cosh l_{k}\sinh l_{j}-\cosh l_{j}\sinh l_{k}\cdot\frac{\cosh l_{i}+\cosh l_{j}\cosh l_{k}}{\sinh l_{j}\sinh l_{k}})$
$\ \ \ \ +\coth d_{ij}\cdot(\cosh l_{j}\sinh l_{k}-\cosh l_{k}\sinh l_{j}\cdot\frac{\cosh l_{i}+\cosh l_{j}\cosh l_{k}}{\sinh l_{j}\sinh l_{k}})]$
$=$ $\frac{1}{A}(-\coth d_{ik}\sinh l_{i}\cosh\theta_{k}-\coth d_{ij}\sinh l_{i}\cosh\theta_{j})$
$=$ $\frac{\partial\theta_{i}}{\partial f_{i}}.$
Q.E.D.
Remark 2.21.
The formula (29) was first obtained by Glickenstein-Thomas [3] for generalized discrete conformal structures on closed surfaces, which has lots of applications. See [9, 10, 12, 13, 14, 15] and others for example. In [11], Xu showed that the formula (29) holds for a special class of discrete conformal structures on surfaces with boundary and conjectured this formula holds for generalized discrete conformal structures on surfaces with boundary. Lemma 2.20 provides an affirmative answer to Xu’s conjecture.
Since the variational formula of generalized angle (15) is not symmetric in $i$ and $j$, we need to obtain a symmetric variational formula of generalized angle by reparameterization.
Lemma 2.22.
There is a change of variables $u_{i}=u_{i}(f_{i})$ such that
$\frac{\partial\theta^{jk}_{i}}{\partial u_{j}}=\frac{\partial\theta^{ik}_{j}}{\partial u_{i}}.$
The function $u$ is also called a discrete conformal factor.
Proof.
Take new coordinates $u_{i}=u_{i}(f_{i})$ such that $\partial\theta^{jk}_{i}/\partial u_{j}=\partial\theta^{ik}_{j}/\partial u_{i}$. By (15), we have
$\frac{\partial\theta^{jk}_{i}}{\partial u_{j}}=\frac{\partial\theta^{jk}_{i}}{\partial f_{j}}\frac{\partial f_{j}}{\partial u_{j}}=\frac{-\partial f_{j}/\partial u_{j}}{\sinh d_{ji}}\cdot\frac{\tanh^{\beta}h_{ij}}{\sinh l_{ij}},$
and
$\frac{\partial\theta^{ik}_{j}}{\partial u_{i}}=\frac{-\partial f_{i}/\partial u_{i}}{\sinh d_{ij}}\cdot\frac{\tanh^{\beta}h_{ij}}{\sinh l_{ij}}.$
This implies
$\frac{\partial f_{i}/\partial u_{i}}{\partial f_{j}/\partial u_{j}}=\frac{\sinh d_{ij}}{\sinh d_{ji}}.$
Recall the following formulas from the proof of [16, Theorem 1.6]
$\frac{\sinh^{2}d_{ij}}{\sinh^{2}d_{ji}}=e^{C_{ij}}e^{2f_{i}-2f_{j}}\quad\text{or}\quad\frac{\sinh^{2}d_{ij}}{\sinh^{2}d_{ji}}=\frac{1+\alpha_{i}e^{2f_{i}}}{1+\alpha_{j}e^{2f_{j}}}.$
(i):
By Remark 1.7, we can set $e^{C_{ij}}=1$. If $\frac{\sinh d_{ij}}{\sinh d_{ji}}=e^{f_{i}-f_{j}}$, then
$\frac{\partial f_{i}}{\partial u_{i}}=e^{f_{i}}\quad\text{and}\ \quad\frac{\partial f_{j}}{\partial u_{j}}=e^{f_{j}}.$
Suppose $\frac{\sinh d_{ij}}{\sinh d_{ji}}=-e^{f_{i}-f_{j}}$. If $d_{ij}<0$, then $d_{ik}<0$ by Lemma 2.17, and thus $d_{ji}>0,\ d_{ki}>0$. We need $d_{jk}>0,\ d_{kj}>0$ to make the condition (1) hold. Hence $\frac{\sinh d_{ik}}{\sinh d_{ki}}=-e^{f_{i}-f_{k}}$ and $\frac{\sinh d_{jk}}{\sinh d_{ki}}=e^{f_{j}-f_{k}}$. Then
(30) $\frac{\partial f_{i}}{\partial u_{i}}=-e^{f_{i}},\ \quad\frac{\partial f_{j}}{\partial u_{j}}=e^{f_{j}}\quad\text{and}\ \quad\frac{\partial f_{k}}{\partial u_{k}}=e^{f_{k}}.$
Similarly, if $d_{ji}<0$, then $\frac{\partial f_{i}}{\partial u_{i}}=e^{f_{i}},\ \frac{\partial f_{j}}{\partial u_{j}}=-e^{f_{j}}$ and $\frac{\partial f_{k}}{\partial u_{k}}=e^{f_{k}}$.
(ii):
If $\frac{\sinh d_{ij}}{\sinh d_{ji}}=\sqrt{\frac{1+\alpha_{i}e^{2f_{i}}}{1+\alpha_{j}e^{2f_{j}}}}$, then
$\frac{\partial f_{i}}{\partial u_{i}}=\sqrt{1+\alpha_{i}e^{2f_{i}}}\quad\text{and}\ \quad\frac{\partial f_{j}}{\partial u_{j}}=\sqrt{1+\alpha_{j}e^{2f_{j}}}.$
By the arguments similar to that in the case $\mathrm{(i)}$, if $\frac{\sinh d_{ij}}{\sinh d_{ji}}=-\sqrt{\frac{1+\alpha_{i}e^{2f_{i}}}{1+\alpha_{j}e^{2f_{j}}}}$, then when $d_{ij}<0$,
$\frac{\partial f_{i}}{\partial u_{i}}=-\sqrt{1+\alpha_{i}e^{2f_{i}}}\quad\text{and}\ \quad\frac{\partial f_{j}}{\partial u_{j}}=\sqrt{1+\alpha_{j}e^{2f_{j}}};$
when $d_{ji}<0$,
$\frac{\partial f_{i}}{\partial u_{i}}=\sqrt{1+\alpha_{i}e^{2f_{i}}}\quad\text{and}\ \quad\frac{\partial f_{j}}{\partial u_{j}}=-\sqrt{1+\alpha_{j}e^{2f_{j}}}.$
Q.E.D.
Remark 2.23.
In fact, the function $u_{i}=u_{i}(f_{i})$ can be computed explicitly. If $\frac{\partial f_{i}}{\partial u_{i}}=e^{f_{i}}$, then $u_{i}=-e^{-f_{i}}$. If $\frac{\partial f_{i}}{\partial u_{i}}=-e^{f_{i}}$, then $u_{i}=e^{-f_{i}}$.
Suppose $\frac{\partial f_{i}}{\partial u_{i}}=\sqrt{1+\alpha_{i}e^{2f_{i}}}$, if $\alpha_{i}=0$, then $u_{i}=f_{i}\in\mathbb{R}$; if $\alpha_{i}\neq 0$, then
$u_{i}=\frac{1}{2}\ln\left|\frac{\sqrt{1+\alpha_{i}e^{2f_{i}}}-1}{\sqrt{1+\alpha_{i}e^{2f_{i}}}+1}\right|.$
Specially, if $\alpha_{i}=-1$, then
(31) $-\tanh u_{i}=\sqrt{1-e^{2f_{i}}},$
which implies $e^{f_{i}}=\frac{1}{\cosh u_{i}}$. Note that $\sqrt{1-e^{2f_{i}}}\neq 0$ and thus $f_{i}\neq 0$. Then $f_{i},u_{i}\in\mathbb{R}_{<0}$. If $\alpha_{i}=1$, then
(32) $-\coth u_{i}=\sqrt{1+e^{2f_{i}}},$
which implies $e^{f_{i}}=-\frac{1}{\sinh u_{i}}$ and $f_{i}\in\mathbb{R}$, $u_{i}\in\mathbb{R}_{<0}$.
Suppose $\frac{\partial f_{i}}{\partial u_{i}}=-\sqrt{1+\alpha_{i}e^{2f_{i}}}$, then if $\alpha_{i}=0$, then $u_{i}=-f_{i}\in\mathbb{R}$; if $\alpha_{i}\neq 0$, then
$u_{i}=-\frac{1}{2}\ln\left|\frac{\sqrt{1+\alpha_{i}e^{2f_{i}}}-1}{\sqrt{1+\alpha_{i}e^{2f_{i}}}+1}\right|.$
Specially, if $\alpha_{i}=-1$, then
$\tanh u_{i}=\sqrt{1-e^{2f_{i}}},$
which implies $e^{f_{i}}=\frac{1}{\cosh u_{i}}$. Similarly, the case that $f_{i}=0$ is removed. Then $f_{i}\in\mathbb{R}_{<0},\ u_{i}\in\mathbb{R}_{>0}$. If $\alpha_{i}=1$, then
$\coth u_{i}=\sqrt{1+e^{2f_{i}}},$
which implies $e^{f_{i}}=\frac{1}{\sinh u_{i}}$ and $f_{i}\in\mathbb{R},\ u_{i}\in\mathbb{R}_{>0}$.
3. Rigidity and existence of the discrete conformal structure (A1)
3.1. Admissible space of the discrete conformal structure (A1)
Suppose $(\Sigma,\mathcal{T},\alpha,\eta)$ is a weighted triangulated surface with boundary, and the weights $\alpha:B\rightarrow\{-1,0,1\}$ and $\eta\in\mathbb{R}^{E}_{>0}$. The admissible space of the discrete conformal factors $f$ for a right-angled hyperbolic hexagon $\{ijk\}\in F$ with edge lengths $l_{ij},l_{jk},l_{ki}$ given by (5) is defined to be
(33) $\Omega_{ijk}(\eta)=\{(f_{i},f_{j},f_{k})\in\mathbb{R}^{3}|$ $\cosh l_{rs}=-\sqrt{(1+\alpha_{r}e^{2f_{r}})(1+\alpha_{s}e^{2f_{s}})}$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +\eta_{rs}e^{f_{r}+f_{s}}>1,\ \forall\{r,s\}\subseteq\{i,j,k\}\}.$
By a change of variables $u_{i}=u_{i}(f_{i})$ in Lemma 2.22, the admissible space $\Omega_{ijk}(\eta)$ of $f$ is transferred to the following admissible space $\mathcal{U}_{ijk}(\eta)$ of the discrete conformal factors $u$ for a right-angled hyperbolic hexagon $\{ijk\}$
$\mathcal{U}_{ijk}(\eta)=u(\Omega_{ijk}(\eta))=\{(u_{i},u_{j},u_{k})\in\mathbb{R}^{N_{1}}\times\mathbb{R}_{<0}^{3-N_{1}}|(f_{i},f_{j},f_{k})\in\Omega_{ijk}(\eta)\},$
where $N_{1}$ is the number of the boundary components $i\in B$ with $\alpha_{i}=0$.
Since $\frac{\sinh d_{rs}}{\sinh d_{sr}}=\sqrt{\frac{1+\alpha_{r}e^{2f_{r}}}{1+\alpha_{s}e^{2f_{s}}}}>0$ in Theorem 1.6, then by the proof of Lemma 2.22, we have
(34) $\frac{\partial f_{r}}{\partial u_{r}}=\sqrt{1+\alpha_{r}e^{2f_{r}}}\quad\text{and}\ \quad\frac{\partial f_{s}}{\partial u_{s}}=\sqrt{1+\alpha_{s}e^{2f_{s}}}.$
According to Remark 2.23, we can describe the admissible space $\mathcal{U}_{ijk}(\eta)$ explicitly as follows:
(i):
If $\alpha\equiv 0$, then $\cosh l_{rs}=-1+\eta_{rs}e^{u_{r}+u_{s}}>1$, which implies $u_{r}+u_{s}>\log\frac{2}{\eta_{rs}}$.
(ii):
If $\alpha\equiv-1$, then $\cosh l_{rs}=-\sqrt{(1-e^{2f_{r}})(1-e^{2f_{s}})}+\eta_{rs}e^{f_{r}+f_{s}}=-\tanh u_{r}\tanh u_{s}+\eta_{rs}\frac{1}{\cosh u_{r}}\frac{1}{\cosh u_{s}}>1$, which implies $\cosh(u_{r}+u_{s})<\eta_{rs}$. Since $u\in\mathbb{R}^{3}_{<0}$, then $u_{r}+u_{s}>-\mathrm{arccosh}\ \eta_{rs}$. It is worth noting that $\eta_{rs}>1$.
(iii):
If $\alpha\equiv 1$, then $\cosh l_{rs}=-\sqrt{(1+e^{2f_{r}})(1+e^{2f_{s}})}+\eta_{rs}e^{f_{r}+f_{s}}=-\coth u_{r}\coth u_{s}+\eta_{rs}\frac{1}{\sinh u_{r}}\frac{1}{\sinh u_{s}}>1$, which implies $\cosh(u_{r}+u_{s})<\eta_{rs}$. Since $u\in\mathbb{R}^{3}_{<0}$, then $u_{r}+u_{s}>-\mathrm{arccosh}\ \eta_{rs}$. It is worth noting that $\eta_{rs}>1$.
(iv):
If $\alpha_{r}=0,\alpha_{s}=-1$, then $\cosh l_{rs}=-\sqrt{1-e^{2f_{s}}}+\eta_{rs}e^{f_{r}+f_{s}}=\tanh u_{s}+\eta_{rs}e^{u_{r}}\frac{1}{\cosh u_{s}}>1$, which implies $u_{r}+u_{s}>\log\frac{1}{\eta_{rs}}$.
(v):
If $\alpha_{r}=0,\alpha_{s}=1$, then $\cosh l_{rs}=-\sqrt{1+e^{2f_{s}}}+\eta_{rs}e^{f_{r}+f_{s}}=\coth u_{s}+\eta_{rs}e^{u_{r}}\frac{-1}{\sinh u_{s}}>1$. Since $u_{s}\in\mathbb{R}_{<0}$, then $u_{r}+u_{s}>\log\frac{1}{\eta_{rs}}$.
(vi):
If $\alpha_{r}=1,\alpha_{s}=-1$, then $\cosh l_{rs}=-\sqrt{(1+e^{2f_{r}})(1-e^{2f_{s}})}+\eta_{rs}e^{f_{r}+f_{s}}=-\coth u_{r}\tanh u_{s}+\eta_{rs}\frac{-1}{\sinh u_{r}}\frac{1}{\cosh u_{s}}>1$. Since $u_{r},u_{s}\in\mathbb{R}_{<0}$, then $\sinh(u_{r}+u_{s})>-\eta_{rs}$ and hence $u_{r}+u_{s}>\mathrm{arcsinh}\ (-\eta_{rs})$.
In summary, the admissible space $\mathcal{U}_{ijk}(\eta)$ of $u$ can be rewritten as
(35) $\mathcal{U}_{ijk}(\eta)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}^{N_{1}}\times\mathbb{R}_{<0}^{3-N_{1}}|u_{r}+u_{s}>C(\eta_{rs}),\ \forall\{r,s\}\subseteq\{i,j,k\}\},$
where $C(\eta_{rs})$ is a constant that depends on $\eta_{rs}$.
The above arguments can imply the following theorem.
Theorem 3.1.
Suppose $(\Sigma,\mathcal{T},\alpha,\eta)$ is a weighted triangulated surface with boundary, and the weights $\alpha:B\rightarrow\{-1,0,1\}$ and $\eta\in\mathbb{R}_{>0}^{E}$ satisfying $\eta_{ij}>\alpha_{i}\alpha_{j}$ if $\alpha_{i}=\alpha_{j}$ for two adjacent boundary components $i,j\in B$. Then the admissible space $\mathcal{U}_{ijk}(\eta)$ is a convex polytope. As a result, the admissible space
$\mathcal{U}(\eta)=\bigcap_{\{ijk\}\in F}\mathcal{U}_{ijk}(\eta)$
is a convex polytope.
3.2. Negative definiteness of Jacobian
For simplicity, we set $l_{r}=l_{st},\ \theta_{r}=\theta^{st}_{r}$ and $A=\sinh l_{r}\sinh l_{s}\sinh\theta_{t}$, where $\{r,s,t\}=\{i,j,k\}$. Define
(36) $Q_{1}=\left(\begin{array}[]{ccc}-1&\cosh\theta_{k}&\cosh\theta_{j}\\ \cosh\theta_{k}&-1&\cosh\theta_{i}\\ \cosh\theta_{j}&\cosh\theta_{i}&-1\\ \end{array}\right).$
Then $\det Q_{1}=-1+2\cosh\theta_{i}\cosh\theta_{j}\cosh\theta_{k}+\cosh^{2}\theta_{i}+\cosh^{2}\theta_{j}+\cosh^{2}\theta_{k}>0$. By calculations similar to (16), we have
$\frac{\partial(\theta_{i},\theta_{j},\theta_{k})}{\partial(l_{i},l_{j},l_{k})}=$ $-\frac{1}{A}\left(\begin{array}[]{ccc}\sinh l_{i}&0&0\\ 0&\sinh l_{j}&0\\ 0&0&\sinh l_{k}\\ \end{array}\right)\times Q_{1}.$
This implies
(37) $\det\bigg{(}\frac{\partial(\theta_{i},\theta_{j},\theta_{k})}{\partial(l_{i},l_{j},l_{k})}\bigg{)}=-\frac{1}{A^{3}}\sinh l_{i}\sinh l_{j}\sinh l_{k}\cdot\det Q_{1}<0.$
Combining (2) and (3), we have
(38) $\frac{\partial(l_{i},l_{j},l_{k})}{\partial(f_{i},f_{j},f_{k})}=\left(\begin{array}[]{ccc}0&\coth d_{jk}&\coth d_{kj}\\ \coth d_{ik}&0&\coth d_{ki}\\ \coth d_{ij}&\coth d_{ji}&0\\ \end{array}\right):=Q_{2}.$
Then
(39) $\det Q_{2}=\coth d_{jk}\coth d_{ki}\coth d_{ij}+\coth d_{kj}\coth d_{ik}\coth d_{ji}.$
For any $\alpha_{i}\in\{-1,0,1\}$, set $S_{i}=e^{f_{i}}$, $C_{i}=\sqrt{1+\alpha_{i}e^{2f_{i}}}$, then $S_{i},C_{i}>0$ and $C_{i}^{2}-\alpha_{i}S_{i}^{2}=1$. By (34), we have
(40) $\frac{\partial(f_{i},f_{j},f_{k})}{\partial(u_{i},u_{j},u_{k})}=\left(\begin{array}[]{ccc}C_{i}&0&0\\ 0&C_{j}&0\\ 0&0&C_{k}\\ \end{array}\right):=Q_{3}.$
Theorem 3.2.
Under the same assumptions as in Theorem 3.1. For a right-angled hyperbolic hexagon $\{ijk\}\in F$ on the admissible space $\mathcal{U}_{ijk}(\eta)$, the Jacobian $\Lambda_{ijk}=\frac{\partial(\theta^{jk}_{i},\theta^{ik}_{j},\theta^{ij}_{k})}{\partial(u_{i},u_{j},u_{k})}$ is non-degenerate, symmetric and negative definite.
Proof.
Since $d>0$ for the discrete conformal structure (A1) in Theorem 1.6, then $\det Q_{2}>0$ by (39). Hence
$\det\bigg{(}\frac{\partial(l_{i},l_{j},l_{k})}{\partial(u_{i},u_{j},u_{k})}\bigg{)}=\det Q_{2}\cdot\det Q_{3}>0.$
Combining with (37) gives
(41) $\det\Lambda_{ijk}=\det\bigg{(}\frac{\partial(\theta^{jk}_{i},\theta^{ik}_{j},\theta^{ij}_{k})}{\partial(l_{i},l_{j},l_{k})}\bigg{)}\cdot\det\bigg{(}\frac{\partial(l_{i},l_{j},l_{k})}{\partial(u_{i},u_{j},u_{k})}\bigg{)}<0.$
This implies the matrix $\Lambda_{ijk}$ is non-degenerate. By Lemma 2.22, the matrix $\Lambda_{ijk}$ is symmetric. Set $a_{rrs}:=C_{r}\cdot\coth d_{rs}>0$. Define
$Q:=Q_{1}Q_{2}Q_{3}=\left(\begin{array}[]{ccc}-1&\cosh\theta_{k}&\cosh\theta_{j}\\ \cosh\theta_{k}&-1&\cosh\theta_{i}\\ \cosh\theta_{j}&\cosh\theta_{i}&-1\\ \end{array}\right)\\ \times\left(\begin{array}[]{ccc}0&a_{jjk}&a_{kkj}\\ a_{iik}&0&a_{kki}\\ a_{iij}&a_{jji}&0\\ \end{array}\right).$
Then
$\Lambda_{ijk}=-\frac{1}{A}\left(\begin{array}[]{ccc}\sinh l_{i}&0&0\\ 0&\sinh l_{j}&0\\ 0&0&\sinh l_{k}\\ \end{array}\right)\times Q.$
To prove the negative definiteness of $\Lambda_{ijk}$, it suffices to prove $Q$ is positive definite. The formula (41) shows that $\det Q>0$. It remains to prove that the determinants of the $1\times 1$ and $2\times 2$ order principal minor of $Q$ are positive. Let $m_{rs}$ be the entry of $Q$ at $r$-th row and $s$-th column. Then
$m_{11}=\cosh\theta_{k}\cdot a_{iik}+\cosh\theta_{j}\cdot a_{iij}$
$m_{22}=\cosh\theta_{k}\cdot a_{jjk}+\cosh\theta_{i}\cdot a_{jji}$
$m_{12}=-a_{jjk}+\cosh\theta_{j}\cdot a_{jji}$
$m_{21}=-a_{iik}+\cosh\theta_{i}\cdot a_{iij}.$
Thus the determinants of the $1\times 1$ principal submatrices of $Q$ is $m_{11}>0$ by $a_{iik}>0,a_{iij}>0$. And the determinants of the $2\times 2$ principal submatrices of $Q$ is
$m_{11}m_{22}-m_{12}m_{21}=$ $\sinh^{2}\theta_{k}a_{iik}a_{jjk}+(\cosh\theta_{i}\cosh\theta_{k}+\cosh\theta_{j})a_{iik}a_{jji}$
$+(\cosh\theta_{j}\cosh\theta_{k}+\cosh\theta_{i})a_{iij}a_{jjk}>0.$
This completes the proof. Q.E.D.
As a result, we have the following corollary.
Corollary 3.3.
Under the same assumptions as in Theorem 3.1. The Jacobian $\Lambda=\frac{\partial(K_{i},...,K_{N})}{\partial(u_{i},...,u_{N})}$ is symmetric and negative definite on the admissible space $\mathcal{U}(\eta)$.
3.3. Rigidity of the discrete conformal structure (A1)
The following theorem gives the rigidity of discrete conformal structure (A1), which is the case $\mathrm{(i)}$ in Theorem 1.9.
Theorem 3.4.
Under the same assumptions as in Theorem 3.1. The function $f$ is determined by its generalized combinatorial curvature $K\in\mathbb{R}^{N}_{>0}$. In particular, the map from $f$ to $K$ is a smooth embedding.
Proof.
Theorem 3.1 and Theorem 3.2 imply that the following energy function
[…]
for a right-angled hyperbolic hexagon $\{ijk\}\in F$ is a well-defined smooth function on $\mathcal{U}_{ijk}(\eta)$. Furthermore, $\mathcal{F}_{ijk}(u_{i},u_{j},u_{k})$ is a strictly concave function on $\mathcal{U}_{ijk}(\eta)$ with $\nabla_{u_{i}}\mathcal{F}_{ijk}=\theta^{jk}_{i}$. Define the function ~$\widetilde{\mathcal{F}}:\mathcal{U}(\eta)\rightarrow\mathbb{R}$ by
~$\widetilde{\mathcal{F}}(u)=\sum_{\{ijk\}\in F}\mathcal{F}_{ijk}(u_{i},u_{j},u_{k}).$
By Corollary 3.3, ~$\widetilde{\mathcal{F}}(u)$ is a strictly concave function on $\mathcal{U}(\eta)$ with $\nabla_{u_{i}}\widetilde{\mathcal{F}}(u)=K_{i}$. According to Lemma 1.11, the map ~$\nabla\widetilde{\mathcal{F}}:\mathcal{U}(\eta)\rightarrow\mathbb{R}^{N}$ is a smooth embedding. Therefore, the map from $\{f\in\Omega(\eta)=\bigcap_{\{ijk\}\in F}\Omega_{ijk}(\eta)\}$ to $\{K\in\mathbb{R}^{N}_{>0}\}$ is a smooth injective map. Q.E.D.
3.4. Existence of discrete conformal structure (A1)
The following theorem characterizes the image of $K$.
Theorem 3.5.
Under the same assumptions as in Theorem 3.1. Suppose $\alpha:B\rightarrow\{0,1\}$. Then the image of $K$ is >$\mathbb{R}^{N}_{>0}$.
To prove Theorem 3.5, we need the following two lemmas.
Lemma 3.6.
Under the same assumptions as in Theorem 3.1. For a right-angled hyperbolic hexagon $\{ijk\}\in F$ on $\Omega_{ijk}(\eta)$, if one of the following conditions is satisfied
(i):
$\lim f_{i}=+\infty,\ \lim f_{j}=+\infty,\ \lim f_{k}=+\infty$;
(ii):
$\lim f_{i}=+\infty,\ \lim f_{j}=+\infty,\ \lim f_{k}=c$;
(iii):
$\lim f_{i}=+\infty,\ \lim f_{j}=a,\ \lim f_{k}=b$.
Then $\lim\theta^{jk}_{i}(f_{i},f_{j},f_{k})=0$, where $a,\ b,\ c$ are constants.
Proof.
For case $\mathrm{(i)}$, $\lim f_{i}=+\infty,\ \lim f_{j}=+\infty,\ \lim f_{k}=+\infty$. By (5), we have
(42) $\cosh l_{ij}=$ $\bigg{[}\eta_{ij}-\sqrt{(e^{-2f_{i}}+\alpha_{i})(e^{-2f_{j}}+\alpha_{j})}\bigg{]}e^{f_{i}+f_{j}},$
$\cosh l_{jk}=$ $\bigg{[}\eta_{jk}-\sqrt{(e^{-2f_{j}}+\alpha_{j})(e^{-2f_{k}}+\alpha_{k})}\bigg{]}e^{f_{j}+f_{k}},$
$\cosh l_{ik}=$ $\bigg{[}\eta_{ik}-\sqrt{(e^{-2f_{i}}+\alpha_{i})(e^{-2f_{k}}+\alpha_{k})}\bigg{]}e^{f_{i}+f_{k}}.$
Then $\lim\cosh l_{ij}:=\lim c_{k}e^{f_{i}+f_{j}}$$\lim\cosh l_{jk}:=\lim c_{i}e^{f_{j}+f_{k}}$ and $\lim\cosh l_{ik}:=\lim c_{j}e^{f_{i}+f_{k}}$, where $c_{i},c_{j},c_{k}$ are positive constants. Set $L_{1}=\frac{\cosh l_{jk}}{\sinh l_{ij}\sinh l_{ik}}$ and $L_{2}=\frac{\cosh l_{ij}\cosh l_{ik}}{\sinh l_{ij}\sinh l_{ik}}$, then $\cosh\theta^{jk}_{i}=L_{1}+L_{2}$. Since
$L_{1}\rightarrow\frac{c_{i}e^{f_{j}+f_{k}}}{c_{k}e^{f_{i}+f_{j}}\cdot c_{j}e^{f_{i}+f_{k}}}\rightarrow 0\quad\ \text{and}\ \quad L_{2}\rightarrow 1,$
then $\cosh\theta^{jk}_{i}\rightarrow 1$ and hence $\theta_{i}^{jk}\rightarrow 0$. The same proofs are suitable for the cases $\mathrm{(ii)}$ and $\mathrm{(iii)}$, and we omit them here. Q.E.D.
Lemma 3.7.
Under the same assumptions as in Theorem 3.1. For a right-angled hyperbolic hexagon $\{ijk\}\in F$ on $\Omega_{ijk}(\eta)$, if $\lim f_{i}=-\infty$, then $\lim f_{j}=\lim f_{k}=+\infty$. Furthermore, $\lim\theta^{jk}_{i}(f_{i},f_{j},f_{k})=+\infty$.
Proof.
According to the definition of the admissible space $\mathcal{U}_{ijk}(\eta)$ by (35), there is no subset $\{r,s\}\subseteq\{i,j,k\}$ such that $\lim u_{r}=-\infty$ and $\lim u_{s}=-\infty$ or $\lim u_{s}=c$, where $c$ is a constant. Since $\frac{\partial f_{r}}{\partial u_{r}}=\sqrt{1+\alpha_{r}e^{2f_{r}}}$ for $r\in\{i,j,k\}$, by Remark 2.23, if $\lim f_{i}=-\infty$, then $\lim u_{i}=-\infty$; if $\lim f_{j}=+\infty$, then $\lim u_{j}=+\infty$ with $\alpha_{j}=0$ and $\lim u_{j}=0^{-}$ with $\alpha_{j}=1$. Therefore, if $\lim f_{i}=-\infty$, then $\lim u_{i}=-\infty$ and hence $\lim u_{j}=\lim u_{k}=+\infty$. This implies $\alpha_{j}=\alpha_{k}=0$ and $\lim f_{j}=\lim f_{k}=+\infty$.
If $\alpha_{i}=0$, then $\cosh l_{ij}=-1+\eta_{ij}e^{u_{i}+u_{j}}$$\cosh l_{jk}=-1+\eta_{jk}e^{u_{j}+u_{k}}$ and $\cosh l_{ik}=-1+\eta_{ik}e^{u_{i}+u_{k}}$. Thus
$L_{1}=$ $\frac{-1+\eta_{jk}e^{u_{j}+u_{k}}}{\sqrt{(-1+\eta_{ij}e^{u_{i}+u_{j}})^{2}-1}\cdot\sqrt{(-1+\eta_{ik}e^{u_{i}+u_{k}})^{2}-1}}$
$=$ $\frac{-1+\eta_{jk}e^{u_{j}+u_{k}}}{\sqrt{\eta_{ij}\eta_{ik}}e^{u_{j}+u_{k}}}\cdot\frac{1}{\sqrt{\eta_{ij}e^{2u_{i}}-2e^{u_{i}-u_{j}}}}\cdot\frac{1}{\sqrt{\eta_{ik}e^{2u_{i}}-2e^{u_{i}-u_{k}}}}\rightarrow+\infty.$
And $L_{2}>0$. Then $\cosh\theta^{jk}_{i}=L_{1}+L_{2}\rightarrow+\infty$, which implies $\theta^{jk}_{i}\rightarrow+\infty$.
If $\alpha_{i}=1$, then $\sqrt{1+e^{2f_{i}}}=-\coth u_{i}$ and $e^{f_{i}}=-\frac{1}{\sinh u_{i}}$. Hence,
$\cosh l_{ij}=-\sqrt{1+e^{2f_{i}}}+\eta_{ij}e^{f_{i}+f_{j}}=\frac{-\cosh u_{i}+\eta_{ij}e^{u_{j}}}{-\sinh u_{i}}\rightarrow-1+2\eta_{ij}e^{u_{i}+u_{j}}.$
Similarly, $\cosh l_{ik}\rightarrow-1+2\eta_{ik}e^{u_{i}+u_{k}}$ and $\cosh l_{jk}=-1+\eta_{jk}e^{u_{j}+u_{k}}$. Similar to the proof of the case $\alpha_{i}=0$, we have $\cosh\theta^{jk}_{i}=L_{1}+L_{2}\rightarrow+\infty$, then $\theta^{jk}_{i}\rightarrow+\infty$.
If $\alpha_{i}=-1$, then $\sqrt{1-e^{2f_{i}}}=-\tanh u_{i}$ and $e^{f_{i}}=\frac{1}{\cosh u_{i}}$. Hence,
$\cosh l_{ij}=-\sqrt{1-e^{2f_{i}}}+\eta_{ij}e^{f_{i}+f_{j}}=\frac{\sinh u_{i}+\eta_{ij}e^{u_{j}}}{\cosh u_{i}}\rightarrow-1+2\eta_{ij}e^{u_{i}+u_{j}}.$
Similarly, $\cosh l_{ik}\rightarrow-1+2\eta_{ik}e^{u_{i}+u_{k}}$ and $\cosh l_{jk}=-1+\eta_{jk}e^{u_{j}+u_{k}}$. Similar to the proof of the case $\alpha_{i}=0$, we have $\cosh\theta^{jk}_{i}=L_{1}+L_{2}\rightarrow+\infty$, then $\theta^{jk}_{i}\rightarrow+\infty$. Q.E.D.
Proof of Theorem 3.5: Let the image of $K$ be $X$. To prove $X=\mathbb{R}^{N}_{>0}$, we will show that $X$ is both open and closed in >$\mathbb{R}^{N}_{>0}$. By the definition of curvature map $K$, we have $X\subseteq\mathbb{R}^{N}_{>0}$. Since the map from $f\in\Omega(\eta)$ to >$\mathbb{R}^{N}_{>0}$ is a smooth injective map, then $X$ is open in >$\mathbb{R}^{N}_{>0}$. It remains to prove that $X$ is closed in >$\mathbb{R}^{N}_{>0}$. We show that if a sequence $f^{(m)}$ in $\Omega(\eta)$ satisfies $\lim_{m\rightarrow+\infty}K^{(m)}\in\mathbb{R}^{N}_{>0}$, then there exists a subsequence, say $f^{(m)}$, such that $\lim_{m\rightarrow+\infty}f^{(m)}$ is in $\Omega(\eta)$.
Suppose otherwise, there exists a subsequence $f^{(m)}$, such that $\lim_{m\rightarrow+\infty}f^{(m)}$ is in the boundary of $\Omega(\eta)$. By (33), the boundary of the admissible space $\Omega(\eta)$ in $[-\infty,+\infty]^{N}$ consists of the following three parts
(i):
There exists an ideal edge $\{rs\}\in E$ such that $l_{rs}=0$, i.e., $u_{r}+u_{s}=-M(\eta_{rs})$ for the admissible space $\mathcal{U}_{ijk}(\eta)$ defined by (35).
(ii):
For $\alpha\in\{0,1\}$, there exists $i\in B$ satisfying $f_{i}^{(m)}\rightarrow+\infty$.
(iii):
For $\alpha\in\{0,1\}$, there exists $i\in B$ satisfying $f_{i}^{(m)}\rightarrow-\infty$.
For case $\mathrm{(i)}$, in a right-angled hyperbolic hexagon $\{ijk\}\in F$ with edge lengths $l_{ij},l_{jk},l_{ik}$ and opposite lengths of boundaries $\theta^{ij}_{k},\theta^{jk}_{i},\theta^{ik}_{j}$. Without loss of generality, we assume $l_{ij}\rightarrow 0$. By the cosine law,
$\cosh\theta^{ik}_{j}=\frac{\cosh l_{ik}+\cosh l_{ij}\cosh l_{jk}}{\sinh l_{ij}\sinh l_{jk}}>\frac{\cosh l_{ij}\cosh l_{jk}}{\sinh l_{ij}\sinh l_{jk}}\geq\frac{\cosh l_{ij}}{\sinh l_{ij}}\rightarrow+\infty,$
which contradicts the assumption that $\lim_{m\rightarrow+\infty}K^{(m)}\in\mathbb{R}^{N}_{>0}$.
The case $\mathrm{(ii)}$ follows from Lemma 3.6. In fact, each generalized angle incident to the boundary component $i\in B$ converges to 0. Hence $\lim_{m\rightarrow+\infty}K^{(m)}=0$. This contradicts with the assumption that $\lim_{m\rightarrow+\infty}K^{(m)}\in\mathbb{R}^{N}_{>0}$.
The case $\mathrm{(iii)}$ follows from Lemma 3.7. In fact, each generalized angle incident to the boundary component $i\in B$ converges to +$\infty$+ ∞. Hence $\lim_{m\rightarrow+\infty}K^{(m)}=+\infty$. This contradicts with the assumption that $\lim_{m\rightarrow+\infty}K^{(m)}\in\mathbb{R}^{N}_{>0}$. Q.E.D.
Remark 3.8.
Note that Lemma 3.6 and Lemma 3.7 hold for $\alpha:B\rightarrow\{-1,0,1\}$. But Theorem 3.5 holds for $\alpha:B\rightarrow\{0,1\}$. If $\alpha_{i}=-1$, then the boundary of the admissible space $\Omega(\eta)$ in $[-\infty,+\infty]^{N}$ will include the case that there exists $i\in B$ satisfying $f_{i}^{(m)}\rightarrow 0^{-}$. We are unable to prove the closeness of $X$ in this case.
4. Rigidity and existence of the discrete conformal structure (A2)
By Remark 1.7, we can set $C\equiv 0$ in the discrete conformal structure (A2).By Remark 1.7, we can set $C\equiv 0$ in the discrete conformal structure (A2). In this case, the discrete conformal structure (A2) is the $(-1,-1,0)$ type generalized circle packings in Guo-Luo [5]. Its rigidity and existence have been proved by Guo-Luo [5], which is as follows.
Theorem 4.1.
([5]) Suppose $(\Sigma,\mathcal{T},\eta)$ is a weighted triangulated surface with boundary and the weight $\eta\in\mathbb{R}_{>0}^{E}$, the discrete conformal structure $f\in\Omega(\eta)$ is determined by its generalized combinatorial curvature $K\in\mathbb{R}^{N}_{>0}$. Furthermore, the image of $K$ is $\mathbb{R}_{>0}^{N}$.
In this section, we reprove the rigidity part of Theorem 4.1. In [5], Guo-Luo spent lots of space to prove the symmetry and negative definiteness of Jacobian. By (2) and (3), we simplify Guo-Luo’s proof, which is similar to that of Theorem 3.2.
The admissible space of the discrete conformal factor $f$ for a right-angled hyperbolic hexagon $\{ijk\}\in F$ with edge lengths $l_{ij},l_{jk},l_{ki}$ given by (7) is defined to be
$\Omega_{ijk}(\eta)=\{(f_{i},f_{j},f_{k})\in\mathbb{R}^{3}|\cosh l_{rs}=-\cosh(f_{r}-f_{s})+\eta_{rs}e^{f_{r}+f_{s}}>1,\ \forall\{r,s\}\subseteq\{i,j,k\}\}.$
By Theorem 1.6, we have $d>0$ for the discrete conformal structure (A2). By a change of variables in Remark 2.23, i.e., $\frac{df_{r}}{du_{r}}=e^{f_{r}}$ and $u_{r}=-e^{-f_{r}}$. Thus
$\eta_{rs}=\frac{\cosh l_{rs}+\cosh(f_{r}-f_{s})}{e^{f_{r}+f_{s}}}>\frac{1+\frac{1}{2}(e^{f_{r}-f_{s}}+e^{f_{s}-f_{r}})}{e^{f_{r}+f_{s}}}=\frac{1}{2}(e^{-f_{r}}+e^{-f_{s}})^{2}=\frac{1}{2}(u_{r}+u_{s})^{2}.$
Since $u_{r},u_{s}\in\mathbb{R}_{<0}$, then $\eta_{rs}>0$ and $u_{r}+u_{s}>-\sqrt{2\eta_{rs}}$. Therefore, the admissible space $\Omega_{ijk}(\eta)$ of $f$ is transferred to the following admissible space $\mathcal{U}_{ijk}(\eta)$ of $u$ for a right-angled hyperbolic hexagon $\{ijk\}\in F$,
$\mathcal{U}_{ijk}(\eta)=u(\Omega_{ijk}(\eta))=\{(u_{i},u_{j},u_{k})\in\mathbb{R}_{<0}^{3}|u_{r}+u_{s}>-\sqrt{2\eta_{rs}},\ \forall\{r,s\}\subseteq\{i,j,k\}\},$
which is a convex polytope. Then the admissible space of $u$ on a weighted triangulated surface $(\Sigma,\mathcal{T},\eta)$ can be defined as
$\mathcal{U}(\eta)=\bigcap_{\{ijk\}\in F}\mathcal{U}_{ijk}(\eta),$
which is also a convex polytope.
Theorem 4.2.
Suppose $(\Sigma,\mathcal{T},\eta)$ is a weighted triangulated surface with boundary and the weight $\eta\in\mathbb{R}_{>0}^{E}$. For a right-angled hyperbolic hexagon $\{ijk\}\in F$ on $\mathcal{U}_{ijk}(\eta)$, the Jacobian $\Lambda_{ijk}=\frac{\partial(\theta^{jk}_{i},\theta^{ik}_{j},\theta^{ij}_{k})}{\partial(u_{i},u_{j},u_{k})}$ is non-degenerate, symmetric and negative definite. As a result, the Jacobian $\Lambda=\frac{\partial(K_{i},...,K_{N})}{\partial(u_{i},...,u_{N})}$ is symmetric and negative definite on $\mathcal{U}(\eta)$.
Proof.
For simplicity, we set $l_{r}=l_{st},\ \theta_{r}=\theta^{st}_{r}$ and $A=\sinh l_{r}\sinh l_{s}\sinh\theta_{t}$, where $\{r,s,t\}=\{i,j,k\}$. Since $\frac{df_{r}}{du_{r}}=e^{f_{r}}$, then
$\frac{\partial(f_{i},f_{j},f_{k})}{\partial(u_{i},u_{j},u_{k})}=\left(\begin{array}[]{ccc}e^{f_{i}}&0&0\\ 0&e^{f_{j}}&0\\ 0&0&e^{f_{k}}\\ \end{array}\right):=Q_{4}.$
Hence
$\Lambda_{ijk}=\frac{\partial(\theta_{i},\theta_{j},\theta_{k})}{\partial(u_{i},u_{j},u_{k})}=-\frac{1}{A}\left(\begin{array}[]{ccc}\sinh l_{i}&0&0\\ 0&\sinh l_{j}&0\\ 0&0&\sinh l_{k}\\ \end{array}\right)Q_{1}Q_{2}Q_{4},$
where $Q_{1}$ is defined by (36) and $Q_{2}$ is defined by (38). Moreover, $\det Q_{1}>0$.
Since $d>0$ for the discrete conformal structure (A2) in Theorem 1.6, then $\det Q_{2}>0$ by (39). Hence,
$\det\Lambda_{ijk}=-\frac{1}{A^{3}}\sinh l_{i}\sinh l_{j}\sinh l_{k}\cdot\det Q_{1}\cdot\det Q_{2}\cdot\det Q_{4}<0,$
which implies $\Lambda_{ijk}$ is non-degenerate. The symmetry of $\Lambda_{ijk}$ follows from Lemma 2.22. Set $a_{rrs}=e^{f_{r}}\cdot\coth d_{rs}>0$. Define
$Q:=Q_{1}Q_{2}Q_{4}=\left(\begin{array}[]{ccc}-1&\cosh\theta_{k}&\cosh\theta_{j}\\ \cosh\theta_{k}&-1&\cosh\theta_{i}\\ \cosh\theta_{j}&\cosh\theta_{i}&-1\\ \end{array}\right)\\ \times\left(\begin{array}[]{ccc}0&a_{jjk}&a_{kkj}\\ a_{iik}&0&a_{kki}\\ a_{iij}&a_{jji}&0\\ \end{array}\right).$
The rest of the proof is similar to that of Theorem 3.2, so we omit it. Q.E.D.
The following theorem gives the rigidity of discrete conformal structure (A2), which is the case $\mathrm{(ii)}$ in Theorem 1.9. As the proof is similar to that of Theorem 3.4, we omit it here.
Theorem 4.3.
Suppose $(\Sigma,\mathcal{T},\eta)$ is a weighted triangulated surface with boundary and the weight $\eta\in\mathbb{R}_{>0}^{E}$. For the discrete conformal structure (A2), the function $f$ is determined by its generalized combinatorial curvature $K\in\mathbb{R}^{E}_{>0}$. In particular, the map from $f$ to $K$ is a smooth embedding.
5. Rigidity and existence of the mixed discrete conformal structure I
In this section, we consider the rigidity and existence for the mixed discrete conformal structure $\mathrm{I}$, i.e., the discrete hyperbolic metrics $l_{ij},l_{jk},l_{ki}$ of a right-angled hyperbolic hexagon $\{ijk\}\in F$ are induced by (7) and (10). By Remark 1.7, we can set $C\equiv 0$ for (7) and (10). As mentioned in Remark 1.8, if $d_{rs}<0$, then $l_{rs},l_{rt}$ are given by (10) and $l_{st}$ is given by (7), where $\{r,s,t\}=\{i,j,k\}$. Without loss of generality, we can assume $d_{ij}<0$ and $d_{ik}<0$. Then for a right-angled hyperbolic hexagon $\{ijk\}\in F$, the edge lengths $l_{ij},l_{ik}$ are given by (10) and the edge length $l_{jk}$ is given by (7), i.e.,
(43) $\cosh l_{ij}$ $=\cosh(f_{j}-f_{i})+\zeta_{ij}e^{f_{i}+f_{j}},$
$\cosh l_{ik}$ $=\cosh(f_{k}-f_{i})+\zeta_{ik}e^{f_{i}+f_{k}},$
$\cosh l_{jk}$ $=-\cosh(f_{j}-f_{k})+\eta_{jk}e^{f_{j}+f_{k}}.$
Here $\eta_{ij}\in\mathbb{R}_{>0}$ and $\zeta_{ij},\zeta_{ik}\in\mathbb{R}$.
5.1. Rigidity of the mixed discrete conformal structure $\mathrm{I}$
The admissible space of the discrete conformal factor $f$ for a right-angled hyperbolic hexagon $\{ijk\}\in F$ with edge lengths given by (43) is defined to be
$\Omega_{ijk}(\zeta_{ij},\eta_{jk},\zeta_{ik})=\{(f_{i},f_{j},f_{k})\in\mathbb{R}^{3}|l_{ij}>0,\ l_{jk}>0,\ l_{ik}>0\}.$
Combining (30) and Remark 2.23 gives $u_{i}=e^{-f_{i}}\in\mathbb{R}_{>0}$, $u_{j}=-e^{-f_{j}}\in\mathbb{R}_{<0}$ and $u_{k}=-e^{-f_{k}}\in\mathbb{R}_{<0}$. Thus
$\eta_{jk}=\frac{\cosh l_{jk}+\cosh(f_{j}-f_{k})}{e^{f_{j}+f_{k}}}>\frac{1+\frac{1}{2}(e^{f_{j}-f_{k}}+e^{f_{k}-f_{j}})}{e^{f_{j}+f_{k}}}=\frac{1}{2}(e^{-f_{j}}+e^{-f_{k}})^{2}=\frac{1}{2}(u_{j}+u_{k})^{2}.$
Since $u_{j},u_{k}\in\mathbb{R}_{<0}$, then $\eta_{jk}>0$ and $u_{j}+u_{k}>-\sqrt{2\eta_{jk}}$. Similarly,
$\zeta_{ij}=\frac{\cosh l_{ij}-\cosh(f_{j}-f_{i})}{e^{f_{i}+f_{j}}}>\frac{1-\frac{1}{2}(e^{f_{j}-f_{i}}+e^{f_{i}-f_{j}})}{e^{f_{i}+f_{j}}}=-\frac{1}{2}(e^{-f_{i}}-e^{-f_{j}})^{2}=-\frac{1}{2}(u_{i}+u_{j})^{2},$
which implies $(u_{i}+u_{j})^{2}>-2\zeta_{ij}$. Moreover, if $\zeta_{ij}>0$, then $u_{i}+u_{j}\in\mathbb{R}$; if $\zeta_{ij}<0$, then $u_{i}+u_{j}>\sqrt{-2\zeta_{ij}}$ or $u_{i}+u_{j}<-\sqrt{-2\zeta_{ij}}$. To remove the solution $u_{i}+u_{j}<-\sqrt{-2\zeta_{ij}}$, we add the condition that $\eta_{jk}+\zeta_{ij}\leq 0$. In fact, $u_{i}+u_{j}<-\sqrt{-2\zeta_{ij}}\leq-\sqrt{2\eta_{jk}}<u_{j}+u_{k}$, which contradicts with the fact $u_{i}>0$ and $u_{k}<0$. Similar arguments are also suitable for $\zeta_{ik}$. Hence, we have the following theorem, which characterizes the admissible space of $u$.
Theorem 5.1.
Suppose $(\Sigma,\mathcal{T},\zeta,\eta)$ is a weighted triangulated surface with boundary, and the weights $\zeta\in\mathbb{R}^{E}$ and $\eta\in\mathbb{R}_{>0}^{E}$. Let $\{ijk\}\in F$ be a right-angled hyperbolic hexagon with edge lengths given by (43). If one of the following conditions is satisfied
(i):
$\zeta_{ij}>0$ and $\zeta_{ik}>0$;
(ii):
$\zeta_{ij}>0$, $\zeta_{ik}<0$ and $\eta_{jk}+\zeta_{ik}\leq 0$;
(iii):
$\zeta_{ij}<0$, $\zeta_{ik}>0$ and $\eta_{jk}+\zeta_{ij}\leq 0$;
(iv):
$\zeta_{ij}<0$, $\zeta_{ik}<0$, $\eta_{jk}+\zeta_{ij}\leq 0$ and $\eta_{jk}+\zeta_{ij}\leq 0$.
Then the admissible space
(44) $\mathcal{U}_{ijk}(\zeta_{ij},\eta_{jk},\zeta_{ik})=$ $u(\Omega_{ijk}(\zeta_{ij},\eta_{jk},\zeta_{ik}))$
$=$ $\{(u_{i},u_{j},u_{k})\in\mathbb{R}_{>0}\times\mathbb{R}_{<0}^{2}|u_{j}+u_{k}>-\sqrt{2\eta_{jk}},$
$\ \ (u_{i}+u_{j})^{2}>-2\zeta_{ij},(u_{i}+u_{k})^{2}>-2\zeta_{ik}\}$
of $u$ is a convex polytope. As a result, the admissible space
$\mathcal{U}(\zeta,\eta)=\bigcap_{\{ijk\}\in F}\mathcal{U}_{ijk}(\zeta_{ij},\eta_{jk},\zeta_{ik})$
is also a convex polytope.
Proof.
According to the arguments above Theorem 5.1, we can describe the admissible space $\mathcal{U}_{ijk}(\zeta_{ij},\eta_{jk},\zeta_{ik})$ explicitly. For the case $\mathrm{(i)}$, the admissible space (44) is reduced to
$\mathcal{U}_{ijk}(\zeta_{ij},\eta_{jk},\zeta_{ik})=\{(u_{i},u_{j},u_{k})\in\mathbb{R}_{>0}\times\mathbb{R}_{<0}^{2}|u_{j}+u_{k}>-\sqrt{2\eta_{jk}}\}.$
For the case $\mathrm{(ii)}$, the admissible space (44) is reduced to
$\mathcal{U}_{ijk}(\zeta_{ij},\eta_{jk},\zeta_{ik})=\{(u_{i},u_{j},u_{k})\in\mathbb{R}_{>0}\times\mathbb{R}_{<0}^{2}|u_{j}+u_{k}>-\sqrt{2\eta_{jk}},u_{i}+u_{k}>\sqrt{-2\zeta_{ik}}\}.$
For the case $\mathrm{(iii)}$, the admissible space (44) is reduced to
$\mathcal{U}_{ijk}(\zeta_{ij},\eta_{jk},\zeta_{ik})=\{(u_{i},u_{j},u_{k})\in\mathbb{R}_{>0}\times\mathbb{R}_{<0}^{2}|u_{j}+u_{k}>-\sqrt{2\eta_{jk}},u_{i}+u_{j}>\sqrt{-2\zeta_{ij}}\}.$
For the case $\mathrm{(iv)}$, the admissible space (44) is reduced to
$\mathcal{U}_{ijk}(\zeta_{ij},\eta_{jk},\zeta_{ik})=\{(u_{i},u_{j},u_{k})\in\mathbb{R}_{>0}\times\mathbb{R}_{<0}^{2}|$ $u_{j}+u_{k}>-\sqrt{2\eta_{jk}},u_{i}+u_{j}>\sqrt{-2\zeta_{ij}},$
$u_{i}+u_{k}>\sqrt{-2\zeta_{ik}}\}.$
Each of these four cases shows that $\mathcal{U}_{ijk}(\zeta_{ij},\eta_{jk},\zeta_{ik})$ is a convex polytope. Q.E.D.
Theorem 5.2.
Under the same assumptions as in Theorem 5.1. The Jacobian $\Lambda_{ijk}=\frac{\partial(\theta^{jk}_{i},\theta^{ik}_{j},\theta^{ij}_{k})}{\partial(u_{i},u_{j},u_{k})}$ is non-degenerate, symmetric and negative definite on $\mathcal{U}_{ijk}(\zeta_{ij},\eta_{jk},\zeta_{ik})$. As a result, the Jacobian $\Lambda=\frac{\partial(K_{i},...,K_{N})}{\partial(u_{i},...,u_{N})}$ is symmetric and negative definite on $\mathcal{U}(\zeta,\eta)$.
Proof.
The formula (30) implies
$\frac{\partial(f_{i},f_{j},f_{k})}{\partial(u_{i},u_{j},u_{k})}=\left(\begin{array}[]{ccc}-e^{f_{i}}&0&0\\ 0&e^{f_{j}}&0\\ 0&0&e^{f_{k}}\\ \end{array}\right):=Q_{5}.$
Then
$\Lambda_{ijk}=\frac{\partial(\theta_{i},\theta_{j},\theta_{k})}{\partial(u_{i},u_{j},u_{k})}=-\frac{1}{A}\left(\begin{array}[]{ccc}\sinh l_{i}&0&0\\ 0&\sinh l_{j}&0\\ 0&0&\sinh l_{k}\\ \end{array}\right)Q_{1}Q_{2}Q_{5},$
where $Q_{1}$ is defined by (36) and $Q_{2}$ is defined by (38). Moreover, $\det Q_{1}>0$.
Since $d_{ij}<0,d_{ik}<0$ and the others are positive for any right-angled hyperbolic hexagon $\{ijk\}\in F$ with edge lengths given by (43), then
$\det Q_{2}=$ $\coth d_{jk}\coth d_{ki}\coth d_{ij}+\coth d_{kj}\coth d_{ik}\coth d_{ji}$
$=$ $\frac{\cosh d_{jk}\cosh d_{ki}\cosh d_{ij}+\cosh d_{kj}\cosh d_{ik}\cosh d_{ji}}{\sinh d_{ij}\sinh d_{jk}\sinh d_{ki}}<0,$
where (1) is used in the second line. Thus
(45) $\det\Lambda_{ijk}=-\frac{1}{A^{3}}\sinh l_{i}\sinh l_{j}\sinh l_{k}\cdot\det Q_{1}\cdot\det Q_{2}\cdot\det Q_{5}<0,$
which implies the matrix $\Lambda_{ijk}$ is non-degenerate. The symmetry of the matrix $\Lambda_{ijk}$ follows from Lemma 2.22. Set $a_{rrs}=e^{f_{r}}\cdot\coth d_{rs}$. Then $a_{iij}<0,\ a_{iik}<0,\ a_{jjk}>0,\ a_{kkj}>0,\ a_{jji}>0,\ a_{kki}>0$. Define
$Q:=Q_{1}Q_{2}Q_{5}=\left(\begin{array}[]{ccc}-1&\cosh\theta_{k}&\cosh\theta_{j}\\ \cosh\theta_{k}&-1&\cosh\theta_{i}\\ \cosh\theta_{j}&\cosh\theta_{i}&-1\\ \end{array}\right)\times\left(\begin{array}[]{ccc}0&a_{jjk}&a_{kkj}\\ -a_{iik}&0&a_{kki}\\ -a_{iij}&a_{jji}&0\\ \end{array}\right).$
By similar arguments in the proof of Theorem 3.2, we will show that the determinants of the $1\times 1$ and $2\times 2$ order principal minor of $Q$ are positive. Let $m_{rs}$ be the entry of $Q$ at $r$-th row and $s$-th column. Then
$m_{11}=-\cosh\theta_{k}\cdot a_{iik}-\cosh\theta_{j}\cdot a_{iij}$
$m_{22}=\cosh\theta_{k}\cdot a_{jjk}+\cosh\theta_{i}\cdot a_{jji}$
$m_{12}=-a_{jjk}+\cosh\theta_{j}\cdot a_{jji}$
$m_{21}=a_{iik}-\cosh\theta_{i}\cdot a_{iij}.$
Then the determinants of the $1\times 1$ principal submatrices of $Q$ is $m_{11}>0$ by $a_{iik}<0,\ a_{iij}<0$. And the determinants of the $2\times 2$ principal submatrices of $Q$ is
$m_{11}m_{22}-m_{12}m_{21}=$ $-[\sinh^{2}\theta_{k}a_{iik}a_{jjk}+(\cosh\theta_{i}\cosh\theta_{k}+\cosh\theta_{j})a_{iik}a_{jji}$
$+(\cosh\theta_{j}\cosh\theta_{k}+\cosh\theta_{i})a_{iij}a_{jjk}]>0.$
This completes the proof. Q.E.D.
The following theorem gives the rigidity of the mixed discrete conformal structure $\mathrm{I}$, which generalizes the case $\mathrm{(iii)}$ in Theorem 1.9. As the proof is similar to that of Theorem 3.4, we omit it here.
Theorem 5.3.
Under the same assumptions as in Theorem 5.1. The function $f$ is determined by its generalized combinatorial curvature $K\in\mathbb{R}_{>0}^{E}$. In particular, the map from $f$ to $K$ is a smooth embedding.
5.2. Existence of the mixed discrete conformal structure $\mathrm{I}$
This proof is similar to that of Theorem 3.5 for the discrete conformal structure (A1).
Theorem 5.4.
Under the same assumptions as in Theorem 5.1. The image of $K$ is >$\mathbb{R}^{N}_{>0}$.
To prove Theorem 5.4, we need the following two lemmas.
Lemma 5.5.
Under the same assumptions as in Theorem 5.1. For any constants $c_{1},...,c_{6}\in\mathbb{R}$, if one of the following conditions is satisfied
(i):
$\lim f_{i}=+\infty,\ \lim f_{j}=+\infty,\ \lim f_{k}=+\infty$;
(ii):
$\lim f_{i}=+\infty,\ \lim f_{j}=+\infty,\ \lim f_{k}=c_{1}$;
(iii):
$\lim f_{i}=+\infty,\ \lim f_{j}=c_{2},\ \lim f_{k}=c_{3}$;
(iv):
$\lim f_{i}=-\infty,\ \lim f_{j}=+\infty,\ \lim f_{k}=+\infty$;
(v):
$\lim f_{i}=-\infty,\ \lim f_{j}=+\infty,\ \lim f_{k}=c_{4}$;
(vi):
$\lim f_{i}=-\infty,\ \lim f_{j}=c_{5},\ \lim f_{k}=c_{6}$.
Then $\lim\theta^{jk}_{i}(f_{i},f_{j},f_{k})=0$.
Proof.
By (43), we have
$\cosh l_{ij}$ $=e^{f_{i}+f_{j}}(\zeta_{ij}+\frac{1}{2}e^{-2f_{i}}+\frac{1}{2}e^{-2f_{j}}),$
$\cosh l_{ik}$ $=e^{f_{i}+f_{k}}(\zeta_{ik}+\frac{1}{2}e^{-2f_{i}}+\frac{1}{2}e^{-2f_{k}}),$
$\cosh l_{jk}$ $=e^{f_{j}+f_{k}}(\eta_{jk}-\frac{1}{2}e^{-2f_{j}}-\frac{1}{2}e^{-2f_{k}}).$
For the case $\mathrm{(i)}$, if $\lim f_{i}=+\infty,\ \lim f_{j}=+\infty,\ \lim f_{k}=+\infty$, then $\lim\cosh l_{ij}:=\lim c_{k}e^{f_{i}+f_{j}}$$\lim\cosh l_{jk}:=\lim c_{i}e^{f_{j}+f_{k}}$ and $\lim\cosh l_{ik}:=\lim c_{j}e^{f_{i}+f_{k}}$, where $c_{i},c_{j},c_{k}$ are positive constants. By the cosine law,
$\cosh\theta^{jk}_{i}=\frac{\cosh l_{jk}+\cosh l_{ij}\cosh l_{ik}}{\sinh l_{ij}\sinh l_{ik}}\rightarrow\frac{c_{i}e^{f_{j}+f_{k}}}{c_{k}e^{f_{i}+f_{j}}\cdot c_{j}e^{f_{i}+f_{k}}}+1\rightarrow 1,$
which implies $\theta^{jk}_{i}\rightarrow 0$. The proofs of the cases $\mathrm{(ii)}$ and $\mathrm{(iii)}$ are similar, so we omit them here.
For the case $\mathrm{(iv)}$, if $\lim f_{i}=-\infty,\ \lim f_{j}=+\infty,\ \lim f_{k}=+\infty$. By (43), we have
$\cosh l_{ij}=e^{f_{j}-f_{i}}(\zeta_{ij}e^{2f_{i}}+\frac{1}{2}e^{2f_{i}-2f_{j}}+\frac{1}{2}),$
$\cosh l_{ik}=e^{f_{k}-f_{i}}(\zeta_{ik}e^{2f_{i}}+\frac{1}{2}e^{2f_{i}-2f_{k}}+\frac{1}{2}),$
$\cosh l_{jk}=e^{f_{j}+f_{k}}(\eta_{jk}-\frac{1}{2}e^{-2f_{j}}-\frac{1}{2}e^{-2f_{k}}).$
Then $\lim\cosh l_{ij}:=\lim c_{k}e^{f_{j}-f_{i}},\ \lim\cosh l_{ik}:=\lim c_{j}e^{f_{k}-f_{i}}$ and $\lim\cosh l_{jk}:=\lim c_{i}e^{f_{j}+f_{k}}$, where $c_{i},\ c_{j},\ c_{k}$ are positive constants. By the cosine law,
$\cosh\theta^{jk}_{i}=\frac{\cosh l_{jk}+\cosh l_{ij}\cosh l_{ik}}{\sinh l_{ij}\sinh l_{ik}}\rightarrow\frac{c_{i}e^{f_{j}+f_{k}}}{c_{k}e^{f_{j}-f_{i}}\cdot c_{j}e^{f_{k}-f_{i}}}+1\rightarrow 1,$
which implies $\theta^{jk}_{i}\rightarrow 0$. The same proofs are suitable for the cases $\mathrm{(v)}$ and $\mathrm{(vi)}$, and we omit them here. Q.E.D.
Lemma 5.6.
Under the same assumptions as in Theorem 5.1. For any constants $c_{1},...,c_{5}\in\mathbb{R}$, if one of the following conditions is satisfied
(i):
$\lim f_{j}=+\infty,\ \lim f_{i}=+\infty,\ \lim f_{k}=+\infty$;
(ii):
$\lim f_{j}=+\infty,\ \lim f_{i}=+\infty,\ \lim f_{k}=c_{1}$;
(iii):
$\lim f_{j}=+\infty,\ \lim f_{i}=-\infty,\ \lim f_{k}=+\infty$;
(iv):
$\lim f_{j}=+\infty,\ \lim f_{i}=-\infty,\ \lim f_{k}=c_{2}$;
(v):
$\lim f_{j}=+\infty,\ \lim f_{i}=c_{3},\ \lim f_{k}=+\infty$;
(vi):
$\lim f_{j}=+\infty,\ \lim f_{i}=c_{4},\ \lim f_{k}=c_{5}$.
Then $\lim\theta^{ik}_{j}(f_{i},f_{j},f_{k})=0$.
Proof.
For the case $\mathrm{(i)}$, if $\lim f_{j}=+\infty,\ \lim f_{i}=+\infty,\ \lim f_{k}=+\infty$, then $\lim\cosh l_{ij}:=\lim c_{k}e^{f_{i}+f_{j}}$$\lim\cosh l_{jk}:=\lim c_{i}e^{f_{j}+f_{k}}$ and $\lim\cosh l_{ik}:=\lim c_{j}e^{f_{i}+f_{k}}$, where $c_{i},c_{j},c_{k}$ are positive constants. By the cosine law,
$\cosh\theta^{ik}_{j}=\frac{\cosh l_{ik}+\cosh l_{ij}\cosh l_{jk}}{\sinh l_{ij}\sinh l_{jk}}\rightarrow\frac{c_{j}e^{f_{i}+f_{k}}}{c_{k}e^{f_{i}+f_{j}}\cdot c_{i}e^{f_{i}+f_{j}}}+1\rightarrow 1,$
which implies $\theta^{ik}_{j}\rightarrow 0$. The same proofs are suitable for the cases $\mathrm{(ii)}$, $\mathrm{(v)}$ and $\mathrm{(vi)}$, and we omit them.
For the case $\mathrm{(iii)}$, if $\lim f_{j}=+\infty,\ \lim f_{i}=-\infty,\ \lim f_{k}=+\infty$, then $\lim\cosh l_{ij}:=\lim c_{k}e^{f_{j}-f_{i}}$$\lim\cosh l_{ik}:=\lim c_{j}e^{f_{k}-f_{i}}$ and $\lim\cosh l_{jk}:=\lim c_{i}e^{f_{j}+f_{k}}$, where $c_{i},\ c_{j},\ c_{k}$ are positive constants. By the cosine law,
$\cosh\theta^{ik}_{j}=\frac{\cosh l_{ik}+\cosh l_{ij}\cosh l_{jk}}{\sinh l_{ij}\sinh l_{jk}}\rightarrow\frac{c_{j}e^{f_{k}-f_{i}}}{c_{k}e^{f_{j}-f_{i}}\cdot c_{i}e^{f_{j}+f_{k}}}+1\rightarrow 1,$
which implies $\theta^{ik}_{j}\rightarrow 0$. The same proof is suitable for the case $\mathrm{(iv)}$, so we omit it. Q.E.D.
Proof of Theorem 5.4: Let the image of $K$ is $X$. The openness of $X$ in >$\mathbb{R}^{N}_{>0}$ follows from the injectivity of the map $f$. To prove the closeness of $X$ in >$\mathbb{R}^{N}_{>0}$, we prove that if a sequence $f^{(m)}$ in $\Omega(\zeta,\eta)=\cap_{\{ijk\}\in F}\Omega_{ijk}(\zeta_{ij},\eta_{jk},\zeta_{ik})$ satisfies $\lim_{m\rightarrow+\infty}K^{(m)}\in\mathbb{R}^{N}_{>0}$, then there exists a subsequence, say $f^{(m)}$, such that $\lim_{m\rightarrow+\infty}f^{(m)}$ is in $\Omega(\zeta,\eta)$.
Suppose otherwise, there exists a subsequence $f^{(m)}$, such that $\lim_{m\rightarrow+\infty}f^{(m)}$ is in the boundary of $\Omega(\zeta,\eta)$. The boundary of the admissible space $\Omega(\zeta,\eta)$ in $[-\infty,+\infty]^{N}$ consists of the following three parts:
(i):
There exists an ideal edge $\{rs\}\in E$ such that $l_{rs}=0$.
(ii):
There exists $i\in B$ in a right-angled hyperbolic hexagon $\{ijk\}\in F$ with edge lengths given by (43), such that $f_{i}^{(m)}\rightarrow\pm\infty$.
(iii):
There exists $j\in B$ in a right-angled hyperbolic hexagon $\{ijk\}\in F$ with edge lengths given by (43), such that $f_{j}^{(m)}\rightarrow+\infty$.
For case $\mathrm{(i)}$, in a right-angled hyperbolic hexagon $\{ijk\}\in F$ with lengths $l_{ij},l_{jk},l_{ik}$ and opposite lengths of boundaries $\theta^{ij}_{k},\theta^{jk}_{i},\theta^{ik}_{j}$. Without loss of generality, we assume $l_{ij}\rightarrow 0$. By the cosine law,
$\cosh\theta^{ik}_{j}=\frac{\cosh l_{ik}+\cosh l_{ij}\cosh l_{jk}}{\sinh l_{ij}\sinh l_{jk}}>\frac{\cosh l_{ij}\cosh l_{jk}}{\sinh l_{ij}\sinh l_{jk}}\geq\frac{\cosh l_{ij}}{\sinh l_{ij}}\rightarrow+\infty,$
which contradicts the assumption that $\lim_{m\rightarrow+\infty}K^{(m)}\in\mathbb{R}^{N}_{>0}$.
The case $\mathrm{(ii)}$ follows from Lemma 5.5 and the case $\mathrm{(iii)}$ follows from Lemma 5.6, which both contradict the assumption that $\lim_{m\rightarrow+\infty}K^{(m)}\in\mathbb{R}^{N}_{>0}$. Q.E.D.
Remark 5.7.
Note that in a right-angled hyperbolic hexagon $\{ijk\}\in F$ with edge lengths given by (43), there is no $j\in B$ such that $f_{j}^{(m)}\rightarrow-\infty$. Otherwise, $u_{j}=-e^{-f_{j}}\rightarrow-\infty$. According to the admissible space $\mathcal{U}_{ijk}(\zeta_{ij},\eta_{jk},\zeta_{ik})$ defined by (44), the equation $u_{j}+u_{k}>-2\sqrt{\eta_{jk}}$ implies $u_{k}\rightarrow+\infty$, which contradicts with the fact that $u_{k}\in\mathbb{R}_{<0}$.
6. Rigidity and existence of the mixed discrete conformal structure II
In this section, we consider the rigidity and the image of $K$ for the mixed discrete conformal structure $\mathrm{II}$, i.e., the discrete hyperbolic metrics $l_{ij},l_{jk},l_{ki}$ of a right-angled hyperbolic hexagon $\{ijk\}\in F$ are induced by (5) and (9). As mentioned in Remark 1.8, if $d_{rs}<0$, then $l_{rs},l_{rt}$ are given by (9) and $l_{st}$ is given by (5), where $\{r,s,t\}=\{i,j,k\}$. Note that $\eta\in\mathbb{R}_{>0}^{E}$ by (5) and $\zeta\in\mathbb{R}^{E}$ by (9). To simplify the notation, we use a new parameter $\xi$ to replace $\eta$ and $\zeta$.
Without loss of generality, we can always assume $d_{ij}<0$ and $d_{ik}<0$. Then for any right-angled hyperbolic hexagon $\{ijk\}\in F$, the edge lengths $l_{ij}$ and $l_{ik}$ are given by (9) and the edge length $l_{jk}$ is given by (5), i.e.,
(46) $\cosh l_{ij}$ $=\sqrt{(1+\alpha_{i}e^{2f_{i}})(1+\alpha_{j}e^{2f_{j}})}+\xi_{ij}e^{f_{i}+f_{j}},$
$\cosh l_{ik}$ $=\sqrt{(1+\alpha_{i}e^{2f_{i}})(1+\alpha_{k}e^{2f_{k}})}+\xi_{ik}e^{f_{i}+f_{k}},$
$\cosh l_{jk}$ $=-\sqrt{(1+\alpha_{j}e^{2f_{j}})(1+\alpha_{k}e^{2f_{k}})}+\xi_{jk}e^{f_{j}+f_{k}}.$
6.1. Admissible space of mixed discrete conformal structure $\mathrm{II}$
The admissible space of the discrete conformal factor $f$ for a right-angled hyperbolic hexagon $\{ijk\}\in F$ with edge lengths given by (46) is defined to be
$\Omega_{ijk}(\xi_{ij},\xi_{jk},\xi_{ik})=\{(f_{i},f_{j},f_{k})\in\mathbb{R}^{3}|l_{ij}>0,\ l_{jk}>0,\ l_{ik}>0\}.$
Since $\frac{\sinh d_{ij}}{\sinh d_{ji}}=-\sqrt{\frac{1+\alpha_{i}e^{2f_{i}}}{1+\alpha_{j}e^{2f_{j}}}}$, $\frac{\sinh d_{jk}}{\sinh d_{kj}}=\sqrt{\frac{1+\alpha_{j}e^{2f_{j}}}{1+\alpha_{k}e^{2f_{k}}}}$, $\frac{\sinh d_{ik}}{\sinh d_{ki}}=-\sqrt{\frac{1+\alpha_{i}e^{2f_{i}}}{1+\alpha_{k}e^{2f_{k}}}}$, then by by the proof of Lemma 2.22, we have
(47) $\frac{\partial f_{i}}{\partial u_{i}}=-\sqrt{1+\alpha_{i}e^{2f_{i}}},\ \frac{\partial f_{j}}{\partial u_{j}}=\sqrt{1+\alpha_{j}e^{2f_{j}}},\ \frac{\partial f_{k}}{\partial u_{k}}=\sqrt{1+\alpha_{k}e^{2f_{k}}}.$
To obtain the admissible space $\mathcal{U}_{ijk}(\xi_{ij},\xi_{jk},\xi_{ik})$ with respect to the discrete conformal factors $u$, we need to characterize $l_{ij}>0$ and $l_{jk}>0$. Since the case that $l_{jk}>0$ has been already discussed in Subsection 3.1, we just need to discuss the case $l_{ij}>0$. By Remark 2.23, we have
(i):
If $\alpha_{i}=0$ and $\alpha_{j}=0$, then $\cosh l_{ij}=1+\xi_{ij}e^{u_{j}-u_{i}}>1$, which implies $\xi_{ij}>0$ and $(u_{j}-u_{i})\in\mathbb{R}$.
(ii):
If $\alpha_{i}=1$ and $\alpha_{j}=0$, then $\cosh l_{ij}=\sqrt{1+e^{2f_{i}}}+\xi_{ij}e^{f_{i}+f_{j}}=\coth u_{i}+\xi_{ij}e^{u_{j}}\frac{1}{\sinh u_{i}}>1$, which implies $\xi_{ij}e^{u_{i}+u_{j}}+1>0$ by $u_{i}>0$. Thus if $\xi_{ij}\geq 0$, then $(u_{i}+u_{j})\in\mathbb{R}$; if $\xi_{ij}<0$, then $u_{i}+u_{j}<\log(-\frac{1}{\xi_{ij}})$. For simplicity, we only consider the case $\xi_{ij}\geq 0$ in the following.
(iii):
If $\alpha_{i}=-1$ and $\alpha_{j}=0$, then $\cosh l_{ij}=\sqrt{1-e^{2f_{i}}}+\xi_{ij}e^{f_{i}+f_{j}}=\tanh u_{i}+\xi_{ij}e^{u_{j}}\frac{1}{\cosh u_{i}}>1$, which implies $\xi_{ij}e^{u_{i}+u_{j}}>1$. Thus $\xi_{ij}>0$ and $u_{i}+u_{j}>\log\frac{1}{\xi_{ij}}$.
(iv):
If $\alpha_{i}=0$ and $\alpha_{j}=1$, then $\cosh l_{ij}=\sqrt{1+e^{2f_{j}}}+\xi_{ij}e^{f_{i}+f_{j}}=-\coth u_{j}+\xi_{ij}e^{-u_{i}}\frac{-1}{\sinh u_{j}}>1$, which implies $e^{u_{i}+u_{j}}+\xi_{ij}>0$ by $u_{j}<0$. Thus if $\xi_{ij}\geq 0$, then $(u_{i}+u_{j})\in\mathbb{R}$; if $\xi_{ij}<0$, then $u_{i}+u_{j}>\log(-\xi_{ij})$.
(v):
If $\alpha_{i}=1$ and $\alpha_{j}=1$, then $\cosh l_{ij}=\sqrt{(1+e^{2f_{i}})(1+e^{2f_{j}})}+\xi_{ij}e^{f_{i}+f_{j}}=-\coth u_{i}\coth u_{j}+\xi_{ij}\frac{1}{\sinh u_{i}}\frac{-1}{\sinh u_{j}}>1$, which implies $\cosh(u_{i}+u_{j})+\xi_{ij}>0$ by $u_{i}>0$ and $u_{j}<0$. Thus if $\xi_{ij}>-1$, then $(u_{i}+u_{j})\in\mathbb{R}$; if $\xi_{ij}<-1$, then $u_{i}+u_{j}>\mathrm{arccosh}\ (-\xi_{ij})$ or $u_{i}+u_{j}<-\mathrm{arccosh}\ (-\xi_{ij})$. For simplicity, we only consider the case $\xi_{ij}>-1$ in the following.
(vi):
If $\alpha_{i}=-1$ and $\alpha_{j}=1$, then $\cosh l_{ij}=\sqrt{(1-e^{2f_{i}})(1+e^{2f_{j}})}+\xi_{ij}e^{f_{i}+f_{j}}=-\tanh u_{i}\coth u_{j}+\xi_{ij}\frac{1}{\cosh u_{i}}\frac{-1}{\sinh u_{j}}>1$, which implies $\sinh(u_{i}+u_{j})+\xi_{ij}>0$ by $u_{j}<0$. Thus $\xi_{ij}\in\mathbb{R}$ and $u_{i}+u_{j}>\mathrm{arcsinh}\ (-\xi_{ij})$.
(vii):
If $\alpha_{i}=0$ and $\alpha_{j}=-1$, then $\cosh l_{ij}=\sqrt{1-e^{2f_{j}}}+\xi_{ij}e^{f_{i}+f_{j}}=-\tanh u_{j}+\xi_{ij}e^{-u_{i}}\frac{1}{\cosh u_{j}}>1$, which implies $e^{u_{i}+u_{j}}<\xi_{ij}$. Thus $\xi_{ij}>0$ and $u_{i}+u_{j}<\log\xi_{ij}$.
(viii):
If $\alpha_{i}=1$ and $\alpha_{j}=-1$, then $\cosh l_{ij}=\sqrt{(1+e^{2f_{i}})(1-e^{2f_{j}})}+\xi_{ij}e^{f_{i}+f_{j}}=-\coth u_{i}\tanh u_{j}+\xi_{ij}\frac{1}{\sinh u_{i}}\frac{1}{\cosh u_{j}}>1$, which implies $\sinh(u_{i}+u_{j})<\xi_{ij}$. Thus $\xi_{ij}\in\mathbb{R}$ and $u_{i}+u_{j}<\mathrm{arcsinh}\ \xi_{ij}$.
(ix):
If $\alpha_{i}=-1$ and $\alpha_{j}=-1$, then $\cosh l_{ij}=\sqrt{(1-e^{2f_{i}})(1-e^{2f_{j}})}+\xi_{ij}e^{f_{i}+f_{j}}=-\tanh u_{i}\tanh u_{j}+\xi_{ij}\frac{1}{\cosh u_{i}}\frac{1}{\cosh u_{j}}>1$, which implies $\cosh(u_{i}+u_{j})<\xi_{ij}$. Thus $\xi_{ij}>1$ and $-\mathrm{arccosh}\ \xi_{ij}<u_{i}+u_{j}<\mathrm{arccosh}\ \xi_{ij}$.
For simplicity, we use triples $(\alpha_{i},\alpha_{j},\alpha_{k})$ to represent different types of the admissible spaces of $u$ for a right-angled hyperbolic hexagon $\{ijk\}\in F$ with edge lengths given by (46). By permutations and combinations, the admissible space $\Omega_{ijk}(\xi_{ij},\xi_{jk},\xi_{ik})$ of $f$ can be transferred to 27 types of the admissible spaces of $u$. Note that the positions of $j$ and $k$ are inter-changeable in (46), we regard $(\alpha_{i},\alpha_{j},\alpha_{k})$ and $(\alpha_{i},\alpha_{k},\alpha_{j})$ as the same type. Therefore, there exist 18 types of the admissible spaces $\mathcal{U}^{(n)}_{ijk}(\xi_{ij},\xi_{jk},\xi_{ik})$ of $u$, as shown below. Here $n=1,2,...,18$. For simplicity, we write $\mathcal{U}^{(n)}_{ijk}(\xi_{ij},\xi_{jk},\xi_{ik})$ as $\mathcal{U}^{(n)}_{ijk}(\xi)$.
(I):
Set $(\alpha_{i},\alpha_{j},\alpha_{k})=(0,0,0)$. Then $u_{j}+u_{k}>\log\frac{2}{\xi_{jk}}$ and $\xi_{jk}>0$ in Subsection 3.1. By the above discussion, we have $\xi_{ij}>0,\ (u_{j}-u_{i})\in\mathbb{R}$ and $\xi_{ik}>0,\ (u_{k}-u_{i})\in\mathbb{R}$. Therefore, the admissible space of $u$ is defined as
(48) $\mathcal{U}^{(1)}_{ijk}(\xi)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}^{3}|u_{j}+u_{k}>\log\frac{2}{\xi_{jk}}\},$
where $\xi_{jk}>0,\ \xi_{ij}>0,\ \xi_{ik}>0$.
(II):
Set $(\alpha_{i},\alpha_{j},\alpha_{k})=(1,0,0)$. Then $u_{j}+u_{k}>\log\frac{2}{\xi_{jk}}$ and $\xi_{jk}>0$. Moreover, $(u_{i}+u_{j})\in\mathbb{R}$ for $\xi_{ij}\geq 0$ and $(u_{i}+u_{k})\in\mathbb{R}$ for $\xi_{ik}\geq 0$. Then
[…]
where $\xi_{jk}>0,\ \xi_{ij}\geq 0,\ \xi_{ik}\geq 0$.
(III):
Set $(\alpha_{i},\alpha_{j},\alpha_{k})=(-1,0,0)$. Then $u_{j}+u_{k}>\log\frac{2}{\xi_{jk}},\ \xi_{jk}>0$. Moreover, $\xi_{ij}>0$$u_{i}+u_{j}>\log\frac{1}{\xi_{ij}}$ and $\xi_{ik}>0,\ u_{i}+u_{k}>\log\frac{1}{\xi_{ik}}$. Then
$\mathcal{U}^{(3)}_{ijk}(\xi)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}_{>0}\times\mathbb{R}^{2}|u_{j}+u_{k}>\log\frac{2}{\xi_{jk}},\ u_{i}+u_{j}>\log\frac{1}{\xi_{ij}},\ u_{i}+u_{k}>\log\frac{1}{\xi_{ik}}\},$
where $\xi_{jk}>0,\ \xi_{ij}>0,\ \xi_{ik}>0$.
(IV):
Set $(\alpha_{i},\alpha_{j},\alpha_{k})=(0,0,1)$. Then $u_{j}+u_{k}>\log\frac{1}{\xi_{jk}}$ and $\xi_{jk}>0$ in Subsection 3.1. Moreover, $\xi_{ij}>0$ and $(u_{j}-u_{i})\in\mathbb{R}$. Similarly, $(u_{i}+u_{k})\in\mathbb{R}$ for $\xi_{ik}\geq 0$ and $u_{i}+u_{k}>\log(-\xi_{ik})$ for $\xi_{ik}<0$. Then
(50) $\mathcal{U}^{(4)}_{ijk}(\xi)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}^{2}\times\mathbb{R}_{<0}|u_{j}+u_{k}>\log\frac{1}{\xi_{jk}},\ e^{u_{i}+u_{k}}>-\xi_{ik}\},$
where $\xi_{jk}>0,\ \xi_{ij}>0,\ \xi_{ik}\in\mathbb{R}$.
The admissible space (50) can be described explicitly as follows:
(i):
For $\xi_{jk}>0,\ \xi_{ij}>0,\ \xi_{ik}\geq 0$, then
(51) $\mathcal{U}^{(4)}_{ijk}(\xi)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}^{2}\times\mathbb{R}_{<0}|u_{j}+u_{k}>\log\frac{1}{\xi_{jk}}\}.$
(ii):
For $\xi_{jk}>0,\ \xi_{ij}>0,\ \xi_{ik}<0$, then
(52) $\mathcal{U}^{(4)}_{ijk}(\xi)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}^{2}\times\mathbb{R}_{<0}|u_{j}+u_{k}>\log\frac{1}{\xi_{jk}},\ u_{i}+u_{k}>\log(-\xi_{ik})\}.$
(V):
Set $(\alpha_{i},\alpha_{j},\alpha_{k})=(1,0,1)$. Then $u_{j}+u_{k}>\log\frac{1}{\xi_{jk}}$ and $\xi_{jk}>0$. Moreover, $(u_{i}+u_{j})\in\mathbb{R}$ for $\xi_{ij}\geq 0$ and $(u_{i}+u_{k})\in\mathbb{R}$ for $\xi_{ik}>-1$. Then
(53) $\mathcal{U}^{(5)}_{ijk}(\xi)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}\times\mathbb{R}_{>0}\times\mathbb{R}_{<0}|u_{j}+u_{k}>\log\frac{1}{\xi_{jk}}\},$
where $\xi_{jk}>0,\ \xi_{ij}\geq 0,\ \xi_{ik}>-1$.
(VI):
Set $(\alpha_{i},\alpha_{j},\alpha_{k})=(-1,0,1)$. Then $u_{j}+u_{k}>\log\frac{1}{\xi_{jk}}$ and $\xi_{jk}>0$. Moreover, $\xi_{ij}>0$$u_{i}+u_{j}>\log\frac{1}{\xi_{ij}}$. Similarly, $\xi_{ik}\in\mathbb{R}$ and $u_{i}+u_{k}>\mathrm{arcsinh}\ (-\xi_{ik})$. Then
$\mathcal{U}^{(6)}_{ijk}(\xi)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}_{>0}\times\mathbb{R}\times\mathbb{R}_{<0}|$ $u_{j}+u_{k}>\log\frac{1}{\xi_{jk}},\ u_{i}+u_{j}>\log\frac{1}{\xi_{ij}},$
$u_{i}+u_{k}>\mathrm{arcsinh}\ (-\xi_{ik})\},$
where $\xi_{jk}>0,\ \xi_{ij}>0,\ \xi_{ik}\in\mathbb{R}$.
(VII):
Set $(\alpha_{i},\alpha_{j},\alpha_{k})=(0,0,-1)$. Then $u_{j}+u_{k}>\log\frac{1}{\xi_{jk}}$ and $\xi_{jk}>0$ in Subsection 3.1. Moreover, $\xi_{ij}>0$$(u_{j}-u_{i})\in\mathbb{R}$. Similarly, $\xi_{ik}>0$ and $u_{i}+u_{k}<\log\xi_{ik}$. Then
$\mathcal{U}^{(7)}_{ijk}(\xi)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}^{2}\times\mathbb{R}_{<0}|u_{j}+u_{k}>\log\frac{1}{\xi_{jk}},\ u_{i}+u_{k}<\log\xi_{ik}\},$
where $\xi_{jk}>0,\ \xi_{ij}>0,\ \xi_{ik}>0$.
(VIII):
Set $(\alpha_{i},\alpha_{j},\alpha_{k})=(1,0,-1)$. Then $u_{j}+u_{k}>\log\frac{1}{\xi_{jk}}$ and $\xi_{jk}>0$. Moreover, $(u_{i}+u_{j})\in\mathbb{R}$ for $\xi_{ij}\geq 0$. Similarly, $\xi_{ik}\in\mathbb{R}$ and $u_{i}+u_{k}<\mathrm{arcsinh}\ \xi_{ik}$. Then
[…]
where $\xi_{jk}>0,\ \xi_{ij}\geq 0,\ \xi_{ik}\in\mathbb{R}$.
(IX):
Set $(\alpha_{i},\alpha_{j},\alpha_{k})=(-1,0,-1)$. Then $u_{j}+u_{k}>\log\frac{1}{\xi_{jk}}$ and $\xi_{jk}>0$. Moreover, $\xi_{ij}>0$ and $u_{i}+u_{j}>\log\frac{1}{\xi_{ij}}$. Similarly, $\xi_{ik}>1$ and $-\mathrm{arccosh}\ \xi_{ik}<u_{i}+u_{k}<\mathrm{arccosh}\ \xi_{ik}$. Then
$\mathcal{U}^{(9)}_{ijk}(\xi)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}_{>0}\times\mathbb{R}\times\mathbb{R}_{<0}|$ $u_{j}+u_{k}>\log\frac{1}{\xi_{jk}},\ u_{i}+u_{j}>\log\frac{1}{\xi_{ij}},$
$-\mathrm{arccosh}\ \xi_{ik}<u_{i}+u_{k}<\mathrm{arccosh}\ \xi_{ik}\},$
where $\xi_{jk}>0,\ \xi_{ij}>0,\ \xi_{ik}>1$.
(X):
Set $(\alpha_{i},\alpha_{j},\alpha_{k})=(0,1,1)$. Then $u_{j}+u_{k}>-\mathrm{arccosh}\ \xi_{jk}$ and $\xi_{jk}>1$ in Subsection 3.1. Moreover, $(u_{i}+u_{j})\in\mathbb{R}$ for $\xi_{ij}\geq 0$ and $u_{i}+u_{j}>\log(-\xi_{ij})$ for $\xi_{ij}<0$. Similarly, $(u_{i}+u_{k})\in\mathbb{R}$ for $\xi_{ik}\geq 0$ and $u_{i}+u_{k}>\log(-\xi_{ik})$ for $\xi_{ik}<0$. Then
(54) $\mathcal{U}^{(10)}_{ijk}(\xi)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}\times\mathbb{R}^{2}_{<0}|$ $u_{j}+u_{k}>-\mathrm{arccosh}\ \xi_{jk},\ e^{u_{i}+u_{j}}>-\xi_{ij},$
$e^{u_{i}+u_{k}}>-\xi_{ik}\},$
where $\xi_{jk}>1,\ \xi_{ij}\in\mathbb{R},\ \xi_{ik}\in\mathbb{R}$.
The admissible space (54) can also be described explicitly as follows:
(i):
For $\xi_{jk}>1,\ \xi_{ij}\geq 0,\ \xi_{ik}\geq 0$, then
$\mathcal{U}^{(10)}_{ijk}(\xi)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}\times\mathbb{R}^{2}_{<0}|u_{j}+u_{k}>-\mathrm{arccosh}\ \xi_{jk}\}.$
(ii):
For $\xi_{jk}>1,\ \xi_{ij}\geq 0,\ \xi_{ik}<0$, then
$\mathcal{U}^{(10)}_{ijk}(\xi)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}\times\mathbb{R}^{2}_{<0}|u_{j}+u_{k}>-\mathrm{arccosh}\ \xi_{jk},\ u_{i}+u_{k}>\log(-\xi_{ik})\}.$
(iii):
For $\xi_{jk}>1,\ \xi_{ij}<0,\ \xi_{ik}\geq 0$, then
$\mathcal{U}^{(10)}_{ijk}(\xi)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}\times\mathbb{R}^{2}_{<0}|u_{j}+u_{k}>-\mathrm{arccosh}\ \xi_{jk},\ u_{i}+u_{j}>\log(-\xi_{ij})\},$
(iv):
For $\xi_{jk}>1,\ \xi_{ij}<0,\ \xi_{ik}<0$, then
[…]
$u_{i}+u_{j}>\log(-\xi_{ij}),\ u_{i}+u_{k}>\log(-\xi_{ik})\}.$
(XI):
Set $(\alpha_{i},\alpha_{j},\alpha_{k})=(1,1,1)$. Then $u_{j}+u_{k}>-\mathrm{arccosh}\ \xi_{jk}$ and $\xi_{jk}>1$. Moreover, $\xi_{ij}>-1$$(u_{i}+u_{j})\in\mathbb{R}$ and $\xi_{ik}>-1,\ (u_{i}+u_{k})\in\mathbb{R}$. Then
(55) $\mathcal{U}^{(11)}_{ijk}(\xi)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}_{>0}\times\mathbb{R}^{2}_{<0}|u_{j}+u_{k}>-\mathrm{arccosh}\ \xi_{jk}\},$
where $\xi_{jk}>1,\ \xi_{ij}>-1,\ \xi_{ik}>-1$.
(XII):
Set $(\alpha_{i},\alpha_{j},\alpha_{k})=(-1,1,1)$. Then $u_{j}+u_{k}>-\mathrm{arccosh}\ \xi_{jk}$ and $\xi_{jk}>1$. Moreover, $\xi_{ij}\in\mathbb{R}$$u_{i}+u_{j}>\mathrm{arcsinh}\ (-\xi_{ij})$ and $\xi_{ik}\in\mathbb{R},\ u_{i}+u_{k}>\mathrm{arcsinh}\ (-\xi_{ik})$. Then
$\mathcal{U}^{(12)}_{ijk}(\xi)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}_{>0}\times\mathbb{R}^{2}_{<0}|$ $u_{j}+u_{k}>-\mathrm{arccosh}\ \xi_{jk},u_{i}+u_{j}>\mathrm{arcsinh}\ (-\xi_{ij}),$
$u_{i}+u_{k}>\mathrm{arcsinh}\ (-\xi_{ik})\},$
where $\xi_{jk}>1,\ \xi_{ij}\in\mathbb{R},\ \xi_{ik}\in\mathbb{R}$.
(XIII):
Set $(\alpha_{i},\alpha_{j},\alpha_{k})=(0,1,-1)$. Then $u_{j}+u_{k}>\mathrm{arcsinh}\ (-\xi_{jk})$ and $\xi_{jk}>0$ in Subsection 3.1. Moreover, $(u_{i}+u_{j})\in\mathbb{R}$ for $\xi_{ij}\geq 0$ and $u_{i}+u_{j}>\log(-\xi_{ij})$ for $\xi_{ij}<0$. Similarly, $\xi_{ik}>0$ and $u_{i}+u_{k}<\log\xi_{ik}$. If $\xi_{ij}<0$, then $\log(-\xi_{ij})+\mathrm{arcsinh}\ (-\xi_{jk})<u_{i}+2u_{j}+u_{k}<2u_{j}+\log\xi_{ik}<\log\xi_{ik}$, which implies $(\xi_{ik}-\xi_{ij})(\xi_{ik}+\xi_{ij})>2\xi_{jk}\xi_{ij}\xi_{ik}$. Thus we assume $\xi_{ik}+\xi_{ij}>0$ for simplicity. Then
$\mathcal{U}^{(13)}_{ijk}(\xi)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}\times\mathbb{R}^{2}_{<0}|$ $u_{j}+u_{k}>\mathrm{arcsinh}\ (-\xi_{jk}),\ e^{u_{i}+u_{j}}>-\xi_{ij},$
$u_{i}+u_{k}<\log\xi_{ik}\},$
where $\xi_{jk}>0,\ \xi_{ik}>0,\ \xi_{ij}\geq 0$ or $\xi_{ij}<0$$\xi_{ij}+\xi_{ik}>0$.
The admissible space $\mathcal{U}^{(13)}_{ijk}(\xi)$ can be described explicitly as follows:
(i):
For $\xi_{jk}>0,\ \xi_{ik}>0,\ \xi_{ij}\geq 0$, then
$\mathcal{U}_{ijk}^{(13)}(\xi)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}\times\mathbb{R}^{2}_{<0}|u_{j}+u_{k}>\mathrm{arcsinh}\ (-\xi_{jk}),u_{i}+u_{k}<\log\xi_{ik}\}.$
(ii):
For $\xi_{jk}>0,\ \xi_{ik}>0,\ \xi_{ij}<0$ and $\xi_{ij}+\xi_{ik}>0$, then
$\mathcal{U}_{ijk}^{(13)}(\xi)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}\times\mathbb{R}^{2}_{<0}|$ $u_{j}+u_{k}>\mathrm{arcsinh}\ (-\xi_{jk}),\ u_{i}+u_{k}<\log\xi_{ik},$
$u_{i}+u_{j}>\log(-\xi_{ij})\}.$
(XIV):
Set $(\alpha_{i},\alpha_{j},\alpha_{k})=(1,1,-1)$. Then $u_{j}+u_{k}>\mathrm{arcsinh}\ (-\xi_{jk})$ and $\xi_{jk}>0$. Moreover, $\xi_{ij}>-1$ and $(u_{i}+u_{j})\in\mathbb{R}$. Similarly, $\xi_{ik}\in\mathbb{R}$ and $u_{i}+u_{k}<\mathrm{arcsinh}\ \xi_{ik}$. Since $u_{i}>0,u_{j}<0$, then $\mathrm{arcsinh}\ (-\xi_{jk})<u_{j}+u_{k}<u_{i}+u_{k}<\mathrm{arcsinh}\ \xi_{ik}$ and hence $\xi_{ik}+\xi_{jk}>0$. Then
$\mathcal{U}^{(14)}_{ijk}(\xi)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}_{>0}\times\mathbb{R}^{2}_{<0}|u_{j}+u_{k}>\mathrm{arcsinh}\ (-\xi_{jk}),u_{i}+u_{k}<\mathrm{arcsinh}\ \xi_{ik}\},$
where $\xi_{jk}>0,\ \xi_{ij}>-1,\ \xi_{ik}\in\mathbb{R}$ and $\xi_{ik}+\xi_{jk}>0$.
(XV):
Set $(\alpha_{i},\alpha_{j},\alpha_{k})=(-1,1,-1)$. Then $u_{j}+u_{k}>\mathrm{arcsinh}\ (-\xi_{jk})$ and $\xi_{jk}>0$. Moreover, $\xi_{ij}\in\mathbb{R}$ and $u_{i}+u_{j}>\mathrm{arcsinh}\ (-\xi_{ij})$. Similarly, $\xi_{ik}>1$ and $-\mathrm{arccosh}\ \xi_{ik}<u_{i}+u_{k}<\mathrm{arccosh}\ \xi_{ik}$. Then $\mathrm{arcsinh}\ (-\xi_{jk})+\mathrm{arcsinh}(-\xi_{ij})<u_{i}+2u_{j}+u_{k}<2u_{j}+\mathrm{arccosh}\xi_{ik}<\mathrm{arccosh}\xi_{ik}$. For simplicity, we require $\xi_{ij}>0$. Then
$\mathcal{U}^{(15)}_{ijk}(\xi)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}_{>0}\times\mathbb{R}^{2}_{<0}|$ $u_{j}+u_{k}>\mathrm{arcsinh}\ (-\xi_{jk}),\ u_{i}+u_{j}>\mathrm{arcsinh}\ (-\xi_{ij}),$
$-\mathrm{arccosh}\ \xi_{ik}<u_{i}+u_{k}<\mathrm{arccosh}\ \xi_{ik}\},$
where $\xi_{jk}>0,\ \xi_{ij}>0,\ \xi_{ik}>1$.
(XVI):
Set $(\alpha_{i},\alpha_{j},\alpha_{k})=(0,-1,-1)$. Then $u_{j}+u_{k}>-\mathrm{arccosh}\ \xi_{jk}$ and $\xi_{jk}>1$ in Subsection 3.1. Moreover, $\xi_{ij}>0,\ u_{i}+u_{j}<\log\xi_{ij}$ and $\xi_{ik}>0,\ u_{i}+u_{k}<\log\xi_{ik}$. Then
$\mathcal{U}^{(16)}_{ijk}(\xi)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}\times\mathbb{R}^{2}_{<0}|$ $u_{j}+u_{k}>-\mathrm{arccosh}\ \xi_{jk},\ u_{i}+u_{j}<\log\xi_{ij},$
$u_{i}+u_{k}<\log\xi_{ik}\},$
where $\xi_{jk}>1,\ \xi_{ij}>0,\ \xi_{ik}>0$.
(XVII):
Set $(\alpha_{i},\alpha_{j},\alpha_{k})=(1,-1,-1)$. Then $u_{j}+u_{k}>-\mathrm{arccosh}\ \xi_{jk}$ and $\xi_{jk}>1$. Moreover, $\xi_{ij}\in\mathbb{R}$ and $u_{i}+u_{j}<\mathrm{arcsinh}\ \xi_{ij}$. Since $u_{i}>0$ and $u_{k}<0$, then $-\mathrm{arccosh}\ \xi_{jk}<u_{j}+u_{k}<u_{i}+u_{j}<\mathrm{arcsinh}\ \xi_{ij}$. For simplicity, we assume $\xi_{ij}>0$. Similarly, $\xi_{ik}\in\mathbb{R}$ and $u_{i}+u_{k}<\mathrm{arcsinh}\ \xi_{ik}$, then we assume $\xi_{ik}>0$ for simplicity. Then
[…]
$u_{i}+u_{j}<\mathrm{arcsinh}\ \xi_{ij},\ u_{i}+u_{k}<\mathrm{arcsinh}\ \xi_{ik}\},$
where $\xi_{jk}>1,\ \xi_{ij}>0,\ \xi_{ik}>0$.
(XVIII):
Set $(\alpha_{i},\alpha_{j},\alpha_{k})=(-1,-1,-1)$. Then $u_{j}+u_{k}>-\mathrm{arccosh}\ \xi_{jk}$ and $\xi_{jk}>1$. Moreover, $\xi_{ij}>1$ and $-\mathrm{arccosh}\ \xi_{ij}<u_{i}+u_{j}<\mathrm{arccosh}\ \xi_{ij}$. Similarly, $\xi_{ik}>1$ and $-\mathrm{arccosh}\ \xi_{ik}<u_{i}+u_{k}<\mathrm{arccosh}\ \xi_{ik}$. Then
[…]
$-\mathrm{arccosh}\ \xi_{ij}<u_{i}+u_{j}<\mathrm{arccosh}\ \xi_{ij},$
$-\mathrm{arccosh}\ \xi_{ik}<u_{i}+u_{k}<\mathrm{arccosh}\ \xi_{ik}\},$
where $\xi_{jk}>1,\ \xi_{ij}>1,\ \xi_{ik}>1$.
In the above different admissible spaces $\mathcal{U}^{(n)}_{ijk}(\xi)$, the weight $\xi$ has different range of values. We are unable to use a uniform formula to express them. For simplicity, we use a notation $\Xi$ to denote the set of the range of weight $\xi$ in the whole admissible spaces $\mathcal{U}^{(n)}_{ijk}(\xi)$.
Theorem 6.1.
Suppose $(\Sigma,\mathcal{T},\alpha,\xi)$ is a weighted triangulated surface with boundary, and the weights $\alpha:B\rightarrow\{-1,0,1\}$ and $\xi\in\Xi$. Let $\{ijk\}\in F$ be a right-angled hyperbolic hexagon with edge lengths given by (46). Then the admissible space $\mathcal{U}^{(n)}_{ijk}(\xi)$ of $u$ has 18 types, each of which is convex polytope. As a result, the admissible space
$\mathcal{U}(\xi)=\bigcap_{\{ijk\}\in F}\mathcal{U}_{ijk}(\xi),\ \mathcal{U}_{ijk}(\xi)\in\{\mathcal{U}^{(n)}_{ijk}(\xi),\ n=1,2,...,18\}$
is also a convex polytope.
Remark 6.2.
The admissible space $\mathcal{U}(\xi)$ is a finite intersection of 18 types of admissible spaces $\mathcal{U}^{(n)}_{ijk}(\xi)$ for $n=1,2,...,18$. But this finite intersection is not arbitrary because of the connectedness of the admissible space $\mathcal{U}(\xi)$. For example, if there exist the admissible space $\mathcal{U}^{(1)}_{ijk}(\xi)$ and $\mathcal{U}^{(18)}_{ijk}(\xi)$ for some right-angled hyperbolic hexagons $\{ijk\}\in F$ with edge lengths given by (46), then there must be the third type of the admissible space $\mathcal{U}^{(n)}_{ijk}(\xi)$ for $n\neq 1,18$.
6.1.1. Negative definiteness of Jacobian
Theorem 6.3.
Under the same assumptions as in Theorem 6.1. The Jacobian $\Lambda_{ijk}=\frac{\partial(\theta^{jk}_{i},\theta^{ik}_{j},\theta^{ij}_{k})}{\partial(u_{i},u_{j},u_{k})}$ is non-degenerate, symmetric and negative definite. As a consequence, the Jacobian $\Lambda=\frac{\partial(K_{i},...,K_{N})}{\partial(u_{i},...,u_{N})}$ is symmetric and negative definite on the admissible space $\mathcal{U}(\xi)$.
Proof.
For simplicity, we set $l_{r}=l_{st},\ \theta_{r}=\theta^{st}_{r}$ and $A=\sinh l_{r}\sinh l_{s}\sinh\theta_{t}$, where $\{r,s,t\}=\{i,j,k\}$. By (47), we have
$\frac{\partial(f_{i},f_{j},f_{k})}{\partial(u_{i},u_{j},u_{k})}=\left(\begin{array}[]{ccc}-C_{i}&0&0\\ 0&C_{j}&0\\ 0&0&C_{k}\\ \end{array}\right):=Q_{6},$
where $C_{r}=\sqrt{1+\alpha_{r}e^{2f_{r}}}$. Then
$\Lambda_{ijk}=\frac{\partial(\theta_{i},\theta_{j},\theta_{k})}{\partial(u_{i},u_{j},u_{k})}=-\frac{1}{A}\left(\begin{array}[]{ccc}\sinh l_{i}&0&0\\ 0&\sinh l_{j}&0\\ 0&0&\sinh l_{k}\\ \end{array}\right)Q_{1}Q_{2}Q_{6},$
where $Q_{1}$ is defined by (36) and $Q_{2}$ is defined by (38). Note that $d_{ij}<0,d_{ik}<0$ and the others are positive for any right-angled hyperbolic hexagon $\{ijk\}\in F$ with edge lengths given by (46). The rest of the proof is similar to that of Theorem 5.2, so we omit it. Q.E.D.
6.2. Rigidity of the mixed discrete conformal structure $\mathrm{II}$
The following theorem gives the rigidity of mixed discrete conformal structure $\mathrm{II}$, which generalizes the case $\mathrm{(iv)}$ in Theorem 1.9. As the proof is similar to that of Theorem 3.4, we omit it.
Theorem 6.4.
Under the same assumptions as in Theorem 6.1. The function $f$ is determined by its generalized combinatorial curvature $K\in\mathbb{R}^{E}_{>0}$. In particular, the map from $f$ to $K$ is a smooth embedding.
6.3. Existence of the mixed discrete conformal structure $\mathrm{II}$
We consider the image of $K$ for four types of the admissible spaces $(0,0,0),(1,0,0),(0,0,1),(1,0,1)$. In other words, we do not consider the case that $\alpha_{j}=\alpha_{k}=1$ for $\alpha\in\{0,1\}$. For simplicity, we use a new notation $\widetilde{\Xi}$ to denote the set of the range of weights $\xi$ in the admissible spaces $\mathcal{U}^{(n)}_{ijk}$, i.e., $n=1,2,4,5.$
Theorem 6.5.
Suppose $(\Sigma,\mathcal{T},\alpha,\xi)$ is a weighted triangulated surface with boundary, and the weights $\alpha:B\rightarrow\{0,1\}$ and $\xi\in\widetilde{\Xi}$. Let $\{ijk\}\in F$ be a right-angled hyperbolic hexagon with edge lengths given by (46). If $\alpha_{j}$ and $\alpha_{k}$ can not take 1 simultaneously, then the image of $K$ is >$\mathbb{R}^{N}_{>0}$.
To prove Theorem 6.5, we need the following four lemmas.
Lemma 6.6.
Under the same assumptions as in Theorem 6.5. For any constants $c_{1},c_{2},c_{3}\in\mathbb{R}$, if one of the following conditions is satisfied
(i):
$\lim f_{i}=+\infty,\ \lim f_{j}=+\infty,\ \lim f_{k}=+\infty$;
(ii):
$\lim f_{i}=+\infty,\ \lim f_{j}=+\infty,\ \lim f_{k}=c_{1}$;
(iii):
$\lim f_{i}=+\infty,\ \lim f_{j}=c_{2},\ \lim f_{k}=c_{3}$.
Then $\lim\theta^{jk}_{i}(f_{i},f_{j},f_{k})=0$.
Proof.
By (46), we have
(56) $\cosh l_{ij}=\bigg{[}\xi_{ij}+\sqrt{(e^{-2f_{i}}+\alpha_{i})(e^{-2f_{j}}+\alpha_{j})}\bigg{]}e^{f_{i}+f_{j}},$
$\cosh l_{ik}=\bigg{[}\xi_{ik}+\sqrt{(e^{-2f_{i}}+\alpha_{i})(e^{-2f_{k}}+\alpha_{k})}\bigg{]}e^{f_{i}+f_{k}},$
$\cosh l_{jk}=\bigg{[}\xi_{jk}-\sqrt{(e^{-2f_{j}}+\alpha_{j})(e^{-2f_{k}}+\alpha_{k})}\bigg{]}e^{f_{j}+f_{k}}.$
For the case $\mathrm{(i)}$, $\lim f_{i}=+\infty,\ \lim f_{j}=+\infty,\ \lim f_{k}=+\infty$. Then $\lim\cosh l_{ij}:=\lim(c_{k}e^{f_{i}+f_{j}})$$\lim\cosh l_{ik}:=\lim(c_{j}e^{f_{i}+f_{k}})$ and $\lim\cosh l_{jk}:=\lim(c_{i}e^{f_{j}+f_{k}})$, where $c_{i},\ c_{j},\ c_{k}$ are positive constants. By the cosine law,
$\cosh\theta^{jk}_{i}=\frac{\cosh l_{jk}+\cosh l_{ij}\cosh l_{ik}}{\sinh l_{ij}\sinh l_{ik}}\rightarrow\frac{c_{i}e^{f_{j}+f_{k}}}{c_{k}e^{f_{i}+f_{j}}\cdot c_{j}e^{f_{i}+f_{k}}}+1\rightarrow 1,$
which implies $\theta^{jk}_{i}\rightarrow 0$. The proofs of the cases $\mathrm{(ii)}$ and $\mathrm{(iii)}$ are similar, so we omit them here.
Q.E.D.
Lemma 6.7.
Under the same assumptions as in Theorem 6.5. If $\lim f_{k}=-\infty$, then $\lim f_{j}=+\infty$. Moreover, if one of the following conditions is satisfied
(i):
$\lim f_{i}=+\infty,\ \lim f_{j}=+\infty,\ \lim f_{k}=-\infty$;
(ii):
$\lim f_{i}=c,\ \lim f_{j}=+\infty,\ \lim f_{k}=-\infty$,
then $\lim\theta^{ij}_{k}(f_{i},f_{j},f_{k})=+\infty$, where $c\in\mathbb{R}$ is a constant.
Proof.
Combining (47) and Remark 2.23, if $\lim f_{k}=-\infty$, then $\lim u_{k}=-\infty$. Note that $l_{jk}>0$ implies that $u_{j}+u_{k}>C(\xi_{jk})$, where $C(\xi_{jk})$ is a constant that depends on $\xi_{jk}$. Then $\lim u_{j}=+\infty$. This implies $\alpha_{j}=0$ and $\lim f_{j}=+\infty$ by Remark 2.23.
Set $\widetilde{L}_{1}=\frac{\cosh l_{ij}}{\sinh l_{ik}\sinh l_{jk}}$ and $\widetilde{L}_{2}=\frac{\cosh l_{ik}\cosh l_{jk}}{\sinh l_{ik}\sinh l_{jk}}$, then $\cosh\theta_{k}^{ij}=\widetilde{L}_{1}+\widetilde{L}_{2}$. Moreover, (46) is reduced to
$\cosh l_{ij}$ $=\sqrt{1+\alpha_{i}e^{2f_{i}}}+\xi_{ij}e^{f_{i}+f_{j}},$
$\cosh l_{ik}$ $\rightarrow\sqrt{1+\alpha_{i}e^{2f_{i}}}+\xi_{ik}e^{f_{i}+f_{k}},$
$\cosh l_{jk}$ $\rightarrow-1+\xi_{jk}e^{f_{j}+f_{k}}.$
For the case $\mathrm{(i)}$, $\lim f_{i}=+\infty,\lim f_{j}=+\infty,\ \lim f_{k}=-\infty$. Then $\lim\cosh l_{ij}:=\lim c_{k}e^{f_{i}+f_{j}}$$\lim\cosh l_{jk}=\lim(-1+\xi_{jk}e^{f_{j}+f_{k}})$ and
$\lim\cosh l_{ik}=\begin{cases}\lim(1+\xi_{ik}e^{f_{i}+f_{k}}),\ &\text{if}\ \alpha_{i}=0,\\ \lim(c_{j}e^{f_{i}}),\ &\text{if}\ \alpha_{i}=1,\end{cases}$
where $c_{j},c_{k}$ are positive constants. If $\alpha_{i}=0$, we have
$\widetilde{L}_{1}\rightarrow$ $\frac{c_{k}e^{f_{i}+f_{j}}}{\sqrt{(-1+\xi_{jk}e^{f_{j}+f_{k}})^{2}-1}\cdot\sqrt{(1+\xi_{ik}e^{f_{i}+f_{k}})^{2}-1}}$
$=$ $\frac{c_{k}e^{f_{i}+f_{j}}}{\sqrt{\xi_{jk}\xi_{ik}}e^{f_{i}+f_{j}}}\cdot\frac{1}{\sqrt{\xi_{jk}e^{2f_{k}}-2e^{f_{k}-f_{j}}}}\cdot\frac{1}{\sqrt{\xi_{ik}e^{2f_{k}}+2e^{f_{k}-f_{i}}}}\rightarrow+\infty.$
And $\widetilde{L}_{2}>0$. Then $\cosh\theta^{ij}_{k}=\widetilde{L}_{1}+\widetilde{L}_{2}\rightarrow+\infty$, which implies $\theta^{ij}_{k}\rightarrow+\infty$. The proof of the case that $\alpha_{i}=1$ is similar, so we omit it here.
For the case $\mathrm{(ii)}$, $\lim f_{i}=c,\lim f_{j}=+\infty,\ \lim f_{k}=-\infty$. Then $\lim\cosh l_{ij}:=\lim c^{\prime}_{k}e^{f_{j}}$$\lim\cosh l_{jk}=\lim(-1+\xi_{jk}e^{f_{j}+f_{k}})$ and
$\lim\cosh l_{ik}=\begin{cases}1,\ &\text{if}\ \alpha_{i}=0,\\ c^{\prime}_{j},\ &\text{if}\ \alpha_{i}=1,\end{cases}$
where $c^{\prime}_{j},c^{\prime}_{k}$ are positive constants. If $\alpha_{i}=0$, then $\cosh l_{ik}\rightarrow 1$ and hence $l_{ik}\rightarrow 0$. Then $\cosh\theta^{ij}_{k}>\widetilde{L}_{2}>\frac{\cosh l_{ik}}{\sinh l_{ik}}\rightarrow+\infty$ and hence $\theta^{ij}_{k}\rightarrow+\infty$. If $\alpha_{i}=1$, similar to the proof of the case $\mathrm{(i)}$, we still have $\widetilde{L}_{1}\rightarrow+\infty$. This implies $\theta^{ij}_{k}\rightarrow+\infty$. This completes the proof. Q.E.D.
Lemma 6.8.
Under the same assumptions as in Theorem 6.5. For any constants $c_{1},c_{2},c_{3}\in\mathbb{R}$, if one of the following conditions is satisfied
(i):
$\lim f_{i}=-\infty,\ \lim f_{j}=+\infty,\ \lim f_{k}=+\infty$;
(ii):
$\lim f_{i}=-\infty,\ \lim f_{j}=+\infty,\ \lim f_{k}=c_{1}$;
(iii):
$\lim f_{i}=-\infty,\ \lim f_{j}=c_{2},\ \lim f_{k}=c_{3}$;
(iv):
$\lim f_{i}=-\infty,\ \lim f_{j}=+\infty,\ \lim f_{k}=-\infty$.
Then $\lim\theta^{jk}_{i}(f_{i},f_{j},f_{k})=+\infty$.
Proof.
Set $L_{1}=\frac{\cosh l_{jk}}{\sinh l_{ij}\sinh l_{ik}}$ and $L_{2}=\frac{\cosh l_{ij}\cosh l_{ik}}{\sinh l_{ij}\sinh l_{ik}}$, then $\cosh\theta^{jk}_{i}=L_{1}+L_{2}$. Since $\lim f_{i}=-\infty$, then (46) is reduced to
(57) $\cosh l_{ij}$ $\rightarrow\sqrt{1+\alpha_{j}e^{2f_{j}}}+\xi_{ij}e^{f_{i}+f_{j}},$
$\cosh l_{ik}$ $\rightarrow\sqrt{1+\alpha_{k}e^{2f_{k}}}+\xi_{ik}e^{f_{i}+f_{k}},$
$\cosh l_{jk}$ $=-\sqrt{(1+\alpha_{j}e^{2f_{j}})(1+\alpha_{k}e^{2f_{k}})}+\xi_{jk}e^{f_{j}+f_{k}}.$
For the case $\mathrm{(i)}$, $\lim f_{j}=+\infty,\ \lim f_{k}=+\infty$. By (57), we have $\lim\cosh l_{jk}:=\lim(c_{i}e^{f_{j}+f_{k}})$,
$\lim\cosh l_{ij}=\begin{cases}\lim(1+\xi_{ij}e^{f_{i}+f_{j}}),\ &\text{if}\ \alpha_{j}=0,\\ \lim(c_{k}e^{f_{j}}),\ &\text{if}\ \alpha_{j}=1,\end{cases}$
and
$\lim\cosh l_{ik}=\begin{cases}\lim(1+\xi_{ik}e^{f_{i}+f_{k}}),\ &\text{if}\ \alpha_{k}=0,\\ \lim(c_{j}e^{f_{k}}),\ &\text{if}\ \alpha_{k}=1,\end{cases}$
where $c_{i},c_{j},c_{k}$ are positive constants. If $\alpha_{j}=\alpha_{k}=0$, then
$L_{1}\rightarrow$ $\frac{c_{i}e^{f_{j}+f_{k}}}{\sqrt{(1+\xi_{ij}e^{f_{i}+f_{j}})^{2}-1}\cdot\sqrt{(1+\xi_{ik}e^{f_{i}+f_{k}})^{2}-1}}$
$=$ $\frac{c_{i}e^{f_{j}+f_{k}}}{\sqrt{\xi_{ij}\xi_{ik}}e^{f_{j}+f_{k}}}\cdot\frac{1}{\sqrt{\xi_{ij}e^{2f_{i}}+2e^{f_{i}-f_{j}}}}\cdot\frac{1}{\sqrt{\xi_{ik}e^{2f_{i}}+2e^{f_{i}-f_{k}}}}\rightarrow+\infty.$
And $L_{2}>0$. Thus $\cosh\theta^{jk}_{i}=L_{1}+L_{2}\rightarrow+\infty$, so $\theta^{jk}_{i}\rightarrow+\infty$. For the cases that $\alpha_{j}=0,\ \alpha_{k}=1$ and $\alpha_{j}=1,\ \alpha_{k}=0$, we still have $\theta^{jk}_{i}\rightarrow+\infty$ by similar proofs.
For the case $\mathrm{(ii)}$, $\lim f_{j}=+\infty,\ \lim f_{k}=c_{1}$. By (57), we have $\lim\cosh l_{jk}:=\lim(c^{\prime}_{i}e^{f_{j}})$ and
$\lim\cosh l_{ik}=\begin{cases}1,\ &\text{if}\ \alpha_{k}=0,\\ c^{\prime}_{j},\ &\text{if}\ \alpha_{k}=1,\end{cases}$
where $c^{\prime}_{i},c^{\prime}_{j}$ are positive constants. If $\alpha_{k}=0$, then $l_{ik}\rightarrow 0$ and hence $L_{2}>\frac{\cosh l_{ik}}{\sinh l_{ik}}\rightarrow+\infty$. This implies $\cosh\theta^{jk}_{i}\rightarrow+\infty$. If $\alpha_{k}=1$, then by the assumption, $\alpha_{j}=0$. Hence, $\lim\cosh l_{ij}=\lim(1+\xi_{ij}e^{f_{i}+f_{j}})$, $\lim\cosh l_{jk}=\lim(c^{\prime}_{i}e^{f_{j}})$ and $\lim\cosh l_{ik}=c^{\prime}_{j}$. Then
$L_{1}\rightarrow$ $\frac{c^{\prime}_{i}e^{f_{j}}}{\sqrt{(1+\xi_{ij}e^{f_{i}+f_{j}})^{2}-1}\cdot\sqrt{(c^{\prime}_{j})^{2}-1}}$
$=$ $\frac{c^{\prime}_{i}e^{f_{j}}}{\sqrt{\xi_{ij}}\sqrt{(c^{\prime}_{j})^{2}-1}e^{f_{j}}}\cdot\frac{1}{\sqrt{\xi_{ij}e^{2f_{i}}+2e^{f_{i}-f_{j}}}}\rightarrow+\infty.$
And $L_{2}>0$. Thus $\cosh\theta^{jk}_{i}=L_{1}+L_{2}\rightarrow+\infty$, so $\theta^{jk}_{i}\rightarrow+\infty$.
For the case $\mathrm{(iii)}$, $\lim f_{j}=c_{2},\ \lim f_{k}=c_{3}$. By (57), we have $\lim\cosh l_{jk}:=c^{\prime\prime}_{i}$ and
$\lim\cosh l_{ij}=\begin{cases}1,\ &\text{if}\ \alpha_{j}=0,\\ c^{\prime\prime}_{k},\ &\text{if}\ \alpha_{j}=1,\end{cases}$
where $c^{\prime\prime}_{i},c^{\prime\prime}_{k}$ are positive constants. If $\alpha_{j}=0$, then $l_{ij}\rightarrow 0$ and hence $L_{2}\rightarrow+\infty$. This implies $\theta^{jk}_{i}\rightarrow+\infty$. If $\alpha_{j}=1$, then by the assumption, $\alpha_{k}=0$. According to the arguments in the case $\mathrm{(ii)}$, we still have $\theta^{jk}_{i}\rightarrow+\infty$.
For the case $\mathrm{(iv)}$, $\lim f_{j}=+\infty,\ \lim f_{k}=-\infty$. By (57), we have $\lim\cosh l_{ik}=1$ and hence $l_{ik}\rightarrow 0$. Then $\cosh\theta^{jk}_{i}>L_{2}\geq\frac{\cosh l_{ik}}{\sinh l_{ik}}\rightarrow+\infty$ and then $\theta^{jk}_{i}\rightarrow+\infty$. This completes the proof. Q.E.D.
Remark 6.9.
According to the proof of the case $\mathrm{(iii)}$ in Lemma 6.8, if $\alpha_{j}=\alpha_{k}=1$, $\lim\cosh l_{ij}=c^{\prime\prime}_{k}$, $\lim\cosh l_{jk}=c^{\prime\prime}_{i}$ and $\lim\cosh l_{ik}=c^{\prime}_{j}$, then $\cosh\theta^{jk}_{i}$, $\cosh\theta^{ik}_{j}$ and $\cosh\theta^{ij}_{k}$ all tend to finite constants, which are far away from 1 and +$\infty$+ ∞.
Lemma 6.10.
Under the same assumptions as in Theorem 6.5. For any constants $c_{1},c_{2},c_{3}\in\mathbb{R}$, if one of the following conditions is satisfied
(i):
$\lim f_{j}=+\infty,\ \lim f_{i}=c_{1},\ \lim f_{k}=+\infty$;
(ii):
$\lim f_{j}=+\infty,\ \lim f_{i}=c_{2},\ \lim f_{k}=c_{3}$.
Then $\lim\theta^{ik}_{j}(f_{i},f_{j},f_{k})=0$.
Proof.
For the case $\mathrm{(i)}$, by (56), we have $\lim\cosh l_{ij}:=\lim(c_{k}e^{f_{j}})$, $\lim\cosh l_{ik}:=c_{j}e^{f_{k}}$ and $\lim\cosh l_{jk}:=\lim(c_{i}e^{f_{j}+f_{k}})$, where $c_{i},c_{j},c_{k}$ are positive constants. Then
$\cosh\theta^{ik}_{j}=\frac{\cosh l_{ik}+\cosh l_{ij}\cosh l_{jk}}{\sinh l_{ij}\sinh l_{jk}}\rightarrow\frac{c_{j}e^{f_{k}}}{c_{k}e^{f_{j}}\cdot c_{i}e^{f_{j}+f_{k}}}+1\rightarrow 1,$
which implies $\theta^{ik}_{j}\rightarrow 0$. The proof of the case $\mathrm{(ii)}$ is similar, so we omit it here. Q.E.D.
Proof of Theorem 6.5: Let the image of $K$ is $X$. The openness of $X$ in >$\mathbb{R}^{N}_{>0}$ follows from the injectivity of the map $f$. To prove the closeness of $X$ in >$\mathbb{R}^{N}_{>0}$, we prove that if a sequence $f^{(m)}$ in $\Omega(\xi)=\bigcap_{\{ijk\}\in F}\Omega_{ijk}(\xi)$ satisfies $\lim_{m\rightarrow+\infty}K^{(m)}\in\mathbb{R}^{N}_{>0}$, then there exists a subsequence, say $f^{(m)}$, such that $\lim_{m\rightarrow+\infty}f^{(m)}$ is in $\Omega(\xi)$.
Suppose otherwise, there exists a subsequence $f^{(m)}$, such that $\lim_{m\rightarrow+\infty}f^{(m)}$ is in the boundary of $\Omega(\xi)$. The boundary of the admissible space $\Omega(\xi)$ in $[-\infty,+\infty]^{N}$ consists of the following three parts:
(i):
There exists an ideal edge $\{rs\}\in E$ such that $l_{rs}=0$.
(ii):
There exists $i\in B$ in a right-angled hyperbolic hexagon with edge lengths given by (46), such that $f_{i}^{(m)}\rightarrow\pm\infty$.
(iii):
There exists $j\in B$ in a right-angled hyperbolic hexagon with edge lengths given by (46), such that $f_{j}^{(m)}\rightarrow\pm\infty$.
(iv):
There exists $k\in B$ in a right-angled hyperbolic hexagon with edge lengths given by (46), such that $f_{k}^{(m)}\rightarrow\pm\infty$.
For the case $\mathrm{(i)}$, in a right-angled hyperbolic hexagon $\{ijk\}\in F$ with lengths $l_{ij},l_{jk},l_{ik}$ and opposite lengths of boundaries $\theta^{ij}_{k},\theta^{jk}_{i},\theta^{ik}_{j}$. Without loss of generality, we assume $l_{ij}\rightarrow 0$. By the cosine law,
$\cosh\theta^{ik}_{j}=\frac{\cosh l_{ik}+\cosh l_{ij}\cosh l_{jk}}{\sinh l_{ij}\sinh l_{jk}}>\frac{\cosh l_{ij}\cosh l_{jk}}{\sinh l_{ij}\sinh l_{jk}}\geq\frac{\cosh l_{ij}}{\sinh l_{ij}}\rightarrow+\infty,$
which contradicts the assumption that $\lim_{m\rightarrow+\infty}K^{(m)}\in\mathbb{R}^{N}_{>0}$.
We will prove that the above cases $\mathrm{(ii)}$, $\mathrm{(iii)}$ and $\mathrm{(iv)}$ can not hold on $\mathcal{U}^{(n)}_{ijk}(\xi)$, $n=1,2,4,5.$
(1): For $(0,0,0)$ type admissible space $\mathcal{U}^{(1)}_{ijk}(\xi)$ in (48), we have $u_{j}+u_{k}>\log\frac{2}{\xi_{jk}}$. By Remark 2.23, $u_{j}=f_{j}$ and $u_{k}=f_{k}$, then $f_{j}+f_{k}$ has a lower bound. Hence, $f_{j},\ f_{k}$ can not tend to $\infty$- ∞ simultaneously.
[…]
(2): For $(1,0,0)$ type admissible space $\mathcal{U}^{(2)}_{ijk}(\xi)$ in (49), we have $u_{j}+u_{k}>\log\frac{2}{\xi_{jk}}$. The proof is the same as that of $(0,0,0)$ type admissible space $\mathcal{U}^{(1)}_{ijk}(\xi)$, so we omit it here.
(3): For $(0,0,1)$ type admissible space $\mathcal{U}^{(4)}_{ijk}(\xi)$ in (50), the positions of $j$ and $k$ are not interchangeable.
(i):
In the admissible space (51), we have $u_{j}+u_{k}>\log\frac{1}{\xi_{jk}}$. By (47) and Remark 2.23, $u_{j}=f_{j}\in\mathbb{R}$ and $u_{k}=-\mathrm{arcsinh}(e^{-f_{k}})\in\mathbb{R}_{0}$, then $f_{j}$ can not tend to $\infty$- ∞.
[…]
If $f_{j}^{(m)}\rightarrow+\infty$ and $f_{k}^{(m)}\rightarrow-\infty$, by Lemma 6.7 (ii), then $(\theta^{ij}_{k})^{(m)}\rightarrow+\infty$ and hence $K^{(m)}_{k}\rightarrow+\infty$. If $f_{j}^{(m)}\rightarrow+\infty$ and $f_{k}^{(m)}\nrightarrow-\infty$, by Lemma 6.10, then $(\theta^{ik}_{j})^{(m)}\rightarrow 0$ and hence $K^{(m)}_{j}\rightarrow 0$. Hence, the subsequence $f^{(m)}_{j}$ can not tend to +$\infty$+ ∞ in the admissible space (51).
[…]
(ii):
In the admissible space (52), we have $u_{j}+u_{k}>\log\frac{1}{\xi_{jk}},\ u_{i}+u_{k}>\log(-\xi_{ik})$. According to the case $\mathrm{(i)}$, $f_{j}$ can not tend to $\infty$- ∞. Similarly, combining (47) and Remark 2.23 gives $u_{i}=-f_{i}$. Then $f_{i}$ can not tend to +$\infty$+ ∞.
If $f_{i}^{(m)}\rightarrow-\infty$, by Lemma 6.8, then $(\theta^{jk}_{i})^{(m)}\rightarrow+\infty$ and hence $K^{(m)}_{i}\rightarrow+\infty$. Hence, the subsequence $f^{(m)}_{i}$ can not tend to $\infty$- ∞ in the admissible space (52).
If $f_{j}^{(m)}\rightarrow+\infty$ and $f_{k}^{(m)}\rightarrow-\infty$, then $f_{i}^{(m)}\rightarrow-\infty$ to ensure $u_{i}+u_{k}>\log(-\xi_{ik})$. This case has been included in the above arguments. If $f_{j}^{(m)}\rightarrow+\infty$ and $f_{k}^{(m)}\nrightarrow-\infty$, by Lemma 6.10, then $(\theta^{ik}_{j})^{(m)}\rightarrow 0$ and hence $K^{(m)}_{j}\rightarrow 0$. Hence, the subsequence $f^{(m)}_{j}$ can not tend to +$\infty$+ ∞ in the admissible space (52).
[…]
(4): For $(1,0,1)$ type admissible space $\mathcal{U}^{(5)}_{ijk}(\xi)$ in (53), we have $u_{j}+u_{k}>\log\frac{1}{\xi_{jk}}$. The proof is the same as that of $(0,0,1)$ type admissible space (51), so we omit it here. Q.E.D.
Remark 6.11.
For $(0,1,1)$ type admissible space $\mathcal{U}^{(10)}_{ijk}(\xi)$ in (54) and $(1,1,1)$ type admissible space $\mathcal{U}^{(11)}_{ijk}(\xi)$ in (55), we are unable to prove the image of $K$ is >$\mathbb{R}^{N}_{>0}$. The reasons follow from Remark 6.9.
Appendix A Proofs of Lemmas
For simplicity, we set $h_{t}=h_{rs},\ q_{t}=q_{rs}$ and assume $||v_{i}||=||v_{j}||=||v_{k}||=1$. Please refer to Figure 3.
Lemma A.1.
If $c_{ijk}$ is time-like, then
$\sinh q_{r}=\cosh h_{s}\cdot\sinh d_{rt}=\cosh h_{t}\cdot\sinh d_{rs},$
$\sinh h_{r}=\cosh q_{s}\cdot\sinh\theta_{rt}=\cosh q_{t}\cdot\sinh\theta_{rs}.$
Proof.
Without loss of generality, we assume $r=i,s=j,t=k$.
In the generalized right-angled hyperbolic triangle $c_{ijk}c_{ik}v_{i}$, by Proposition 2.1 we have $-v_{i}\ast c_{ijk}=(c_{ik}\ast c_{ijk})(v_{i}\ast c_{ik})$. By Proposition 2.2 $\mathrm{(i)}$, $c_{ik}\ast c_{ijk}=-\cosh h_{j}$. Moreover, by Proposition 2.2 $\mathrm{(ii)}$, if $c_{ik}$ and $v_{i}$ are on same sides of the hyperplane $v_{i}^{\bot}$, then
$v_{i}\ast c_{ik}=+\sinh(d_{\mathbb{H}}(c_{ik},v_{i}^{\bot}))=+\sinh(-d_{ik})=-\sinh d_{ik};$
if $c_{ik}$ and $v_{i}$ are on opposite sides of the hyperplane $v_{i}^{\bot}$, then
$v_{i}\ast c_{ik}=-\sinh(d_{\mathbb{H}}(c_{ik},v_{i}^{\bot}))=-\sinh(+d_{ik})=-\sinh d_{ik}.$
Thus $v_{i}\ast c_{ik}=-\sinh d_{ik}$. Similarly, if $c_{ijk}$ and $v_{i}$ are on same sides of the hyperplane $v_{i}^{\bot}$, then $v_{i}\ast c_{ijk}=+\sinh(d_{\mathbb{H}}(c_{ijk},v_{i}^{\bot}))=+\sinh(-q_{i})=-\sinh q_{i}$; if $c_{ijk}$ and $v_{i}$ are on opposite sides of the hyperplane $v_{i}^{\bot}$, then $v_{i}\ast c_{ijk}=-\sinh(d_{\mathbb{H}}(c_{ijk},v_{i}^{\bot}))=-\sinh(+q_{i})=-\sinh q_{i}$. Thus $v_{i}\ast c_{ijk}=-\sinh q_{i}$. Then $\sinh q_{i}=\cosh h_{j}\cdot\sinh d_{ik}$.
In the generalized right-angled hyperbolic triangle $c_{ijk}c_{ij}v_{i}$, we have $-v_{i}\ast c_{ijk}=(c_{ij}\ast c_{ijk})(v_{i}\ast c_{ij})=-\cosh h_{k}\cdot(v_{i}\ast c_{jk})$. Similar arguments can imply $v_{i}\ast c_{ij}=-\sinh d_{ij}$. Then $\sinh q_{i}=\cosh h_{k}\cdot\sinh d_{ij}$.
In the generalized right-angled hyperbolic triangle $c_{ijk}c^{\prime}_{ik}v^{\prime}_{i}$, we have $-v^{\prime}_{i}\ast c_{ijk}=(c^{\prime}_{ik}\ast c_{ijk})(v^{\prime}_{i}\ast c^{\prime}_{ik})=-\cosh q_{j}\cdot(v^{\prime}_{i}\ast c^{\prime}_{ik})$. Similarly, $v^{\prime}_{i}\ast c_{ijk}=-\sinh h_{i}$ and $v^{\prime}_{i}\ast c^{\prime}_{ik}=-\sinh\theta_{ik}$. Then $\sinh h_{i}=\cosh q_{j}\cdot\sinh\theta_{ik}$.
In the generalized right-angled hyperbolic triangle $c_{ijk}c^{\prime}_{ij}v^{\prime}_{i}$, we have $-v^{\prime}_{i}\ast c_{ijk}=(c^{\prime}_{ij}\ast c_{ijk})(v^{\prime}_{i}\ast c^{\prime}_{ij})=-\cosh q_{k}\cdot(v^{\prime}_{i}\ast c^{\prime}_{ij})$. Similarly, $v^{\prime}_{i}\ast c_{ijk}=-\sinh h_{i}$ and $v^{\prime}_{i}\ast c^{\prime}_{ij}=-\sinh\theta_{ij}$. Then $\sinh h_{i}=\cosh q_{k}\cdot\sinh\theta_{ij}$. Q.E.D.
Proposition A.2.
If $c_{ijk}\in\mathbb{R}^{3}$ lies in the open domain $D_{\mathrm{V}}$, then
$\cosh q_{i}=-\sinh h_{j}\cdot\sinh d_{ik}=-\sinh h_{k}\cdot\sinh d_{ij},$
$\cosh q_{j}=\sinh h_{i}\cdot\sinh d_{jk}=\sinh h_{k}\cdot\sinh d_{ji},$
$\cosh q_{k}=\sinh h_{i}\cdot\sinh d_{kj}=\sinh h_{j}\cdot\sinh d_{ki},$
$\cosh h_{i}=\sinh q_{j}\cdot\sinh\theta_{ik}=\sinh q_{k}\cdot\sinh\theta_{ij},$
$\cosh h_{j}=-\sinh q_{i}\cdot\sinh\theta_{jk}=\sinh q_{k}\cdot\sinh\theta_{ji},$
$\cosh h_{k}=-\sinh q_{i}\cdot\sinh\theta_{kj}=\sinh q_{j}\cdot\sinh\theta_{ki}.$
If $c_{ijk}\in\mathbb{R}^{3}$ lies in the open domain $D_{\mathrm{II}}$, then we swap $h$ and $q$, and $\theta$ and $d$.
Proof.
(1):
Suppose $c_{ijk}$ lies in the domain $D_{\mathrm{V}}$. In the generalized right-angled hyperbolic triangle $c_{ijk}c_{ik}v_{i}$, by Proposition 2.1, we have $-v_{i}\ast c_{ijk}=(c_{ik}\ast c_{ijk})(v_{i}\ast c_{ik})$. By Proposition 2.2 $\mathrm{(iii)}$, $v_{i}\ast c_{ijk}=\cosh(d_{\mathbb{H}}(v_{i}^{\bot},c_{ijk}^{\bot}))=\cosh(-q_{i})=\cosh q_{i}$. Moreover, by Proposition 2.2 $\mathrm{(ii)}$, $v_{i}\ast c_{ik}=+\sinh d_{\mathbb{H}}(v_{i}^{\bot},c_{ik})=+\sinh(-d_{ik})=-\sinh d_{ik}$. If $c_{ik}$ and $c_{ijk}$ are on same sides of the hyperplane $c_{ijk}^{\bot}$, then
$c_{ik}\ast c_{ijk}=+\sinh(d_{\mathbb{H}}(c_{ik},c_{ijk}^{\bot}))=+\sinh(-h_{j})=-\sinh h_{j}.$
If $c_{ik}$ and $c_{ijk}$ are on opposite sides of the hyperplane $c_{ijk}^{\bot}$, then
$c_{ik}\ast c_{ijk}=-\sinh(d_{\mathbb{H}}(c_{ik},c_{ijk}^{\bot}))=-\sinh h_{j}.$
Thus $c_{ik}\ast c_{ijk}=-\sinh h_{j}$. Then $\cosh q_{i}=-\sinh h_{j}\cdot\sinh d_{ik}$. By Lemma 2.5, it is easy to check that $d_{ik}<0$. Hence $h_{j}>0$. Then $d_{\mathbb{H}}(c_{ik},c_{ijk}^{\bot})=h_{j}$, which implies that $c_{ik}$ and $c_{ijk}$ are on opposite sides of the hyperplane $c_{ijk}^{\bot}$.
In the generalized right-angled hyperbolic triangle $c_{ijk}c_{ij}v_{i}$, we have $-v_{i}\ast c_{ijk}=(c_{ij}\ast c_{ijk})(v_{i}\ast c_{ij})=-\cosh q_{i}$. Similarly, $v_{i}\ast c_{ij}=-\sinh d_{ij}$. If $c_{ij}$ and $c_{ijk}$ are on same sides of the hyperplane $c_{ijk}^{\bot}$, then $c_{ij}\ast c_{ijk}=+\sinh(d_{\mathbb{H}}(c_{ij},c_{ijk}^{\bot}))=+\sinh(-h_{k})=-\sinh h_{k}$. If $c_{ij}$ and $c_{ijk}$ are on opposite sides of the hyperplane $c_{ijk}^{\bot}$, then $c_{ij}\ast c_{ijk}=-\sinh(d_{\mathbb{H}}(c_{ij},c_{ijk}^{\bot}))=-\sinh h_{k}$. Thus $c_{ij}\ast c_{ijk}=-\sinh h_{k}$. Then $\cosh q_{i}=-\sinh h_{k}\cdot\sinh d_{ij}$. Similarly, $d_{ij}<0$ by Lemma 2.5 and hence $h_{k}>0$. So $d_{\mathbb{H}}(c_{ij},c_{ijk}^{\bot})=h_{k}$, which implies that $c_{ij}$ and $c_{ijk}$ are on opposite sides of the hyperplane $c_{ijk}^{\bot}$.
In the generalized right-angled hyperbolic triangle $c_{ijk}c_{jk}v_{j}$, we have $-v_{j}\ast c_{ijk}=(c_{jk}\ast c_{ijk})(v_{j}\ast c_{jk})$. By Proposition 2.2 $\mathrm{(iii)}$, $v_{j}\ast c_{ijk}=-\cosh(d_{\mathbb{H}}(v_{j}^{\bot},c_{ijk}^{\bot}))=-\cosh q_{j}$. Moreover, by Proposition 2.2 $\mathrm{(ii)}$, $c_{jk}\ast c_{ijk}=-\sinh(d_{\mathbb{H}}(c_{jk},c_{ijk}^{\bot}))=-\sinh h_{i}$. Similarly, $v_{j}\ast c_{jk}=-\sinh d_{jk}$. Then $\cosh q_{j}=\sinh h_{i}\cdot\sinh d_{jk}$.
In the generalized right-angled hyperbolic triangle $c_{ijk}c_{ij}v_{j}$, we have $-v_{j}\ast c_{ijk}=(c_{ij}\ast c_{ijk})(v_{j}\ast c_{ij})$. Since $v_{j}\ast c_{ijk}=-\cosh q_{j}$, $v_{j}\ast c_{ij}=-\sinh d_{ji}$ and $c_{ij}\ast c_{ijk}=-\sinh h_{k}$, then $\cosh q_{j}=\sinh h_{k}\cdot\sinh d_{ji}$.
In the generalized right-angled hyperbolic triangle $c_{ijk}c_{jk}v_{k}$, we have $-v_{k}\ast c_{ijk}=(c_{jk}\ast c_{ijk})(v_{k}\ast c_{jk})$. Since $v_{k}\ast c_{ijk}=-\cosh q_{k}$, $v_{k}\ast c_{jk}=-\sinh d_{kj}$ and $c_{jk}\ast c_{ijk}=-\sinh h_{i}$, then $\cosh q_{k}=\sinh h_{i}\cdot\sinh d_{kj}$.
In the generalized right-angled hyperbolic triangle $c_{ijk}c_{ik}v_{k}$, we have $-v_{k}\ast c_{ijk}=(c_{ik}\ast c_{ijk})(v_{k}\ast c_{ik})$. Since $v_{k}\ast c_{ijk}=-\cosh q_{k}$, $v_{k}\ast c_{ik}=-\sinh d_{ki}$ and $c_{ik}\ast c_{ijk}=-\sinh h_{j}$, then $\cosh q_{k}=\sinh h_{j}\cdot\sinh d_{ki}$.
In the generalized right-angled hyperbolic triangle $c_{ijk}c^{\prime}_{ik}v^{\prime}_{i}$, we have $-v^{\prime}_{i}\ast c_{ijk}=(c^{\prime}_{ik}\ast c_{ijk})(v^{\prime}_{i}\ast c^{\prime}_{ik})$. Since $v^{\prime}_{i}\ast c_{ijk}=-\cosh h_{i}$, $v^{\prime}_{i}\ast c^{\prime}_{ik}=-\sinh\theta_{ik}$ and $c^{\prime}_{ik}\ast c_{ijk}=-\sinh q_{j}$, then $\cosh h_{i}=\sinh q_{j}\cdot\sinh\theta_{ik}$.
In the generalized right-angled hyperbolic triangle $c_{ijk}c^{\prime}_{ij}v^{\prime}_{i}$, we have $-v^{\prime}_{i}\ast c_{ijk}=(c^{\prime}_{ij}\ast c_{ijk})(v^{\prime}_{i}\ast c^{\prime}_{ij})$. Since $v^{\prime}_{i}\ast c_{ijk}=-\cosh h_{i}$, $v^{\prime}_{i}\ast c^{\prime}_{ij}=-\sinh\theta_{ij}$ and $c^{\prime}_{ij}\ast c_{ijk}=-\sinh q_{k}$, then $\cosh h_{i}=\sinh q_{k}\cdot\sinh\theta_{ij}$.
In the generalized right-angled hyperbolic triangle $c_{ijk}c^{\prime}_{jk}v^{\prime}_{j}$, we have $-v^{\prime}_{j}\ast c_{ijk}=(c^{\prime}_{jk}\ast c_{ijk})(v^{\prime}_{j}\ast c^{\prime}_{jk})$. By Proposition 2.2 $\mathrm{(iii)}$, $v^{\prime}_{j}\ast c_{ijk}=-\cosh(d_{\mathbb{H}}((v^{\prime}_{j})^{\bot},c_{ijk}^{\bot}))=-\cosh h_{j}$. Since $c^{\prime}_{jk}$ and $c_{ijk}$ are on opposite sides of the hyperplane $c_{ijk}^{\bot}$, then $c^{\prime}_{jk}\ast c_{ijk}=-\sinh(d_{\mathbb{H}}(c^{\prime}_{jk},c_{ijk}^{\bot}))=-\sinh(-q_{i})=+\sinh q_{i}$. Due to $v^{\prime}_{j}\ast c^{\prime}_{jk}=-\sinh\theta_{jk}$, then $\cosh h_{j}=-\sinh q_{i}\cdot\sinh\theta_{jk}$.
In the generalized right-angled hyperbolic triangle $c_{ijk}c^{\prime}_{ij}v^{\prime}_{j}$, we have $-v^{\prime}_{j}\ast c_{ijk}=(c^{\prime}_{ij}\ast c_{ijk})(v^{\prime}_{j}\ast c^{\prime}_{ij})$. Note that $v^{\prime}_{j}\ast c_{ijk}=-\cosh h_{j}$, $v^{\prime}_{j}\ast c^{\prime}_{ij}=-\sinh\theta_{ji}$ and $c^{\prime}_{ij}\ast c_{ijk}=-\sinh q_{k}$, then $\cosh h_{j}=\sinh q_{k}\cdot\sinh\theta_{ji}$.
In the generalized right-angled hyperbolic triangle $c_{ijk}c^{\prime}_{jk}v^{\prime}_{k}$, we have $-v^{\prime}_{k}\ast c_{ijk}=(c^{\prime}_{jk}\ast c_{ijk})(v^{\prime}_{k}\ast c^{\prime}_{jk})$. Since $v^{\prime}_{k}\ast c_{ijk}=-\cosh h_{k}$, $v^{\prime}_{k}\ast c^{\prime}_{jk}=-\sinh\theta_{kj}$ and $c^{\prime}_{jk}\ast c_{ijk}=\sinh q_{i}$, then $\cosh h_{k}=-\sinh q_{i}\cdot\sinh\theta_{kj}$.
In the generalized right-angled hyperbolic triangle $c_{ijk}c^{\prime}_{ik}v^{\prime}_{k}$, we have $-v^{\prime}_{k}\ast c_{ijk}=(c^{\prime}_{ik}\ast c_{ijk})(v^{\prime}_{k}\ast c^{\prime}_{ik})$. Since $v^{\prime}_{k}\ast c_{ijk}=-\cosh h_{k}$, $v^{\prime}_{k}\ast c^{\prime}_{ik}=-\sinh\theta_{ki}$ and $c^{\prime}_{ik}\ast c_{ijk}=-\sinh q_{j}$, then $\cosh h_{k}=\sinh q_{j}\cdot\sinh\theta_{ki}$.
(2):
Suppose $c_{ijk}$ lies in the domain $D_{\mathrm{II}}$. In the generalized right-angled hyperbolic triangle $c_{ijk}c_{ik}v_{i}$, we have $-v_{i}\ast c_{ijk}=(c_{ik}\ast c_{ijk})(v_{i}\ast c_{ik})$. By Proposition 2.2 $\mathrm{(iii)}$, $v_{i}\ast c_{ijk}=-\cosh q_{i}$ and $c_{ik}\ast c_{ijk}=-\sinh h_{j}$. Since $v_{i}\ast c_{ik}=-\sinh d_{\mathbb{H}}(v_{i}^{\bot},c_{ik})=-\sinh(+d_{ik})=-\sinh d_{ik}$, then $\cosh q_{i}=\sinh h_{j}\cdot\sinh d_{ik}$.
In the generalized right-angled hyperbolic triangle $c_{ijk}c_{ij}v_{i}$, we have $-v_{i}\ast c_{ijk}=(c_{ij}\ast c_{ijk})(v_{i}\ast c_{ij})=\cosh q_{i}$. Similarly, $v_{i}\ast c_{ij}=-\sinh d_{ij}$ and $c_{ij}\ast c_{ijk}=-\sinh h_{k}$. Then $\cosh q_{i}=\sinh h_{k}\cdot\sinh d_{ij}$.
In the generalized right-angled hyperbolic triangle $c_{ijk}c_{jk}v_{j}$, we have $-v_{j}\ast c_{ijk}=(c_{jk}\ast c_{ijk})(v_{j}\ast c_{jk})$. By Proposition 2.2 $\mathrm{(iii)}$, then $v_{j}\ast c_{ijk}=-\cosh(d_{\mathbb{H}}(v_{j}^{\bot},c_{ijk}^{\bot}))=-\cosh q_{j}$. Since $c_{jk}$ and $c_{ijk}$ are on opposite sides of the hyperplane $c_{ijk}^{\bot}$, then $c_{jk}\ast c_{ijk}=-\sinh(d_{\mathbb{H}}(c_{jk},c_{ijk}^{\bot}))=-\sinh(-h_{i})=+\sinh h_{i}$. Due to $v_{j}\ast c_{jk}=-\sinh d_{jk}$, then $\cosh q_{j}=-\sinh h_{i}\cdot\sinh d_{jk}$.
In the generalized right-angled hyperbolic triangle $c_{ijk}c_{ij}v_{j}$, we have $-v_{j}\ast c_{ijk}=(c_{ij}\ast c_{ijk})(v_{j}\ast c_{ij})$. Since $v_{j}\ast c_{ijk}=-\cosh q_{j}$, $v_{j}\ast c_{ij}=-\sinh d_{ji}$ and $c_{ij}\ast c_{ijk}=-\sinh h_{k}$, then $\cosh q_{j}=\sinh h_{k}\cdot\sinh d_{ji}$.
In the generalized right-angled hyperbolic triangle $c_{ijk}c_{jk}v_{k}$, we have $-v_{k}\ast c_{ijk}=(c_{jk}\ast c_{ijk})(v_{k}\ast c_{jk})$. Since $v_{k}\ast c_{ijk}=-\cosh q_{k}$, $v_{k}\ast c_{jk}=-\sinh d_{kj}$ and $c_{jk}\ast c_{ijk}=\sinh h_{i}$, then $\cosh q_{k}=-\sinh h_{i}\cdot\sinh d_{kj}$.
In the generalized right-angled hyperbolic triangle $c_{ijk}c_{ik}v_{k}$, we have $-v_{k}\ast c_{ijk}=(c_{ik}\ast c_{ijk})(v_{k}\ast c_{ik})$. Since $v_{k}\ast c_{ijk}=-\cosh q_{k}$, $v_{k}\ast c_{ik}=-\sinh d_{ki}$ and $c_{ik}\ast c_{ijk}=-\sinh h_{j}$, then $\cosh q_{k}=\sinh h_{j}\cdot\sinh d_{ki}$.
In the generalized right-angled hyperbolic triangle $c_{ijk}c^{\prime}_{ik}v^{\prime}_{i}$, we have $-v^{\prime}_{i}\ast c_{ijk}=(c^{\prime}_{ik}\ast c_{ijk})(v^{\prime}_{i}\ast c^{\prime}_{ik})$. Since $v^{\prime}_{i}\ast c_{ijk}=\cosh h_{i}$ and $v^{\prime}_{i}\ast c^{\prime}_{ik}=-\sinh\theta_{ik}$. If $c^{\prime}_{ik}$ and $c_{ijk}$ are on same sides of the hyperplane $c_{ijk}^{\bot}$, then $c^{\prime}_{ik}\ast c_{ijk}=+\sinh(d_{\mathbb{H}}(c^{\prime}_{ik},c_{ijk}^{\bot}))=+\sinh(-q_{j})=-\sinh q_{j}$. If $c^{\prime}_{ik}$ and $c_{ijk}$ are on opposite sides of the hyperplane $c_{ijk}^{\bot}$, then $c^{\prime}_{ij}\ast c_{ijk}=-\sinh(d_{\mathbb{H}}(c^{\prime}_{ik},c_{ijk}^{\bot}))=-\sinh(+q_{j})=-\sinh q_{j}$. Thus $c^{\prime}_{ik}\ast c_{ijk}=-\sinh q_{j}$. Then $\cosh h_{i}=-\sinh q_{j}\cdot\sinh\theta_{ik}$. By Lemma 2.6, it is easy to check that $\theta_{ik}<0$. Hence $q_{j}>0$. Then $d_{\mathbb{H}}(c^{\prime}_{ik},c_{ijk}^{\bot})=q_{j}$, which implies that $c^{\prime}_{ik}$ and $c_{ijk}$ are on opposite sides of the hyperplane $c_{ijk}^{\bot}$.
In the generalized right-angled hyperbolic triangle $c_{ijk}c^{\prime}_{ij}v^{\prime}_{i}$, we have $-v^{\prime}_{i}\ast c_{ijk}=(c^{\prime}_{ij}\ast c_{ijk})(v^{\prime}_{i}\ast c^{\prime}_{ij})$. Since $v^{\prime}_{i}\ast c_{ijk}=\cosh h_{i}$ and $v^{\prime}_{i}\ast c^{\prime}_{ij}=-\sinh\theta_{ij}$. If $c^{\prime}_{ij}$ and $c_{ijk}$ are on same sides of the hyperplane $c_{ijk}^{\bot}$, then $c^{\prime}_{ij}\ast c_{ijk}=+\sinh(d_{\mathbb{H}}(c^{\prime}_{ij},c_{ijk}^{\bot}))=+\sinh(-q_{k})=-\sinh q_{k}$. If $c^{\prime}_{ij}$ and $c_{ijk}$ are on opposite sides of the hyperplane $c_{ijk}^{\bot}$, then $c^{\prime}_{ij}\ast c_{ijk}=-\sinh(d_{\mathbb{H}}(c^{\prime}_{ij},c_{ijk}^{\bot}))=-\sinh(+q_{k})=-\sinh q_{k}$. Thus $c^{\prime}_{ij}\ast c_{ijk}=-\sinh q_{k}$. Then $\cosh h_{i}=-\sinh q_{k}\cdot\sinh\theta_{ij}$. Similarly, $\theta_{ij}<0$ by Lemma 2.6 and hence $q_{k}>0$. Then $d_{\mathbb{H}}(c^{\prime}_{ij},c_{ijk}^{\bot})=q_{k}$, which implies that $c^{\prime}_{ij}$ and $c_{ijk}$ are on opposite sides of the hyperplane $c_{ijk}^{\bot}$.
In the generalized right-angled hyperbolic triangle $c_{ijk}c^{\prime}_{jk}v^{\prime}_{j}$, we have $-v^{\prime}_{j}\ast c_{ijk}=(c^{\prime}_{jk}\ast c_{ijk})(v^{\prime}_{j}\ast c^{\prime}_{jk})$. Since $v^{\prime}_{j}\ast c_{ijk}=-\cosh h_{j}$, $v^{\prime}_{j}\ast c^{\prime}_{jk}=-\sinh\theta_{jk}$ and $c^{\prime}_{jk}\ast c_{ijk}=-\sinh q_{i}$. Then $\cosh h_{j}=\sinh q_{i}\cdot\sinh\theta_{jk}$.
In the generalized right-angled hyperbolic triangle $c_{ijk}c^{\prime}_{ij}v^{\prime}_{j}$, we have $-v^{\prime}_{j}\ast c_{ijk}=(c^{\prime}_{ij}\ast c_{ijk})(v^{\prime}_{j}\ast c^{\prime}_{ij})$. Since $v^{\prime}_{j}\ast c_{ijk}=-\cosh h_{j}$, $v^{\prime}_{j}\ast c^{\prime}_{ij}=-\sinh\theta_{ji}$ and $c^{\prime}_{ij}\ast c_{ijk}=-\sinh q_{k}$, then $\cosh h_{j}=\sinh q_{k}\cdot\sinh\theta_{ji}$.
In the generalized right-angled hyperbolic triangle $c_{ijk}c^{\prime}_{jk}v^{\prime}_{k}$, we have $-v^{\prime}_{k}\ast c_{ijk}=(c^{\prime}_{jk}\ast c_{ijk})(v^{\prime}_{k}\ast c^{\prime}_{jk})$. Since $v^{\prime}_{k}\ast c_{ijk}=-\cosh h_{k}$, $v^{\prime}_{k}\ast c^{\prime}_{jk}=-\sinh\theta_{kj}$ and $c^{\prime}_{jk}\ast c_{ijk}=-\sinh q_{i}$, then $\cosh h_{k}=\sinh q_{i}\cdot\sinh\theta_{kj}$.
In the generalized right-angled hyperbolic triangle $c_{ijk}c^{\prime}_{ik}v^{\prime}_{k}$, we have $-v^{\prime}_{k}\ast c_{ijk}=(c^{\prime}_{ik}\ast c_{ijk})(v^{\prime}_{k}\ast c^{\prime}_{ik})$. Since $v^{\prime}_{k}\ast c_{ijk}=-\cosh h_{k}$, $v^{\prime}_{k}\ast c^{\prime}_{ik}=-\sinh\theta_{ki}$ and $c^{\prime}_{ik}\ast c_{ijk}=-\sinh q_{j}$, then $\cosh h_{k}=\sinh q_{j}\cdot\sinh\theta_{ki}$.
Q.E.D.
Proposition A.3.
Let $c_{ijk}\in\mathbb{R}^{3}$ lie in the open domain $D_{\mathrm{vi}}$, then
$\cosh q_{i}=\sinh h_{j}\cdot\sinh d_{ik}=-\sinh h_{k}\cdot\sinh d_{ij},$
$\cosh q_{j}=\sinh h_{i}\cdot\sinh d_{jk}=\sinh h_{k}\cdot\sinh d_{ji},$
$\cosh q_{k}=\sinh h_{i}\cdot\sinh d_{kj}=-\sinh h_{j}\cdot\sinh d_{ki},$
$\cosh h_{i}=\sinh q_{j}\cdot\sinh\theta_{ik}=\sinh q_{k}\cdot\sinh\theta_{ij},$
$\cosh h_{j}=\sinh q_{i}\cdot\sinh\theta_{jk}=-\sinh q_{k}\cdot\sinh\theta_{ji},$
$\cosh h_{k}=-\sinh q_{i}\cdot\sinh\theta_{kj}=\sinh q_{j}\cdot\sinh\theta_{ki}.$
Proof.
In the generalized right-angled hyperbolic triangle $c_{ijk}c_{ik}v_{i}$, we have $-v_{i}\ast c_{ijk}=(c_{ik}\ast c_{ijk})(v_{i}\ast c_{ik})$. By Proposition 2.2 $\mathrm{(iii)}$, $v_{i}\ast c_{ijk}=\cosh d_{\mathbb{H}}(v_{i}^{\bot},c_{ijk}^{\bot})=\cosh q_{i}$. By Proposition 2.2 $\mathrm{(ii)}$, $v_{i}\ast c_{ik}=+\sinh d_{\mathbb{H}}(v_{i}^{\bot},c_{ik})=+\sinh(-d_{ik})=-\sinh d_{ik}$. If $c_{ik}$ and $c_{ijk}$ are on same sides of the hyperplane $c_{ijk}^{\bot}$, then $c_{ik}\ast c_{ijk}=+\sinh(d_{\mathbb{H}}(c_{ik},c_{ijk}^{\bot}))=+\sinh h_{j}$; If $c_{ik}$ and $c_{ijk}$ are on opposite sides of the hyperplane $c_{ijk}^{\bot}$, then $c_{ik}\ast c_{ijk}=-\sinh(d_{\mathbb{H}}(c_{ik},c_{ijk}^{\bot}))=-\sinh(-h_{j})=+\sinh h_{j}$. Thus $c_{ik}\ast c_{ijk}=\sinh h_{j}$. Then $\cosh q_{i}=\sinh h_{j}\cdot\sinh d_{ik}$. By Lemma 2.5, it is easy to check that $d_{ik}<0$. Then $h_{j}<0$ and hence $d_{\mathbb{H}}(c_{ik},c_{ijk}^{\bot})=-h_{j}$, which implies that $c_{ik}$ and $c_{ijk}$ are on opposite sides of the hyperplane $c_{ijk}^{\bot}$.
In the generalized right-angled hyperbolic triangle $c_{ijk}c_{ij}v_{i}$, we have $-v_{i}\ast c_{ijk}=(c_{ij}\ast c_{ijk})(v_{i}\ast c_{ij})$. Similarly, $v_{i}\ast c_{ij}=-\sinh d_{ij}$ and $c_{ij}\ast c_{ijk}=-\sinh h_{k}$. Then $\cosh q_{i}=-\sinh h_{k}\cdot\sinh d_{ij}$.
In the generalized right-angled hyperbolic triangle $c_{ijk}c_{jk}v_{j}$, we have $-v_{j}\ast c_{ijk}=(c_{jk}\ast c_{ijk})(v_{j}\ast c_{jk})$. Since $v_{j}\ast c_{ijk}=-\cosh q_{j}$, $v_{j}\ast c_{jk}=-\sinh d_{jk}$ and $c_{jk}\ast c_{ijk}=-\sinh h_{i}$, then $\cosh q_{j}=\sinh h_{i}\cdot\sinh d_{jk}$.
In the generalized right-angled hyperbolic triangle $c_{ijk}c_{ij}v_{j}$, we have $-v_{j}\ast c_{ijk}=(c_{ij}\ast c_{ijk})(v_{j}\ast c_{ij})$. Since $v_{j}\ast c_{ijk}=-\cosh q_{j}$, $v_{j}\ast c_{ij}=-\sinh d_{ji}$ and $c_{ij}\ast c_{ijk}=-\sinh h_{k}$, then $\cosh q_{j}=\sinh h_{k}\cdot\sinh d_{ji}$.
In the generalized right-angled hyperbolic triangle $c_{ijk}c_{jk}v_{k}$, we have $-v_{k}\ast c_{ijk}=(c_{jk}\ast c_{ijk})(v_{k}\ast c_{jk})$. Since $v_{k}\ast c_{ijk}=-\cosh q_{k}$, $v_{k}\ast c_{jk}=-\sinh d_{kj}$ and $c_{jk}\ast c_{ijk}=-\sinh h_{i}$, then $\cosh q_{k}=\sinh h_{i}\cdot\sinh d_{kj}$.
In the generalized right-angled hyperbolic triangle $c_{ijk}c_{ik}v_{k}$, we have $-v_{k}\ast c_{ijk}=(c_{ik}\ast c_{ijk})(v_{k}\ast c_{ik})$. Since $v_{k}\ast c_{ijk}=-\cosh q_{k}$, $v_{k}\ast c_{ik}=-\sinh d_{ki}$ and $c_{ik}\ast c_{ijk}=\sinh h_{j}$, then $\cosh q_{k}=-\sinh h_{j}\cdot\sinh d_{ki}$.
In the generalized right-angled hyperbolic triangle $c_{ijk}c^{\prime}_{ik}v^{\prime}_{i}$, we have $-v^{\prime}_{i}\ast c_{ijk}=(c^{\prime}_{ik}\ast c_{ijk})(v^{\prime}_{i}\ast c^{\prime}_{ik})$. Since $v^{\prime}_{i}\ast c_{ijk}=-\cosh h_{i}$, $v^{\prime}_{i}\ast c^{\prime}_{ik}=-\sinh\theta_{ik}$ and $c^{\prime}_{ik}\ast c_{ijk}=-\sinh q_{j}$, then $\cosh h_{i}=\sinh q_{j}\cdot\sinh\theta_{ik}$.
In the generalized right-angled hyperbolic triangle $c_{ijk}c^{\prime}_{ij}v^{\prime}_{i}$, we have $-v^{\prime}_{i}\ast c_{ijk}=(c^{\prime}_{ij}\ast c_{ijk})(v^{\prime}_{i}\ast c^{\prime}_{ij})$. Since $v^{\prime}_{i}\ast c_{ijk}=-\cosh h_{i}$, $v^{\prime}_{i}\ast c^{\prime}_{ij}=-\sinh\theta_{ij}$ and $c^{\prime}_{ij}\ast c_{ijk}=-\sinh q_{k}$, then $\cosh h_{i}=\sinh q_{k}\cdot\sinh\theta_{ij}$.
In the generalized right-angled hyperbolic triangle $c_{ijk}c^{\prime}_{jk}v^{\prime}_{j}$, we have $-v^{\prime}_{j}\ast c_{ijk}=(c^{\prime}_{jk}\ast c_{ijk})(v^{\prime}_{j}\ast c^{\prime}_{jk})$. Since $v^{\prime}_{j}\ast c_{ijk}=\cosh h_{j}$ and $v^{\prime}_{j}\ast c^{\prime}_{jk}=-\sinh\theta_{jk}$. If $c^{\prime}_{jk}$ and $c_{ijk}$ are on same sides of the hyperplane $c_{ijk}^{\bot}$, then $c^{\prime}_{jk}\ast c_{ijk}=+\sinh(d_{\mathbb{H}}(c^{\prime}_{jk},c_{ijk}^{\bot}))=+\sinh q_{i}$; If $c^{\prime}_{jk}$ and $c_{ijk}$ are on opposite sides of the hyperplane $c_{ijk}^{\bot}$, then $c^{\prime}_{jk}\ast c_{ijk}=-\sinh(d_{\mathbb{H}}(c^{\prime}_{jk},c_{ijk}^{\bot}))=-\sinh(-q_{i})=+\sinh q_{i}$. Thus $c^{\prime}_{jk}\ast c_{ijk}=\sinh q_{i}$. Then $\cosh h_{j}=\sinh q_{i}\cdot\sinh\theta_{jk}$. By Lemma 2.6, it is easy to check that $\theta_{jk}<0$. Then $q_{i}<0$ and hence $d_{\mathbb{H}}(c^{\prime}_{jk},c_{ijk}^{\bot})=-q_{i}$, which implies that $c^{\prime}_{jk}$ and $c_{ijk}$ are on opposite sides of the hyperplane $c_{ijk}^{\bot}$.
In the generalized right-angled hyperbolic triangle $c_{ijk}c^{\prime}_{ij}v^{\prime}_{j}$, we have $-v^{\prime}_{j}\ast c_{ijk}=(c^{\prime}_{ij}\ast c_{ijk})(v^{\prime}_{j}\ast c^{\prime}_{ij})$. Since $v^{\prime}_{j}\ast c_{ijk}=\cosh h_{j}$, $v^{\prime}_{j}\ast c^{\prime}_{ij}=-\sinh\theta_{ji}$ and $c^{\prime}_{ij}\ast c_{ijk}=-\sinh q_{k}$, then $\cosh h_{j}=-\sinh q_{k}\cdot\sinh\theta_{ji}$.
In the generalized right-angled hyperbolic triangle $c_{ijk}c^{\prime}_{jk}v^{\prime}_{k}$, we have $-v^{\prime}_{k}\ast c_{ijk}=(c^{\prime}_{jk}\ast c_{ijk})(v^{\prime}_{k}\ast c^{\prime}_{jk})$. Since $v^{\prime}_{k}\ast c_{ijk}=-\cosh h_{k}$, $v^{\prime}_{k}\ast c^{\prime}_{jk}=-\sinh\theta_{kj}$ and $c^{\prime}_{jk}\ast c_{ijk}=\sinh q_{i}$, then $\cosh h_{k}=-\sinh q_{i}\cdot\sinh\theta_{kj}$.
In the generalized right-angled hyperbolic triangle $c_{ijk}c^{\prime}_{ik}v^{\prime}_{k}$, we have $-v^{\prime}_{k}\ast c_{ijk}=(c^{\prime}_{ik}\ast c_{ijk})(v^{\prime}_{k}\ast c^{\prime}_{ik})$. Since $v^{\prime}_{k}\ast c_{ijk}=-\cosh h_{k}$, $v^{\prime}_{k}\ast c^{\prime}_{ik}=-\sinh\theta_{ki}$ and $c^{\prime}_{ik}\ast c_{ijk}=-\sinh q_{j}$, then $\cosh h_{k}=\sinh q_{j}\cdot\sinh\theta_{ki}$. Q.E.D.

### v3

@@
 Abstract.
-This is a continuation of [16] studying the discrete conformal structures on surfaces with boundary, in which we gave a classification of the discrete conformal structures on surfaces with boundary.
-In this paper, we prove the rigidity and existence of these discrete conformal structures on surfaces with boundary.
+In [14], we introduced discrete conformal structures on surfaces with boundary via an axiomatic framework, and provided a classification of such discrete conformal structures.
+The present work focuses on the rigidity and existence of these discrete conformal structures on surfaces with boundary.
+As a direct consequence, the results by Guo-Luo in [5] and Guo in [4], which deal with the rigidity and existence of discrete conformal structures on surfaces with boundary, are extended to a very general context.
 Key words and phrases:
@@
 1. Introduction
-1.1. Basic definition and notations
-Let $\widetilde{\mathcal{T}}={(V,\widetilde{E},\widetilde{F})}$ be a triangulation of a closed surface $\widetilde{\Sigma}$, where $V,\widetilde{E},\widetilde{F}$ represent the sets of vertices, edges and faces respectively and $V$ is a finite subset of $\widetilde{\Sigma}$ with $|V|=N$.
-Denote $N(V)$ as a small open regular disjoint neighborhood of the union of all vertices $V$.
-Then $\Sigma=\widetilde{\Sigma}-N(V)$ is a compact surface with $N$ boundary components.
-The intersection $\mathcal{T}=\widetilde{\mathcal{T}}\cap\Sigma$ is called an ideal triangulation of the surface $\Sigma$.
-The intersections $E=\widetilde{E}\cap\Sigma$ and $F=\widetilde{F}\cap\Sigma$ are the sets of the ideal edges and ideal faces of $\Sigma$ in the ideal triangulation $\mathcal{T}$ respectively.
-The intersection of an ideal face and the boundary $\partial\Sigma$ is called a boundary arc.
-For simplicity, we use $B=\{1,2,...,N\}$ to denote the boundary components and label them as $i\in B$.
-We use $E_{+}$ to denote the set of oriented edges.
-Denote $\{ij\}\in E$ as the ideal edge between two adjacent boundary components $i,j\in B$.
-Label the ordered ideal edge as $(i,j)$.
-Denote $\{ijk\}\in F$ as the ideal face adjacent to the boundary components $i,j,k\in B$.
-In fact, the ideal face $\{ijk\}$ is a hexagon, which corresponds to the triangle $v_{i}v_{j}v_{k}\in\widetilde{F}$ in $\widetilde{\mathcal{T}}$.
-The sets of real valued functions on $B,E$ and $E_{+}$ are denoted by $\mathbb{R}^{N}$, $\mathbb{R}^{E}$ and $\mathbb{R}^{E_{+}}$ respectively.
-An edge length function associated to $\mathcal{T}$ is a vector $l\in\mathbb{R}^{E}_{>0}$ assigning each ideal edge $\{ij\}\in E$ a positive number $l_{ij}=l_{ji}$.
-For any three positive numbers, there exists a unique right-angled hyperbolic hexagon (up to isometry) with the lengths of three non-pairwise adjacent edges given by these three positive numbers [7].
-Hence, for any ideal face $\{ijk\}\in F$, there exists a unique right-angled hyperbolic hexagon whose three non-pairwise adjacent edges have lengths $l_{ij},l_{ik},l_{jk}$.
-Gluing all such right-angled hyperbolic hexagons isomorphically along the ideal edges in pairs, one can construct a hyperbolic surface with totally geodesic boundary from the ideal triangulation $\mathcal{T}$.
-Conversely, any ideally triangulated hyperbolic surface $(\Sigma,\mathcal{T})$ with totally geodesic boundary produces a function $l\in\mathbb{R}^{E}_{>0}$ with $l_{ij}$ given by the length of the shortest geodesic connecting the boundary components $i,j\in B$.
-The edge length function $l\in\mathbb{R}^{E}_{>0}$ is called a discrete hyperbolic metric on $(\Sigma,\mathcal{T})$.
+1.1. Basic definition and previous results
+Let $\widetilde{\mathcal{T}}=(V,\widetilde{E},\widetilde{F})$ be a triangulation of a closed surface $\widetilde{\Sigma}$, where $V$, $\widetilde{E}$, and $\widetilde{F}$ denote the sets of vertices, unoriented edges, and faces, respectively.
+Here $V$ is a finite subset of $\widetilde{\Sigma}$ with $|V|=N$.
+Let $N(V)$ be a small open regular neighborhood of the union of all vertices, chosen such that the neighborhoods around distinct vertices are disjoint.
+Then $\Sigma=\widetilde{\Sigma}\setminus N(V)$ is a compact surface with $N$ boundary components.
+The intersection $\mathcal{T}=\widetilde{\mathcal{T}}\cap\Sigma$ is referred to as an ideal triangulation of $\Sigma$.
+Let $E=\widetilde{E}\cap\Sigma$ and $F=\widetilde{F}\cap\Sigma$ denote the sets of unoriented ideal edges and ideal faces of $\Sigma$ with respect to $\mathcal{T}$, respectively.
+A boundary arc is defined as a connected component of the intersection between an ideal face and the boundary $\partial\Sigma$.
+For simplicity, we index the boundary components of $\Sigma$ by the set $B=\{1,2,\dots,N\}$, where each element $i\in B$ corresponds to a unique boundary component.
+Let $E_{+}$ denote the set of oriented ideal edges.
+For any two adjacent boundary components $i,j\in B$, the unoriented ideal edge connecting them is denoted by $\{ij\}\in E$, with the corresponding oriented ideal edge given by $(i,j)$.
+An ideal face $\{ijk\}\in F$ is adjacent to the boundary components $i,j,k\in B$.
+Specifically, such an ideal face is a hexagon, which corresponds to the triangle $v_{i}v_{j}v_{k}\in\widetilde{F}$ in $\widetilde{\mathcal{T}}$.
+The sets of real-valued functions on $B$, $E$, and $E_{+}$ are denoted by $\mathbb{R}^{N}$, $\mathbb{R}^{E}$, and $\mathbb{R}^{E_{+}}$, respectively.
+A edge length function associated to $\mathcal{T}$ is a vector $l\in\mathbb{R}^{E}_{>0}$ that assigns to each ideal edge $\{ij\}\in E$ a positive number $l_{ij}=l_{ji}$.
+For any ideal face $\{ijk\}\in F$, there exists a unique right-angled hyperbolic hexagon (up to isometry) whose three non-adjacent edges have lengths $l_{ij},l_{ik},l_{jk}$ (see [7, Theorem 3.5.14]).
+By gluing all such right-angled hyperbolic hexagons isometrically along the ideal edges in pairs, one can construct a hyperbolic surface with totally geodesic boundary from the ideal triangulation $\mathcal{T}$.
+Conversely, any ideally triangulated hyperbolic surface $(\Sigma,\mathcal{T})$ with totally geodesic boundary induces a function $l\in\mathbb{R}^{E}_{>0}$, where $l_{ij}$ is the length of the shortest geodesic connecting the boundary components $i,j\in B$.
+Such an edge length function $l\in\mathbb{R}^{E}_{>0}$ is referred to as a discrete hyperbolic metric on $(\Sigma,\mathcal{T})$.
 The length $K_{i}$ of the boundary component $i\in B$ is called the generalized combinatorial curvature of the discrete hyperbolic metric $l$ at $i$.
-Namely, the generalized combinatorial curvature $K_{i}$ is defined to be
-$K_{i}=\sum_{\{ijk\}\in F}\theta^{jk}_{i},$
-where the summation is taken over all the right-angled hyperbolic hexagons adjacent to $i$ and the generalized angle $\theta^{jk}_{i}$ is the length of the boundary arc of the right-angled hyperbolic hexagon $\{ijk\}$ at $i$.
-Motivated by Glickenstein’s work in [2] and Glickenstein-Thomas’ work in [3] for triangulated closed surfaces, we introduced the following two definitions of partial edge length and discrete conformal structures on an ideally triangulated surface with boundary.
+Specifically, the generalized combinatorial curvature $K_{i}$ is defined by
+(1) $K_{i}=\sum_{\{ijk\}\in F}\theta^{jk}_{i},$
+where the summation is taken over all right-angled hyperbolic hexagons adjacent to $i$, and the generalized angle $\theta^{jk}_{i}$ denotes the length of the boundary arc of the right-angled hyperbolic hexagon $\{ijk\}$ at $i$.
+Motivated by the work of Glickenstein [2] and that of Glickenstein-Thomas [3] on discrete conformal structures for triangulated closed surfaces, we introduced in [14] the following two definitions regarding partial edge lengths and discrete conformal structures on ideally triangulated surfaces with boundary.
 Definition 1.1 ([14], Definition 1.1).
 Let $(\Sigma,\mathcal{T})$ be an ideally triangulated surface with boundary.
-An assignment of partial edge lengths is a map $d\in\mathbb{R}^{E_{+}}$ such that $l_{ij}=d_{ij}+d_{ji}>0$ for every edge $\{ij\}\in E$ and
+An assignment of partial edge lengths is a map $d\in\mathbb{R}^{E_{+}}$ satisfying $l_{ij}=d_{ij}+d_{ji}>0$ for every edge $\{ij\}\in E$, and
 (2) $\sinh d_{ij}\sinh d_{jk}\sinh d_{ki}=\sinh d_{ji}\sinh d_{kj}\sinh d_{ik}$
@@
 Let $(\Sigma,\mathcal{T})$ be an ideally triangulated surface with boundary.
-A discrete conformal structure $d=d(f)$ on $(\Sigma,\mathcal{T})$ is a smooth map, sending a function $f\in\mathbb{R}^{N}$ defined on boundary components $B$ to a partial edge length function $d\in\mathbb{R}^{E_{+}}$, such that
+A discrete conformal structure $d=d(f)$ on $(\Sigma,\mathcal{T})$ is a smooth map that sends a function $f\in\mathbb{R}^{N}$ defined on the boundary components $B$ to a partial edge length function $d\in\mathbb{R}^{E_{+}}$, satisfying
 (3) $\frac{\partial l_{ij}}{\partial f_{i}}=\coth d_{ij}$
-for each $(i,j)\in E_{+}$ and
+for each $(i,j)\in E_{+}$, and
 (4) $\frac{\partial d_{ij}}{\partial f_{k}}=0$
@@
 Remark 1.3.
-In general, the partial edge lengths $d\in\mathbb{R}^{E_{+}}$ do not satisfy the symmetry condition, i.e., $d_{ij}\neq d_{ji}$.
-Denote $E_{ij}$ as the hyperbolic geodesic in the hyperbolic plane $\mathbb{H}^{2}$ extending the ideal edge $\{ij\}\in E$.
-The partial edge length $d_{ij}$ represents the signed distance from a point $c_{ij}$ in $E_{ij}$ to $i\in B$ and $d_{ji}$ represents the signed distance from $c_{ij}$ to $j\in B$.
-This point $c_{ij}$ is called the edge center of the ideal edge $\{ij\}$.
-Note that the signed distance $d_{ij}$ of $c_{ij}$ to $i$ is positive if $c_{ij}$ is on the same side of $i$ as $j$ along the hyperbolic geodesic $E_{ij}$, and negative otherwise.
+In general, the partial edge lengths $d\in\mathbb{R}^{E_{+}}$ may take negative values and do not satisfy the symmetry condition, i.e., $d_{ij}\neq d_{ji}$.
+Let $E_{ij}$ denote the hyperbolic geodesic in the hyperbolic plane $\mathbb{H}^{2}$ that extends the ideal edge $\{ij\}\in E$.
+The partial edge length $d_{ij}$ represents the signed distance from a point $c_{ij}\in E_{ij}$ to the boundary component $i\in B$, while $d_{ji}$ represents the signed distance from $c_{ij}$ to $j\in B$.
+This point $c_{ij}$ is referred to as the edge center of the ideal edge $\{ij\}$.
+Note that the signed distance $d_{ij}$ from $c_{ij}$ to $i$ is positive if $c_{ij}$ lies on the same side of $i$ as $j$ along the hyperbolic geodesic $E_{ij}$, and negative otherwise.
 Please refer to Figure 1.
@@
 Poincaré dual to the ideal edge $\{ij\}\in E$ with the position determined by $d_{ij}$ and $d_{ji}$ in the Klein model (dotted).
-The red lines represent the boundary arcs of the ideal face $\{ijk\}$.
-The black triangle $v_{i}v_{j}v_{k}$ is the corresponding hyper-ideal hyperbolic triangle.
+The red lines represent the boundary arcs of the ideal face $\{ijk\}$, and the black triangle $v_{i}v_{j}v_{k}$ is the corresponding hyper-ideal hyperbolic triangle.
 Remark 1.4.
-We briefly explain the motivation of Definition 1.1 for the partial edge length on surfaces with boundary.
-One can refer to [16] for more details.
-The hyperbolic geodesic orthogonal to the hyperbolic geodesic $E_{ij}$ through $c_{ij}$ is called the edge perpendicular, denoted by $P_{ij}$.
-We can identify the perpendicular $P_{ij}$ with a unique 2-dimensional time-like subspace $\widetilde{P}_{ij}$ of the Lorentzian space $\mathbb{R}^{3}$ such that $\widetilde{P}_{ij}\cap\mathbb{H}^{2}=P_{ij}$.
-Since the 2-dimensional time-like subspaces $\widetilde{P}_{ij}$ and $\widetilde{P}_{jk}$ associated to $P_{ij}$ and $P_{jk}$ intersect in a 1-dimensional subspace of $\mathbb{R}^{3}$, then the intersection $\widetilde{P}_{ij}\cap\widetilde{P}_{jk}$ corresponds to a point $q$ in the Klein model of hyperbolic plane $\mathbb{H}^{2}$.
-We view this point $q$ as the intersection point of the perpendiculars $P_{ij}$ and $P_{jk}$.
-Note that the point $q$ may be not in $\mathbb{H}^{2}$, i.e., the perpendiculars $P_{ij},P_{jk}$ may not intersect within $\mathbb{H}^{2}$.
-Lemma 2.4 in [16] shows that the perpendiculars $P_{ij},P_{jk},P_{ki}$ of $\{ijk\}$ intersect at a common point $c_{ijk}$ (called face center) if and only if the condition (1) holds.
-Hence, the condition (1) in Definition 1.1 ensures the existence of a dual geometric structure on the Poincaré dual of an ideally triangulated surface with boundary.
-Remark 1.5.
-The geometric motivation for Definition 1.2 is that the vertex $v_{k}$ and face center $c_{ijk}$ of a hyper-ideal hyperbolic triangle should be in the 2-dimensional vector subspace of $\mathbb{R}^{3}$ including the vertex $\overline{v}_{k}$ of the new hyper-ideal hyperbolic triangle under the $f_{k}$-conformal variation.
-Please refer to [16, Remark 1.3] for more details.
-The main result in [16] is the following theorem, which gives the explicit forms of the discrete conformal structures in Definition 1.2.
+We briefly elaborate on the motivation behind Definition 1.1 concerning partial edge lengths on surfaces with boundary; further details can be found in [14].
+The hyperbolic geodesic passing through $c_{ij}$ and Lorentz orthogonal to $E_{ij}$ is referred to as the edge perpendicular, denoted by $P_{ij}$.
+This edge perpendicular $P_{ij}$ can be identified with a unique 2-dimensional time-like subspace $\widetilde{P}_{ij}$ of the Lorentzian space $\mathbb{R}^{3}$, satisfying $\widetilde{P}_{ij}\cap\mathbb{H}^{2}=P_{ij}$.
+Since the 2-dimensional subspaces $\widetilde{P}_{ij}$ and $\widetilde{P}_{jk}$ (associated to $P_{ij}$ and $P_{jk}$, respectively) intersect in a 1-dimensional subspace of $\mathbb{R}^{3}$, their intersection $\widetilde{P}_{ij}\cap\widetilde{P}_{jk}$ corresponds to a point $q$ in the Klein model of $\mathbb{H}^{2}$.
+This point $q$ is viewed as the intersection of the perpendiculars $P_{ij}$ and $P_{jk}$.
+Note that $q$ may be not in $\mathbb{H}^{2}$, meaning the perpendiculars $P_{ij},P_{jk}$ may not intersect within $\mathbb{H}^{2}$.
+Lemma 2.4 in [14] demonstrates that the perpendiculars $P_{ij}$, $P_{jk}$, and $P_{ki}$ of the ideal face $\{ijk\}$ intersect at a common point $c_{ijk}$ (termed the face center) if and only if the condition (2) is satisfied.
+Consequently, the condition (2) in Definition 1.1 guarantees the existence of a geometric structure on the Poincaré dual of an ideally triangulated surface with boundary.
+Moreover, the geometric motivation behind Definition 1.2 is discussed in detail in [14, Remark 1.3].
+The main result in [14] is the following theorem, which provides explicit forms and a classification of the discrete conformal structures (as defined in Definition 1.2) on surfaces with boundary.
 Theorem 1.5 ([14], Theorem 1.4).
-Let $(\Sigma,\mathcal{T})$ be an ideally triangulated surface with boundary and let $d=d(f)$ be a discrete conformal structure on $(\Sigma,\mathcal{T})$.
-There exist constant vectors $\alpha\in\mathbb{R}^{N}$, $\eta\in\mathbb{R}_{>0}^{E}$ and $\zeta\in\mathbb{R}^{E}$ satisfying $\eta_{ij}=\eta_{ji}$ and $\zeta_{ij}=\zeta_{ji}$, such that for any right-angled hyperbolic hexagon $\{ijk\}\in F$,
+Let $(\Sigma,\mathcal{T})$ be an ideally triangulated surface with boundary, and let $d=d(f)$ be a discrete conformal structure on $(\Sigma,\mathcal{T})$.
+There exist constant vectors $\alpha\in\mathbb{R}^{N}$ and $\eta\in\mathbb{R}^{E}$ satisfying $\eta_{ij}=\eta_{ji}$, such that for any right-angled hyperbolic hexagon $\{ijk\}\in F$, the value of $\frac{\sinh d_{ij}}{\sinh d_{ji}}$ can only take one of the following four forms:
+$\sqrt{\frac{1+\alpha_{i}e^{2f_{i}}}{1+\alpha_{j}e^{2f_{j}}}},\ -\sqrt{\frac{1+\alpha_{i}e^{2f_{i}}}{1+\alpha_{j}e^{2f_{j}}}},\ e^{\frac{1}{2}C_{ij}(f_{i}-f_{j})},\ -e^{\frac{1}{2}C_{ij}(f_{i}-f_{j})},$
+where $C\in\mathbb{R}^{E_{+}}$ is a constant vector satisfying $C_{ij}+C_{jk}+C_{ki}=0$ for any $\{ijk\}\in F$, and $C_{rs}+C_{sr}=0$ for any subset $\{r,s\}\subseteq\{i,j,k\}$. Furthermore,
 (A1):
-if $\frac{\sinh d_{ij}}{\sinh d_{ji}}=\sqrt{\frac{1+\alpha_{i}e^{2f_{i}}}{1+\alpha_{j}e^{2f_{j}}}}>0$, then the discrete conformal structure $d=d(f)$ is
-(4) $\coth d_{ij}=-\frac{\alpha_{i}e^{2f_{i}}}{\sinh l_{ij}}\sqrt{\frac{1+\alpha_{j}e^{2f_{j}}}{1+\alpha_{i}e^{2f_{i}}}}+\frac{\eta_{ij}e^{f_{i}+f_{j}}}{\sinh l_{ij}}$
-with
+if $\frac{\sinh d_{ij}}{\sinh d_{ji}}=\sqrt{\frac{1+\alpha_{i}e^{2f_{i}}}{1+\alpha_{j}e^{2f_{j}}}}>0$ with $01+\alpha_{i}e^{2f_{i}}>$ 0 and $01+\alpha_{j}e^{2f_{j}}>$ 0, then
 (5) $\cosh l_{ij}=-\sqrt{(1+\alpha_{i}e^{2f_{i}})(1+\alpha_{j}e^{2f_{j}})}+\eta_{ij}e^{f_{i}+f_{j}};$
 (A2):
-if $\frac{\sinh d_{ij}}{\sinh d_{ji}}=e^{\frac{1}{2}C_{ij}(f_{i}-f_{j})}>0$, then the discrete conformal structure $d=d(f)$ is
-(6) $\coth d_{ij}=\frac{\sinh(f_{j}-f_{i}-C_{ij})}{\sinh l_{ij}}+\frac{\eta_{ij}e^{f_{i}+f_{j}}}{\sinh l_{ij}}$
-with
-(7) $\cosh l_{ij}=-\cosh(f_{j}-f_{i}-C_{ij})+\eta_{ij}e^{f_{i}+f_{j}},$
-where $C\in\mathbb{R}^{E_{+}}$ is a constant vector satisfying $C_{ij}+C_{jk}+C_{ki}=0$ for any $\{ijk\}\in F$ and $C_{rs}+C_{sr}=0$ for any $\{r,s\}\subseteq\{i,j,k\}$;
+if $\frac{\sinh d_{ij}}{\sinh d_{ji}}=\sqrt{\frac{1+\alpha_{i}e^{2f_{i}}}{1+\alpha_{j}e^{2f_{j}}}}>0$ with $01+\alpha_{i}e^{2f_{i}}<$ 0 and $01+\alpha_{j}e^{2f_{j}}<$ 0, then
+(6) $\cosh l_{ij}=\sqrt{(1+\alpha_{i}e^{2f_{i}})(1+\alpha_{j}e^{2f_{j}})}+\eta_{ij}e^{f_{i}+f_{j}};$
+(A3):
+if $\frac{\sinh d_{ij}}{\sinh d_{ji}}=e^{\frac{1}{2}C_{ij}(f_{i}-f_{j})}>0$, then
+(7) $\cosh l_{ij}=-\cosh(f_{j}-f_{i}-C_{ij})+\eta_{ij}e^{f_{i}+f_{j}};$
 (B1):
-if $\frac{\sinh d_{ij}}{\sinh d_{ji}}=-\sqrt{\frac{1+\alpha_{i}e^{2f_{i}}}{1+\alpha_{j}e^{2f_{j}}}}<0$, then the discrete conformal structure $d=d(f)$ is
-(8) $\coth d_{ij}=\frac{\alpha_{i}e^{2f_{i}}}{\sinh l_{ij}}\sqrt{\frac{1+\alpha_{j}e^{2f_{j}}}{1+\alpha_{i}e^{2f_{i}}}}+\frac{\zeta_{ij}e^{f_{i}+f_{j}}}{\sinh l_{ij}}$
-with
-(9) $\cosh l_{ij}=\sqrt{(1+\alpha_{i}e^{2f_{i}})(1+\alpha_{j}e^{2f_{j}})}+\zeta_{ij}e^{f_{i}+f_{j}};$
+if $\frac{\sinh d_{ij}}{\sinh d_{ji}}=-\sqrt{\frac{1+\alpha_{i}e^{2f_{i}}}{1+\alpha_{j}e^{2f_{j}}}}<0$ with $01+\alpha_{i}e^{2f_{i}}>$ 0 and $01+\alpha_{j}e^{2f_{j}}>$ 0, then
+(8) $\cosh l_{ij}=\sqrt{(1+\alpha_{i}e^{2f_{i}})(1+\alpha_{j}e^{2f_{j}})}+\eta_{ij}e^{f_{i}+f_{j}};$
 (B2):
-if $\frac{\sinh d_{ij}}{\sinh d_{ji}}=-e^{\frac{1}{2}C_{ij}(f_{i}-f_{j})}<0$, then the discrete conformal structure $d=d(f)$ is
-$\coth d_{ij}=-\frac{\sinh(f_{j}-f_{i}-C_{ij})}{\sinh l_{ij}}+\frac{\zeta_{ij}e^{f_{i}+f_{j}}}{\sinh l_{ij}}$
-with
-(10) $\cosh l_{ij}=\cosh(f_{j}-f_{i}-C_{ij})+\zeta_{ij}e^{f_{i}+f_{j}},$
-where $C\in\mathbb{R}^{E_{+}}$ is a constant vector satisfying $C_{ij}+C_{jk}+C_{ki}=0$ for any $\{ijk\}\in F$ and $C_{rs}+C_{sr}=0$ for any $\{r,s\}\subseteq\{i,j,k\}$.
+if $\frac{\sinh d_{ij}}{\sinh d_{ji}}=-\sqrt{\frac{1+\alpha_{i}e^{2f_{i}}}{1+\alpha_{j}e^{2f_{j}}}}<0$ with $01+\alpha_{i}e^{2f_{i}}<$ 0 and $01+\alpha_{j}e^{2f_{j}}<$ 0, then
+(9) $\cosh l_{ij}=-\sqrt{(1+\alpha_{i}e^{2f_{i}})(1+\alpha_{j}e^{2f_{j}})}+\eta_{ij}e^{f_{i}+f_{j}};$
+(B3):
+if $\frac{\sinh d_{ij}}{\sinh d_{ji}}=-e^{\frac{1}{2}C_{ij}(f_{i}-f_{j})}<0$, then
+(10) $\cosh l_{ij}=\cosh(f_{j}-f_{i}-C_{ij})+\eta_{ij}e^{f_{i}+f_{j}}.$
+Moreover, there are exactly six types of combinations of such discrete conformal structures that can exist on the surface with boundary, namely (A1), (A2), (A3), the combination of (A1) and (B1), the combination of (A2) and (B2), and the combination of (A3) and (B3).
+The latter three mixed combinations are referred to as the mixed discrete conformal structure I, mixed discrete conformal structure II, and mixed discrete conformal structure III, respectively.
 Remark 1.6.
-As mentioned in [16, Remark 1.5], we can always assume $\alpha:B\rightarrow\{-1,0,1\}$ in the discrete conformal structures (A1) and (B1) and $C\equiv 0$ in the discrete conformal structures (A2) and (B2) for simplicity.
+As noted in [14, Remark 1.5], for the discrete conformal structures (A1) and (B1), the discrete conformal structure $d=d(f)$ allows a reparameterization such that $\alpha:B\to\{-1,0,1\}$, with the induced discrete hyperbolic metric $l$ remaining invariant.
+Thus we always assume $\alpha:B\to\{-1,0,1\}$ in (A1) and (B1) .
+For the discrete conformal structures (A2) and (B2), the conditions $01+\alpha_{i}e^{2f_{i}}<$ 0 and $01+\alpha_{j}e^{2f_{j}}<$ 0 imply $\alpha_{i}<0$ and $\alpha_{j}<0$.
+Via reparameterization, we can assume $\alpha\equiv-1$ in (A2) and (B2).
+For the discrete conformal structures (A3) and (B3), if $(\Sigma,\mathcal{T})$ has genus zero, a reparameterization allows us to set $C\equiv 0$.
+For surfaces of non-zero genus, we take $C\equiv 0$ in (A3) and (B3) as special cases.
 Remark 1.7.
-According to [16, Remark 1.7], the discrete conformal structures (A1) and (A2) can exist alone on an ideally triangulated surface with boundary, but they can not exist simultaneously on the same ideally triangulated surface with boundary.
-Moreover, the discrete conformal structures (B1) and (B2) can not exist alone on an ideally triangulated surface with boundary.
-Therefore, we consider the mixed type of discrete conformal structures, i.e., the discrete hyperbolic metrics $l_{ij},l_{jk},l_{ki}$ of a right-angled hyperbolic hexagon $\{ijk\}\in F$ are induced by (7) and (10) or (5) and (9), such that the condition (1) holds.
-We call the former the mixed discrete conformal structure I and the latter the mixed discrete conformal structure II.
-For example, set $\{r,s,t\}=\{i,j,k\}$, if $d_{rs}<0$, then $d_{rt}<0$ by Lemma 2.17, and $d_{sr}>0,\ d_{tr}>0$ by the conditions $l_{rs}=d_{rs}+d_{sr}>0$ and $l_{rt}=d_{rt}+d_{tr}>0$ in Definition 1.1.
-We need $d_{st}>0,\ d_{ts}>0$ to make the condition (1) hold.
-Thus $l_{rs},l_{rt}$ are defined by (10) and $l_{st}$ is defined by (7) or $l_{rs},l_{rt}$ are defined by (9) and $l_{st}$ is defined by (5).
-1.2. Rigidity and existence of discrete conformal structures
-An important problem in discrete conformal geometry is to obtain the relationships between the discrete conformal structures and their combinatorial curvatures.
-The main result in this paper is the following theorem, which gives the global rigidity and existence of discrete conformal structures on ideally triangulated surfaces with boundary.
-For convenience of description, we use the new parameters $\xi^{(1)}$ and $\xi^{(2)}$ to replace parameters $\eta$ and $\zeta$ in the mixed discrete conformal structures $\mathrm{I}$ and $\mathrm{II}$, respectively.
+The right-angled hyperbolic hexagon $\{ijk\}\in F$ with edge lengths $l_{ij},l_{jk},l_{ki}$ induced by (5) and (8), (6) and (9), and (7) and (10) is referred to as Type-I, Type-II, and Type-III, respectively.
+Let $\{r,s,t\}=\{i,j,k\}$.
+For a Type-I right-angled hyperbolic hexagon, if $d_{rs}<0$, then $d_{rt}<0$ by Lemma 2.16.
+Furthermore, from the conditions $l_{rs}=d_{rs}+d_{sr}>0$ and $l_{rt}=d_{rt}+d_{tr}>0$ in Definition 1.1, it follows that $d_{sr}>0$ and $d_{tr}>0$.
+For the condition (2) to hold, $d_{st}>0$ and $d_{ts}>0$ are required.
+Thus $l_{rs}$ and $l_{rt}$ are defined by (8), while $l_{st}$ is defined by (5).
+Analogous reasoning applies to Type-II and Type-III right-angled hyperbolic hexagons.
+Consequently, for the mixed discrete conformal structure I, all right-angled hyperbolic hexagons are either Type-I or a combination of Type-I right-angled hyperbolic hexagons and those with edge lengths induced by (5).
+Similar arguments also apply for the mixed discrete conformal structures II and III.
+1.2. Main results
+A central problem in discrete conformal geometry concerns the relationship between discrete conformal structures and their combinatorial curvatures.
+The main result in this paper is the following theorem, which establishes the global rigidity and existence of discrete conformal structures on ideally triangulated surfaces with boundary.
 Theorem 1.8.
-Let $(\Sigma,\mathcal{T})$ be an ideally triangulated surface with boundary and let $d=d(f)$ be a discrete conformal structure on $(\Sigma,\mathcal{T})$.
+Let $(\Sigma,\mathcal{T})$ be an ideally triangulated surface with boundary, and let $d=d(f)$ be a discrete conformal structure on $(\Sigma,\mathcal{T})$.
 (i):
-For the discrete conformal structure (A1), let $\alpha:B\rightarrow\{-1,0,1\}$ and $\eta\in\mathbb{R}_{>0}^{E}$ be the weights on $(\Sigma,\mathcal{T})$ satisfying $\eta_{ij}>\alpha_{i}\alpha_{j}$ if $\alpha_{i}=\alpha_{j}$ for any two adjacent boundary components $i,j\in B$;
+For the discrete conformal structure (A1), let $\alpha:B\rightarrow\{-1,0,1\}$ and $\eta\in\mathbb{R}_{>0}^{E}$ be the weights on $(\Sigma,\mathcal{T})$ satisfying $\eta_{ij}>\alpha_{i}\alpha_{j}$ for any two adjacent boundary components $i,j\in B$.
+Then the discrete conformal factor $f$ is uniquely determined by its generalized combinatorial curvature $K\in\mathbb{R}^{N}_{>0}$.
+Furthermore, if $\alpha:B\rightarrow\{0,1\}$, then the image of $K$ is >$\mathbb{R}^{N}_{>0}$.
 (ii):
-For the discrete conformal structure (A2), let $\eta\in\mathbb{R}_{>0}^{E}$ be the weight on $(\Sigma,\mathcal{T})$;
+For the discrete conformal structure (A2), let $\eta\in(-1,+\infty)^{E}$ be the weight on $(\Sigma,\mathcal{T})$.
+Then the discrete conformal factor $f$ is uniquely determined by its generalized combinatorial curvature $K\in\mathbb{R}^{N}_{>0}$.
+Furthermore, if $\eta\in(-1,0]^{E}$, then the image of $K$ is >$\mathbb{R}^{N}_{>0}$.
 (iii):
-For the mixed discrete conformal structure $\mathrm{I}$, i.e., the discrete hyperbolic metrics $l_{ij},l_{jk},l_{ki}$ of a right-angled hyperbolic hexagon $\{ijk\}$ are induced by (7) and (10), let $\xi^{(1)}\in\mathbb{R}_{>0}^{E}$ be the weights on $(\Sigma,\mathcal{T})$;
+For the discrete conformal structure (A3), let $\eta\in\mathbb{R}_{>0}^{E}$ be the weight on $(\Sigma,\mathcal{T})$.
+Then the discrete conformal factor $f$ is uniquely determined by its generalized combinatorial curvature $K\in\mathbb{R}^{N}_{>0}$.
+Furthermore, the image of $K$ is >$\mathbb{R}^{N}_{>0}$.
 (iv):
-For the mixed discrete conformal structure $\mathrm{II}$, i.e., the discrete hyperbolic metrics $l_{ij},l_{jk},l_{ki}$ of a right-angled hyperbolic hexagon $\{ijk\}$ are induced by (5) and (9), let $\alpha:B\rightarrow\{-1,0,1\}$ and $\xi^{(2)}:E\rightarrow(1,+\infty)$ be the weights on $(\Sigma,\mathcal{T})$;
-Then the discrete conformal factor $f$ is determined by its generalized combinatorial curvature $K\in\mathbb{R}^{N}_{>0}$.
-In particular, the map from $f$ to $K$ is a smooth embedding.
-Furthermore, the image of $K$ is >$\mathbb{R}^{N}_{>0}$ for the cases $\mathrm{(ii)}$ and $\mathrm{(iii)}$, for the case $\mathrm{(i)}$ if $\alpha:B\rightarrow\{0,1\}$, and for the case $\mathrm{(iv)}$ if $\alpha:B\rightarrow\{0,1\}$ and satisfies a restrictive condition.
+For the mixed discrete conformal structure I, let $\alpha:B\rightarrow\{-1,0,1\}$ and $\eta\in(1,+\infty)^{E}$ be the weights on $(\Sigma,\mathcal{T})$.
+Then the discrete conformal factor $f$ is uniquely determined by its generalized combinatorial curvature $K\in\mathbb{R}^{N}_{>0}$.
+Furthermore, if $\alpha:B\rightarrow\{0,1\}$, and $\alpha_{j}$ and $\alpha_{k}$ cannot both be 1 simultaneously for any right-angled hyperbolic hexagon with edge lengths $l_{ij},l_{ik}$ given by (8) and $l_{jk}$ given by (5), then the image of $K$ is >$\mathbb{R}^{N}_{>0}$.
+(v):
+For the mixed discrete conformal structure II, let $\eta\in[1,+\infty)^{E}$ be the weight on $(\Sigma,\mathcal{T})$.
+Then the discrete conformal factor $f$ is uniquely determined by its generalized combinatorial curvature $K\in\mathbb{R}^{N}_{>0}$.
+(vi):
+For the mixed discrete conformal structure III, let $\eta\in\mathbb{R}_{>0}^{E}$ be the weight on $(\Sigma,\mathcal{T})$.
+Then the discrete conformal factor $f$ is uniquely determined by its generalized combinatorial curvature $K\in\mathbb{R}^{N}_{>0}$.
+Furthermore, the image of $K$ is >$\mathbb{R}^{N}_{>0}$.
 Remark 1.9.
-If $\alpha\equiv 0$, the case $\mathrm{(i)}$ in Theorem 1.9 generalizes Guo’s results [4] and Li-Xu-Zhou’s results [6].
-If $\alpha\equiv 1$, the case $\mathrm{(i)}$ in Theorem 1.9 was proved by Guo-Luo [5].
-Moreover, the case $\mathrm{(ii)}$ in Theorem 1.9 was also proved by Guo-Luo [5].
-We include them in Theorem 1.9 for completeness.
-For the case $\mathrm{(iii)}$ in Theorem 1.9, the weight $\xi^{(1)}\in\mathbb{R}_{>0}^{E}$ can be extended to a wider range, please see Theorem 5.1.
-For the case $\mathrm{(iv)}$ in Theorem 1.9, the weight $\xi^{(2)}:E\rightarrow(1,+\infty)$ can also be extended to a wider range, and we will use a new notation $\Xi$ to denote the set of the range of weight because of its complexity, please see Subsection 6.1.
-Moreover, for the image of $K$ in the mixed discrete conformal structure $\mathrm{II}$, we not only require $\alpha:B\rightarrow\{0,1\}$, but also $\alpha$ satisfies a restrictive condition.
-Here the restrictive condition means to exclude the case that $\alpha_{j}=\alpha_{k}=1$ for a right-angled hyperbolic hexagon $\{ijk\}\in F$ with the edge lengths $l_{ij}$ and $l_{ik}$ are given by (9) and the edge length $l_{jk}$ is given by (5), please see Subsection 6.3.
+If $\alpha\equiv 0$, then Theorem 1.8 (i) generalizes the results obtained by Guo [4] and Li-Xu-Zhou [6].
+For $\alpha\equiv 1$, Theorem 1.8 (i) was established by Guo-Luo [5].
+Theorem 1.8 (i) unifies these existing results and further incorporates mixed-type cases where $\alpha_{i}=1$ for some non-empty subset $V_{1}\subseteq V$ and $\alpha_{j}=0$ for $j\in V\setminus V_{1}$.
+Theorem 1.8 (ii) generalizes the results of Guo-Luo [5], while Theorem 1.8 (iii), which was proven by Guo-Luo [5], is included here for completeness.
+In Theorem 1.8 (iv), the weight $\eta\in(1,+\infty)^{E}$ admits an extension to a broader range, with details provided in Subsection 8.1.
+For Theorem 1.8 (v), the weight $\eta\in[1,+\infty)^{E}$ can also be extended to a broader range (see Theorem 7.5).
+Note that only a rigidity result is presented here, without any existence result; the reason is elaborated in Remark 7.6.
+In Theorem 1.8 (vi), the weight $\eta\in\mathbb{R}_{>0}^{E}$ similarly allows for extension to a broader range, as discussed in Theorem 6.1.
 1.3. Basic ideas of the proof of Theorem 1.8
-The proof for the existence of discrete conformal structures uses the continuity method.
-We will show that the image of the curvature map is both open and closed subset of >$\mathbb{R}^{N}_{>0}$, and thus is the whole space >$\mathbb{R}^{N}_{>0}$.
-The proof for the rigidity of discrete conformal structures involves a variational principle, which was pioneered by Colin de Verdière’s work [1] for tangential circle packings on triangulated surfaces.
-The proof of the rigidity consists of the following three steps.
-The first step is to give a characterization of the admissible space of discrete conformal factor $f$ for a right-angled hyperbolic hexagon.
-By a change of variables $u_{i}=u_{i}(f_{i})$ described in the proof of Lemma 2.22, the admissible space $\Omega_{ijk}$ of $f$ is transferred to the admissible space $\mathcal{U}_{ijk}$ of $u$, which is proved to be convex and simply connected.
-The second step is to prove the Jacobian of the generalized angles $\theta$ with respect to $u$ for a right-angled hyperbolic hexagon is symmetric and negative definite.
-As a consequence, the Jacobian of the generalized combinatorial curvature $K$ with respect to $u$ is symmetric and negative definite on the admissible space $\mathcal{U}=\cap\ \mathcal{U}_{ijk}$.
-The first step and second step enable us to define a strictly concave function on the admissible space $\mathcal{U}$.
-The third step is to use the following well-known result from analysis:
+The proof of the existence of discrete conformal structures relies on the continuity method.
+Specifically, we will show that the image of the curvature map $K$ is both an open and a closed subset of >$\mathbb{R}^{N}_{>0}$; consequently, it coincides with the entire space >$\mathbb{R}^{N}_{>0}$.
+The proof of the rigidity of discrete conformal structures involves a variational principle introduced by Colin de Verdière [1] for tangential circle packings on triangulated closed surfaces.
+The rigidity proof proceeds in three steps: First, we characterize the admissible space of discrete conformal factors $f$ for a right-angled hyperbolic hexagon.
+By a variable transformation $u_{i}=u_{i}(f_{i})$ (as detailed in the proof of Lemma 2.21), the admissible space $\Omega_{ijk}$ of $f$ is transformed into the admissible space $\mathcal{U}_{ijk}$ of $u$, which is shown to be convex and simply connected.
+Second, we prove that the Jacobian of the generalized angles $\theta$ with respect to $u$ for a right-angled hyperbolic hexagon is symmetric and negative definite.
+Consequently, the Jacobian of the generalized combinatorial curvature $K$ with respect to $u$ is symmetric and negative definite on the admissible space $\mathcal{U}=\cap\,\mathcal{U}_{ijk}$.
+These two steps collectively enable the definition of a strictly concave function on $\mathcal{U}$.
+Third, we invoke the following well-known result from analysis to establish rigidity.
 Lemma 1.10.
@@
 Furthermore, $\nabla f$ is a smooth embedding.
-The key step is how to get a change of variables $u_{i}=u_{i}(f_{i})$. For this, we need to calculate the variation formulas of the generalized angles, which is as follows.
+A critical component of the proof involves constructing the necessary variable transformation $u_{i}=u_{i}(f_{i})$.
+To this end, we require the following result concerning the variation of generalized angles.
 Theorem 1.11.
-Let $(\Sigma,\mathcal{T})$ be an ideally triangulated surface with boundary and let $d=d(f)$ be a discrete conformal structure on $(\Sigma,\mathcal{T})$.
+Let $(\Sigma,\mathcal{T})$ be an ideally triangulated surface with boundary, and let $d=d(f)$ be a discrete conformal structure on $(\Sigma,\mathcal{T})$.
 (i):
-For any right-angled hyperbolic hexagon $\{ijk\}\in F$,
+For any right-angled hyperbolic hexagon $\{ijk\}\in F$, the following holds:
 $\frac{\partial\theta^{jk}_{i}}{\partial f_{j}}=\frac{-1}{\sinh d_{ji}}\frac{\tanh^{\beta}h_{ij}}{\sinh l_{ij}},$
-where $\beta=1$ if $c_{ijk}$ is time-like and $\beta=-1$ if $c_{ijk}$ is space-like.
-If $c_{ijk}$ is light-like, we interpret the formula as $\tanh^{\beta}h_{ij}=\tanh^{\beta}h_{jk}=\tanh^{\beta}h_{ki}=1$.
+where $\beta=1$ if the face center $c_{ijk}$ is time-like, and $\beta=-1$ if $c_{ijk}$ is space-like.
+If $c_{ijk}$ is light-like, the formula is interpreted as $\tanh^{\beta}h_{ij}=\tanh^{\beta}h_{jk}=\tanh^{\beta}h_{ki}=1$.
+Here, $h_{ij}$ denotes the signed distance from the face center $c_{ijk}$ to the hyperbolic geodesic line $E_{ij}$, which is positive if $c_{ijk}$ lies on the same side of $E_{ij}$ as the right-angled hyperbolic hexagon $\{ijk\}$, negative otherwise, and zero if $c_{ijk}$ lies on $E_{ij}$.
 (ii):
-Furthermore, we have the following Glickenstein-Thomas formula for $\frac{\partial\theta^{jk}_{i}}{\partial f_{i}}$, $\frac{\partial\theta^{ik}_{j}}{\partial f_{i}}$ and $\frac{\partial\theta^{ij}_{k}}{\partial f_{i}}$, i.e.,
+For a right-angled hyperbolic hexagon $\{ijk\}\in F$, the variation of generalized angles satisfies the following relation:
 $\frac{\partial\theta^{jk}_{i}}{\partial f_{i}}=\cosh l_{ij}\frac{\partial\theta^{ik}_{j}}{\partial f_{i}}+\cosh l_{ik}\frac{\partial\theta^{ij}_{k}}{\partial f_{i}}.$
 (iii):
-Moreover, there is a change of variables $u_{i}=u_{i}(f_{i})$ such that
-(11) $\frac{\partial\theta^{jk}_{i}}{\partial u_{j}}=\frac{\partial\theta^{ik}_{j}}{\partial u_{i}}.$
-The function $u$ is also called a discrete conformal factor.
-Here $h_{ij}$ is the signed distance of the face center $c_{ijk}$ to the hyperbolic geodesic $E_{ij}$, which is positive if $c_{ijk}$ is on the same side of the geodesic $E_{ij}$ as the right-angled hyperbolic hexagon $\{ijk\}$ and negative otherwise (or zero if $c_{ijk}$ is on the geodesic $E_{ij}$).
+There exists a change of variables given by $u_{i}=u_{i}(f_{i})$ such that
+$\frac{\partial\theta^{jk}_{i}}{\partial u_{j}}=\frac{\partial\theta^{ik}_{j}}{\partial u_{i}}.$
+The function $u$ is also referred to as a discrete conformal factor.
 1.4. Organization of the paper
-In Section 2, we prove Theorem 1.12 by some useful lemmas.
-In Section 3, we prove Theorem 1.9 (i).
-In Section 4, we prove Theorem 1.9 (ii).
-In Section 5, we prove Theorem 1.9 (iii).
-In Section 6, we prove Theorem 1.9 (iv).
+In Section 2, we establish Theorem 1.11.
+In Section 3, we investigate the discrete conformal structure (A1) and prove Theorem 1.8 (i).
+Section 4 is devoted to the discrete conformal structure (A2), where Theorem 1.8 (ii) is proven.
+In Section 5, we study the discrete conformal structure (A3) and establish Theorem 1.8 (iii).
+Section 6 focuses on the mixed discrete conformal structure III, with a proof of a generalization of Theorem 1.8 (vi).
+Similarly, Section 7 investigates the mixed discrete conformal structure II and proves a generalization of Theorem 1.8 (v).
+Finally, Section 8 studies the mixed discrete conformal structure I and establishes a generalization of Theorem 1.8 (iv).
+Acknowledgements
+The research of X. Xu is supported by National Natural Science Foundation of China under grant no. 12471057.
 2. Discrete conformal variations of generalized angles
-In this section, we first present some basic facts in the 2-dimensional hyperbolic geometry, which are discussed in [7, Chapter 3 and 6].
-Then we prove Theorem 1.12.
-Different from the proof of Theorem 7 in Glickenstein-Thomas [3], we are unable to get the variational formulas of generalized angle by similar geometric properties directly from the figure under the $f_{k}$-conformal variation, and we obtain them by complex calculations.
-2.1. Some basic knowledge
-For any vectors $x,y$ in $\mathbb{R}^{3}$, the Lorentzian inner product * of $x$ and $y$ is given by $x*y=xJy^{T}$, where $J=\mathrm{diag}\{1,1,-1\}$.
-If $x*y=0$, then we say $x,y\in\mathbb{R}^{3}$ are Lorentz orthogonal.
-The Lorentzian cross product $\otimes$ of $x$ and $y$ is defined as $x\otimes y:=J(x\times y)$, where $\times$ is the Euclidean cross product.
-The Lorentzian norm of $x\in\mathbb{R}^{3}$ is defined to be the complex number $||x||=\sqrt{x*x}$.
-Here $||x||$ is either positive, zero or positive imaginary.
-If $||x||$ is positive imaginary, we denote by $|||x|||$ its absolute value.
-Furthermore, $x\in\mathbb{R}^{3}$ is called space-like if $x*x>0$, light-like if $x*x=0$ and time-like if $x*x<0$.
-The hyperboloid model of $\mathbb{H}^{2}$ is defined to be
-$\mathbb{H}^{2}=\{x=(x_{1},x_{2},x_{3})\in\mathbb{R}^{3}|\ x*x=-1\ \text{and}\ x_{3}>0\},$
-which can be embedded in $(\mathbb{R}^{3},*)$.
-A geodesic (or hyperbolic line) in $\mathbb{H}^{2}$ is the non-empty intersection of $\mathbb{H}^{2}$ with a 2-dimensional vector subspace of $\mathbb{R}^{3}$.
-The Klein model of $\mathbb{H}^{2}$ is the center projection of $\mathbb{H}^{2}$ into the plane $x_{3}=1$ of $\mathbb{R}^{3}$.
-In the Klein model, the interior, boundary and external points of the unit disk represent the time-like, light-like and space-like vectors respectively, and they are called hyperbolic points, ideal points and hyper-ideal points respectively.
-A geodesic in $\mathbb{H}^{2}$ is a chord on the unit disk.
-In the Klein model, a triangle $xyz$ with its three vertices $x,y,z$ being hyperbolic is called a hyperbolic triangle.
-If at least one of $x,y,z$ is ideal or hyper-ideal, then we say the triangle $xyz$ is a generalized hyperbolic triangle.
-Specially, if $x,y,z$ are all hyper-ideal, we call this triangle as a hyper-ideal hyperbolic triangle for simplicity.
-The following proposition follows from the properties of Lorentzian cross product, which is used extensively in this paper.
+In this section, we first recall some fundamental results from 2-dimensional hyperbolic geometry (see [7, Chapters 3 and 6]).
+We then provide a geometric interpretation along with several useful lemmas.
+Finally, we proceed to prove Theorem 1.11.
+2.1. Some fundamental results
+For any vectors $x,y$ in the Lorentzian space $\mathbb{R}^{3}$, the Lorentzian inner product * of $x$ and $y$ is defined as $x*y=xJy^{\mathrm{T}}$, where $J=\mathrm{diag}\{1,1,-1\}$.
+Two vectors $x,y\in\mathbb{R}^{3}$ are said to be Lorentz orthogonal if $x*y=0$.
+The Lorentzian cross product $\otimes$ of $x$ and $y$ is defined as $x\otimes y=J(x\times y)$, where $\times$ denotes the Euclidean cross product.
+The Lorentzian norm of a vector $x\in\mathbb{R}^{3}$ is defined as the complex number $\|x\|=\sqrt{x*x}$.
+Here $\|x\|$ is either positive, zero, or positive imaginary.
+When $\|x\|$ is positive imaginary, its absolute value is denoted by $|\|x\||$.
+Furthermore, a vector $x\in\mathbb{R}^{3}$ is termed space-like if $x*x>0$, light-like if $x*x=0$, and time-like if $x*x<0$.
+The hyperboloid model of $\mathbb{H}^{2}$ is defined as
+$\mathbb{H}^{2}=\{x=(x_{1},x_{2},x_{3})\in\mathbb{R}^{3}\mid x*x=-1\ \text{and}\ x_{3}>0\},$
+which is embedded in the Lorentzian space $(\mathbb{R}^{3},*)$.
+A vector subspace of $\mathbb{R}^{3}$ is termed time-like if it contains at least one time-like vector.
+A geodesic (or hyperbolic line) in $\mathbb{H}^{2}$ is the non-empty intersection of $\mathbb{H}^{2}$ with a 2-dimensional time-like vector subspace of $\mathbb{R}^{3}$.
+The Klein model of $\mathbb{H}^{2}$ is defined as the central projection of $\mathbb{H}^{2}$ onto the plane $x_{3}=1$ in $\mathbb{R}^{3}$.
+In this model, the interior, boundary, and exterior points of the unit disk correspond to time-like, light-like, and space-like vectors, respectively, and are referred to as hyperbolic points, ideal points, and hyper-ideal points, correspondingly.
+In the Klein model, a geodesic in $\mathbb{H}^{2}$ is represented as an open chord of the unit disk.
+A triangle $xyz$ with all three vertices $x,y,z$ being hyperbolic is called a hyperbolic triangle.
+If at least one of $x,y,z$ is ideal or hyper-ideal, then the triangle $xyz$ is termed a generalized hyperbolic triangle.
+Specifically, if $x,y,z$ are hyper-ideal, this triangle is referred to as a hyper-ideal hyperbolic triangle for brevity.
+It should be noted that, in this paper, each edge of a generalized hyperbolic triangle $xyz$ is required to intersect the unit disk.
+The following proposition, which follows from properties of the Lorentzian cross product, is used extensively in this paper.
 Proposition 2.1.
@@
 $-(z*y)=(z*x)(x*y).$
-For two vectors $x,y\in\mathbb{H}^{2}$, we denoted by $d_{\mathbb{H}}(x,y)$ the hyperbolic distance between $x$ and $y$, which is the arc length of the geodesic between $x$ and $y$.
-We use $\mathrm{Span}(x)$ or $\mathrm{Span}(x,y)$ to denote the vector subspace spanned $x$ or $x,y$.
-If $x\in\mathbb{R}^{3}$ is a space-like vector, then the Lorentzian complement $x^{\bot}:=\{w\in\mathbb{R}^{3}|x*w=0\}$ of $\mathrm{Span}(x)$ is a 2-dimension vector subspace of $\mathbb{R}^{3}$ and $x^{\bot}\cap\mathbb{H}^{2}$ is non-empty.
-This implies that $x^{\bot}\cap\mathbb{H}^{2}$ is a geodesic in $\mathbb{H}^{2}$.
-And the hyperbolic distance between $x$ and $y\in\mathbb{H}^{2}$ is defined to be $d_{\mathbb{H}}(y,x^{\bot})=\inf\{d_{\mathbb{H}}(y,z)|z\in x^{\bot}\cap\mathbb{H}^{2}\}$.
+For any two vectors $x,y\in\mathbb{H}^{2}$, there exists a unique geodesic passing through $x$ and $y$.
+The arc length of the geodesic segment between $x$ and $y$ is defined as the hyperbolic distance between $x$ and $y$, denoted by $d_{\mathbb{H}}(x,y)$.
+We use $\textrm{Span}(x)$ and $\textrm{Span}(x,y)$ to denote the vector subspace spanned by $x$ and by $x,y$, respectively.
+For a space-like vector $x\in\mathbb{R}^{3}$, the Lorentzian complement of $\textrm{Span}(x)$, defined as $x^{\bot}:=\{w\in\mathbb{R}^{3}\mid x*w=0\}$, is a 2-dimensional time-like vector subspace of $\mathbb{R}^{3}$, and thus $x^{\bot}\cap\mathbb{H}^{2}$ is a geodesic in $\mathbb{H}^{2}$.
+The hyperbolic distance between $x$ and $y\in\mathbb{H}^{2}$ is defined as $d_{\mathbb{H}}(y,x^{\bot})=\inf\{d_{\mathbb{H}}(y,z)\mid z\in x^{\bot}\cap\mathbb{H}^{2}\}$.
 Let $x,y\in\mathbb{R}^{3}$ be space-like vectors.
-If $\mathrm{Span}(x,y)\cap\mathbb{H}^{2}$ is non-empty, then the hyperbolic distance between $x$ and $y$ is defined by $d_{\mathbb{H}}(x^{\bot},y^{\bot})=\inf\{d_{\mathbb{H}}(z,w)|z\in x^{\bot}\cap\mathbb{H}^{2},w\in y^{\bot}\cap\mathbb{H}^{2}\}$.
-If $\mathrm{Span}(x,y)\cap\mathbb{H}^{2}$ is empty, then $x^{\bot}$ and $y^{\bot}$ intersect in an angle $\angle(x^{\bot},y^{\bot})$ in $\mathbb{H}^{2}$.
+If $\textrm{Span}(x,y)\cap\mathbb{H}^{2}$ is non-empty, then the hyperbolic distance between $x$ and $y$ is defined by $d_{\mathbb{H}}(x^{\bot},y^{\bot})=\inf\{d_{\mathbb{H}}(z,w)\mid z\in x^{\bot}\cap\mathbb{H}^{2},w\in y^{\bot}\cap\mathbb{H}^{2}\}$.
+If $\textrm{Span}(x,y)\cap\mathbb{H}^{2}$ is empty, then $x^{\bot}$ and $y^{\bot}$ intersect at an angle $\angle(x^{\bot},y^{\bot})$ in $\mathbb{H}^{2}$.
 Proposition 2.2 ([7], Chapter 3).
-Suppose $x,y\in\mathbb{H}^{2}$ are time-like vectors and $z,w\in\mathbb{R}^{3}$ are space-like vectors. Then
+Suppose $x,y\in\mathbb{H}^{2}$ are time-like vectors, and $z,w\in\mathbb{R}^{3}$ are space-like vectors. Then
 (i):
-$x*y=-|||x|||\cdot|||y|||\cosh d_{\mathbb{H}}(x,y)$;
+$x*y=-|\|x\||\cdot|\|y\||\cosh d_{\mathbb{H}}(x,y)$;
 (ii):
-$|x*z|=|||x|||\cdot||z||\sinh d_{\mathbb{H}}(x,z^{\bot})$, and $x*z<0$ if and only if $x$ and $z$ are on opposite sides of the hyperplane $z^{\bot}$;
+$|x*z|=|\|x\||\cdot\|z\|\sinh d_{\mathbb{H}}(x,z^{\bot})$, and $x*z<0$ if and only if $x$ and $z$ lie on opposite sides of the hyperplane $z^{\bot}$;
 (iii):
-if $\mathrm{Span}(z,w)\cap\mathbb{H}^{2}$ is non-empty, then $|z*w|=||z||\cdot||w||\cosh d_{\mathbb{H}}(z^{\bot},w^{\bot})$, and $z*w<0$ if and only if $z$ and $w$ are oppositely oriented tangent vectors of $Q$, where $Q$ is the hyperbolic line Lorentz orthogonal to $z^{\bot}$ and $w^{\bot}$;
+if $\textrm{Span}(z,w)\cap\mathbb{H}^{2}\neq\emptyset$, then $|z*w|=\|z\|\cdot\|w\|\cosh d_{\mathbb{H}}(z^{\bot},w^{\bot})$, and $z*w<0$ if and only if $z$ and $w$ are oppositely oriented tangent vectors of $Q$, where $Q$ is the hyperbolic line Lorentz orthogonal to $z^{\bot}$ and $w^{\bot}$;
 (iv):
-if $\mathrm{Span}(z,w)\cap\mathbb{H}^{2}$ is empty, then $z*w=||z||\cdot||w||\cos\angle(z^{\bot},w^{\bot})$.
-The following proposition is useful for identifying perpendicular geodesics in the Klein model.
+if $\textrm{Span}(z,w)\cap\mathbb{H}^{2}=\emptyset$, then $z*w=\|z\|\cdot\|w\|\cos\angle(z^{\bot},w^{\bot})$.
+The following two propositions are useful for characterizing perpendicular geodesics in the Klein model.
 Proposition 2.3.
-([8], Proposition 4) Let $x\in\mathbb{R}^{3}$ be a space-like vector and let $\gamma=x^{\bot}\cap\mathbb{H}^{2}$ be the geodesic corresponding to $x^{\bot}$.
-Suppose $\omega$ is another geodesic that intersects $\gamma$, and $P$ is the plane such that $\omega=P\cap\mathbb{H}^{2}$.
-Then $\gamma$ and $\omega$ meet at a right angle if and only if $x\in P$.
+([7], Theorem 3.2.11) Let $P$ be a geodesic in $\mathbb{H}^{2}$, and let $y\in\mathbb{H}^{2}$ be a time-like vector.
+There exists a unique geodesic in $\mathbb{H}^{2}$ passing through $y$ and Lorentz orthogonal to $P$.
 Proposition 2.4.
-([7], Theorem 3.2.11) Let $P$ be a geodesic in $\mathbb{H}^{2}$ and let $y\in\mathbb{H}^{2}$ be a time-like vector.
-Then there exists a unique geodesic of $\mathbb{H}^{2}$ through $y$ and Lorentz orthogonal to $P$.
-2.2. A geometric explanation
-Consider a hyper-ideal hyperbolic triangle $v_{i}v_{j}v_{k}$ in the Kelin model.
-Set $\{r,s,t\}=\{i,j,k\}$.
-Set $E_{rs}=\mathrm{Span}(v_{r},v_{s})\cap\mathbb{H}^{2}$, then $E_{rs}$ is a geodesic in $\mathbb{H}^{2}$.
-Set $v_{r}^{\bot}\cap\mathbb{H}^{2}=E^{\prime}_{st}$, then $E^{\prime}_{st}$ is also a geodesic in $\mathbb{H}^{2}$ and Lorentz orthogonal to the geodesics $E_{rs},E_{rt}$.
-The geodesic arc in $E^{\prime}_{st}$ bounded by $E_{rs}$ and $E_{rt}$ is called a boundary arc, denoted by $r\in B$.
-The geodesic arc between two adjacent boundary arcs $r$ and $s$ is denoted by $\{rs\}$, which is a part of $E_{rs}$ and called an ideal edge.
-Moreover, the six geodesic arcs form a right-angled hyperbolic hexagon $\{ijk\}$.
+([8], Proposition 4) Let $x\in\mathbb{R}^{3}$ be a space-like vector, and let $\gamma=x^{\bot}\cap\mathbb{H}^{2}$ be the geodesic corresponding to $x^{\bot}$.
+Suppose $\omega$ is another geodesic intersecting $\gamma$, and $P$ is the plane such that $\omega=P\cap\mathbb{H}^{2}$.
+Then $\gamma$ and $\omega$ intersect at a right angle if and only if $x\in P$.
+2.2. A geometric interpretation
+Let $v_{i}v_{j}v_{k}$ be a hyper-ideal hyperbolic triangle in the Klein model, and let $\{r,s,t\}=\{i,j,k\}$.
+By the notations in Remark 1.3, $E_{rs}=\textrm{Span}(v_{r},v_{s})\cap\mathbb{H}^{2}$ is a geodesic in $\mathbb{H}^{2}$.
+Define $E^{\prime}_{st}=v_{r}^{\bot}\cap\mathbb{H}^{2}$.
+Then $E^{\prime}_{st}$ is also a geodesic in $\mathbb{H}^{2}$, and is Lorentz orthogonal to both $E_{rs}$ and $E_{rt}$.
+The geodesic arc in $E^{\prime}_{st}$ bounded by $E_{rs}$ and $E_{rt}$ is a boundary arc, denoted by $r$.
+An ideal edge $\{rs\}$ is the geodesic arc between two adjacent boundary arcs $r$ and $s$, lying on $E_{rs}$.
+Furthermore, the six geodesic arcs form a right-angled hyperbolic hexagon $\{ijk\}$.
 Conversely, for any geodesic $E_{rs}$ in $\mathbb{H}^{2}$, there exists a unique space-like vector $v^{\prime}_{t}$ such that $(v^{\prime}_{t})^{\bot}\cap\mathbb{H}^{2}=E_{rs}$.
-Then the three points $v^{\prime}_{i},v^{\prime}_{j}$ and $v^{\prime}_{k}$ are hyper-ideal and form a hyper-ideal hyperbolic triangle $v^{\prime}_{i}v^{\prime}_{j}v^{\prime}_{k}$, which is called the polar triangle of the hyper-ideal hyperbolic triangle $v_{i}v_{j}v_{k}$.
-Please refer to Figure 2.
+The three points $v^{\prime}_{i},v^{\prime}_{j},v^{\prime}_{k}$ are hyper-ideal and form a hyper-ideal hyperbolic triangle $v^{\prime}_{i}v^{\prime}_{j}v^{\prime}_{k}$, referred to as the polar triangle of $v_{i}v_{j}v_{k}$.
+See Figure 2 for illustration.
 Figure 2.
-A hyper-ideal hyperbolic triangle $v_{i}v_{j}v_{k}$ and its polar triangle $v^{\prime}_{i}v^{\prime}_{j}v^{\prime}_{k}$ in the Kelin model (left), and a right-angled hyperbolic hexagon $\{ijk\}\in F$ in the Poincaré model (right).
-By the construction of the polar triangle $v^{\prime}_{i}v^{\prime}_{j}v^{\prime}_{k}$, the vertices $v^{\prime}_{i},v^{\prime}_{j},v^{\prime}_{k}$ are determined by the vertices $v_{i},v_{j},v_{k}$ and vice versa.
-Without loss of generality, we assume $||v_{i}||=||v_{j}||=||v_{k}||=1$, then
-(12) $v^{\prime}_{i}=\frac{v_{k}\otimes v_{j}}{||v_{k}\otimes v_{j}||},\ v^{\prime}_{j}=\frac{v_{i}\otimes v_{k}}{||v_{i}\otimes v_{k}||},\ v^{\prime}_{k}=\frac{v_{j}\otimes v_{i}}{||v_{j}\otimes v_{i}||}.$
-Therefore, given a hyper-ideal hyperbolic triangle $v_{i}v_{j}v_{k}$, there exists a unique polar triangle $v^{\prime}_{i}v^{\prime}_{j}v^{\prime}_{k}$ with its vertices satisfying (12).
-The length of the ideal edge $\{rs\}\in E$ is denoted by $l_{rs}$ and the length of the boundary arc $t\in B$ is denoted by $\theta^{rs}_{t}$.
+A hyper-ideal hyperbolic triangle $v_{i}v_{j}v_{k}$ and its polar triangle $v^{\prime}_{i}v^{\prime}_{j}v^{\prime}_{k}$ in the Klein model (left), and a right-angled hyperbolic hexagon $\{ijk\}\in F$ in the Poincaré disk model (right).
+By the construction of the polar triangle $v^{\prime}_{i}v^{\prime}_{j}v^{\prime}_{k}$, the vertices $v^{\prime}_{i},v^{\prime}_{j},v^{\prime}_{k}$ are determined by the vertices $v_{i},v_{j},v_{k}$, and vice versa.
+Without loss of generality, we assume $\|v_{i}\|=\|v_{j}\|=\|v_{k}\|=1$.
+Then
+(11) $v^{\prime}_{i}=\frac{v_{k}\otimes v_{j}}{\|v_{k}\otimes v_{j}\|},\ v^{\prime}_{j}=\frac{v_{i}\otimes v_{k}}{\|v_{i}\otimes v_{k}\|},\ v^{\prime}_{k}=\frac{v_{j}\otimes v_{i}}{\|v_{j}\otimes v_{i}\|}.$
+Therefore, given a hyper-ideal hyperbolic triangle $v_{i}v_{j}v_{k}$, there exists a unique polar triangle $v^{\prime}_{i}v^{\prime}_{j}v^{\prime}_{k}$ whose vertices satisfy (11).
+Let $l_{rs}$ denote the length of the ideal edge $\{rs\}\in E$, and let $\theta^{rs}_{t}$ denote the length of the boundary arc $t\in B$.
 Lemma 2.5.
-The vectors $c_{ijk},c_{rs},v^{\prime}_{t}$ lie in the same 2-dimensional vector subspace of $\mathbb{R}^{3}$, where the notations of $c_{ijk},c_{rs}$ follows from Remark 1.4.
+The vectors $c_{ijk}$, $c_{rs}$, and $v^{\prime}_{t}$ lie in the same 2-dimensional time-like vector subspace of the Lorentzian space $\mathbb{R}^{3}$, where the notations $c_{ijk}$ and $c_{rs}$ are defined in Remark 1.4 and Remark 1.3, respectively.
 Proof.
 Without loss of generality, we assume $r=i,\ s=j,\ t=k$.
-Since $c_{ij}\in E_{ij}$, by Proposition 2.4, there exists a unique geodesic of $\mathbb{H}^{2}$ through $c_{ij}$ and Lorentz orthogonal to $E_{ij}$, which is $P_{ij}=\mathrm{Span}(c_{ij},c_{ijk})\cap\mathbb{H}^{2}$.
+Since $c_{ij}\in E_{ij}$, by Proposition 2.3, there exists a unique geodesic in $\mathbb{H}^{2}$ passing through $c_{ij}$ and Lorentz orthogonal to $E_{ij}$, which is $P_{ij}=\textrm{Span}(c_{ij},c_{ijk})\cap\mathbb{H}^{2}$.
 Note that $E_{ij}=(v^{\prime}_{k})^{\bot}\cap\mathbb{H}^{2}$.
-By Proposition 2.3, since $E_{ij}$ and $P_{ij}$ meet at a right angle, then $v^{\prime}_{k}\in\mathrm{Span}(c_{ij},c_{ijk})$.
+By Proposition 2.4, since $E_{ij}$ and $P_{ij}$ intersect at a right angle, it follows that $v^{\prime}_{k}\in\textrm{Span}(c_{ij},c_{ijk})$.
 Q.E.D.
-Combining Remark 1.4 and Lemma 2.5, we have $c_{ijk}=\mathrm{Span}(v^{\prime}_{i},c_{jk})\cap\mathrm{Span}(v^{\prime}_{j},c_{ik})\cap\mathrm{Span}(v^{\prime}_{k},c_{ij})$.
-Due to the equivalence of $v_{i}v_{j}v_{k}$ and $v^{\prime}_{i}v^{\prime}_{j}v^{\prime}_{k}$, it is natural to consider a dual model.
-Note that for any $c_{ijk}\in\mathbb{R}^{3}$, $\mathrm{Span}(v_{r},c_{ijk})\cap v_{r}^{\bot}$ may not intersect within $\mathbb{H}^{2}$.
-For the convenience of calculations, we focus on the case that
-$\mathrm{Span}(v_{r},c_{ijk})\cap v_{r}^{\bot}\cap\mathbb{H}^{2}=\mathrm{Span}(v_{r},c_{ijk})\cap E^{\prime}_{st}\neq\emptyset,\quad\forall r\in\{i,j,k\}.$
-This condition limits the positions of $c_{ijk}$ when $c_{ijk}$ is space-like, i.e., $c_{ijk}$ can only be in some certain open domains.
-Denote $c^{\prime}_{st}=\mathrm{Span}(v_{r},c_{ijk})\cap E^{\prime}_{st}$ as the dual edge center, which is the unique point along the geodesic $E^{\prime}_{st}$ such that $c^{\prime}_{st}$ is of (signed) distance $\theta_{st}$ from $\{rt\}$ and $\theta_{ts}$ from $\{rs\}$.
-Thus we have
+Combining Remark 1.4 and Lemma 2.5, we obtain $c_{ijk}=\textrm{Span}(v^{\prime}_{i},c_{jk})\cap\textrm{Span}(v^{\prime}_{j},c_{ik})\cap\textrm{Span}(v^{\prime}_{k},c_{ij})$.
+Given the equivalence between $v_{i}v_{j}v_{k}$ and $v^{\prime}_{i}v^{\prime}_{j}v^{\prime}_{k}$, it is natural to consider a dual model.
+Note that for any $c_{ijk}\in\mathbb{R}^{3}$, the intersection $\textrm{Span}(v_{r},c_{ijk})\cap v_{r}^{\bot}$ may not lie within $\mathbb{H}^{2}$.
+In this paper, we focus on the case that
+(12) $\textrm{Span}(v_{r},c_{ijk})\cap v_{r}^{\bot}\cap\mathbb{H}^{2}=\textrm{Span}(v_{r},c_{ijk})\cap E^{\prime}_{st}\neq\emptyset,\ \forall r\in\{i,j,k\}.$
+The condition (12) constrains the position of $c_{ijk}$ when it is space-like, namely, $c_{ijk}$ can only lie within certain specific open domains.
+Let $c^{\prime}_{st}=\textrm{Span}(v_{r},c_{ijk})\cap E^{\prime}_{st}$ denote the dual edge center, which is the unique point on the geodesic $E^{\prime}_{st}$ such that the signed distance from $c^{\prime}_{st}$ to $\{rt\}$ is $\theta_{st}$, and the signed distance from $c^{\prime}_{st}$ to $\{rs\}$ is $\theta_{ts}$.
+Thus
 (13) $\theta_{r}^{st}=\theta_{st}+\theta_{ts}.$
-Note that the signed distance $\theta_{st}$ is positive if $c^{\prime}_{st}$ is on the same side of $\{rt\}$ as $\{rs\}$ along the hyperbolic geodesic $E^{\prime}_{st}$, and negative otherwise.
-The above construction shows the following lemma.
+Note that the signed distance $\theta_{st}$ is positive if $c^{\prime}_{st}$ and $\{rs\}$ lie on the same side of $\{rt\}$ along the hyperbolic geodesic $E^{\prime}_{st}$, and negative otherwise.
+The above construction yields the following lemma.
 Lemma 2.6.
-The vectors $c_{ijk},c^{\prime}_{rs},v_{t}$ lie in the same 2-dimensional vector subspace of $\mathbb{R}^{3}$.
-By Proposition 2.4, the edge perpendicular $P^{\prime}_{ij}=\mathrm{Span}(v_{k},c_{ijk})\cap\mathbb{H}^{2}$ is through $c^{\prime}_{ij}$ and Lorentz orthogonal to $E^{\prime}_{ij}$.
-Hence, the three edge perpendiculars $P^{\prime}_{ij},P^{\prime}_{jk},P^{\prime}_{ki}$ of $\{ijk\}$ also intersect in a point $c_{ijk}$.
-Similar to [16, Lemma 2.4], we further have that the three perpendiculars $P^{\prime}_{ij},P^{\prime}_{jk},P^{\prime}_{ki}$ of $\{ijk\}$ intersect in a point if and only if
-(14) $\sinh\theta_{ij}\sinh\theta_{jk}\sinh\theta_{ki}=\sinh\theta_{ji}\sinh\theta_{kj}\sinh\theta_{ik}.$
-Hence, Definition 1.1 is equivalent to the following definition with respect to the parameter $\theta$.
-For precision, we rewrite the boundary component $i\in B$ as $\{j^{\prime}k^{\prime}\}\in B$.
-We use $B_{+}$ to denote the sets of oriented boundary components and label them with ordered pairs $(j^{\prime},k^{\prime})\in B_{+}$.
-Definition 2.7.
-Let $(\Sigma,\mathcal{T})$ be an ideally triangulated surface with boundary.
-An assignment of partial edge lengths is a map $\theta:B_{+}\rightarrow\mathbb{R}$, such that $\theta_{i}^{jk}=\theta_{jk}+\theta_{kj}>0$ for each oriented boundary arc $(j^{\prime},k^{\prime})\in B_{+}$ and formula (14) holds for every right-angled hyperbolic hexagon $\{ijk\}\in F$.
-Recall the definition of $h_{rs}$, which is the signed distance of the face center $c_{ijk}$ to the hyperbolic geodesic $E_{rs}$ and is positive if $c_{ijk}$ is on the same side of the geodesic $E_{rs}$ as the hyper-ideal hyperbolic triangle $v_{i}v_{j}v_{k}$ and negative otherwise (or zero if $c_{ijk}$ is on the geodesic $E_{ij}$).
-Similarly, we denote $q_{rs}$ as the signed distance from $c_{ijk}$ to $c^{\prime}_{rs}$, which is defined to be positive if $c_{ijk}$ is on the same side of the geodesic $E^{\prime}_{rs}$ as the polar triangle $v^{\prime}_{i}v^{\prime}_{j}v^{\prime}_{k}$ and negative otherwise (or zero if $c_{ijk}$ is on the geodesic $E^{\prime}_{rs}$).
+The vectors $c_{ijk}$, $c^{\prime}_{rs}$, and $v_{t}$ lie in the same 2-dimensional time-like vector subspace of $\mathbb{R}^{3}$.
+By Proposition 2.3, let $P^{\prime}_{ij}$ denote the edge perpendicular passing through $c^{\prime}_{ij}$ and Lorentz orthogonal to $E^{\prime}_{ij}$.
+Then $P^{\prime}_{ij}=\textrm{Span}(v_{k},c_{ijk})\cap\mathbb{H}^{2}$.
+Consequently, the three edge perpendiculars $P^{\prime}_{ij},P^{\prime}_{jk},P^{\prime}_{ki}$ of $\{ijk\}$ intersect in a common point $c_{ijk}$.
+Recall the definition of the signed distance $h_{rs}$ from the face center $c_{ijk}$ to the hyperbolic geodesic $E_{rs}$.
+It is positive if $c_{ijk}$ lies on the same side of $E_{rs}$ as the right-angled hyperbolic hexagon $\{ijk\}$, negative otherwise, and zero if $c_{ijk}$ lies on $E_{rs}$.
+Similarly, let $q_{rs}$ denote the signed distance from $c_{ijk}$ to $E^{\prime}_{rs}$, which is defined to be positive if $c_{ijk}$ lies on the same side of the geodesic $E^{\prime}_{rs}$ as the polar triangle $v^{\prime}_{i}v^{\prime}_{j}v^{\prime}_{k}$, negative otherwise, and zero if $c_{ijk}$ lies on $E^{\prime}_{rs}$.
 Please refer to Figure 3.
 Figure 3. Data for two hyper-ideal hyperbolic triangles in the Klein model
-In the Klein model, the hyper-ideal hyperbolic triangle $v_{i}v_{j}v_{k}$ and the polar triangle $v^{\prime}_{i}v^{\prime}_{j}v^{\prime}_{k}$ divide the unit disk into 13 open domains, denoted by $D_{1},D_{2},...,D_{13}$ respectively.
+In the Klein model, the hyper-ideal hyperbolic triangle $v_{i}v_{j}v_{k}$ and its polar triangle $v^{\prime}_{i}v^{\prime}_{j}v^{\prime}_{k}$ partition the unit disk into 13 open domains, denoted by $D_{1},D_{2},...,D_{13}$ respectively.
 Please refer to Figure 4.
-Figure 4. Division for a right-angled hyperbolic hexagon $\{ijk\}$ in the Klein model. If $c_{ijk}$ is space-like, then it can be in red and blue open domains.
-By the definition of $h$ and $q$, one can directly obtain the relationships between the position of $c_{ijk}$ and the signs of $h$ and $q$ in these 13 open domains, as shown in Table 1.
-For example, suppose $c_{ijk}$ is in the domain $D_{1}$.
-Then $c_{ijk}$ is on the opposite side of $E_{ik}$ as $v_{i}v_{j}v_{k}$ and hence $h_{ik}<0$.
-And $c_{ijk}$ is on the same side of $E_{ij},E_{jk}$ as $v_{i}v_{j}v_{k}$ and hence $h_{ij}>0,h_{jk}>0$.
-Similarly, $c_{ijk}$ is also on the opposite side of $E^{\prime}_{ij}$ as $v^{\prime}_{i}v^{\prime}_{j}v^{\prime}_{k}$ and hence $q_{ij}<0$.
-And $c_{ijk}$ is on the same side of $E^{\prime}_{ik},E^{\prime}_{jk}$ as $v^{\prime}_{i}v^{\prime}_{j}v^{\prime}_{k}$ and hence $q_{jk}>0,q_{ik}>0$.
-If $c_{ijk}$ is space-like, then $c_{ijk}$ can be only in one of 12 open domains and each domain is bounded by two tangential lines and unit circle, denoted by $D_{\mathrm{I}},D_{\mathrm{II}},D_{\mathrm{III}},D_{\mathrm{IV}},D_{\mathrm{V}},D_{\mathrm{VI}}$ and $D_{\mathrm{i}},D_{\mathrm{ii}},D_{\mathrm{iii}},D_{\mathrm{iv}},D_{\mathrm{v}},D_{\mathrm{vi}}$ respectively.
+Figure 4. Partition of the right-angled hyperbolic hexagon $\{ijk\}$ in the Klein model. If $c_{ijk}$ is space-like, it may lie in the open domains bounded by two red tangents or two blue tangents.
+By the definitions of $h$ and $q$, one can directly obtain the relationships between the position of $c_{ijk}$ and the signs of $h$ and $q$ in these 13 open domains, as shown in Table 1.
+For example, if $c_{ijk}\in D_{1}$, then $c_{ijk}$ lies on the opposite side of $E_{ik}$ as $v_{i}v_{j}v_{k}$, so $h_{ik}<0$.
+Moreover, $c_{ijk}$ lies on the same side of $E_{ij}$ and $E_{jk}$ as $v_{i}v_{j}v_{k}$, which implies $h_{ij}>0$ and $h_{jk}>0$.
+Similarly, $c_{ijk}$ lies on the opposite side of $E^{\prime}_{ij}$, and the same side of $E^{\prime}_{ik}$ and $E^{\prime}_{jk}$ as $v^{\prime}_{i}v^{\prime}_{j}v^{\prime}_{k}$, which implies $q_{ij}<0$, $q_{jk}>0$, and $q_{ik}>0$.
+If $c_{ijk}$ is space-like, it can lie only in one of 12 open domains, each bounded by two tangent lines and the unit circle.
+These domains are denoted by $D_{\mathrm{I}},D_{\mathrm{II}},D_{\mathrm{III}},D_{\mathrm{IV}},D_{\mathrm{V}},D_{\mathrm{VI}}$ and $D_{\mathrm{i}},D_{\mathrm{ii}},D_{\mathrm{iii}},D_{\mathrm{iv}},D_{\mathrm{v}},D_{\mathrm{vi}}$ respectively.
+Specifically, the domain $D_{\mathrm{i}}$ is enclosed by the tangents from vertices $v_{k}$ and $v_{j}^{\prime}$ to the unit circle, together with the unit circle itself.
+The domain $D_{\mathrm{I}}$ is enclosed by the tangents from vertices $v_{i}^{\prime}$ and $v_{j}^{\prime}$ to the unit circle, together with the unit circle itself.
+The remaining domains are defined analogously.
 Please refer to Figure 4.
-If $c_{ijk}$ is not in one of 12 open domains, then either $\mathrm{Span}(c_{ijk},v_{r})\cap\mathbb{H}^{2}=\emptyset$ or $\mathrm{Span}(c_{ijk},v^{\prime}_{r})\cap\mathbb{H}^{2}=\emptyset$, which contradicts with the assumption that the edge centers $c_{st}\in E_{rs}$ and $c^{\prime}_{st}\in E^{\prime}_{rs}$ are in $\mathbb{H}^{2}$ for $\{r,s,t\}=\{i,j,k\}$.
+If $c_{ijk}$ does not lie in any of these 12 open domains, then either $\textrm{Span}(c_{ijk},v_{r})\cap\mathbb{H}^{2}=\emptyset$ or $\textrm{Span}(c_{ijk},v^{\prime}_{r})\cap\mathbb{H}^{2}=\emptyset$, which contradicts the assumption that the edge centers $c_{st}\in E_{rs}$ and $c^{\prime}_{st}\in E^{\prime}_{rs}$ are in $\mathbb{H}^{2}$ for $\{r,s,t\}=\{i,j,k\}$.
 Similar to the case that $c_{ijk}$ is time-like, one can also directly obtain the relationships between the position of $c_{ijk}$ and the signs of $h$ and $q$ in these 12 open domains, as shown in Table 1.
-By moving $c_{ijk}$ to the light cone continuously, one can obtain similar results.
-Note that $c_{ijk}$ can only be on the unit disk that removes 12 points.
+By continuously moving $c_{ijk}$ to the light cone, analogous results are obtained.
+Note that if $c_{ijk}$ is light-like, then it can lie only on the unit disk with 12 points removed.
 Domains $D_{1}$($D_{\mathrm{i}}$) $D_{2}$($D_{\mathrm{I}}$) $D_{3}$($D_{\mathrm{ii}}$) $D_{4}$($D_{\mathrm{II}}$) $D_{5}$($D_{\mathrm{iii}}$) $D_{6}$($D_{\mathrm{III}}$) $D_{7}$($D_{\mathrm{iv}}$) $D_{8}$($D_{\mathrm{IV}}$) $D_{9}$($D_{\mathrm{v}}$) $D_{10}$($D_{\mathrm{V}}$) $D_{11}$($D_{\mathrm{vi}}$) $D_{12}$($D_{\mathrm{VI}}$) $D_{13}$
@@
 Table 1. Relationships between the position of $c_{ijk}$ and the signs of $h$ and $q$
-In the following, we provide another way to obtain Table 1.
+In the following, we present an alternative derivation of Table 1.
 Suppose $c_{ijk}$ is time-like.
-Without loss of generality, we assume $c_{ijk}$ is in the domain $D_{1}$.
-By Lemma 2.5, the vectors $c_{ijk},c_{ik},v^{\prime}_{j}$ lie in the same 2-dimensional vector subspace of $\mathbb{R}^{3}$.
-Due to $c_{ijk}$ is not in the polar triangle $v^{\prime}_{i}v^{\prime}_{j}v^{\prime}_{k}$, then $c_{ik}\in E_{ik}$ is not in $\{ik\}$.
-Furthermore, since $c_{ijk}$ is in $D_{1}$, then $c_{ik}$ is on the opposite side of $k$ as $i$.
-By the definition of $d$ in Remark 1.3, we have $d_{ki}<0,d_{ik}>0$.
-Similarly, we have $d_{kj}<0,d_{jk}>0,d_{ij}>0,d_{ji}>0$.
-By Lemma 2.6, the vectors $c_{ijk},c^{\prime}_{ij},v_{k}$ lie in the same 2-dimensional vector subspace of $\mathbb{R}^{3}$.
-Due to $c_{ijk}$ is not in the hyper-ideal hyperbolic triangle $v_{i}v_{j}v_{k}$, then $c^{\prime}_{ij}\in E^{\prime}_{ij}$ is not in $k$.
-Furthermore, since $c_{ijk}$ is in $D_{1}$, then $c^{\prime}_{ij}$ is on the opposite side of $\{ik\}$ as $\{jk\}$ and hence $\theta_{ji}<0,\theta_{ij}>0$.
-Similarly, we have $\theta_{jk}<0,\theta_{kj}>0,\theta_{ik}>0,\theta_{ki}>0$.
-By Corollary 2.11, we have $h_{ik}<0,h_{jk}>0,h_{ij}>0$ and $q_{ij}<0,q_{ik}>0,q_{jk}>0$.
-The case that $c_{ijk}$ is space-like follows from the following two lemmas.
+Without loss of generality, we assume $c_{ijk}\in D_{1}$.
+By Lemma 2.5, the vectors $c_{ijk}$, $c_{ik}$, and $v^{\prime}_{j}$ lie in the same 2-dimensional vector subspace of $\mathbb{R}^{3}$.
+Since $c_{ijk}\notin v^{\prime}_{i}v^{\prime}_{j}v^{\prime}_{k}$, then $c_{ik}\notin\{ik\}$.
+Furthermore, since $c_{ik}$ lies on the opposite side of $k$ as $i$.
+by the definition of $d$ in Remark 1.3, we have $d_{ki}<0$ and $d_{ik}>0$.
+Similarly, $d_{kj}<0$, $d_{jk}>0$, $d_{ij}>0$, and $d_{ji}>0$.
+By Lemma 2.6, the vectors $c_{ijk}$, $c^{\prime}_{ij}$, and $v_{k}$ lie in the same 2-dimensional vector subspace of $\mathbb{R}^{3}$.
+Since $c_{ijk}\notin v_{i}v_{j}v_{k}$, it follows that $c^{\prime}_{ij}\notin k$.
+Furthermore, since $c^{\prime}_{ij}$ is on the opposite side of $\{ik\}$ as $\{jk\}$, it follows that $\theta_{ji}<0$ and $\theta_{ij}>0$.
+Similarly, $\theta_{jk}<0$, $\theta_{kj}>0$, $\theta_{ik}>0$, and $\theta_{ki}>0$.
+By Corollary 2.10, we have $h_{ik}<0$, $h_{jk}>0$, $h_{ij}>0$, $q_{ij}<0$, $q_{ik}>0$, and $q_{jk}>0$.
+The case that $c_{ijk}$ is space-like is established via the two lemmas below.
 Lemma 2.7.
-The signs of $q$ and $h$ in the domains $D_{\mathrm{I}},D_{\mathrm{II}},D_{\mathrm{III}},D_{\mathrm{IV}},D_{\mathrm{V}},D_{\mathrm{VI}}$ are the same as those in the domains $D_{2},D_{4},D_{6},D_{8},D_{10},D_{12}$, respectively.
+The signs of $q$ and $h$ in the domains $D_{\mathrm{I}},D_{\mathrm{II}},D_{\mathrm{III}},D_{\mathrm{IV}},D_{\mathrm{V}},D_{\mathrm{VI}}$ coincide with those in the domains $D_{2},D_{4},D_{6},D_{8},D_{10},D_{12}$, respectively.
 Proof.
-We only prove the signs of $q$ and $h$ in the domain $D_{\mathrm{V}}$ are the same as those in the domain $D_{10}$, and the remaining cases are similar.
-If $c_{ijk}$ is in the domain $D_{10}$, then $q_{jk}<0$ and the rest is positive by Table 1.
-Suppose $c_{ijk}$ is in the domain $D_{\mathrm{V}}$.
-By Lemma 2.5, the vectors $c_{ijk},c_{ij},v^{\prime}_{k}$ lie in the same 2-dimensional vector subspace of $\mathbb{R}^{3}$.
-Due to $c_{ijk}$ is not in the polar triangle $v^{\prime}_{i}v^{\prime}_{j}v^{\prime}_{k}$, then $c_{ij}\in E_{ij}$ is not in $\{ij\}$.
-Furthermore, since $c_{ijk}$ is in $D_{\mathrm{V}}$, then $c_{ij}$ is on the opposite side of $j$ as $i$ and hence $d_{ij}<0,d_{ji}>0$.
-Similarly, we have $d_{ik}<0,d_{ki}>0,d_{jk}>0,d_{kj}>0$.
-By Proposition 2.12, we have $h_{ij}>0,h_{jk}>0,h_{ik}>0$.
-Note that $c_{ijk}$ is in the polar triangle $v^{\prime}_{i}v^{\prime}_{j}v^{\prime}_{k}$.
-By Lemma 2.6, the vectors $c_{ijk},c^{\prime}_{rs},v_{t}$ lie in the same 2-dimensional vector subspace of $\mathbb{R}^{3}$ for $\{r,s,t\}=\{i,j,k\}$.
-Then $c^{\prime}_{rs}\in E^{\prime}_{rs}$ is in $t$.
-Hence, $\theta_{ij}>0,\theta_{ji}>0,\theta_{jk}<0,\theta_{kj}>0,\theta_{ik}>0,\theta_{ki}>0$.
-By Proposition 2.12, we have $q_{ij}>0,q_{jk}<0,q_{ik}>0$.
+We only prove the signs of $q$ and $h$ in the domain $D_{\mathrm{V}}$ coincide with those in the domain $D_{10}$, and the remaining cases follow analogously.
+If $c_{ijk}\in D_{10}$, then $q_{jk}<0$ and the rest are positive by Table 1.
+Suppose $c_{ijk}\in D_{\mathrm{V}}$.
+By Lemma 2.5, the vectors $c_{ijk}$, $c_{ij}$, and $v^{\prime}_{k}$ lie in the same 2-dimensional vector subspace of $\mathbb{R}^{3}$.
+Since $c_{ijk}\notin v^{\prime}_{i}v^{\prime}_{j}v^{\prime}_{k}$, it follows that $c_{ij}\notin\{ij\}$.
+Furthermore, since $c_{ij}$ lies on the opposite side of $j$ as $i$, it follows that $d_{ij}<0$ and $d_{ji}>0$.
+Similarly, $d_{ik}<0$, $d_{ki}>0$, $d_{jk}>0$, and $d_{kj}>0$.
+By Proposition 2.11, we have $h_{ij}>0$, $h_{jk}>0$, and $h_{ik}>0$.
+Note that $c_{ijk}\in v_{i}v_{j}v_{k}$.
+By Lemma 2.6, the vectors $c_{ijk}$, $c^{\prime}_{rs}$, and $v_{t}$ lie in the same 2-dimensional vector subspace of $\mathbb{R}^{3}$ for $\{r,s,t\}=\{i,j,k\}$.
+Consequently, $c^{\prime}_{rs}\in t$.
+Hence, $\theta_{ij}>0$, $\theta_{ji}>0$, $\theta_{jk}<0$, $\theta_{kj}>0$, $\theta_{ik}>0$, and $\theta_{ki}>0$.
+By Proposition 2.11, we have $q_{ij}>0$, $q_{jk}<0$, and $q_{ik}>0$.
 Q.E.D.
 Lemma 2.8.
-The signs of $q$ and $h$ in the domains $D_{\mathrm{i}},D_{\mathrm{ii}},D_{\mathrm{iii}},D_{\mathrm{iv}},D_{\mathrm{v}},D_{\mathrm{vi}}$ are the same as those in the domains $D_{1},D_{3},D_{5},D_{7},D_{9},D_{11}$, respectively.
+The signs of $q$ and $h$ in the domains $D_{\mathrm{i}},D_{\mathrm{ii}},D_{\mathrm{iii}},D_{\mathrm{iv}},D_{\mathrm{v}},D_{\mathrm{vi}}$ coincide with those in the domains $D_{1},D_{3},D_{5},D_{7},D_{9},D_{11}$, respectively.
 Proof.
-We only prove the signs of $q$ and $h$ in the domain $D_{\mathrm{vi}}$ are the same as those in the domain $D_{11}$, and the remaining cases are similar.
-If $c_{ijk}$ is in the domain $D_{11}$, then $q_{i}<0,h_{j}<0$ and the rest is positive by Table 1.
-Suppose $c_{ijk}$ is in the domain $D_{\mathrm{vi}}$.
-By Lemma 2.5, the vectors $c_{ijk},c_{ij},v^{\prime}_{k}$ lie in the same 2-dimensional vector subspace of $\mathbb{R}^{3}$.
-Due to $c_{ijk}$ is not in the polar triangle $v^{\prime}_{i}v^{\prime}_{j}v^{\prime}_{k}$, then $c_{ij}\in E_{ij}$ is not in $\{ij\}$.
-Furthermore, since $c_{ijk}$ is in $D_{\mathrm{vi}}$, then $c_{ij}$ is on the opposite side of $j$ as $i$ and hence $d_{ij}<0,d_{ji}>0$.
-Similarly, we have $d_{ik}<0,d_{ki}>0,d_{jk}>0,d_{kj}>0$.
-By Proposition 2.14, we have $h_{ij}<0,h_{jk}>0,h_{ik}>0$.
-By Lemma 2.6, the vectors $c_{ijk},c^{\prime}_{jk},v_{i}$ lie in the same 2-dimensional vector subspace of $\mathbb{R}^{3}$.
-Due to $c_{ijk}$ is not in the hyper-ideal hyperbolic triangle $v_{i}v_{j}v_{k}$, then $c^{\prime}_{jk}\in E^{\prime}_{jk}$ is not in $i$.
-Furthermore, since $c_{ijk}$ is in $D_{\mathrm{vi}}$, then $c^{\prime}_{jk}$ is on the opposite side of $\{ik\}$ as $\{ij\}$ and hence $\theta_{jk}<0,\theta_{kj}>0$.
-Similarly, we have $\theta_{ji}<0,\theta_{ij}>0,\theta_{ik}>0,\theta_{ki}>0$.
-By Proposition 2.14, we have $q_{ij}>0,q_{jk}<0,q_{ik}>0$.
+We only prove the signs of $q$ and $h$ in the domain $D_{\mathrm{vi}}$ coincide with those in the domain $D_{11}$, and the remaining cases follow analogously.
+If $c_{ijk}\in D_{11}$, then $q_{jk}<0$, $h_{ik}<0$, and the rest are positive by Table 1.
+Suppose $c_{ijk}\in D_{\mathrm{vi}}$.
+By Lemma 2.5, the vectors $c_{ijk}$, $c_{ij}$, and $v^{\prime}_{k}$ lie in the same 2-dimensional vector subspace of $\mathbb{R}^{3}$.
+Since $c_{ijk}\notin v^{\prime}_{i}v^{\prime}_{j}v^{\prime}_{k}$, it follows that $c_{ij}\notin\{ij\}$.
+Furthermore, since $c_{ij}$ lies on the opposite side of $j$ as $i$, it follows that $d_{ij}<0$ and $d_{ji}>0$.
+Similarly, $d_{ik}<0$, $d_{ki}>0$, $d_{jk}>0$, and $d_{kj}>0$.
+By Proposition 2.13, we have $h_{ij}<0$, $h_{jk}>0$, and $h_{ik}>0$.
+By Lemma 2.6, the vectors $c_{ijk}$, $c^{\prime}_{jk}$, and $v_{i}$ lie in the same 2-dimensional vector subspace of $\mathbb{R}^{3}$.
+Since $c_{ijk}\notin v_{i}v_{j}v_{k}$, it follows that $c^{\prime}_{jk}\notin i$.
+Furthermore, since $c^{\prime}_{jk}$ lies on the opposite side of $\{ik\}$ as $\{ij\}$, it follows that $\theta_{jk}<0$ and $\theta_{kj}>0$.
+Similarly, $\theta_{ji}<0$, $\theta_{ij}>0$, $\theta_{ik}>0$, and $\theta_{ki}>0$.
+By Proposition 2.13, we have $q_{ij}>0$, $q_{jk}<0$, and $q_{ik}>0$.
 Q.E.D.
-By direct calculations, we can get the following results.
-As the proofs of the Lemma 2.10, Proposition 2.12 and Proposition 2.14 are too long, we defer them to Appendix A.
-For simplicity, we set $h_{t}=h_{rs},\ q_{t}=q_{rs}$ for $\{r,s,t\}=\{i,j,k\}$.
+By direct calculations, we obtain the following results.
+As the proofs of Lemma 2.9, Proposition 2.11 and Proposition 2.13 are too long, we defer them to Appendix A.
+For simplicity, we set $h_{t}=h_{rs}$ and $q_{t}=q_{rs}$ for $\{r,s,t\}=\{i,j,k\}$.
 Lemma 2.9.
@@
 $\sinh h_{r}=\cosh q_{s}\cdot\sinh\theta_{rt}=\cosh q_{t}\cdot\sinh\theta_{rs}.$
-As a direct corollary of Lemma 2.10, we have
+As a direct corollary of Lemma 2.9, we have the following result.
 Corollary 2.10.
-If $c_{ijk}$ is time-like, then the signs of $q_{r}$, $d_{rt}$ and $d_{rs}$ are the same and the signs of $h_{r}$, $\theta_{rs}$ and $\theta_{rt}$ are the same.
+If $c_{ijk}$ is time-like, then the signs of $q_{r}$, $d_{rt}$, and $d_{rs}$ coincide, and the signs of $h_{r}$, $\theta_{rs}$ and $\theta_{rt}$ coincide.
 Proposition 2.11.
@@
 $\cosh h_{k}=-\sinh q_{i}\cdot\sinh\theta_{kj}=\sinh q_{j}\cdot\sinh\theta_{ki}.$
-If $c_{ijk}\in\mathbb{R}^{3}$ lies in the open domain $D_{\mathrm{II}}$, then we swap $h$ and $q$, and $\theta$ and $d$.
-If $c_{ijk}$ lies in the domains $D_{\mathrm{I}},D_{\mathrm{III}},D_{\mathrm{IV}}$ or $D_{\mathrm{VI}}$, we have similar propositions.
-Using these propositions, we can prove Lemma 2.8 and further obtain Table 1.
-Moreover, from Table 1, we see that $c_{ijk}$ lies in the open domains $D_{\mathrm{I}},D_{\mathrm{II}},D_{\mathrm{III}},D_{\mathrm{IV}},D_{\mathrm{V}}$ or $D_{\mathrm{VI}}$ if and only if one of $q_{i},q_{j},q_{k},h_{i},h_{j},h_{k}$ is negative and the others are positive.
-Hence, these propositions can be written as a unified form.
+If $c_{ijk}\in\mathbb{R}^{3}$ lies in the open domain $D_{\mathrm{II}}$, then we swap $h$ with $q$, and $\theta$ with $d$.
+If $c_{ijk}$ lies in the domains $D_{\mathrm{I}},D_{\mathrm{III}},D_{\mathrm{IV}}$ or $D_{\mathrm{VI}}$, analogous propositions hold.
+These propositions enable us to prove Lemma 2.7 and further derive Table 1.
+Moreover, Table 1 shows that $c_{ijk}$ lies in the open domains $D_{\mathrm{I}},D_{\mathrm{II}},D_{\mathrm{III}},D_{\mathrm{IV}},D_{\mathrm{V}}$ or $D_{\mathrm{VI}}$ if and only if one of $q_{i},q_{j},q_{k},h_{i},h_{j},h_{k}$ is negative and the rest are positive.
+Consequently, these propositions can be written in a unified form.
 Lemma 2.12.
@@
 $\cosh h_{t}=-\sinh q_{r}\cdot\sinh\theta_{ts}=\sinh q_{s}\cdot\sinh\theta_{tr}.$
-If $h_{r}<0$, then we swap $h,\ \theta$ and $q,\ d$.
+If $h_{r}<0$, then we swap $h$ with $q$, and $\theta$ with $d$.
 Proposition 2.13.
@@
 $\cosh h_{k}=-\sinh q_{i}\cdot\sinh\theta_{kj}=\sinh q_{j}\cdot\sinh\theta_{ki}.$
-Similarly, if $c_{ijk}$ lies in the domains $D_{\mathrm{i}},D_{\mathrm{ii}},D_{\mathrm{iii}},D_{\mathrm{iv}}$ or $D_{\mathrm{v}}$, we have the corresponding propositions.
-Using these propositions, we can prove Lemma 2.9 and further obtain Table 1.
-Moreover, from Table 1, we see that $c_{ijk}$ lies in the open domains $D_{\mathrm{i}},D_{\mathrm{ii}},D_{\mathrm{iii}},D_{\mathrm{iv}},D_{\mathrm{v}}$ or $D_{\mathrm{vi}}$ if and only if $q_{r}<0,h_{s}<0$ and the rest are positive.
-Hence, these propositions can also be written as a unified form.
+Similarly, if $c_{ijk}$ lies in the domains $D_{\mathrm{i}},D_{\mathrm{ii}},D_{\mathrm{iii}},D_{\mathrm{iv}}$ or $D_{\mathrm{v}}$, corresponding propositions hold.
+These propositions allow us to prove Lemma 2.8 and further derive Table 1.
+Moreover, Table 1 shows that $c_{ijk}$ lies in the open domains $D_{\mathrm{i}},D_{\mathrm{ii}},D_{\mathrm{iii}},D_{\mathrm{iv}},D_{\mathrm{v}}$ or $D_{\mathrm{vi}}$ if and only if $q_{r}<0$, $h_{s}<0$, and the rest are positive.
+Consequently, these propositions also admit a unified form.
 Lemma 2.14.
-Let $c_{ijk}\in\mathbb{R}^{3}$ lie in one of the open domains $D_{\mathrm{i}},D_{\mathrm{ii}},D_{\mathrm{iii}},D_{\mathrm{iv}},D_{\mathrm{v}},D_{\mathrm{vi}}$. If $q_{r}<0,h_{s}<0$, then
+Let $c_{ijk}\in\mathbb{R}^{3}$ lie in one of the open domains $D_{\mathrm{i}},D_{\mathrm{ii}},D_{\mathrm{iii}},D_{\mathrm{iv}},D_{\mathrm{v}},D_{\mathrm{vi}}$. If $q_{r}<0$ and $h_{s}<0$, then
 $\cosh q_{r}=\sinh h_{s}\cdot\sinh d_{rt}=-\sinh h_{t}\cdot\sinh d_{rs},$
@@
 $\cosh h_{t}=-\sinh q_{r}\cdot\sinh\theta_{ts}=\sinh q_{s}\cdot\sinh\theta_{tr}.$
-As a direct corollary of Lemma 2.13 and Lemma 2.15, we have
+As a direct corollary of Lemma 2.12 and Lemma 2.14, we have the following result.
 Corollary 2.15.
-If $c_{ijk}$ is space-like, then the signs of $q_{r}$, $d_{rt}$ and $d_{rs}$ are the same and the signs of $h_{r}$, $\theta_{rs}$ and $\theta_{rt}$ are the same.
+If $c_{ijk}$ is space-like, then the signs of $q_{r}$, $d_{rt}$, and $d_{rs}$ coincide, and the signs of $h_{r}$, $\theta_{rs}$, and $\theta_{rt}$ coincide.
 Combining Corollary 2.10 and Corollary 2.15 gives the following lemma.
 Lemma 2.16.
-For any right-angled hyperbolic hexagon $\{ijk\}\in F$, the signs of $q_{r}$, $d_{rt}$ and $d_{rs}$ are the same and the signs of $h_{r}$, $\theta_{rs}$ and $\theta_{rt}$ are the same.
+For any right-angled hyperbolic hexagon $\{ijk\}\in F$, the signs of $q_{r}$, $d_{rt}$, and $d_{rs}$ coincide, and the signs of $h_{r}$, $\theta_{rs}$, and $\theta_{rt}$ coincide.
 Proof.
 If $c_{ijk}$ is time-like, the conclusion follows from Corollary 2.10.
-If $c_{ijk}$ is space-like, the conclusion follows from Corollary 2.16.
-If $c_{ijk}$ is light-like, the conclusion follows from continuity by moving the center to the light cone.
+If $c_{ijk}$ is space-like, the conclusion is derived from Corollary 2.15.
+If $c_{ijk}$ is light-like, the conclusion holds by continuity upon moving $c_{ijk}$ to the light cone.
 Q.E.D.
 2.3. Variational formulas of generalized angle
-We divide Theorem 1.12 into Lemma 2.18, Lemma 2.20 and Lemma 2.22, and prove them respectively.
+We decompose Theorem 1.11 into Lemma 2.17, Lemma 2.19, and Lemma 2.21, and prove them respectively.
 Lemma 2.17.
-For a right-angled hyperbolic hexagon $\{ijk\}\in F$, we have
+For any right-angled hyperbolic hexagon $\{ijk\}\in F$, we have
 (14) $\frac{\partial\theta^{jk}_{i}}{\partial f_{j}}=\frac{-1}{\sinh d_{ji}}\frac{\tanh^{\beta}h_{ij}}{\sinh l_{ij}},$
-where $\beta=1$ if $c_{ijk}$ is time-like and $\beta=-1$ if $c_{ijk}$ is space-like.
-If $c_{ijk}$ is light-like, we interpret the formula as $\tanh^{\beta}h_{ij}=\tanh^{\beta}h_{jk}=\tanh^{\beta}h_{ki}=1$.
+where $\beta=1$ if $c_{ijk}$ is time-like, and $\beta=-1$ if $c_{ijk}$ is space-like.
+If $c_{ijk}$ is light-like, the formula is interpreted as $\tanh^{\beta}h_{ij}=\tanh^{\beta}h_{jk}=\tanh^{\beta}h_{ki}=1$.
 Proof.
-Without loss of generality, we assume $||v_{i}||=||v_{j}||=||v_{k}||=1$ and $||c_{ijk}||^{2}=\pm 1$.
-For simplicity, we set $l_{r}=l_{st},\ \theta_{r}=\theta^{st}_{r},\ h_{r}=h_{st},\ q_{r}=q_{st}$ and $A=\sinh l_{r}\sinh l_{s}\sinh\theta_{t}$, where $\{r,s,t\}=\{i,j,k\}$.
-By generalized hyperbolic cosine law, i.e., $\cosh l_{r}=\cosh\theta_{r}\sinh l_{s}\sinh l_{t}-\cosh l_{s}\cosh l_{t}$, we have
+Without loss of generality, we assume that $\|v_{i}\|=\|v_{j}\|=\|v_{k}\|=1$ and $\|c_{ijk}\|^{2}=\pm 1$.
+For simplicity, we set $l_{r}=l_{st},\ \theta_{r}=\theta^{st}_{r},\ h_{r}=h_{st},\ q_{r}=q_{st}$, and $A=\sinh l_{r}\sinh l_{s}\sinh\theta_{t}$, where $\{r,s,t\}=\{i,j,k\}$.
+By generalized hyperbolic cosine law, i.e., $\cosh l_{r}=\cosh\theta_{r}\sinh l_{s}\sinh l_{t}-\cosh l_{s}\cosh l_{t}$, we derive
 (15) $\frac{\partial\theta_{i}}{\partial l_{i}}=\frac{\sinh l_{i}}{A},\ \frac{\partial\theta_{i}}{\partial l_{j}}=-\frac{\sinh l_{i}\cosh\theta_{k}}{A},\ \frac{\partial\theta_{i}}{\partial l_{k}}=-\frac{\sinh l_{i}\cosh\theta_{j}}{A}.$
-By (2) and (3), we have
+Using (3) and (4), we obtain
 (16) $\frac{\partial\theta_{i}}{\partial f_{j}}$ $=\frac{\partial\theta_{i}}{\partial l_{i}}\frac{\partial l_{i}}{\partial f_{j}}+\frac{\partial\theta_{i}}{\partial l_{j}}\frac{\partial l_{j}}{\partial f_{j}}+\frac{\partial\theta_{i}}{\partial l_{k}}\frac{\partial l_{k}}{\partial f_{j}}$
@@
 $\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ -\sinh\theta_{ki}\sinh\theta_{ik}\coth d_{ji}),$
-where (13) is used in the forth line. Set
+where (13) is used in the fourth line. Define
 (17) $L_{1}=\coth d_{jk}-\cosh\theta_{ki}\cosh\theta_{ik}\coth d_{ji},\ L_{2}=-\sinh\theta_{ki}\sinh\theta_{ik}\coth d_{ji}.$
-Then
+It follows from (16) that
 (18) $\frac{\partial\theta_{i}}{\partial f_{j}}=\frac{1}{\sinh l_{k}\sinh\theta_{j}}\cdot(L_{1}+L_{2}).$
-To calculate $L_{1}$ and $L_{2}$, we first give some formulas. Please refer to Figure 3. Applying Proposition 2.1 to the generalized right-angled hyperbolic triangle $c_{ijk}c_{ij}v_{j}$ gives
-(20) $v_{j}\ast c_{ijk}=-(c_{ij}\ast c_{ijk})(v_{j}\ast c_{ij})=\sinh d_{ji}\cdot(c_{ij}\ast c_{ijk}),$
-where $v_{j}\ast c_{ij}=-\sinh d_{ji}$ follows from Proposition 2.2 (ii). Similarly, in the generalized right-angled hyperbolic triangle $c_{ijk}c_{jk}v_{j}$, we have
+To compute $L_{1}$ and $L_{2}$, some preliminary results are required.
+We begin by referring to Figure 3.
+By Proposition 2.2 (ii), we have $v_{j}\ast c_{ij}=+\sinh d_{\mathbb{H}}(v_{j}^{\bot},c_{ij})=+\sinh(-d_{ji})=-\sinh d_{ji}$, or $v_{j}\ast c_{ij}=-\sinh d_{\mathbb{H}}(v_{j}^{\bot},c_{ij})=-\sinh(+d_{ji})=-\sinh d_{ji}$.
+In both cases, it follows that $v_{j}\ast c_{ij}=-\sinh d_{ji}$.
+Applying Proposition 2.1 to the generalized right-angled hyperbolic triangle $c_{ijk}c_{ij}v_{j}$ yields
+(19) $v_{j}\ast c_{ijk}=-(c_{ij}\ast c_{ijk})(v_{j}\ast c_{ij})=\sinh d_{ji}\cdot(c_{ij}\ast c_{ijk}).$
+Similarly, in the generalized right-angled hyperbolic triangle $c_{ijk}c_{jk}v_{j}$, we derive
 (20) $v_{j}\ast c_{ijk}=-(v_{j}\ast c_{jk})(c_{jk}\ast c_{ijk})=\sinh d_{jk}\cdot(c_{jk}\ast c_{ijk}),$
@@
 (21) $\sinh d_{ji}\cdot(c_{ij}\ast c_{ijk})=\sinh d_{jk}\cdot(c_{jk}\ast c_{ijk}).$
-By Proposition 2.2 $\mathrm{(iii)}$, we have $(c_{ijk}\otimes v_{j})\ast(v_{j}\otimes c_{jk})=||v_{j}\otimes c_{ijk}||\cdot||v_{j}\otimes c_{jk}||(-\cosh\theta_{ik}).$ Moreover,
-$(c_{ijk}\otimes v_{j})\ast(v_{j}\otimes c_{jk})=(c_{jk}\ast c_{ijk})||v_{j}||^{2}-(v_{j}\ast c_{ijk})(v_{j}\ast c_{jk})=(c_{jk}\ast c_{ijk})\cdot\cosh^{2}d_{jk},$
-where (21) is used in the second equality. Since $||v_{j}\otimes c_{jk}||^{2}=(v_{j}\ast c_{jk})^{2}+1=\cosh^{2}d_{jk}$, then
-(23) $\cosh\theta_{ik}=-\frac{(c_{ijk}\otimes v_{j})\ast(v_{j}\otimes c_{jk})}{||v_{j}\otimes c_{ijk}||\cdot||v_{j}\otimes c_{jk}||}=-\frac{(c_{jk}\ast c_{ijk})\cdot\cosh d_{jk}}{||v_{j}\otimes c_{ijk}||}.$
-Similarly, $(c_{ijk}\otimes v_{j})\ast(v_{j}\otimes c_{ij})=||v_{j}\otimes c_{ijk}||\cdot||v_{j}\otimes c_{ij}||(-\cosh\theta_{ki}).$ Moreover,
-$(c_{ijk}\otimes v_{j})\ast(v_{j}\otimes c_{ij})=(c_{ij}\ast c_{ijk})||v_{j}||^{2}-(v_{j}\ast c_{ijk})(v_{j}\ast c_{ij})=(c_{ij}\ast c_{ijk})\cdot\cosh^{2}d_{ji},$
-where (20) is used in the second equality. Since $||v_{j}\otimes c_{ij}||^{2}=(v_{j}\ast c_{ij})^{2}+1=\cosh^{2}d_{ji}$, then
-(24) $\cosh\theta_{ki}=-\frac{(c_{ijk}\otimes v_{j})\ast(v_{j}\otimes c_{ij})}{||v_{j}\otimes c_{ijk}||\cdot||v_{j}\otimes c_{ij}||}=-\frac{(c_{ij}\ast c_{ijk})\cdot\cosh d_{ji}}{||v_{j}\otimes c_{ijk}||}.$
+By Proposition 2.2 (iii), we have $(c_{ijk}\otimes v_{j})\ast(v_{j}\otimes c_{jk})=\|v_{j}\otimes c_{ijk}\|\cdot\|v_{j}\otimes c_{jk}\|(-\cosh\theta_{ik}).$ Moreover, direct calculation yields
+$(c_{ijk}\otimes v_{j})\ast(v_{j}\otimes c_{jk})=(c_{jk}\ast c_{ijk})\|v_{j}\|^{2}-(v_{j}\ast c_{ijk})(v_{j}\ast c_{jk})=(c_{jk}\ast c_{ijk})\cdot\cosh^{2}d_{jk},$
+where (20) is used in the second equality. Since $\|v_{j}\otimes c_{jk}\|^{2}=(v_{j}\ast c_{jk})^{2}+1=\cosh^{2}d_{jk}$, it follows that
+(22) $\cosh\theta_{ik}=-\frac{(c_{ijk}\otimes v_{j})\ast(v_{j}\otimes c_{jk})}{\|v_{j}\otimes c_{ijk}\|\cdot\|v_{j}\otimes c_{jk}\|}=-\frac{(c_{jk}\ast c_{ijk})\cdot\cosh d_{jk}}{\|v_{j}\otimes c_{ijk}\|}.$
+Similarly, $(c_{ijk}\otimes v_{j})\ast(v_{j}\otimes c_{ij})=\|v_{j}\otimes c_{ijk}\|\cdot\|v_{j}\otimes c_{ij}\|(-\cosh\theta_{ki}).$ Moreover,
+$(c_{ijk}\otimes v_{j})\ast(v_{j}\otimes c_{ij})=(c_{ij}\ast c_{ijk})\|v_{j}\|^{2}-(v_{j}\ast c_{ijk})(v_{j}\ast c_{ij})=(c_{ij}\ast c_{ijk})\cdot\cosh^{2}d_{ji},$
+where (19) is used in the second equality. Since $\|v_{j}\otimes c_{ij}\|^{2}=(v_{j}\ast c_{ij})^{2}+1=\cosh^{2}d_{ji}$, it follows that
+(23) $\cosh\theta_{ki}=-\frac{(c_{ijk}\otimes v_{j})\ast(v_{j}\otimes c_{ij})}{\|v_{j}\otimes c_{ijk}\|\cdot\|v_{j}\otimes c_{ij}\|}=-\frac{(c_{ij}\ast c_{ijk})\cdot\cosh d_{ji}}{\|v_{j}\otimes c_{ijk}\|}.$
 By the properties of Lorentzian cross product, we have
-(25) $||v_{j}\otimes c_{ijk}||^{2}=(v_{j}\otimes c_{ijk})*(v_{j}\otimes c_{ijk})=(v_{j}\ast c_{ijk})^{2}-||c_{ijk}||^{2},$
+(24) $\|v_{j}\otimes c_{ijk}\|^{2}=(v_{j}\otimes c_{ijk})*(v_{j}\otimes c_{ijk})=(v_{j}\ast c_{ijk})^{2}-\|c_{ijk}\|^{2},$
 and
-(26) $||c_{ij}\otimes c_{ijk}||^{2}=(c_{ij}\otimes c_{ijk})*(c_{ij}\otimes c_{ijk})=(c_{ij}\ast c_{ijk})^{2}+||c_{ijk}||^{2}.$
+(25) $\|c_{ij}\otimes c_{ijk}\|^{2}=(c_{ij}\otimes c_{ijk})*(c_{ij}\otimes c_{ijk})=(c_{ij}\ast c_{ijk})^{2}+\|c_{ijk}\|^{2}.$
 Adding (24) and (25) gives
-(27) $||v_{j}\otimes c_{ijk}||^{2}+||c_{ij}\otimes c_{ijk}||^{2}=$ $(v_{j}\ast c_{ijk})^{2}+(c_{ij}\ast c_{ijk})^{2}$
+(26) $\|v_{j}\otimes c_{ijk}\|^{2}+\|c_{ij}\otimes c_{ijk}\|^{2}$ $=(v_{j}\ast c_{ijk})^{2}+(c_{ij}\ast c_{ijk})^{2}$
 $=(c_{ij}\ast c_{ijk})^{2}\cdot\sinh^{2}d_{ji}+(c_{ij}\ast c_{ijk})^{2}$
@@
 (27) $\sinh^{2}\theta_{ki}$ $=\cosh^{2}\theta_{ki}-1$
-$=$ $\frac{1}{||v_{j}\otimes c_{ijk}||^{2}}[(c_{ij}\ast c_{ijk})^{2}\cosh^{2}d_{ji}-||v_{j}\otimes c_{ijk}||^{2}]$
-$=$ $\frac{||c_{ij}\otimes c_{ijk}||^{2}}{||v_{j}\otimes c_{ijk}||^{2}},$
+$=\frac{1}{\|v_{j}\otimes c_{ijk}\|^{2}}[(c_{ij}\ast c_{ijk})^{2}\cosh^{2}d_{ji}-\|v_{j}\otimes c_{ijk}\|^{2}]$
+$=\frac{\|c_{ij}\otimes c_{ijk}\|^{2}}{\|v_{j}\otimes c_{ijk}\|^{2}},$
 where (23) is used in the second line and (26) is used in the last line.
-Now we begin to calculate $L_{1}$ and $L_{2}$ in (18).
+We are now ready to compute $L_{1}$ and $L_{2}$ in (17). For $L_{2}$, we proceed as follows:
 $L_{2}$ $=-\sinh\theta_{ki}\sinh\theta_{ik}\cdot\coth d_{ji}$
-$=$ $-\sinh\theta_{ki}\cdot\sinh\theta_{ik}\cdot\frac{\cosh d_{ji}}{\sinh d_{ji}}\cdot\frac{||v_{j}\otimes c_{ijk}||}{c_{ij}\ast c_{ijk}}\cdot\frac{c_{ij}\ast c_{ijk}}{||v_{j}\otimes c_{ijk}||}$
-$=$ $\sinh\theta_{ki}\sinh\theta_{ik}\cdot\frac{||v_{j}\otimes c_{ijk}||}{(c_{ij}\ast c_{ijk})\sinh d_{ji}}\cdot\cosh\theta_{ki},$
-where (24) is used in the last line.
+$=-\sinh\theta_{ki}\cdot\sinh\theta_{ik}\cdot\frac{\cosh d_{ji}}{\sinh d_{ji}}\cdot\frac{\|v_{j}\otimes c_{ijk}\|}{c_{ij}\ast c_{ijk}}\cdot\frac{c_{ij}\ast c_{ijk}}{\|v_{j}\otimes c_{ijk}\|}$
+$=\sinh\theta_{ki}\sinh\theta_{ik}\cdot\frac{\|v_{j}\otimes c_{ijk}\|}{(c_{ij}\ast c_{ijk})\sinh d_{ji}}\cdot\cosh\theta_{ki},$
+where (23) is used in the last line. For $L_{1}$, we have
 $L_{1}$ $=\coth d_{jk}-\cosh\theta_{ki}\cosh\theta_{ik}\cdot\coth d_{ji}$
-$=$ $\frac{\cosh d_{jk}\cdot(c_{jk}\ast c_{ijk})}{\sinh d_{ji}\cdot(c_{ij}\ast c_{ijk})}+\frac{(c_{ij}\ast c_{ijk})\cosh d_{ji}}{||v_{j}\otimes c_{ijk}||}\cdot\cosh\theta_{ik}\cdot\frac{\cosh d_{ji}}{\sinh d_{ji}}$
-$=$ $\frac{\cosh\theta_{ik}}{\sinh d_{ji}\cdot(c_{ij}\ast c_{ijk})\cdot||v_{j}\otimes c_{ijk}||}\big{(}-||v_{j}\otimes c_{ijk}||^{2}+(c_{ij}\ast c_{ijk})^{2}\cosh^{2}d_{ji}\big{)}$
-$=$ $\frac{\cosh\theta_{ik}\cdot||c_{ij}\otimes c_{ijk}||^{2}}{\sinh d_{ji}\cdot(c_{ij}\ast c_{ijk})\cdot||v_{j}\otimes c_{ijk}||}$
-$=$ $\frac{\cosh\theta_{ik}\sinh^{2}\theta_{ki}||v_{j}\otimes c_{ijk}||}{\sinh d_{ji}(c_{ij}\ast c_{ijk})},$
-where (22) and (24) are used in the second line, (23) is used in the third line, (27) is used in the forth line, and (28) is used in the last line. Therefore, the formula (19) can be written as
+$=\frac{\cosh d_{jk}\cdot(c_{jk}\ast c_{ijk})}{\sinh d_{ji}\cdot(c_{ij}\ast c_{ijk})}+\frac{(c_{ij}\ast c_{ijk})\cosh d_{ji}}{\|v_{j}\otimes c_{ijk}\|}\cdot\cosh\theta_{ik}\cdot\frac{\cosh d_{ji}}{\sinh d_{ji}}$
+$=\frac{\cosh\theta_{ik}}{\sinh d_{ji}\cdot(c_{ij}\ast c_{ijk})\cdot\|v_{j}\otimes c_{ijk}\|}\big{(}-\|v_{j}\otimes c_{ijk}\|^{2}+(c_{ij}\ast c_{ijk})^{2}\cosh^{2}d_{ji}\big{)}$
+$=\frac{\cosh\theta_{ik}\cdot\|c_{ij}\otimes c_{ijk}\|^{2}}{\sinh d_{ji}\cdot(c_{ij}\ast c_{ijk})\cdot\|v_{j}\otimes c_{ijk}\|}$
+$=\frac{\cosh\theta_{ik}\sinh^{2}\theta_{ki}\|v_{j}\otimes c_{ijk}\|}{\sinh d_{ji}(c_{ij}\ast c_{ijk})},$
+where (21) and (23) are used in the second line, (22) is used in the third line, (26) is used in the fourth line, and (27) is used in the last line. Substituting $L_{1}$ and $L_{2}$ into (18) yields
 $\frac{\partial\theta_{i}}{\partial f_{j}}$ $=\frac{1}{\sinh l_{k}\sinh\theta_{j}}\cdot(L_{1}+L_{2})$
-$=$ $\frac{1}{\sinh l_{k}\sinh\theta_{j}}\cdot\frac{\sinh\theta_{ki}\cdot||v_{j}\otimes c_{ijk}||}{(c_{ij}\ast c_{ijk})\sinh d_{ji}}\cdot(\cosh\theta_{ik}\sinh\theta_{ki}+\sinh\theta_{ik}\cosh\theta_{ki})$
-$=$ $\frac{1}{\sinh l_{k}\sinh\theta_{j}}\cdot\frac{\sinh\theta_{ki}\cdot||v_{j}\otimes c_{ijk}||}{(c_{ij}\ast c_{ijk})\sinh d_{ji}}\cdot\sinh\theta_{j}$
-$=$ $\frac{1}{\sinh l_{k}\sinh d_{ji}}\cdot\frac{\sinh\theta_{ki}\cdot||v_{j}\otimes c_{ijk}||}{c_{ij}\ast c_{ijk}}.$
-(i): Suppose $c_{ijk}$ is time-like, then $||c_{ijk}||^{2}=-1$ and $v_{j}*c_{ijk}=-\sinh q_{j}$.
-By (25), we have $||v_{j}\otimes c_{ijk}||=\cosh q_{j}$.
-Then $\sinh\theta_{ki}\cdot||v_{j}\otimes c_{ijk}||=\sinh\theta_{ki}\cdot\cosh q_{j}=\sinh h_{k}$ by Lemma 2.10.
-Since $c_{ij}\ast c_{ijk}=-\cosh h_{k}$, then
+$=\frac{1}{\sinh l_{k}\sinh\theta_{j}}\cdot\frac{\sinh\theta_{ki}\cdot\|v_{j}\otimes c_{ijk}\|}{(c_{ij}\ast c_{ijk})\sinh d_{ji}}\cdot(\cosh\theta_{ik}\sinh\theta_{ki}+\sinh\theta_{ik}\cosh\theta_{ki})$
+$=\frac{1}{\sinh l_{k}\sinh\theta_{j}}\cdot\frac{\sinh\theta_{ki}\cdot\|v_{j}\otimes c_{ijk}\|}{(c_{ij}\ast c_{ijk})\sinh d_{ji}}\cdot\sinh\theta_{j}$
+$=\frac{1}{\sinh l_{k}\sinh d_{ji}}\cdot\frac{\sinh\theta_{ki}\cdot\|v_{j}\otimes c_{ijk}\|}{c_{ij}\ast c_{ijk}}.$
+(i): Assume that $c_{ijk}$ is time-like.
+Then $\|c_{ijk}\|^{2}=-1$ and $v_{j}*c_{ijk}=-\sinh q_{j}$.
+By (24), we have $\|v_{j}\otimes c_{ijk}\|=\cosh q_{j}$.
+Furthermore, by Lemma 2.9, we obtain $\sinh\theta_{ki}\cdot\|v_{j}\otimes c_{ijk}\|=\sinh\theta_{ki}\cdot\cosh q_{j}=\sinh h_{k}$.
+Since $c_{ij}\ast c_{ijk}=-\cosh h_{k}$, it follows that
 $\frac{\partial\theta_{i}}{\partial f_{j}}=\frac{1}{\sinh d_{ji}\sinh l_{k}}\cdot\frac{\sinh h_{k}}{-\cosh h_{k}}=\frac{-1}{\sinh d_{ji}}\frac{\tanh h_{k}}{\sinh l_{k}}.$
-(ii): Suppose $c_{ijk}$ is space-like, then $||c_{ijk}||^{2}=1$ and $v_{j}*c_{ijk}=\pm\cosh q_{j}$, thus $||v_{j}\otimes c_{ijk}||=\sinh q_{j}$ if $q_{j}>0$ and $||v_{j}\otimes c_{ijk}||=-\sinh q_{j}$ if $q_{j}<0$ by (25).
-If $c_{ijk}$ is in one of the domains $D_{\mathrm{I}},D_{\mathrm{II}},D_{\mathrm{V}},D_{\mathrm{VI}}$, then $q_{j}>0$ and $h_{k}>0$ by Table 1.
-Then $\sinh\theta_{ki}\cdot||v_{j}\otimes c_{ijk}||=\sinh\theta_{ki}\cdot\sinh q_{j}=\cosh h_{k}$ by Lemma 2.13.
-And $c_{ij}\ast c_{ijk}=-\sinh h_{k}$.
-If $c_{ijk}$ is in the domain $D_{\mathrm{III}}$, then $q_{j}<0$ and $h_{k}>0$ by Table 1.
-Then $\sinh\theta_{ki}\cdot||v_{j}\otimes c_{ijk}||=\sinh\theta_{ki}\cdot(-\sinh q_{j})=\cosh h_{k}$ by Lemma 2.13.
-And $c_{ij}\ast c_{ijk}=-\sinh h_{k}$.
-If $c_{ijk}$ is in the domain $D_{\mathrm{IV}}$, then $q_{j}>0$ and $h_{k}<0$ by Table 1.
-Then $\sinh\theta_{ki}\cdot||v_{j}\otimes c_{ijk}||=\sinh\theta_{ki}\cdot\sinh q_{j}=-\cosh h_{k}$ by Lemma 2.13.
+(ii): Assume that $c_{ijk}$ is space-like.
+Then $\|c_{ijk}\|^{2}=1$ and $v_{j}*c_{ijk}=\pm\cosh q_{j}$.
+It follows from (24) that $\|v_{j}\otimes c_{ijk}\|=\sinh q_{j}$ if $q_{j}>0$, and $\|v_{j}\otimes c_{ijk}\|=-\sinh q_{j}$ if $q_{j}<0$.
+We proceed to analyze different cases based on the domain to which $c_{ijk}$ belongs.
+If $c_{ijk}\in D_{\mathrm{I}}\cup D_{\mathrm{II}}\cup D_{\mathrm{V}}\cup D_{\mathrm{VI}}$, then Table 1 implies $q_{j}>0$ and $h_{k}>0$.
+By Lemma 2.12, we have $\sinh\theta_{ki}\cdot\|v_{j}\otimes c_{ijk}\|=\sinh\theta_{ki}\cdot\sinh q_{j}=\cosh h_{k}$.
+Additionally, it is known that $c_{ij}\ast c_{ijk}=-\sinh h_{k}$.
+If $c_{ijk}\in D_{\mathrm{III}}$, then Table 1 implies $q_{j}<0$ and $h_{k}>0$.
+By Lemma 2.12, we have $\sinh\theta_{ki}\cdot\|v_{j}\otimes c_{ijk}\|=\sinh\theta_{ki}\cdot(-\sinh q_{j})=\cosh h_{k}$ by Lemma 2.12.
+Additionally, it is known that $c_{ij}\ast c_{ijk}=-\sinh h_{k}$.
+If $c_{ijk}\in D_{\mathrm{IV}}$, then Table 1 implies $q_{j}>0$ and $h_{k}<0$.
+By Lemma 2.12, we have $\sinh\theta_{ki}\cdot\|v_{j}\otimes c_{ijk}\|=\sinh\theta_{ki}\cdot\sinh q_{j}=-\cosh h_{k}$.
 Note that $c_{ij}\ast c_{ijk}=-\sinh(d_{\mathbb{H}}(c_{ij},(c_{ijk})^{\bot}))=-\sinh(-h_{k})=\sinh h_{k}$.
-If $c_{ijk}$ is in one of the domains $D_{\mathrm{i}},D_{\mathrm{ii}},D_{\mathrm{vi}}$, then $q_{j}>0$ and $h_{k}>0$ by Table 1.
-Then $\sinh\theta_{ki}\cdot||v_{j}\otimes c_{ijk}||=\sinh\theta_{ki}\cdot\sinh q_{j}=\cosh h_{k}$ by Lemma 2.15.
-And $c_{ij}\ast c_{ijk}=-\sinh h_{k}$.
-If $c_{ijk}$ is in the domain $D_{\mathrm{iii}}$, then $q_{j}<0$ and $h_{k}>0$ by Table 1.
-Then $\sinh\theta_{ki}\cdot||v_{j}\otimes c_{ijk}||=\sinh\theta_{ki}\cdot(-\sinh q_{j})=\cosh h_{k}$ Lemma 2.15.
-And $c_{ij}\ast c_{ijk}=-\sinh h_{k}$.
-If $c_{ijk}$ is in the domain $D_{\mathrm{v}}$, then $q_{j}>0$ and $h_{k}<0$ by Table 1.
-Then $\sinh\theta_{ki}\cdot||v_{j}\otimes c_{ijk}||=\sinh\theta_{ki}\cdot\sinh q_{j}=-\cosh h_{k}$ by Lemma 2.15.
-And $c_{ij}\ast c_{ijk}=\sinh h_{k}$.
-If $c_{ijk}$ is in the domain $D_{\mathrm{iv}}$, then $q_{j}<0$ and $h_{k}<0$ by Table 1.
-Then $\sinh\theta_{ki}\cdot||v_{j}\otimes c_{ijk}||=\sinh\theta_{ki}\cdot(-\sinh q_{j})=-\cosh h_{k}$ by Lemma 2.15.
-And $c_{ij}\ast c_{ijk}=\sinh h_{k}$.
-In summary, if $c_{ijk}$ is space-like, then
+If $c_{ijk}\in D_{\mathrm{i}}\cup D_{\mathrm{ii}}\cup D_{\mathrm{vi}}$, then Table 1 implies $q_{j}>0$ and $h_{k}>0$.
+By Lemma 2.14, we have $\sinh\theta_{ki}\cdot\|v_{j}\otimes c_{ijk}\|=\sinh\theta_{ki}\cdot\sinh q_{j}=\cosh h_{k}$.
+Additionally, $c_{ij}\ast c_{ijk}=-\sinh h_{k}$.
+If $c_{ijk}\in D_{\mathrm{iii}}$, then Table 1 implies $q_{j}<0$ and $h_{k}>0$.
+By Lemma 2.14, we have $\sinh\theta_{ki}\cdot\|v_{j}\otimes c_{ijk}\|=\sinh\theta_{ki}\cdot(-\sinh q_{j})=\cosh h_{k}$.
+Additionally, $c_{ij}\ast c_{ijk}=-\sinh h_{k}$.
+If $c_{ijk}\in D_{\mathrm{v}}$, then Table 1 implies $q_{j}>0$ and $h_{k}<0$.
+By Lemma 2.14, we have $\sinh\theta_{ki}\cdot\|v_{j}\otimes c_{ijk}\|=\sinh\theta_{ki}\cdot\sinh q_{j}=-\cosh h_{k}$.
+Additionally, $c_{ij}\ast c_{ijk}=\sinh h_{k}$.
+If $c_{ijk}\in D_{\mathrm{iv}}$, then Table 1 implies $q_{j}<0$ and $h_{k}<0$.
+By Lemma 2.14, we have $\sinh\theta_{ki}\cdot\|v_{j}\otimes c_{ijk}\|=\sinh\theta_{ki}\cdot(-\sinh q_{j})=-\cosh h_{k}$.
+Additionally, $c_{ij}\ast c_{ijk}=\sinh h_{k}$.
+In all the above cases for space-like $c_{ijk}$, we can conclude that
 $\frac{\partial\theta_{i}}{\partial f_{j}}=\frac{-1}{\sinh d_{ji}}\frac{\coth h_{k}}{\sinh l_{k}}.$
-(iii): The case that $c_{ijk}$ is light-like follows from continuity by moving the center to the light cone.
-In fact, if $c_{ijk}$ is light-like, then $||c_{ijk}||^{2}=0$.
+(iii): The case where $c_{ijk}$ is light-like can be derived by continuity, specifically by considering the limit as the face center approaches the light cone.
+In detail, for a light-like $c_{ijk}$, we have $\|c_{ijk}\|^{2}=0$.
 Combining (27) and (25) gives
-$\frac{\sinh^{2}\theta_{ki}\cdot||v_{j}\otimes c_{ijk}||^{2}}{(c_{ij}\ast c_{ijk})^{2}}=\frac{||c_{ij}\otimes c_{ijk}||^{2}}{(c_{ij}\ast c_{ijk})^{2}}=1.$
-Since $||c_{ij}\otimes c_{ijk}||=-(c_{ij}\ast c_{ijk})$, then
+$\frac{\sinh^{2}\theta_{ki}\cdot\|v_{j}\otimes c_{ijk}\|^{2}}{(c_{ij}\ast c_{ijk})^{2}}=\frac{\|c_{ij}\otimes c_{ijk}\|^{2}}{(c_{ij}\ast c_{ijk})^{2}}=1.$
+Furthermore, since $\|c_{ij}\otimes c_{ijk}\|=-(c_{ij}\ast c_{ijk})$, it follows that
 $\frac{\partial\theta_{i}}{\partial f_{j}}=\frac{-1}{\sinh d_{ji}\sinh l_{k}}.$
@@
 Remark 2.18.
-Actually, the formulas in the proof of Lemma 2.18 are the sine and cosine laws of generalized right-angled hyperbolic triangles.
-In order to avoid discussing one by one, we use Lorentzian inner product and Lorentzian cross product to unify calculations.
+The formulas derived in the proof of Lemma 2.17 essentially correspond to the sine and cosine laws for generalized right-angled hyperbolic triangles.
+To avoid the need for cumbersome case-by-case analyses, we adopt the Lorentzian inner product and Lorentzian cross product as a unifying framework for the calculations.
 Lemma 2.19.
@@
 Proof.
-Under the same notations as in the proof of Lemma 2.18, a routine calculation gives
+Under the same notations as used in the proof of Lemma 2.17, analogous calculations yield
 $\frac{\partial\theta_{i}}{\partial f_{i}}$ $=\frac{\partial\theta_{i}}{\partial l_{i}}\frac{\partial l_{i}}{\partial f_{i}}+\frac{\partial\theta_{i}}{\partial l_{j}}\frac{\partial l_{j}}{\partial f_{i}}+\frac{\partial\theta_{i}}{\partial l_{k}}\frac{\partial l_{k}}{\partial f_{i}}$
@@
 $=\frac{-\sinh l_{i}}{A}\cdot\left(\cosh\theta_{k}\cdot\coth d_{ik}+\cosh\theta_{j}\cdot\coth d_{ij}\right).$
-Similar to (17), we have
+Similar to (16), we derive
 $\frac{\partial\theta_{j}}{\partial f_{i}}=\frac{\sinh l_{j}}{A}(\coth d_{ik}-\cosh\theta_{i}\coth d_{ij}),$
@@
 Remark 2.20.
-The formula (29) was first obtained by Glickenstein-Thomas [3] for generalized discrete conformal structures on closed surfaces, which has lots of applications.
-See [9, 10, 12, 13, 14, 15] and others for example.
-In [11], Xu showed that the formula (29) holds for a special class of discrete conformal structures on surfaces with boundary and conjectured this formula holds for generalized discrete conformal structures on surfaces with boundary.
-Lemma 2.20 provides an affirmative answer to Xu’s conjecture.
-Since the variational formula of generalized angle (15) is not symmetric in $i$ and $j$, we need to obtain a symmetric variational formula of generalized angle by reparameterization.
+Formula (28) was first established by Glickenstein-Thomas [3] in the context of discrete conformal structures on closed surfaces, and it has since been widely applied.
+For instance, relevant discussions can be found in [9, 12, 11, 13] and related works.
+In [10], Xu showed that formula (28) remains valid for a specific class of discrete conformal structures on surfaces with boundary, and conjectured that this formula holds for generic discrete conformal structures on surfaces with boundary.
+Lemma 2.19 provides an affirmative resolution to Xu’s conjecture.
+Given that the variational formula for the generalized angle (14) lacks symmetry in $i$ and $j$, we aim to derive a symmetric variational formula for the generalized angle via reparameterization.
 Lemma 2.21.
-There is a change of variables $u_{i}=u_{i}(f_{i})$ such that
+There exists a change of variables $u_{i}=u_{i}(f_{i})$ such that
 $\frac{\partial\theta^{jk}_{i}}{\partial u_{j}}=\frac{\partial\theta^{ik}_{j}}{\partial u_{i}}.$
-The function $u$ is also called a discrete conformal factor.
+The function $u$ is also referred to as a discrete conformal factor.
 Proof.
-Take new coordinates $u_{i}=u_{i}(f_{i})$ such that $\partial\theta^{jk}_{i}/\partial u_{j}=\partial\theta^{ik}_{j}/\partial u_{i}$. By (15), we have
+Suppose there exists a change of variables $u_{i}=u_{i}(f_{i})$ such that $\partial\theta^{jk}_{i}/\partial u_{j}=\partial\theta^{ik}_{j}/\partial u_{i}$. It follows from (14) that
 $\frac{\partial\theta^{jk}_{i}}{\partial u_{j}}=\frac{\partial\theta^{jk}_{i}}{\partial f_{j}}\frac{\partial f_{j}}{\partial u_{j}}=\frac{-\partial f_{j}/\partial u_{j}}{\sinh d_{ji}}\cdot\frac{\tanh^{\beta}h_{ij}}{\sinh l_{ij}},$
@@
 $\frac{\partial f_{i}/\partial u_{i}}{\partial f_{j}/\partial u_{j}}=\frac{\sinh d_{ij}}{\sinh d_{ji}}.$
-Recall the following formulas from the proof of [16, Theorem 1.6]
+From the proof of [14, Theorem 1.5], we recall the identities
 $\frac{\sinh^{2}d_{ij}}{\sinh^{2}d_{ji}}=e^{C_{ij}}e^{2f_{i}-2f_{j}}\quad\text{or}\quad\frac{\sinh^{2}d_{ij}}{\sinh^{2}d_{ji}}=\frac{1+\alpha_{i}e^{2f_{i}}}{1+\alpha_{j}e^{2f_{j}}}.$
 (i):
-By Remark 1.7, we can set $e^{C_{ij}}=1$. If $\frac{\sinh d_{ij}}{\sinh d_{ji}}=e^{f_{i}-f_{j}}$, then
+By Remark 1.6, we set $e^{C_{ij}}=1$. If $\frac{\sinh d_{ij}}{\sinh d_{ji}}=e^{f_{i}-f_{j}}$, then
 $\frac{\partial f_{i}}{\partial u_{i}}=e^{f_{i}}\quad\text{and}\ \quad\frac{\partial f_{j}}{\partial u_{j}}=e^{f_{j}}.$
 Suppose $\frac{\sinh d_{ij}}{\sinh d_{ji}}=-e^{f_{i}-f_{j}}$.
-If $d_{ij}<0$, then $d_{ik}<0$ by Lemma 2.17, and thus $d_{ji}>0,\ d_{ki}>0$.
-We need $d_{jk}>0,\ d_{kj}>0$ to make the condition (1) hold.
-Hence $\frac{\sinh d_{ik}}{\sinh d_{ki}}=-e^{f_{i}-f_{k}}$ and $\frac{\sinh d_{jk}}{\sinh d_{ki}}=e^{f_{j}-f_{k}}$.
-Then
+If $d_{ij}<0$, then $d_{ik}<0$ by Lemma 2.16, which implies $d_{ji}>0$ and $d_{ki}>0$.
+For the condition (2) to hold, we require $d_{jk}>0$ and $d_{kj}>0$.
+It follows that $\frac{\sinh d_{ik}}{\sinh d_{ki}}=-e^{f_{i}-f_{k}}$ and $\frac{\sinh d_{jk}}{\sinh d_{ki}}=e^{f_{j}-f_{k}}$.
+Hence
 (29) $\frac{\partial f_{i}}{\partial u_{i}}=-e^{f_{i}},\ \quad\frac{\partial f_{j}}{\partial u_{j}}=e^{f_{j}}\quad\text{and}\ \quad\frac{\partial f_{k}}{\partial u_{k}}=e^{f_{k}}.$
@@
 (ii):
-If $\frac{\sinh d_{ij}}{\sinh d_{ji}}=\sqrt{\frac{1+\alpha_{i}e^{2f_{i}}}{1+\alpha_{j}e^{2f_{j}}}}$, then
+If $\frac{\sinh d_{ij}}{\sinh d_{ji}}=\sqrt{\frac{1+\alpha_{i}e^{2f_{i}}}{1+\alpha_{j}e^{2f_{j}}}}$ with $01+\alpha_{i}e^{2f_{i}}>$ 0 and $01+\alpha_{j}e^{2f_{j}}>$ 0, then
 $\frac{\partial f_{i}}{\partial u_{i}}=\sqrt{1+\alpha_{i}e^{2f_{i}}}\quad\text{and}\ \quad\frac{\partial f_{j}}{\partial u_{j}}=\sqrt{1+\alpha_{j}e^{2f_{j}}}.$
-By the arguments similar to that in the case $\mathrm{(i)}$, if $\frac{\sinh d_{ij}}{\sinh d_{ji}}=-\sqrt{\frac{1+\alpha_{i}e^{2f_{i}}}{1+\alpha_{j}e^{2f_{j}}}}$, then when $d_{ij}<0$,
-$\frac{\partial f_{i}}{\partial u_{i}}=-\sqrt{1+\alpha_{i}e^{2f_{i}}}\quad\text{and}\ \quad\frac{\partial f_{j}}{\partial u_{j}}=\sqrt{1+\alpha_{j}e^{2f_{j}}};$
+If $\frac{\sinh d_{ij}}{\sinh d_{ji}}=\sqrt{\frac{1+\alpha_{i}e^{2f_{i}}}{1+\alpha_{j}e^{2f_{j}}}}$ with $01+\alpha_{i}e^{2f_{i}}<$ 0 and $01+\alpha_{j}e^{2f_{j}}<$ 0, then $\alpha\equiv-1$ and
+$\frac{\partial f_{i}}{\partial u_{i}}=\sqrt{e^{2f_{i}}-1}\quad\text{and}\ \quad\frac{\partial f_{j}}{\partial u_{j}}=\sqrt{e^{2f_{j}}-1}.$
+(iii):
+If $\frac{\sinh d_{ij}}{\sinh d_{ji}}=-\sqrt{\frac{1+\alpha_{i}e^{2f_{i}}}{1+\alpha_{j}e^{2f_{j}}}}$ with $01+\alpha_{i}e^{2f_{i}}>$ 0 and $01+\alpha_{j}e^{2f_{j}}>$ 0, then by arguments analogous to those in the case (i), when $d_{ij}<0$,
+$\frac{\partial f_{i}}{\partial u_{i}}=-\sqrt{1+\alpha_{i}e^{2f_{i}}},\quad\frac{\partial f_{j}}{\partial u_{j}}=\sqrt{1+\alpha_{j}e^{2f_{j}}}\quad\text{and}\ \quad\frac{\partial f_{k}}{\partial u_{k}}=\sqrt{1+\alpha_{k}e^{2f_{k}}};$
 when $d_{ji}<0$,
-$\frac{\partial f_{i}}{\partial u_{i}}=\sqrt{1+\alpha_{i}e^{2f_{i}}}\quad\text{and}\ \quad\frac{\partial f_{j}}{\partial u_{j}}=-\sqrt{1+\alpha_{j}e^{2f_{j}}}.$
+$\frac{\partial f_{i}}{\partial u_{i}}=\sqrt{1+\alpha_{i}e^{2f_{i}}},\quad\frac{\partial f_{j}}{\partial u_{j}}=-\sqrt{1+\alpha_{j}e^{2f_{j}}}\quad\text{and}\ \quad\frac{\partial f_{k}}{\partial u_{k}}=\sqrt{1+\alpha_{k}e^{2f_{k}}}.$
+If $\frac{\sinh d_{ij}}{\sinh d_{ji}}=-\sqrt{\frac{1+\alpha_{i}e^{2f_{i}}}{1+\alpha_{j}e^{2f_{j}}}}$ with $01+\alpha_{i}e^{2f_{i}}<$ 0 and $01+\alpha_{j}e^{2f_{j}}<$ 0, then when $d_{ij}<0$,
+(30) $\frac{\partial f_{i}}{\partial u_{i}}=-\sqrt{e^{2f_{i}}-1},\quad\frac{\partial f_{j}}{\partial u_{j}}=\sqrt{e^{2f_{j}}-1}\quad\text{and}\ \quad\frac{\partial f_{k}}{\partial u_{k}}=\sqrt{e^{2f_{k}}-1};$
+when $d_{ji}<0$,
+$\frac{\partial f_{i}}{\partial u_{i}}=\sqrt{e^{2f_{i}}-1},\quad\frac{\partial f_{j}}{\partial u_{j}}=-\sqrt{e^{2f_{j}}-1}\quad\text{and}\ \quad\frac{\partial f_{k}}{\partial u_{k}}=\sqrt{e^{2f_{k}}-1}.$
 Q.E.D.
 Remark 2.22.
-In fact, the function $u_{i}=u_{i}(f_{i})$ can be computed explicitly.
-If $\frac{\partial f_{i}}{\partial u_{i}}=e^{f_{i}}$, then $u_{i}=-e^{-f_{i}}$.
-If $\frac{\partial f_{i}}{\partial u_{i}}=-e^{f_{i}}$, then $u_{i}=e^{-f_{i}}$.
+In fact, the function $u_{i}=u_{i}(f_{i})$ admits an explicit expression.
+(i):
+If $\frac{\partial f_{i}}{\partial u_{i}}=e^{f_{i}}$, then $u_{i}=-e^{-f_{i}}$. If $\frac{\partial f_{i}}{\partial u_{i}}=-e^{f_{i}}$, then $u_{i}=e^{-f_{i}}$.
+(ii):
 Suppose $\frac{\partial f_{i}}{\partial u_{i}}=\sqrt{1+\alpha_{i}e^{2f_{i}}}$, if $\alpha_{i}=0$, then $u_{i}=f_{i}\in\mathbb{R}$; if $\alpha_{i}\neq 0$, then
 $u_{i}=\frac{1}{2}\ln\left|\frac{\sqrt{1+\alpha_{i}e^{2f_{i}}}-1}{\sqrt{1+\alpha_{i}e^{2f_{i}}}+1}\right|.$
-Specially, if $\alpha_{i}=-1$, then
+In particular, if $\alpha_{i}=-1$, then
 (31) $-\tanh u_{i}=\sqrt{1-e^{2f_{i}}},$
-which implies $e^{f_{i}}=\frac{1}{\cosh u_{i}}$. Note that $\sqrt{1-e^{2f_{i}}}\neq 0$ and thus $f_{i}\neq 0$. Then $f_{i},u_{i}\in\mathbb{R}_{<0}$. If $\alpha_{i}=1$, then
+which implies $e^{f_{i}}=\frac{1}{\cosh u_{i}}$.
+Note that $\sqrt{1-e^{2f_{i}}}\neq 0$ and thus $f_{i}\neq 0$.
+Consequently, $f_{i}\in\mathbb{R}_{<0}$ and $u_{i}\in\mathbb{R}_{<0}$.
+If $\alpha_{i}=1$, then
 (32) $-\coth u_{i}=\sqrt{1+e^{2f_{i}}},$
-which implies $e^{f_{i}}=-\frac{1}{\sinh u_{i}}$ and $f_{i}\in\mathbb{R}$, $u_{i}\in\mathbb{R}_{<0}$.
-Suppose $\frac{\partial f_{i}}{\partial u_{i}}=-\sqrt{1+\alpha_{i}e^{2f_{i}}}$, then if $\alpha_{i}=0$, then $u_{i}=-f_{i}\in\mathbb{R}$; if $\alpha_{i}\neq 0$, then
+which implies $e^{f_{i}}=-\frac{1}{\sinh u_{i}}$ with $f_{i}\in\mathbb{R}$ and $u_{i}\in\mathbb{R}_{<0}$.
+(iii):
+Suppose $\frac{\partial f_{i}}{\partial u_{i}}=-\sqrt{1+\alpha_{i}e^{2f_{i}}}$, if $\alpha_{i}=0$, then $u_{i}=-f_{i}\in\mathbb{R}$; if $\alpha_{i}\neq 0$, then
 $u_{i}=-\frac{1}{2}\ln\left|\frac{\sqrt{1+\alpha_{i}e^{2f_{i}}}-1}{\sqrt{1+\alpha_{i}e^{2f_{i}}}+1}\right|.$
-Specially, if $\alpha_{i}=-1$, then
+In particular, if $\alpha_{i}=-1$, then
 $\tanh u_{i}=\sqrt{1-e^{2f_{i}}},$
-which implies $e^{f_{i}}=\frac{1}{\cosh u_{i}}$. Similarly, the case that $f_{i}=0$ is removed. Then $f_{i}\in\mathbb{R}_{<0},\ u_{i}\in\mathbb{R}_{>0}$. If $\alpha_{i}=1$, then
+which implies $e^{f_{i}}=\frac{1}{\cosh u_{i}}$. Similarly, the case that $f_{i}=0$ is excluded, so $f_{i}\in\mathbb{R}_{<0}$ and $u_{i}\in\mathbb{R}_{>0}$. If $\alpha_{i}=1$, then
 $\coth u_{i}=\sqrt{1+e^{2f_{i}}},$
-which implies $e^{f_{i}}=\frac{1}{\sinh u_{i}}$ and $f_{i}\in\mathbb{R},\ u_{i}\in\mathbb{R}_{>0}$.
+which implies $e^{f_{i}}=\frac{1}{\sinh u_{i}}$ with $f_{i}\in\mathbb{R}$ and $u_{i}\in\mathbb{R}_{>0}$.
+(iv):
+If $\frac{\partial f_{i}}{\partial u_{i}}=\sqrt{e^{2f_{i}}-1}$, then
+$\tan u_{i}=\sqrt{e^{2f_{i}}-1},\ \text{for}\ u_{i}\in(0,\frac{\pi}{2}),$
+$-\frac{1}{\tan u_{i}}=\sqrt{e^{2f_{i}}-1},\ \text{for}\ u_{i}\in(-\frac{\pi}{2},0).$
+This implies that $e^{f_{i}}=\frac{1}{\cos u_{i}}$ for $u_{i}\in(0,\frac{\pi}{2})$, and $e^{f_{i}}=-\frac{1}{\sin u_{i}}$ for $u_{i}\in(-\frac{\pi}{2},0)$.
+If $\frac{\partial f_{i}}{\partial u_{i}}=-\sqrt{e^{2f_{i}}-1}$, then
+$-\tan u_{i}=\sqrt{e^{2f_{i}}-1},\ \text{for}\ u_{i}\in(-\frac{\pi}{2},0),$
+$\frac{1}{\tan u_{i}}=\sqrt{e^{2f_{i}}-1},\ \text{for}\ u_{i}\in(0,\frac{\pi}{2}).$
+This implies that $e^{f_{i}}=\frac{1}{\cos u_{i}}$ for $u_{i}\in(-\frac{\pi}{2},0)$, and $e^{f_{i}}=\frac{1}{\sin u_{i}}$ for $u_{i}\in(0,\frac{\pi}{2})$.
 3. Rigidity and existence of the discrete conformal structure (A1)
 3.1. Admissible space of the discrete conformal structure (A1)
-Suppose $(\Sigma,\mathcal{T},\alpha,\eta)$ is a weighted triangulated surface with boundary, and the weights $\alpha:B\rightarrow\{-1,0,1\}$ and $\eta\in\mathbb{R}^{E}_{>0}$.
-The admissible space of the discrete conformal factors $f$ for a right-angled hyperbolic hexagon $\{ijk\}\in F$ with edge lengths $l_{ij},l_{jk},l_{ki}$ given by (5) is defined to be
-(33) $\Omega_{ijk}(\eta)=\{(f_{i},f_{j},f_{k})\in\mathbb{R}^{3}|$ $\cosh l_{rs}=-\sqrt{(1+\alpha_{r}e^{2f_{r}})(1+\alpha_{s}e^{2f_{s}})}$
+Suppose $(\Sigma,\mathcal{T},\alpha,\eta)$ is a weighted triangulated surface with boundary, where the weights are given by $\alpha:B\rightarrow\{-1,0,1\}$ and $\eta\in\mathbb{R}^{E}_{>0}$.
+For a right-angled hyperbolic hexagon $\{ijk\}\in F$ with edge lengths $l_{ij},l_{jk},l_{ki}$ given by (5), the admissible space of the discrete conformal factors $f$ is defined as
+(33) $\Omega_{ijk}(\eta)=\{(f_{i},f_{j},f_{k})\in\mathbb{R}^{3}\mid$ $\cosh l_{rs}=-\sqrt{(1+\alpha_{r}e^{2f_{r}})(1+\alpha_{s}e^{2f_{s}})}$
 $\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +\eta_{rs}e^{f_{r}+f_{s}}>1,\ \forall\{r,s\}\subseteq\{i,j,k\}\}.$
-By a change of variables $u_{i}=u_{i}(f_{i})$ in Lemma 2.22, the admissible space $\Omega_{ijk}(\eta)$ of $f$ is transferred to the following admissible space $\mathcal{U}_{ijk}(\eta)$ of the discrete conformal factors $u$ for a right-angled hyperbolic hexagon $\{ijk\}$
-$\mathcal{U}_{ijk}(\eta)=u(\Omega_{ijk}(\eta))=\{(u_{i},u_{j},u_{k})\in\mathbb{R}^{N_{1}}\times\mathbb{R}_{<0}^{3-N_{1}}|(f_{i},f_{j},f_{k})\in\Omega_{ijk}(\eta)\},$
+By a change of variables $u_{i}=u_{i}(f_{i})$ in Lemma 2.21, the admissible space $\Omega_{ijk}(\eta)$ is transformed into the following admissible space $\mathcal{U}_{ijk}(\eta)$ of the discrete conformal factors $u$ for a right-angled hyperbolic hexagon $\{ijk\}$:
+$\mathcal{U}_{ijk}(\eta)=u(\Omega_{ijk}(\eta))=\{(u_{i},u_{j},u_{k})\in\mathbb{R}^{N_{1}}\times\mathbb{R}_{<0}^{3-N_{1}}\mid(f_{i},f_{j},f_{k})\in\Omega_{ijk}(\eta)\},$
 where $N_{1}$ is the number of the boundary components $i\in B$ with $\alpha_{i}=0$.
-Since $\frac{\sinh d_{rs}}{\sinh d_{sr}}=\sqrt{\frac{1+\alpha_{r}e^{2f_{r}}}{1+\alpha_{s}e^{2f_{s}}}}>0$ in Theorem 1.6, then by the proof of Lemma 2.22, we have
+Since $\frac{\sinh d_{rs}}{\sinh d_{sr}}=\sqrt{\frac{1+\alpha_{r}e^{2f_{r}}}{1+\alpha_{s}e^{2f_{s}}}}>0$ with $01+\alpha_{r}e^{2f_{r}}>$ 0 and $01+\alpha_{s}e^{2f_{s}}>$ 0 in Theorem 1.5, it follows from the proof of Lemma 2.21 that
 (34) $\frac{\partial f_{r}}{\partial u_{r}}=\sqrt{1+\alpha_{r}e^{2f_{r}}}\quad\text{and}\ \quad\frac{\partial f_{s}}{\partial u_{s}}=\sqrt{1+\alpha_{s}e^{2f_{s}}}.$
-According to Remark 2.23, we can describe the admissible space $\mathcal{U}_{ijk}(\eta)$ explicitly as follows:
+By Remark 2.22, the following results are obtained:
 (i):
@@
 If $\alpha\equiv-1$, then $\cosh l_{rs}=-\sqrt{(1-e^{2f_{r}})(1-e^{2f_{s}})}+\eta_{rs}e^{f_{r}+f_{s}}=-\tanh u_{r}\tanh u_{s}+\eta_{rs}\frac{1}{\cosh u_{r}}\frac{1}{\cosh u_{s}}>1$, which implies $\cosh\,(u_{r}+u_{s})<\eta_{rs}$.
-Since $u\in\mathbb{R}^{3}_{<0}$, then $u_{r}+u_{s}>-\mathrm{arccosh}\ \eta_{rs}$.
-It is worth noting that $\eta_{rs}>1$.
+Since $u\in\mathbb{R}^{3}_{<0}$, it follows that $u_{r}+u_{s}>-\mathrm{arccosh}\,\eta_{rs}$.
+Note that $\eta_{rs}>1$.
 (iii):
 If $\alpha\equiv 1$, then $\cosh l_{rs}=-\sqrt{(1+e^{2f_{r}})(1+e^{2f_{s}})}+\eta_{rs}e^{f_{r}+f_{s}}=-\coth u_{r}\coth u_{s}+\eta_{rs}\frac{1}{\sinh u_{r}}\frac{1}{\sinh u_{s}}>1$, which implies $\cosh\,(u_{r}+u_{s})<\eta_{rs}$.
-Since $u\in\mathbb{R}^{3}_{<0}$, then $u_{r}+u_{s}>-\mathrm{arccosh}\ \eta_{rs}$.
-It is worth noting that $\eta_{rs}>1$.
+Since $u\in\mathbb{R}^{3}_{<0}$, it follows that $u_{r}+u_{s}>-\mathrm{arccosh}\,\eta_{rs}$.
+Note that $\eta_{rs}>1$.
 (iv):
-If $\alpha_{r}=0,\alpha_{s}=-1$, then $\cosh l_{rs}=-\sqrt{1-e^{2f_{s}}}+\eta_{rs}e^{f_{r}+f_{s}}=\tanh u_{s}+\eta_{rs}e^{u_{r}}\frac{1}{\cosh u_{s}}>1$, which implies $u_{r}+u_{s}>\log\frac{1}{\eta_{rs}}$.
+If $\alpha_{r}=0$ and $\alpha_{s}=-1$, then $\cosh l_{rs}=-\sqrt{1-e^{2f_{s}}}+\eta_{rs}e^{f_{r}+f_{s}}=\tanh u_{s}+\eta_{rs}e^{u_{r}}\frac{1}{\cosh u_{s}}>1$, which implies $u_{r}+u_{s}>\log\frac{1}{\eta_{rs}}$.
 (v):
-If $\alpha_{r}=0,\alpha_{s}=1$, then $\cosh l_{rs}=-\sqrt{1+e^{2f_{s}}}+\eta_{rs}e^{f_{r}+f_{s}}=\coth u_{s}+\eta_{rs}e^{u_{r}}\frac{-1}{\sinh u_{s}}>1$.
-Since $u_{s}\in\mathbb{R}_{<0}$, then $u_{r}+u_{s}>\log\frac{1}{\eta_{rs}}$.
+If $\alpha_{r}=0$ and $\alpha_{s}=1$, then $\cosh l_{rs}=-\sqrt{1+e^{2f_{s}}}+\eta_{rs}e^{f_{r}+f_{s}}=\coth u_{s}+\eta_{rs}e^{u_{r}}\frac{-1}{\sinh u_{s}}>1$.
+Since $u_{s}\in\mathbb{R}_{<0}$, it follows that $u_{r}+u_{s}>\log\frac{1}{\eta_{rs}}$.
 (vi):
-If $\alpha_{r}=1,\alpha_{s}=-1$, then $\cosh l_{rs}=-\sqrt{(1+e^{2f_{r}})(1-e^{2f_{s}})}+\eta_{rs}e^{f_{r}+f_{s}}=-\coth u_{r}\tanh u_{s}+\eta_{rs}\frac{-1}{\sinh u_{r}}\frac{1}{\cosh u_{s}}>1$.
-Since $u_{r},u_{s}\in\mathbb{R}_{<0}$, then $\sinh(u_{r}+u_{s})>-\eta_{rs}$ and hence $u_{r}+u_{s}>\mathrm{arcsinh}\ (-\eta_{rs})$.
-In summary, the admissible space $\mathcal{U}_{ijk}(\eta)$ of $u$ can be rewritten as
-(35) $\mathcal{U}_{ijk}(\eta)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}^{N_{1}}\times\mathbb{R}_{<0}^{3-N_{1}}|u_{r}+u_{s}>C(\eta_{rs}),\ \forall\{r,s\}\subseteq\{i,j,k\}\},$
-where $C(\eta_{rs})$ is a constant that depends on $\eta_{rs}$.
-The above arguments can imply the following theorem.
+If $\alpha_{r}=1$ and $\alpha_{s}=-1$, then $\cosh l_{rs}=-\sqrt{(1+e^{2f_{r}})(1-e^{2f_{s}})}+\eta_{rs}e^{f_{r}+f_{s}}=-\coth u_{r}\tanh u_{s}+\eta_{rs}\frac{-1}{\sinh u_{r}}\frac{1}{\cosh u_{s}}>1$.
+Since $u_{r}\in\mathbb{R}_{<0}$ and $u_{s}\in\mathbb{R}_{<0}$, it follows that $\sinh\,(u_{r}+u_{s})>-\eta_{rs}$, and hence $u_{r}+u_{s}>\mathrm{arcsinh}\,(-\eta_{rs})$.
+In summary, the admissible space $\mathcal{U}_{ijk}(\eta)$ can be rewritten as
+(35) $\mathcal{U}_{ijk}(\eta)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}^{N_{1}}\times\mathbb{R}_{<0}^{3-N_{1}}\mid u_{r}+u_{s}>C(\eta_{rs}),\ \forall\{r,s\}\subseteq\{i,j,k\}\},$
+where $C(\eta_{rs})$ is a constant depending on $\eta_{rs}$.
 Theorem 3.1.
-Suppose $(\Sigma,\mathcal{T},\alpha,\eta)$ is a weighted triangulated surface with boundary, and the weights $\alpha:B\rightarrow\{-1,0,1\}$ and $\eta\in\mathbb{R}_{>0}^{E}$ satisfying $\eta_{ij}>\alpha_{i}\alpha_{j}$ if $\alpha_{i}=\alpha_{j}$ for two adjacent boundary components $i,j\in B$.
+Suppose $(\Sigma,\mathcal{T},\alpha,\eta)$ is a weighted triangulated surface with boundary, where the weighs are given by $\alpha:B\rightarrow\{-1,0,1\}$ and $\eta\in\mathbb{R}_{>0}^{E}$ satisfying $\eta_{ij}>\alpha_{i}\alpha_{j}$ for any two adjacent boundary components $i,j\in B$.
 Then the admissible space $\mathcal{U}_{ijk}(\eta)$ is a convex polytope.
-As a result, the admissible space
-$\mathcal{U}(\eta)=\bigcap_{\{ijk\}\in F}\mathcal{U}_{ijk}(\eta)$
-is a convex polytope.
-3.2. Negative definiteness of Jacobian
-For simplicity, we set $l_{r}=l_{st},\ \theta_{r}=\theta^{st}_{r}$ and $A=\sinh l_{r}\sinh l_{s}\sinh\theta_{t}$, where $\{r,s,t\}=\{i,j,k\}$. Define
+As a result, the admissible space $\mathcal{U}(\eta)=\bigcap_{\{ijk\}\in F}\mathcal{U}_{ijk}(\eta)$ is also a convex polytope.
+3.2. Negative definiteness of the Jacobian
+For a hyper-ideal hyperbolic triangle $\{ijk\}\in F$, we introduce some simplifying notations: $l_{r}=l_{st}$, $\theta_{r}=\theta^{st}_{r}$, and $A=\sinh l_{r}\sinh l_{s}\sinh\theta_{t}$, where $\{r,s,t\}=\{i,j,k\}$.
+Define
 (36) $Q_{1}=\left(\begin{array}[]{ccc}-1&\cosh\theta_{k}&\cosh\theta_{j}\\ \cosh\theta_{k}&-1&\cosh\theta_{i}\\ \cosh\theta_{j}&\cosh\theta_{i}&-1\\ \end{array}\right).$
-Then $\det Q_{1}=-1+2\cosh\theta_{i}\cosh\theta_{j}\cosh\theta_{k}+\cosh^{2}\theta_{i}+\cosh^{2}\theta_{j}+\cosh^{2}\theta_{k}>0$. By calculations similar to (16), we have
+Then $\det Q_{1}=-1+2\cosh\theta_{i}\cosh\theta_{j}\cosh\theta_{k}+\cosh^{2}\theta_{i}+\cosh^{2}\theta_{j}+\cosh^{2}\theta_{k}>0$. Following calculations analogous to those in (15), we obtain
 $\frac{\partial(\theta_{i},\theta_{j},\theta_{k})}{\partial(l_{i},l_{j},l_{k})}=$ $-\frac{1}{A}\left(\begin{array}[]{ccc}\sinh l_{i}&0&0\\ 0&\sinh l_{j}&0\\ 0&0&\sinh l_{k}\\ \end{array}\right)\times Q_{1}.$
@@
 (37) $\det\bigg{(}\frac{\partial(\theta_{i},\theta_{j},\theta_{k})}{\partial(l_{i},l_{j},l_{k})}\bigg{)}=-\frac{1}{A^{3}}\sinh l_{i}\sinh l_{j}\sinh l_{k}\cdot\det Q_{1}<0.$
-Combining (2) and (3), we have
+By combining (3) and (4), we derive
 (38) $\frac{\partial(l_{i},l_{j},l_{k})}{\partial(f_{i},f_{j},f_{k})}=\left(\begin{array}[]{ccc}0&\coth d_{jk}&\coth d_{kj}\\ \coth d_{ik}&0&\coth d_{ki}\\ \coth d_{ij}&\coth d_{ji}&0\\ \end{array}\right):=Q_{2}.$
@@
 (39) $\det Q_{2}=\coth d_{jk}\coth d_{ki}\coth d_{ij}+\coth d_{kj}\coth d_{ik}\coth d_{ji}.$
-For any $\alpha_{i}\in\{-1,0,1\}$, set $S_{i}=e^{f_{i}}$, $C_{i}=\sqrt{1+\alpha_{i}e^{2f_{i}}}$, then $S_{i},C_{i}>0$ and $C_{i}^{2}-\alpha_{i}S_{i}^{2}=1$. By (34), we have
+For any $\alpha_{i}\in\{-1,0,1\}$, set $S_{i}=e^{f_{i}}$ and $C_{i}=\sqrt{1+\alpha_{i}e^{2f_{i}}}$, so $S_{i}>0$, $C_{i}>0$, and $C_{i}^{2}-\alpha_{i}S_{i}^{2}=1$. It follows from (34) that
 (40) $\frac{\partial(f_{i},f_{j},f_{k})}{\partial(u_{i},u_{j},u_{k})}=\left(\begin{array}[]{ccc}C_{i}&0&0\\ 0&C_{j}&0\\ 0&0&C_{k}\\ \end{array}\right):=Q_{3}.$
 Theorem 3.2.
-Under the same assumptions as in Theorem 3.1.
-For a right-angled hyperbolic hexagon $\{ijk\}\in F$ on the admissible space $\mathcal{U}_{ijk}(\eta)$, the Jacobian $\Lambda_{ijk}=\frac{\partial(\theta^{jk}_{i},\theta^{ik}_{j},\theta^{ij}_{k})}{\partial(u_{i},u_{j},u_{k})}$ is non-degenerate, symmetric and negative definite.
+Under the same assumptions as those in Theorem 3.1, for a right-angled hyperbolic hexagon $\{ijk\}\in F$ on $\mathcal{U}_{ijk}(\eta)$, the Jacobian $\Lambda_{ijk}=\frac{\partial(\theta^{jk}_{i},\theta^{ik}_{j},\theta^{ij}_{k})}{\partial(u_{i},u_{j},u_{k})}$ is symmetric and negative definite.
 Proof.
-Since $d>0$ for the discrete conformal structure (A1) in Theorem 1.6, then $\det Q_{2}>0$ by (39). Hence
+Since $d>0$ for the discrete conformal structure (A1) in Theorem 1.5, then $\det Q_{2}>0$ by (39). Consequently,
 $\det\bigg{(}\frac{\partial(l_{i},l_{j},l_{k})}{\partial(u_{i},u_{j},u_{k})}\bigg{)}=\det Q_{2}\cdot\det Q_{3}>0.$
@@
 (41) $\det\Lambda_{ijk}=\det\bigg{(}\frac{\partial(\theta^{jk}_{i},\theta^{ik}_{j},\theta^{ij}_{k})}{\partial(l_{i},l_{j},l_{k})}\bigg{)}\cdot\det\bigg{(}\frac{\partial(l_{i},l_{j},l_{k})}{\partial(u_{i},u_{j},u_{k})}\bigg{)}<0.$
-This implies the matrix $\Lambda_{ijk}$ is non-degenerate. By Lemma 2.22, the matrix $\Lambda_{ijk}$ is symmetric. Set $a_{rrs}:=C_{r}\cdot\coth d_{rs}>0$. Define
+This implies the matrix $\Lambda_{ijk}$ is non-degenerate. The symmetry of $\Lambda_{ijk}$ follows from Lemma 2.21. Set $a_{rrs}:=C_{r}\cdot\coth d_{rs}>0$. Define
 $Q:=Q_{1}Q_{2}Q_{3}=\left(\begin{array}[]{ccc}-1&\cosh\theta_{k}&\cosh\theta_{j}\\ \cosh\theta_{k}&-1&\cosh\theta_{i}\\ \cosh\theta_{j}&\cosh\theta_{i}&-1\\ \end{array}\right)\\ \times\left(\begin{array}[]{ccc}0&a_{jjk}&a_{kkj}\\ a_{iik}&0&a_{kki}\\ a_{iij}&a_{jji}&0\\ \end{array}\right).$
-Then
+Then $\Lambda_{ijk}$ can be expressed as
 $\Lambda_{ijk}=-\frac{1}{A}\left(\begin{array}[]{ccc}\sinh l_{i}&0&0\\ 0&\sinh l_{j}&0\\ 0&0&\sinh l_{k}\\ \end{array}\right)\times Q.$
-To prove the negative definiteness of $\Lambda_{ijk}$, it suffices to prove $Q$ is positive definite.
-The formula (41) shows that $\det Q>0$.
-It remains to prove that the determinants of the $1\times 1$ and $2\times 2$ order principal minor of $Q$ are positive.
-Let $m_{rs}$ be the entry of $Q$ at $r$-th row and $s$-th column.
+To prove the negative definiteness of $\Lambda_{ijk}$, it suffices to show that $Q$ is positive definite.
+Formula (41) shows that $\det Q>0$.
+It remains to verify that the $1\times 1$ and $2\times 2$ leading principal minors of $Q$ are positive.
+Let $m_{rs}$ denote the entry of $Q$ at $r$-th row and $s$-th column.
 Then
@@
 $m_{21}=-a_{iik}+\cosh\theta_{i}\cdot a_{iij}.$
-Thus the determinants of the $1\times 1$ principal submatrices of $Q$ is $m_{11}>0$ by $a_{iik}>0,a_{iij}>0$. And the determinants of the $2\times 2$ principal submatrices of $Q$ is
+Thus the $1\times 1$ leading principal minor of $Q$ is $m_{11}>0$ due to $a_{iik}>0$ and $a_{iij}>0$. And the $2\times 2$ leading principal minor of $Q$ is
 $m_{11}m_{22}-m_{12}m_{21}$ $=\sinh^{2}\theta_{k}a_{iik}a_{jjk}+(\cosh\theta_{i}\cosh\theta_{k}+\cosh\theta_{j})a_{iik}a_{jji}$
@@
 This completes the proof. Q.E.D.
-As a result, we have the following corollary.
+As a consequence, we obtain the following corollary.
 Corollary 3.3.
-Under the same assumptions as in Theorem 3.1.
-The Jacobian $\Lambda=\frac{\partial(K_{i},...,K_{N})}{\partial(u_{i},...,u_{N})}$ is symmetric and negative definite on the admissible space $\mathcal{U}(\eta)$.
+Under the same assumptions as those in Theorem 3.1, the Jacobian $\Lambda=\frac{\partial(K_{i},...,K_{N})}{\partial(u_{i},...,u_{N})}$ is symmetric and negative definite on $\mathcal{U}(\eta)$.
+Proof.
+The conclusion follows directly from Theorem 3.2 and the identity $\Lambda=\sum_{\{ijk\}\in F}\Lambda_{ijk}$.
+Specifically, each $\Lambda_{ijk}$ is extended to an $N\times N$ matrix via zero-padding, such that it acts non-trivially on a vector $(1,...,N)$ exclusively through the coordinates corresponding to the boundary components $i,j,k$ of the right-angled hyperbolic hexagon $\{ijk\}\in F$.
+Q.E.D.
 3.3. Rigidity of the discrete conformal structure (A1)
-The following theorem gives the rigidity of discrete conformal structure (A1), which is the case $\mathrm{(i)}$ in Theorem 1.9.
+The following theorem establishes the rigidity of the discrete conformal structure (A1), which corresponds to the rigidity part of Theorem 1.8 (i).
 Theorem 3.4.
-Under the same assumptions as in Theorem 3.1.
-The function $f$ is determined by its generalized combinatorial curvature $K\in\mathbb{R}^{N}_{>0}$.
-In particular, the map from $f$ to $K$ is a smooth embedding.
+Under the same assumptions as those in Theorem 3.1, the discrete conformal factor $f$ is uniquely determined by its generalized combinatorial curvature $K\in\mathbb{R}^{N}_{>0}$.
 Proof.
-Theorem 3.1 and Theorem 3.2 imply that the following energy function
+Theorem 3.1 and Theorem 3.2 imply that for a right-angled hyperbolic hexagon $\{ijk\}\in F$, the following energy function
 […]
-for a right-angled hyperbolic hexagon $\{ijk\}\in F$ is a well-defined smooth function on $\mathcal{U}_{ijk}(\eta)$.
-Furthermore, $\mathcal{F}_{ijk}(u_{i},u_{j},u_{k})$ is a strictly concave function on $\mathcal{U}_{ijk}(\eta)$ with $\nabla_{u_{i}}\mathcal{F}_{ijk}=\theta^{jk}_{i}$.
+is a well-defined smooth function on $\mathcal{U}_{ijk}(\eta)$.
+Furthermore, $\mathcal{F}_{ijk}(u_{i},u_{j},u_{k})$ is strictly concave on $\mathcal{U}_{ijk}(\eta)$ with $\nabla_{u_{i}}\mathcal{F}_{ijk}=\theta^{jk}_{i}$.
 Define the function ~$\widetilde{\mathcal{F}}:\mathcal{U}(\eta)\rightarrow\mathbb{R}$ by
 ~$\widetilde{\mathcal{F}}(u)=\sum_{\{ijk\}\in F}\mathcal{F}_{ijk}(u_{i},u_{j},u_{k}).$
-By Corollary 3.3, ~$\widetilde{\mathcal{F}}(u)$ is a strictly concave function on $\mathcal{U}(\eta)$ with $\nabla_{u_{i}}\widetilde{\mathcal{F}}(u)=K_{i}$.
-According to Lemma 1.11, the map ~$\nabla\widetilde{\mathcal{F}}:\mathcal{U}(\eta)\rightarrow\mathbb{R}^{N}$ is a smooth embedding.
-Therefore, the map from $\{f\in\Omega(\eta)=\bigcap_{\{ijk\}\in F}\Omega_{ijk}(\eta)\}$ to $\{K\in\mathbb{R}^{N}_{>0}\}$ is a smooth injective map.
+By Corollary 3.3, ~$\widetilde{\mathcal{F}}(u)$ is strictly concave on $\mathcal{U}(\eta)$ with $\nabla_{u_{i}}\widetilde{\mathcal{F}}(u)=K_{i}$.
+By Lemma 1.10, the gradient map ~$\nabla\widetilde{\mathcal{F}}:\mathcal{U}(\eta)\rightarrow\mathbb{R}^{N}$ is a smooth embedding.
+Note that $\mathcal{U}(\eta)$ and $\Omega(\eta)=\bigcap_{\{ijk\}\in F}\Omega_{ijk}(\eta)$ are diffeomorphic.
+Consequently, the map from $\{f\in\Omega(\eta)\}$ to $\{K\in\mathbb{R}^{N}_{>0}\}$ is a smooth injective map.
 Q.E.D.
-3.4. Existence of discrete conformal structure (A1)
-The following theorem characterizes the image of $K$.
+3.4. Existence of the discrete conformal structure (A1)
+The following theorem characterizes the image of $K$ for the discrete conformal structure (A1), which corresponds to the existence part of Theorem 1.8 (i).
 Theorem 3.5.
-Under the same assumptions as in Theorem 3.1. Suppose $\alpha:B\rightarrow\{0,1\}$. Then the image of $K$ is >$\mathbb{R}^{N}_{>0}$.
+Suppose $(\Sigma,\mathcal{T},\alpha,\eta)$ is a weighted triangulated surface with boundary, where the weights are given by $\alpha:B\rightarrow\{0,1\}$ and $\eta\in\mathbb{R}_{>0}^{E}$ satisfying $\eta_{ij}>\alpha_{i}\alpha_{j}$ for any two adjacent boundary components $i,j\in B$.
+The image of $K$ is >$\mathbb{R}^{N}_{>0}$.
 To prove Theorem 3.5, we need the following two lemmas.
 Lemma 3.6.
-Under the same assumptions as in Theorem 3.1. For a right-angled hyperbolic hexagon $\{ijk\}\in F$ on $\Omega_{ijk}(\eta)$, if one of the following conditions is satisfied
+Under the same assumptions as those in Theorem 3.1, for a right-angled hyperbolic hexagon $\{ijk\}\in F$ on $\Omega_{ijk}(\eta)$, if one of the following conditions is satisfied
 (i):
@@
 $\lim f_{i}=+\infty,\ \lim f_{j}=a,\ \lim f_{k}=b$.
-Then $\lim\theta^{jk}_{i}(f_{i},f_{j},f_{k})=0$, where $a,\ b,\ c$ are constants.
+where $a,b,c$ are constants. Then $\lim\theta^{jk}_{i}(f_{i},f_{j},f_{k})=0$.
 Proof.
-For case $\mathrm{(i)}$, $\lim f_{i}=+\infty,\ \lim f_{j}=+\infty,\ \lim f_{k}=+\infty$. By (5), we have
+For the case (i), it follows from (5) that
 (42) $\cosh l_{ij}=$ $\bigg{[}\eta_{ij}-\sqrt{(e^{-2f_{i}}+\alpha_{i})(e^{-2f_{j}}+\alpha_{j})}\bigg{]}e^{f_{i}+f_{j}},$
@@
 $L_{1}\rightarrow\frac{c_{i}e^{f_{j}+f_{k}}}{c_{k}e^{f_{i}+f_{j}}\cdot c_{j}e^{f_{i}+f_{k}}}\rightarrow 0\quad\ \text{and}\ \quad L_{2}\rightarrow 1,$
-then $\cosh\theta^{jk}_{i}\rightarrow 1$ and hence $\theta_{i}^{jk}\rightarrow 0$. The same proofs are suitable for the cases $\mathrm{(ii)}$ and $\mathrm{(iii)}$, and we omit them here. Q.E.D.
+it follows that $\cosh\theta^{jk}_{i}\rightarrow 1$, which implies $\theta_{i}^{jk}\rightarrow 0$.
+The same arguments apply for the cases (ii) and (iii), the details of which are omitted here for brevity.
+Q.E.D.
 Lemma 3.7.
-Under the same assumptions as in Theorem 3.1.
-For a right-angled hyperbolic hexagon $\{ijk\}\in F$ on $\Omega_{ijk}(\eta)$, if $\lim f_{i}=-\infty$, then $\lim f_{j}=\lim f_{k}=+\infty$.
+Under the same assumptions as those in Theorem 3.1, for a right-angled hyperbolic hexagon $\{ijk\}\in F$ on $\Omega_{ijk}(\eta)$, if $\lim f_{i}=-\infty$, then $\lim f_{j}=\lim f_{k}=+\infty$.
 Furthermore, $\lim\theta^{jk}_{i}(f_{i},f_{j},f_{k})=+\infty$.
 Proof.
-According to the definition of the admissible space $\mathcal{U}_{ijk}(\eta)$ by (35), there is no subset $\{r,s\}\subseteq\{i,j,k\}$ such that $\lim u_{r}=-\infty$ and $\lim u_{s}=-\infty$ or $\lim u_{s}=c$, where $c$ is a constant.
-Since $\frac{\partial f_{r}}{\partial u_{r}}=\sqrt{1+\alpha_{r}e^{2f_{r}}}$ for $r\in\{i,j,k\}$, by Remark 2.23, if $\lim f_{i}=-\infty$, then $\lim u_{i}=-\infty$; if $\lim f_{j}=+\infty$, then $\lim u_{j}=+\infty$ with $\alpha_{j}=0$ and $\lim u_{j}=0^{-}$ with $\alpha_{j}=1$.
-Therefore, if $\lim f_{i}=-\infty$, then $\lim u_{i}=-\infty$ and hence $\lim u_{j}=\lim u_{k}=+\infty$.
-This implies $\alpha_{j}=\alpha_{k}=0$ and $\lim f_{j}=\lim f_{k}=+\infty$.
-If $\alpha_{i}=0$, then $\cosh l_{ij}=-1+\eta_{ij}e^{u_{i}+u_{j}}$$\cosh l_{jk}=-1+\eta_{jk}e^{u_{j}+u_{k}}$ and $\cosh l_{ik}=-1+\eta_{ik}e^{u_{i}+u_{k}}$. Thus
+The definition of the admissible space $\mathcal{U}_{ijk}(\eta)$ in (35) implies $u_{r}+u_{s}>C(\eta_{rs})$ for any $\{r,s\}\subseteq\{i,j,k\}$.
+Hence, there exists no subset $\{r,s\}\subseteq\{i,j,k\}$ such that $\lim u_{r}=-\infty$ and either $\lim u_{s}=-\infty$ or $\lim u_{s}=c$, where $c$ is a constant.
+By combining (34) and Remark 2.22, the following hold: if $\lim f_{i}=-\infty$, then $\lim u_{i}=-\infty$; if $\lim f_{j}=+\infty$, then $\lim u_{j}=+\infty$ when $\alpha_{j}=0$ and $\lim u_{j}=0^{-}$ when $\alpha_{j}=1$.
+Therefore, if $\lim f_{i}=-\infty$, then $\lim u_{i}=-\infty$, which implies $\lim u_{j}=\lim u_{k}=+\infty$.
+This further yields $\alpha_{j}=\alpha_{k}=0$ and $\lim f_{j}=\lim f_{k}=+\infty$.
+If $\alpha_{i}=0$, then $\cosh l_{ij}=-1+\eta_{ij}e^{u_{i}+u_{j}}$, $\cosh l_{jk}=-1+\eta_{jk}e^{u_{j}+u_{k}}$, and $\cosh l_{ik}=-1+\eta_{ik}e^{u_{i}+u_{k}}$. Thus
 $L_{1}$ $=\frac{-1+\eta_{jk}e^{u_{j}+u_{k}}}{\sqrt{(-1+\eta_{ij}e^{u_{i}+u_{j}})^{2}-1}\cdot\sqrt{(-1+\eta_{ik}e^{u_{i}+u_{k}})^{2}-1}}$
 $=\frac{-1+\eta_{jk}e^{u_{j}+u_{k}}}{\sqrt{\eta_{ij}\eta_{ik}}e^{u_{j}+u_{k}}}\cdot\frac{1}{\sqrt{\eta_{ij}e^{2u_{i}}-2e^{u_{i}-u_{j}}}}\cdot\frac{1}{\sqrt{\eta_{ik}e^{2u_{i}}-2e^{u_{i}-u_{k}}}}\rightarrow+\infty.$
-And $L_{2}>0$. Then $\cosh\theta^{jk}_{i}=L_{1}+L_{2}\rightarrow+\infty$, which implies $\theta^{jk}_{i}\rightarrow+\infty$.
+Since $L_{2}>0$, it follows that $\cosh\theta^{jk}_{i}=L_{1}+L_{2}\rightarrow+\infty$, which implies $\theta^{jk}_{i}\rightarrow+\infty$.
 If $\alpha_{i}=1$, then $\sqrt{1+e^{2f_{i}}}=-\coth u_{i}$ and $e^{f_{i}}=-\frac{1}{\sinh u_{i}}$. Hence,
@@
 Similarly, $\cosh l_{ik}\rightarrow-1+2\eta_{ik}e^{u_{i}+u_{k}}$ and $\cosh l_{jk}=-1+\eta_{jk}e^{u_{j}+u_{k}}$.
-Similar to the proof of the case $\alpha_{i}=0$, we have $\cosh\theta^{jk}_{i}=L_{1}+L_{2}\rightarrow+\infty$, then $\theta^{jk}_{i}\rightarrow+\infty$.
+By reasoning analogous to the case $\alpha_{i}=0$, we have $\cosh\theta^{jk}_{i}=L_{1}+L_{2}\rightarrow+\infty$, which implies $\theta^{jk}_{i}\rightarrow+\infty$.
 If $\alpha_{i}=-1$, then $\sqrt{1-e^{2f_{i}}}=-\tanh u_{i}$ and $e^{f_{i}}=\frac{1}{\cosh u_{i}}$. Hence,
@@
 Similarly, $\cosh l_{ik}\rightarrow-1+2\eta_{ik}e^{u_{i}+u_{k}}$ and $\cosh l_{jk}=-1+\eta_{jk}e^{u_{j}+u_{k}}$.
-Similar to the proof of the case $\alpha_{i}=0$, we have $\cosh\theta^{jk}_{i}=L_{1}+L_{2}\rightarrow+\infty$, then $\theta^{jk}_{i}\rightarrow+\infty$.
+Similar to the proof of the case $\alpha_{i}=0$, we have $\cosh\theta^{jk}_{i}=L_{1}+L_{2}\rightarrow+\infty$, which implies $\theta^{jk}_{i}\rightarrow+\infty$.
 Q.E.D.
-Proof of Theorem 3.5: Let the image of $K$ be $X$.
-To prove $X=\mathbb{R}^{N}_{>0}$, we will show that $X$ is both open and closed in >$\mathbb{R}^{N}_{>0}$.
-By the definition of curvature map $K$, we have $X\subseteq\mathbb{R}^{N}_{>0}$.
-Since the map from $f\in\Omega(\eta)$ to >$\mathbb{R}^{N}_{>0}$ is a smooth injective map, then $X$ is open in >$\mathbb{R}^{N}_{>0}$.
-It remains to prove that $X$ is closed in >$\mathbb{R}^{N}_{>0}$.
-We show that if a sequence $f^{(m)}$ in $\Omega(\eta)$ satisfies $\lim_{m\rightarrow+\infty}K^{(m)}\in\mathbb{R}^{N}_{>0}$, then there exists a subsequence, say $f^{(m)}$, such that $\lim_{m\rightarrow+\infty}f^{(m)}$ is in $\Omega(\eta)$.
-Suppose otherwise, there exists a subsequence $f^{(m)}$, such that $\lim_{m\rightarrow+\infty}f^{(m)}$ is in the boundary of $\Omega(\eta)$.
-By (33), the boundary of the admissible space $\Omega(\eta)$ in $[-\infty,+\infty]^{N}$ consists of the following three parts
+Proof of Theorem 3.5: Let $X$ be the image of $K$.
+To prove $X=\mathbb{R}^{N}_{>0}$, we aim to show that $X$ is both an open and a closed subset of >$\mathbb{R}^{N}_{>0}$.
+By the definition of the curvature $K$, it follows directly that $X\subseteq\mathbb{R}^{N}_{>0}$.
+Since the map from $f\in\Omega(\eta)$ to >$\mathbb{R}^{N}_{>0}$ is a smooth injective map, its image $X$ must be an open subset of >$\mathbb{R}^{N}_{>0}$.
+It remains to verify that $X$ is closed in >$\mathbb{R}^{N}_{>0}$.
+Suppose $\{f^{(m)}\}$ is a sequence in $\Omega(\eta)$ such that $\lim_{m\rightarrow+\infty}K^{(m)}\in\mathbb{R}^{N}_{>0}$.
+We shall show that there exists a subsequence, still denoted by $\{f^{(m)}\}$, such that $\lim_{m\rightarrow+\infty}f^{(m)}\in\Omega(\eta)$.
+Suppose, for contradiction, that there exists a subsequence $\{f^{(m)}\}$ such that $\lim_{m\rightarrow+\infty}f^{(m)}$ lies on the boundary of $\Omega(\eta)$.
+By (33), the boundary of the admissible space $\Omega(\eta)$ in $[-\infty,+\infty]^{N}$ consists of three distinct components, specified as follows:
 (i):
-There exists an ideal edge $\{rs\}\in E$ such that $l_{rs}=0$, i.e., $u_{r}+u_{s}=-M(\eta_{rs})$ for the admissible space $\mathcal{U}_{ijk}(\eta)$ defined by (35).
+There exists an ideal edge $\{rs\}\in E$ such that $l_{rs}=0$, a condition equivalent to $u_{r}+u_{s}=C(\eta_{rs})$ for the admissible space $\mathcal{U}_{ijk}(\eta)$ defined by (35);
 (ii):
-For $\alpha\in\{0,1\}$, there exists $i\in B$ satisfying $f_{i}^{(m)}\rightarrow+\infty$.
+For $\alpha\in\{0,1\}$, there exists $i\in B$ such that $f_{i}^{(m)}\rightarrow+\infty$;
 (iii):
-For $\alpha\in\{0,1\}$, there exists $i\in B$ satisfying $f_{i}^{(m)}\rightarrow-\infty$.
-For case $\mathrm{(i)}$, in a right-angled hyperbolic hexagon $\{ijk\}\in F$ with edge lengths $l_{ij},l_{jk},l_{ik}$ and opposite lengths of boundaries $\theta^{ij}_{k},\theta^{jk}_{i},\theta^{ik}_{j}$.
+For $\alpha\in\{0,1\}$, there exists $i\in B$ such that $f_{i}^{(m)}\rightarrow-\infty$.
+For the case (i), consider a right-angled hyperbolic hexagon $\{ijk\}\in F$ with edge lengths $l_{ij},l_{jk},l_{ik}$ and opposite boundary lengths $\theta^{ij}_{k},\theta^{jk}_{i},\theta^{ik}_{j}$.
 Without loss of generality, we assume $l_{ij}\rightarrow 0$.
-By the cosine law,
+By the cosine law, we have
 $\cosh\theta^{ik}_{j}=\frac{\cosh l_{ik}+\cosh l_{ij}\cosh l_{jk}}{\sinh l_{ij}\sinh l_{jk}}>\frac{\cosh l_{ij}\cosh l_{jk}}{\sinh l_{ij}\sinh l_{jk}}\geq\frac{\cosh l_{ij}}{\sinh l_{ij}}\rightarrow+\infty,$
 which contradicts the assumption that $\lim_{m\rightarrow+\infty}K^{(m)}\in\mathbb{R}^{N}_{>0}$.
-The case $\mathrm{(ii)}$ follows from Lemma 3.6.
-In fact, each generalized angle incident to the boundary component $i\in B$ converges to 0.
-Hence $\lim_{m\rightarrow+\infty}K^{(m)}=0$.
-This contradicts with the assumption that $\lim_{m\rightarrow+\infty}K^{(m)}\in\mathbb{R}^{N}_{>0}$.
-The case $\mathrm{(iii)}$ follows from Lemma 3.7.
-In fact, each generalized angle incident to the boundary component $i\in B$ converges to +$\infty$+ ∞.
-Hence $\lim_{m\rightarrow+\infty}K^{(m)}=+\infty$.
-This contradicts with the assumption that $\lim_{m\rightarrow+\infty}K^{(m)}\in\mathbb{R}^{N}_{>0}$.
+For the case (ii), by Lemma 3.6, each generalized angle incident to the boundary component $i\in B$ converges to 0.
+Hence, $\lim_{m\rightarrow+\infty}K_{i}^{(m)}=0$.
+For the case (iii), Lemma 3.7 implies that $\lim_{m\rightarrow+\infty}K_{i}^{(m)}=+\infty$.
+Both cases contradict the assumption that $\lim_{m\rightarrow+\infty}K^{(m)}\in\mathbb{R}^{N}_{>0}$.
 Q.E.D.
 Remark 3.8.
-Note that Lemma 3.6 and Lemma 3.7 hold for $\alpha:B\rightarrow\{-1,0,1\}$.
-But Theorem 3.5 holds for $\alpha:B\rightarrow\{0,1\}$.
-If $\alpha_{i}=-1$, then the boundary of the admissible space $\Omega(\eta)$ in $[-\infty,+\infty]^{N}$ will include the case that there exists $i\in B$ satisfying $f_{i}^{(m)}\rightarrow 0^{-}$.
-We are unable to prove the closeness of $X$ in this case.
+Note that Lemma 3.6 and Lemma 3.7 are valid for $\alpha:B\rightarrow\{-1,0,1\}$.
+In contrast, Theorem 3.5 only holds for $\alpha:B\rightarrow\{0,1\}$.
+Specifically, when $\alpha_{i}=-1$, the boundary of the admissible space $\Omega(\eta)$ in $[-\infty,+\infty]^{N}$ will include the case that there exists $i\in B$ such that $f_{i}^{(m)}\rightarrow 0^{-}$.
+For this case, we are unable to prove the closedness of $X$.
 4. Rigidity and existence of the discrete conformal structure (A2)
-By Remark 1.7, we can set $C\equiv 0$ in the discrete conformal structure (A2).By Remark 1.7, we can set $C\equiv 0$ in the discrete conformal structure (A2).
-In this case, the discrete conformal structure (A2) is the $(-1,-1,0)$ type generalized circle packings in Guo-Luo [5].
-Its rigidity and existence have been proved by Guo-Luo [5], which is as follows.
+By Remark 1.6, we set $\alpha\equiv-1$. Then (6) can be written as
+(43) $\cosh l_{ij}=\sqrt{(e^{2f_{i}}-1)(e^{2f_{j}}-1)}+\eta_{ij}e^{f_{i}+f_{j}}.$
+For a right-angled hyperbolic hexagon $\{ijk\}\in F$ with edge lengths $l_{ij},l_{jk},l_{ki}$ given by (43), the admissible space of the discrete conformal factors $f$ is defined as
+(44) $\Omega_{ijk}(\eta)=\{(f_{i},f_{j},f_{k})\in\mathbb{R}_{>0}^{3}\mid l_{ij}>0,\ l_{jk}>0,\ l_{ik}>0\}.$
+Since $\frac{\partial f_{i}}{\partial u_{i}}=\sqrt{e^{2f_{i}}-1}$, it follows from Remark 2.22 (iv) that $\tan u_{i}=\sqrt{e^{2f_{i}}-1}$ for $u_{i}\in(0,\frac{\pi}{2})$ and $-\frac{1}{\tan u_{i}}=\sqrt{e^{2f_{i}}-1}$ for $u_{i}\in(-\frac{\pi}{2},0)$.
+For simplicity, we consider $u\in(-\frac{\pi}{2},0)^{N}$.
+Consequently,
+(45) $\cosh l_{ij}=\frac{1}{\tan u_{i}\tan u_{j}}+\eta_{ij}\frac{1}{\sin u_{i}\sin u_{j}}>1,$
+which implies $\cos(u_{i}+u_{j})>-\eta_{ij}$.
+This further yields $u_{i}+u_{j}\in(-\pi,0)$ when $\eta_{ij}\in[1,+\infty)$, and $u_{i}+u_{j}>-\arccos\eta_{ij}$ when $\eta_{ij}\in(-1,1)$.
+Hence, the admissible space $\Omega_{ijk}(\eta)$ is transformed into the following admissible space $\mathcal{U}_{ijk}(\eta)$ of $u$ for a right-angled hyperbolic hexagon $\{ijk\}\in F$:
+(46) $\mathcal{U}_{ijk}(\eta)=u(\Omega_{ijk}(\eta))=\{(u_{i},u_{j},u_{k})\in(-\frac{\pi}{2},0)^{3}\mid\cos(u_{r}+u_{s})>-\eta_{rs},\,\forall\{r,s\}\subseteq\{i,j,k\}\},$
+where $\eta_{rs}\in(-1,+\infty)$.
+Hence, $\mathcal{U}_{ijk}(\eta)$ is a convex polytope.
+As a result, the admissible space $\mathcal{U}(\eta)=\bigcap_{\{ijk\}\in F}\mathcal{U}_{ijk}(\eta)$ is also a convex polytope on $(\Sigma,\mathcal{T},\eta)$.
 Theorem 4.1.
-([5]) Suppose $(\Sigma,\mathcal{T},\eta)$ is a weighted triangulated surface with boundary and the weight $\eta\in\mathbb{R}_{>0}^{E}$, the discrete conformal structure $f\in\Omega(\eta)$ is determined by its generalized combinatorial curvature $K\in\mathbb{R}^{N}_{>0}$.
+Suppose $(\Sigma,\mathcal{T},\eta)$ is a weighted triangulated surface with boundary, where $\eta\in(-1,+\infty)^{E}$.
+For a right-angled hyperbolic hexagon $\{ijk\}\in F$ on $\mathcal{U}_{ijk}(\eta)$, the Jacobian $\Lambda_{ijk}=\frac{\partial(\theta^{jk}_{i},\theta^{ik}_{j},\theta^{ij}_{k})}{\partial(u_{i},u_{j},u_{k})}$ is symmetric and negative definite.
+Consequently, the Jacobian $\Lambda=\frac{\partial(K_{i},...,K_{N})}{\partial(u_{i},...,u_{N})}$ is symmetric and negative definite on $\mathcal{U}(\eta)$.
+Proof.
+For simplicity, we adopt the notations introduced in Subsection 3.2. Define
+$\frac{\partial(f_{i},f_{j},f_{k})}{\partial(u_{i},u_{j},u_{k})}=\left(\begin{array}[]{ccc}\sqrt{e^{2f_{i}}-1}&0&0\\ 0&\sqrt{e^{2f_{j}}-1}&0\\ 0&0&\sqrt{e^{2f_{k}}-1}\\ \end{array}\right):=Q_{4}.$
+Then $\Lambda_{ijk}$ can be expressed as
+$\Lambda_{ijk}=\frac{\partial(\theta_{i},\theta_{j},\theta_{k})}{\partial(u_{i},u_{j},u_{k})}=-\frac{1}{A}\left(\begin{array}[]{ccc}\sinh l_{i}&0&0\\ 0&\sinh l_{j}&0\\ 0&0&\sinh l_{k}\\ \end{array}\right)Q_{1}Q_{2}Q_{4},$
+where $Q_{1}$ and $Q_{2}$ are defined by (36) and (38), respectively.
+Furthermore, $\det Q_{1}>0$.
+Since $d>0$ for the discrete conformal structure (A2) in Theorem 1.5, it follows from (39) that $\det Q_{2}>0$.
+Consequently, $\det\Lambda_{ijk}<0$, which implies $\Lambda_{ijk}$ is non-degenerate.
+The symmetry of $\Lambda_{ijk}$ follows from Lemma 2.21.
+Set $a_{rrs}=\sqrt{e^{2f_{r}}-1}\cdot\coth d_{rs}>0$.
+Define
+$Q:=Q_{1}Q_{2}Q_{4}=\left(\begin{array}[]{ccc}-1&\cosh\theta_{k}&\cosh\theta_{j}\\ \cosh\theta_{k}&-1&\cosh\theta_{i}\\ \cosh\theta_{j}&\cosh\theta_{i}&-1\\ \end{array}\right)\\ \times\left(\begin{array}[]{ccc}0&a_{jjk}&a_{kkj}\\ a_{iik}&0&a_{kki}\\ a_{iij}&a_{jji}&0\\ \end{array}\right).$
+The remaining part of the proof is analogous to that of Theorem 3.2, and thus is omitted here.
+Q.E.D.
+The following theorem establishes the rigidity of the discrete conformal structure (A2), which corresponds to the rigidity part of Theorem 1.8 (ii).
+As its proof is analogous to that of Theorem 3.4, it is omitted here.
+Theorem 4.2.
+Suppose $(\Sigma,\mathcal{T},\eta)$ is a weighted triangulated surface with boundary, where $\eta\in(-1,+\infty)^{E}$.
+The discrete conformal factor $f$ is uniquely determined by its generalized combinatorial curvature $K\in\mathbb{R}^{N}_{>0}$.
+The following theorem characterizes the image of $K$ for the discrete conformal structure (A2), which corresponds to the existence part of Theorem 1.8 (ii).
+Theorem 4.3.
+Suppose $(\Sigma,\mathcal{T},\eta)$ is a weighted triangulated surface with boundary, where $\eta\in(-1,0]^{E}$. The image of $K$ is >$\mathbb{R}^{N}_{>0}$.
+To prove Theorem 4.3, we need the following lemma. The proof is nearly identical to that of Lemma 3.6, and is thus omitted here.
+Lemma 4.4.
+Under the same assumptions as those in Theorem 4.2, for a right-angled hyperbolic hexagon $\{ijk\}\in F$ on $\Omega_{ijk}(\eta)$, if one of the following conditions is satisfied
+(i):
+$\lim f_{i}=+\infty,\ \lim f_{j}=+\infty,\ \lim f_{k}=+\infty$,
+(ii):
+$\lim f_{i}=+\infty,\ \lim f_{j}=+\infty,\ \lim f_{k}=c$,
+(iii):
+$\lim f_{i}=+\infty,\ \lim f_{j}=a,\ \lim f_{k}=b$,
+where $a,b,c$ are positive constants. Then $\lim\theta^{jk}_{i}(f_{i},f_{j},f_{k})=0$.
+Proof of Theorem 4.3: Let $X$ be the image of $K$.
+The openness of $X$ in >$\mathbb{R}^{N}_{>0}$ is a consequence of the smooth injectivity of the map, a property that plays a pivotal role in establishing the rigidity theorem.
+To prove the closedness of $X$ in >$\mathbb{R}^{N}_{>0}$, we show that for any sequence $\{f^{(m)}\}\in\Omega(\eta)=\bigcap_{\{ijk\}\in F}\Omega_{ijk}(\eta)$ satisfying $\lim_{m\rightarrow+\infty}K^{(m)}\in\mathbb{R}^{N}_{>0}$, there exists a subsequence, still denoted by $\{f^{(m)}\}$, such that $\lim_{m\rightarrow+\infty}f^{(m)}\in\Omega(\eta)$.
+To derive a contradiction, suppose there exists a subsequence $\{f^{(m)}\}$ such that $\lim_{m\rightarrow+\infty}f^{(m)}$ lies on the boundary of $\Omega(\eta)$.
+By (44), the boundary of the admissible space $\Omega(\eta)$ in $[-\infty,+\infty]^{N}$ consists of the following three parts:
+(i):
+There exists an ideal edge $\{rs\}\in E$ such that $l_{rs}=0$;
+(ii):
+There exists $i\in B$ such that $f_{i}^{(m)}\rightarrow+\infty$;
+(iii):
+There exists $i\in B$ such that $f_{i}^{(m)}\rightarrow 0^{+}$.
+The proof for the case (i) is identical to that of Theorem 3.5 and is thus omitted here.
+For the case (ii), by Lemma 4.4, each generalized angle incident to the boundary component $i\in B$ converges to 0.
+Consequently, $\lim_{m\rightarrow+\infty}K_{i}^{(m)}=0$.
+This contradicts the assumption that $\lim_{m\rightarrow+\infty}K^{(m)}\in\mathbb{R}^{N}_{>0}$.
+For the case (iii), if $f_{i}^{(m)}\rightarrow 0^{+}$, then $\cosh l_{ij}=\sqrt{(e^{2f_{i}}-1)(e^{2f_{j}}-1)}+\eta_{ij}e^{f_{i}+f_{j}}\rightarrow\eta_{ij}e^{f_{j}}\geq 1$.
+However, since $\eta_{ij}\in(-1,0]$, it follows that $\eta_{ij}e^{f_{j}}\leq 0$, which is a contradiction.
+Q.E.D.
+Remark 4.5.
+If $\eta=-\cos\Phi\in(-1,1]^{E}$ and $e^{f}=\cosh r$, then the discrete conformal structure (A2) reduces to the $(-1,-1,1)$-type generalized circle packings introduced by Guo-Luo in [5], where the rigidity of such generalized circle packings was also established.
+Consequently, Theorem 4.2 generalizes the rigidity result for Guo-Luo’s $(-1,-1,1)$-type generalized circle packings in [5].
+Guo-Luo further asserted that the existence holds for $\eta\in(-1,1]^{E}$.
+However, their proof in [5] contains a gap when $\eta\in(0,1)^{E}$, because they did not account for the boundary behavior corresponding to $r\rightarrow 0^{+}$.
+As a result, Theorem 4.3 holds for $\eta\in(-1,0]^{E}$ rather than $\eta\in(-1,1]^{E}$.
+5. Rigidity and existence of the discrete conformal structure (A3)
+By Remark 1.6, we set $C\equiv 0$ in the discrete conformal structure (A3).
+In this case, the discrete conformal structure (A3) coincides precisely with the $(-1,-1,0)$-type generalized circle packings introduced by Guo-Luo in [5].
+Guo-Luo [5] have established the rigidity and existence results for such structures, which can be stated as follows.
+Theorem 5.1.
+([5]) Suppose $(\Sigma,\mathcal{T},\eta)$ is a weighted triangulated surface with boundary, where $\eta\in\mathbb{R}_{>0}^{E}$.
+The discrete conformal factor $f$ is uniquely determined by its generalized combinatorial curvature $K\in\mathbb{R}^{N}_{>0}$.
 Furthermore, the image of $K$ is $\mathbb{R}_{>0}^{N}$.
-In this section, we reprove the rigidity part of Theorem 4.1.
-In [5], Guo-Luo spent lots of space to prove the symmetry and negative definiteness of Jacobian.
-By (2) and (3), we simplify Guo-Luo’s proof, which is similar to that of Theorem 3.2.
-The admissible space of the discrete conformal factor $f$ for a right-angled hyperbolic hexagon $\{ijk\}\in F$ with edge lengths $l_{ij},l_{jk},l_{ki}$ given by (7) is defined to be
-$\Omega_{ijk}(\eta)=\{(f_{i},f_{j},f_{k})\in\mathbb{R}^{3}|\cosh l_{rs}=-\cosh(f_{r}-f_{s})+\eta_{rs}e^{f_{r}+f_{s}}>1,\ \forall\{r,s\}\subseteq\{i,j,k\}\}.$
-By Theorem 1.6, we have $d>0$ for the discrete conformal structure (A2). By a change of variables in Remark 2.23, i.e., $\frac{df_{r}}{du_{r}}=e^{f_{r}}$ and $u_{r}=-e^{-f_{r}}$. Thus
+In this section, we present a concise proof for the rigidity part of Theorem 5.1, which simplifies the argument given by Guo-Luo in [5].
+For the proof of the existence part, we refer the reader to [5].
+For a right-angled hyperbolic hexagon $\{ijk\}\in F$ with edge lengths $l_{ij},l_{jk},l_{ki}$ given by (7), the admissible space of the discrete conformal factor $f$ is defined as
+$\Omega_{ijk}(\eta)=\{(f_{i},f_{j},f_{k})\in\mathbb{R}^{3}\mid\cosh l_{rs}=-\cosh\,(f_{r}-f_{s})+\eta_{rs}e^{f_{r}+f_{s}}>1,\,\forall\{r,s\}\subseteq\{i,j,k\}\}.$
+By Theorem 1.5, for the discrete conformal structure (A3), we have $d>0$.
+By the variable transformation in Remark 2.22, specifically, $\frac{df_{r}}{du_{r}}=e^{f_{r}}$ and $u_{r}=-e^{-f_{r}}$, we derive
 $\eta_{rs}=\frac{\cosh l_{rs}+\cosh(f_{r}-f_{s})}{e^{f_{r}+f_{s}}}>\frac{1+\frac{1}{2}(e^{f_{r}-f_{s}}+e^{f_{s}-f_{r}})}{e^{f_{r}+f_{s}}}=\frac{1}{2}(e^{-f_{r}}+e^{-f_{s}})^{2}=\frac{1}{2}(u_{r}+u_{s})^{2}.$
-Since $u_{r},u_{s}\in\mathbb{R}_{<0}$, then $\eta_{rs}>0$ and $u_{r}+u_{s}>-\sqrt{2\eta_{rs}}$.
-Therefore, the admissible space $\Omega_{ijk}(\eta)$ of $f$ is transferred to the following admissible space $\mathcal{U}_{ijk}(\eta)$ of $u$ for a right-angled hyperbolic hexagon $\{ijk\}\in F$,
-$\mathcal{U}_{ijk}(\eta)=u(\Omega_{ijk}(\eta))=\{(u_{i},u_{j},u_{k})\in\mathbb{R}_{<0}^{3}|u_{r}+u_{s}>-\sqrt{2\eta_{rs}},\ \forall\{r,s\}\subseteq\{i,j,k\}\},$
-which is a convex polytope. Then the admissible space of $u$ on a weighted triangulated surface $(\Sigma,\mathcal{T},\eta)$ can be defined as
-$\mathcal{U}(\eta)=\bigcap_{\{ijk\}\in F}\mathcal{U}_{ijk}(\eta),$
-which is also a convex polytope.
+Since $u_{r}\in\mathbb{R}_{<0}$ and $u_{s}\in\mathbb{R}_{<0}$, it follows that $\eta_{rs}>0$ and $u_{r}+u_{s}>-\sqrt{2\eta_{rs}}$.
+Consequently, for a right-angled hyperbolic hexagon $\{ijk\}\in F$, the admissible space $\Omega_{ijk}(\eta)$ of $f$ is transformed into the following admissible space $\mathcal{U}_{ijk}(\eta)$ of $u$:
+(47) $\mathcal{U}_{ijk}(\eta)=u(\Omega_{ijk}(\eta))=\{(u_{i},u_{j},u_{k})\in\mathbb{R}_{<0}^{3}\mid u_{r}+u_{s}>-\sqrt{2\eta_{rs}},\,\forall\{r,s\}\subseteq\{i,j,k\}\},$
+which is a convex polytope. Consequently, the admissible space $\mathcal{U}(\eta)=\bigcap_{\{ijk\}\in F}\mathcal{U}_{ijk}(\eta)$ is also a convex polytope on $(\Sigma,\mathcal{T},\eta)$.
 Theorem 5.2.
-Suppose $(\Sigma,\mathcal{T},\eta)$ is a weighted triangulated surface with boundary and the weight $\eta\in\mathbb{R}_{>0}^{E}$.
-For a right-angled hyperbolic hexagon $\{ijk\}\in F$ on $\mathcal{U}_{ijk}(\eta)$, the Jacobian $\Lambda_{ijk}=\frac{\partial(\theta^{jk}_{i},\theta^{ik}_{j},\theta^{ij}_{k})}{\partial(u_{i},u_{j},u_{k})}$ is non-degenerate, symmetric and negative definite.
-As a result, the Jacobian $\Lambda=\frac{\partial(K_{i},...,K_{N})}{\partial(u_{i},...,u_{N})}$ is symmetric and negative definite on $\mathcal{U}(\eta)$.
+Suppose $(\Sigma,\mathcal{T},\eta)$ is a weighted triangulated surface with boundary, where $\eta\in\mathbb{R}_{>0}^{E}$.
+For a right-angled hyperbolic hexagon $\{ijk\}\in F$ on $\mathcal{U}_{ijk}(\eta)$, the Jacobian $\Lambda_{ijk}=\frac{\partial(\theta^{jk}_{i},\theta^{ik}_{j},\theta^{ij}_{k})}{\partial(u_{i},u_{j},u_{k})}$ is symmetric and negative definite.
+Consequently, the Jacobian $\Lambda=\frac{\partial(K_{i},...,K_{N})}{\partial(u_{i},...,u_{N})}$ is symmetric and negative definite on $\mathcal{U}(\eta)$.
 Proof.
-For simplicity, we set $l_{r}=l_{st},\ \theta_{r}=\theta^{st}_{r}$ and $A=\sinh l_{r}\sinh l_{s}\sinh\theta_{t}$, where $\{r,s,t\}=\{i,j,k\}$. Since $\frac{df_{r}}{du_{r}}=e^{f_{r}}$, then
-$\frac{\partial(f_{i},f_{j},f_{k})}{\partial(u_{i},u_{j},u_{k})}=\left(\begin{array}[]{ccc}e^{f_{i}}&0&0\\ 0&e^{f_{j}}&0\\ 0&0&e^{f_{k}}\\ \end{array}\right):=Q_{4}.$
-Hence
-$\Lambda_{ijk}=\frac{\partial(\theta_{i},\theta_{j},\theta_{k})}{\partial(u_{i},u_{j},u_{k})}=-\frac{1}{A}\left(\begin{array}[]{ccc}\sinh l_{i}&0&0\\ 0&\sinh l_{j}&0\\ 0&0&\sinh l_{k}\\ \end{array}\right)Q_{1}Q_{2}Q_{4},$
-where $Q_{1}$ is defined by (36) and $Q_{2}$ is defined by (38). Moreover, $\det Q_{1}>0$.
-Since $d>0$ for the discrete conformal structure (A2) in Theorem 1.6, then $\det Q_{2}>0$ by (39). Hence,
-$\det\Lambda_{ijk}=-\frac{1}{A^{3}}\sinh l_{i}\sinh l_{j}\sinh l_{k}\cdot\det Q_{1}\cdot\det Q_{2}\cdot\det Q_{4}<0,$
-which implies $\Lambda_{ijk}$ is non-degenerate. The symmetry of $\Lambda_{ijk}$ follows from Lemma 2.22. Set $a_{rrs}=e^{f_{r}}\cdot\coth d_{rs}>0$. Define
-$Q:=Q_{1}Q_{2}Q_{4}=\left(\begin{array}[]{ccc}-1&\cosh\theta_{k}&\cosh\theta_{j}\\ \cosh\theta_{k}&-1&\cosh\theta_{i}\\ \cosh\theta_{j}&\cosh\theta_{i}&-1\\ \end{array}\right)\\ \times\left(\begin{array}[]{ccc}0&a_{jjk}&a_{kkj}\\ a_{iik}&0&a_{kki}\\ a_{iij}&a_{jji}&0\\ \end{array}\right).$
-The rest of the proof is similar to that of Theorem 3.2, so we omit it. Q.E.D.
-The following theorem gives the rigidity of discrete conformal structure (A2), which is the case $\mathrm{(ii)}$ in Theorem 1.9. As the proof is similar to that of Theorem 3.4, we omit it here.
+For simplicity, we adopt the notations introduced in Subsection 3.2. Since $\frac{df_{r}}{du_{r}}=e^{f_{r}}$, it follows that
+$\frac{\partial(f_{i},f_{j},f_{k})}{\partial(u_{i},u_{j},u_{k})}=\left(\begin{array}[]{ccc}e^{f_{i}}&0&0\\ 0&e^{f_{j}}&0\\ 0&0&e^{f_{k}}\\ \end{array}\right):=Q_{5}.$
+Then $\Lambda_{ijk}$ can be expressed as
+$\Lambda_{ijk}=\frac{\partial(\theta_{i},\theta_{j},\theta_{k})}{\partial(u_{i},u_{j},u_{k})}=-\frac{1}{A}\left(\begin{array}[]{ccc}\sinh l_{i}&0&0\\ 0&\sinh l_{j}&0\\ 0&0&\sinh l_{k}\\ \end{array}\right)Q_{1}Q_{2}Q_{5},$
+where $Q_{1}$ and $Q_{2}$ are defined by (36) and (38), respectively.
+Furthermore, $\det Q_{1}>0$.
+Since $d>0$ for the discrete conformal structure (A3) in Theorem 1.5, it follows from (39) that $\det Q_{2}>0$.
+Consequently, $\det\Lambda_{ijk}<0$, which implies $\Lambda_{ijk}$ is non-degenerate.
+The symmetry of $\Lambda_{ijk}$ follows from Lemma 2.21.
+Set $a_{rrs}=e^{f_{r}}\cdot\coth d_{rs}>0$.
+Define
+$Q:=Q_{1}Q_{2}Q_{5}=\left(\begin{array}[]{ccc}-1&\cosh\theta_{k}&\cosh\theta_{j}\\ \cosh\theta_{k}&-1&\cosh\theta_{i}\\ \cosh\theta_{j}&\cosh\theta_{i}&-1\\ \end{array}\right)\\ \times\left(\begin{array}[]{ccc}0&a_{jjk}&a_{kkj}\\ a_{iik}&0&a_{kki}\\ a_{iij}&a_{jji}&0\\ \end{array}\right).$
+The remaining part of the proof is analogous to that of Theorem 3.2, and thus is omitted here.
+Q.E.D.
+The following theorem establishes the rigidity of the discrete conformal structure (A3), which corresponds to the rigidity part of Theorem 1.8 (iii).
+As its proof is similar to that of Theorem 3.4, it is omitted here.
 Theorem 5.3.
-Suppose $(\Sigma,\mathcal{T},\eta)$ is a weighted triangulated surface with boundary and the weight $\eta\in\mathbb{R}_{>0}^{E}$.
-For the discrete conformal structure (A2), the function $f$ is determined by its generalized combinatorial curvature $K\in\mathbb{R}^{E}_{>0}$.
-In particular, the map from $f$ to $K$ is a smooth embedding.
-5. Rigidity and existence of the mixed discrete conformal structure I
-In this section, we consider the rigidity and existence for the mixed discrete conformal structure $\mathrm{I}$, i.e., the discrete hyperbolic metrics $l_{ij},l_{jk},l_{ki}$ of a right-angled hyperbolic hexagon $\{ijk\}\in F$ are induced by (7) and (10).
-By Remark 1.7, we can set $C\equiv 0$ for (7) and (10).
-As mentioned in Remark 1.8, if $d_{rs}<0$, then $l_{rs},l_{rt}$ are given by (10) and $l_{st}$ is given by (7), where $\{r,s,t\}=\{i,j,k\}$.
-Without loss of generality, we can assume $d_{ij}<0$ and $d_{ik}<0$.
-Then for a right-angled hyperbolic hexagon $\{ijk\}\in F$, the edge lengths $l_{ij},l_{ik}$ are given by (10) and the edge length $l_{jk}$ is given by (7), i.e.,
-(43) $\cosh l_{ij}$ $=\cosh(f_{j}-f_{i})+\zeta_{ij}e^{f_{i}+f_{j}},$
-$\cosh l_{ik}$ $=\cosh(f_{k}-f_{i})+\zeta_{ik}e^{f_{i}+f_{k}},$
+Suppose $(\Sigma,\mathcal{T},\eta)$ is a weighted triangulated surface with boundary, where $\eta\in\mathbb{R}_{>0}^{E}$.
+The discrete conformal factor $f$ is uniquely determined by its generalized combinatorial curvature $K\in\mathbb{R}^{N}_{>0}$.
+6. Rigidity and existence of the mixed discrete conformal structure III
+In this section, we investigate the rigidity and existence of the mixed discrete conformal structure III, specifically, an ideally triangulated surface with boundary admitting both discrete conformal structures (A3) and (B3).
+All right-angled hyperbolic hexagons are either Type-III or a combination of Type-III right-angled hyperbolic hexagons and those with edge lengths given by (7).
+Here, a right-angled hyperbolic hexagon is Type-III if its edge lengths are given by both (7) and (10).
+By Remark 1.6, we set $C\equiv 0$ for both (7) and (10).
+As noted in Remark 1.7, if $d_{rs}<0$, then $d_{rt}<0,\ d_{sr}>0,\ d_{tr}>0,\ d_{st}>0,\ d_{ts}>0$, and $l_{rs},l_{rt}$ are given by (10) and $l_{st}$ is given by (7), where $\{r,s,t\}=\{i,j,k\}$.
+Without loss of generality, we assume $d_{ij}<0$ in this section.
+This implies $d_{ik}<0$, $d_{ji}>0$, $d_{ki}>0$, $d_{jk}>0$, and $d_{kj}>0$.
+For the Type-III right-angled hyperbolic hexagon $\{ijk\}\in F$, the edge lengths $l_{ij},l_{ik}$ are given by (10), and $l_{jk}$ is given by (7), namely,
+(48) $\cosh l_{ij}$ $=\cosh(f_{j}-f_{i})+\eta_{ij}e^{f_{i}+f_{j}},$
+$\cosh l_{ik}$ $=\cosh(f_{k}-f_{i})+\eta_{ik}e^{f_{i}+f_{k}},$
 $\cosh l_{jk}$ $=-\cosh(f_{j}-f_{k})+\eta_{jk}e^{f_{j}+f_{k}}.$
-Here $\eta_{ij}\in\mathbb{R}_{>0}$ and $\zeta_{ij},\zeta_{ik}\in\mathbb{R}$.
-5.1. Rigidity of the mixed discrete conformal structure $\mathrm{I}$
-The admissible space of the discrete conformal factor $f$ for a right-angled hyperbolic hexagon $\{ijk\}\in F$ with edge lengths given by (43) is defined to be
-$\Omega_{ijk}(\zeta_{ij},\eta_{jk},\zeta_{ik})=\{(f_{i},f_{j},f_{k})\in\mathbb{R}^{3}|l_{ij}>0,\ l_{jk}>0,\ l_{ik}>0\}.$
-Combining (30) and Remark 2.23 gives $u_{i}=e^{-f_{i}}\in\mathbb{R}_{>0}$, $u_{j}=-e^{-f_{j}}\in\mathbb{R}_{<0}$ and $u_{k}=-e^{-f_{k}}\in\mathbb{R}_{<0}$. Thus
+Here $\eta_{ij}\in\mathbb{R}$, $\eta_{ik}\in\mathbb{R}$, and $\eta_{jk}\in\mathbb{R}_{>0}$.
+6.1. Rigidity of the mixed discrete conformal structure III
+For a right-angled hyperbolic hexagon $\{ijk\}\in F$ with edge lengths given by (48), the admissible space of the discrete conformal factor $f$ is defined as
+$\Omega_{ijk}(\eta)=\{(f_{i},f_{j},f_{k})\in\mathbb{R}^{3}\mid l_{ij}>0,\ l_{jk}>0,\ l_{ik}>0\}.$
+Combining (29) and Remark 2.22 yields $u_{i}=e^{-f_{i}}\in\mathbb{R}_{>0}$, $u_{j}=-e^{-f_{j}}\in\mathbb{R}_{<0}$, and $u_{k}=-e^{-f_{k}}\in\mathbb{R}_{<0}$. By (48), we have
 $\eta_{jk}=\frac{\cosh l_{jk}+\cosh(f_{j}-f_{k})}{e^{f_{j}+f_{k}}}>\frac{1+\frac{1}{2}(e^{f_{j}-f_{k}}+e^{f_{k}-f_{j}})}{e^{f_{j}+f_{k}}}=\frac{1}{2}(e^{-f_{j}}+e^{-f_{k}})^{2}=\frac{1}{2}(u_{j}+u_{k})^{2}.$
-Since $u_{j},u_{k}\in\mathbb{R}_{<0}$, then $\eta_{jk}>0$ and $u_{j}+u_{k}>-\sqrt{2\eta_{jk}}$. Similarly,
-$\zeta_{ij}=\frac{\cosh l_{ij}-\cosh(f_{j}-f_{i})}{e^{f_{i}+f_{j}}}>\frac{1-\frac{1}{2}(e^{f_{j}-f_{i}}+e^{f_{i}-f_{j}})}{e^{f_{i}+f_{j}}}=-\frac{1}{2}(e^{-f_{i}}-e^{-f_{j}})^{2}=-\frac{1}{2}(u_{i}+u_{j})^{2},$
-which implies $(u_{i}+u_{j})^{2}>-2\zeta_{ij}$.
-Moreover, if $\zeta_{ij}>0$, then $u_{i}+u_{j}\in\mathbb{R}$; if $\zeta_{ij}<0$, then $u_{i}+u_{j}>\sqrt{-2\zeta_{ij}}$ or $u_{i}+u_{j}<-\sqrt{-2\zeta_{ij}}$.
-To remove the solution $u_{i}+u_{j}<-\sqrt{-2\zeta_{ij}}$, we add the condition that $\eta_{jk}+\zeta_{ij}\leq 0$.
-In fact, $u_{i}+u_{j}<-\sqrt{-2\zeta_{ij}}\leq-\sqrt{2\eta_{jk}}<u_{j}+u_{k}$, which contradicts with the fact $u_{i}>0$ and $u_{k}<0$.
-Similar arguments are also suitable for $\zeta_{ik}$.
-Hence, we have the following theorem, which characterizes the admissible space of $u$.
+Since $u_{j}\in\mathbb{R}_{<0}$ and $u_{k}\in\mathbb{R}_{<0}$, it follows that $\eta_{jk}>0$ and $u_{j}+u_{k}>-\sqrt{2\eta_{jk}}$. Similarly,
+$\eta_{ij}=\frac{\cosh l_{ij}-\cosh(f_{j}-f_{i})}{e^{f_{i}+f_{j}}}>\frac{1-\frac{1}{2}(e^{f_{j}-f_{i}}+e^{f_{i}-f_{j}})}{e^{f_{i}+f_{j}}}=-\frac{1}{2}(e^{-f_{i}}-e^{-f_{j}})^{2}=-\frac{1}{2}(u_{i}+u_{j})^{2},$
+which implies $(u_{i}+u_{j})^{2}>-2\eta_{ij}$.
+Moreover, if $\eta_{ij}>0$, then $u_{i}+u_{j}\in\mathbb{R}$; if $\eta_{ij}\leq 0$, then $u_{i}+u_{j}>\sqrt{-2\eta_{ij}}$ or $u_{i}+u_{j}<-\sqrt{-2\eta_{ij}}$.
+To exclude the solution $u_{i}+u_{j}<-\sqrt{-2\eta_{ij}}$, we add the condition that $\eta_{jk}+\eta_{ij}\leq 0$.
+In fact, $u_{i}+u_{j}<-\sqrt{-2\eta_{ij}}\leq-\sqrt{2\eta_{jk}}<u_{j}+u_{k}$, which contradicts the fact $u_{i}\in\mathbb{R}_{>0}$ and $u_{k}\in\mathbb{R}_{<0}$.
+Similar arguments also apply for $\eta_{ik}$.
+Consequently, we obtain the following theorem, which characterizes the admissible space of $u$.
 Theorem 6.1.
-Suppose $(\Sigma,\mathcal{T},\zeta,\eta)$ is a weighted triangulated surface with boundary, and the weights $\zeta\in\mathbb{R}^{E}$ and $\eta\in\mathbb{R}_{>0}^{E}$.
-Let $\{ijk\}\in F$ be a right-angled hyperbolic hexagon with edge lengths given by (43).
-If one of the following conditions is satisfied
+Let $\{ijk\}\in F$ be a right-angled hyperbolic hexagon with edge lengths given by (48). If one of the following conditions is satisfied
 (i):
-$\zeta_{ij}>0$ and $\zeta_{ik}>0$;
+$\eta_{jk}>0$, $\eta_{ij}>0$, and $\eta_{ik}>0$;
 (ii):
-$\zeta_{ij}>0$, $\zeta_{ik}<0$ and $\eta_{jk}+\zeta_{ik}\leq 0$;
+$\eta_{jk}>0$, $\eta_{ij}>0$, $\eta_{ik}\leq 0$, and $\eta_{jk}+\eta_{ik}\leq 0$;
 (iii):
-$\zeta_{ij}<0$, $\zeta_{ik}>0$ and $\eta_{jk}+\zeta_{ij}\leq 0$;
+$\eta_{jk}>0$, $\eta_{ij}\leq 0$, $\eta_{ik}>0$, and $\eta_{jk}+\eta_{ij}\leq 0$;
 (iv):
-$\zeta_{ij}<0$, $\zeta_{ik}<0$, $\eta_{jk}+\zeta_{ij}\leq 0$ and $\eta_{jk}+\zeta_{ij}\leq 0$.
+$\eta_{jk}>0$, $\eta_{ij}\leq 0$, $\eta_{ik}\leq 0$, $\eta_{jk}+\eta_{ij}\leq 0$, and $\eta_{jk}+\eta_{ij}\leq 0$.
 Then the admissible space
-(44) $\mathcal{U}_{ijk}(\zeta_{ij},\eta_{jk},\zeta_{ik})=$ $u(\Omega_{ijk}(\zeta_{ij},\eta_{jk},\zeta_{ik}))$
-$=$ $\{(u_{i},u_{j},u_{k})\in\mathbb{R}_{>0}\times\mathbb{R}_{<0}^{2}|u_{j}+u_{k}>-\sqrt{2\eta_{jk}},$
-$\ \ (u_{i}+u_{j})^{2}>-2\zeta_{ij},(u_{i}+u_{k})^{2}>-2\zeta_{ik}\}$
-of $u$ is a convex polytope. As a result, the admissible space
-$\mathcal{U}(\zeta,\eta)=\bigcap_{\{ijk\}\in F}\mathcal{U}_{ijk}(\zeta_{ij},\eta_{jk},\zeta_{ik})$
-is also a convex polytope.
+(49) $\mathcal{U}_{ijk}(\eta)$ $=u(\Omega_{ijk}(\eta))$
+=$=\{(u_{i},u_{j},u_{k})\in\mathbb{R}_{>0}\times\mathbb{R}_{<0}^{2}\mid u_{j}+u_{k}>-\sqrt{2\eta_{jk}},$
+$\ \ \ \ \ \ \ \ (u_{i}+u_{j})^{2}>-2\eta_{ij},(u_{i}+u_{k})^{2}>-2\eta_{ik}\}$
+of $u$ is a convex polytope.
 Proof.
-According to the arguments above Theorem 5.1, we can describe the admissible space $\mathcal{U}_{ijk}(\zeta_{ij},\eta_{jk},\zeta_{ik})$ explicitly.
-For the case $\mathrm{(i)}$, the admissible space (44) is reduced to
-$\mathcal{U}_{ijk}(\zeta_{ij},\eta_{jk},\zeta_{ik})=\{(u_{i},u_{j},u_{k})\in\mathbb{R}_{>0}\times\mathbb{R}_{<0}^{2}|u_{j}+u_{k}>-\sqrt{2\eta_{jk}}\}.$
-For the case $\mathrm{(ii)}$, the admissible space (44) is reduced to
-$\mathcal{U}_{ijk}(\zeta_{ij},\eta_{jk},\zeta_{ik})=\{(u_{i},u_{j},u_{k})\in\mathbb{R}_{>0}\times\mathbb{R}_{<0}^{2}|u_{j}+u_{k}>-\sqrt{2\eta_{jk}},u_{i}+u_{k}>\sqrt{-2\zeta_{ik}}\}.$
-For the case $\mathrm{(iii)}$, the admissible space (44) is reduced to
-$\mathcal{U}_{ijk}(\zeta_{ij},\eta_{jk},\zeta_{ik})=\{(u_{i},u_{j},u_{k})\in\mathbb{R}_{>0}\times\mathbb{R}_{<0}^{2}|u_{j}+u_{k}>-\sqrt{2\eta_{jk}},u_{i}+u_{j}>\sqrt{-2\zeta_{ij}}\}.$
-For the case $\mathrm{(iv)}$, the admissible space (44) is reduced to
-$\mathcal{U}_{ijk}(\zeta_{ij},\eta_{jk},\zeta_{ik})=\{(u_{i},u_{j},u_{k})\in\mathbb{R}_{>0}\times\mathbb{R}_{<0}^{2}|$ $u_{j}+u_{k}>-\sqrt{2\eta_{jk}},u_{i}+u_{j}>\sqrt{-2\zeta_{ij}},$
-$u_{i}+u_{k}>\sqrt{-2\zeta_{ik}}\}.$
-Each of these four cases shows that $\mathcal{U}_{ijk}(\zeta_{ij},\eta_{jk},\zeta_{ik})$ is a convex polytope. Q.E.D.
+According to the arguments above Theorem 6.1, we can explicitly characterize the admissible space $\mathcal{U}_{ijk}(\eta)$. For the case (i), the admissible space (49) reduces to
+$\mathcal{U}_{ijk}(\eta)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}_{>0}\times\mathbb{R}_{<0}^{2}\mid u_{j}+u_{k}>-\sqrt{2\eta_{jk}}\}.$
+For the case (ii), the admissible space (49) reduces to
+$\mathcal{U}_{ijk}(\eta)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}_{>0}\times\mathbb{R}_{<0}^{2}\mid u_{j}+u_{k}>-\sqrt{2\eta_{jk}},u_{i}+u_{k}>\sqrt{-2\eta_{ik}}\}.$
+For the case (iii), the admissible space (49) reduces to
+$\mathcal{U}_{ijk}(\eta)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}_{>0}\times\mathbb{R}_{<0}^{2}\mid u_{j}+u_{k}>-\sqrt{2\eta_{jk}},u_{i}+u_{j}>\sqrt{-2\eta_{ij}}\}.$
+For the case (iv), the admissible space (49) reduces to
+$\mathcal{U}_{ijk}(\eta)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}_{>0}\times\mathbb{R}_{<0}^{2}\mid$ $\ u_{j}+u_{k}>-\sqrt{2\eta_{jk}},u_{i}+u_{j}>\sqrt{-2\eta_{ij}},$
+$\ u_{i}+u_{k}>\sqrt{-2\eta_{ik}}\}.$
+Each of these four cases shows that $\mathcal{U}_{ijk}(\eta)$ is a convex polytope. Q.E.D.
+Corollary 6.2.
+Suppose $(\Sigma,\mathcal{T},\eta)$ is a weighted triangulated surface with boundary. If one of the following conditions holds:
+(1):
+All right-angled hyperbolic hexagons are Type-III (i.e., the edge lengths of any right-angled hyperbolic hexagons are given by (48)), and the weight $\eta$ satisfies one of the conditions (i)–(iv) in Theorem 6.1;
+(2):
+Not all right-angled hyperbolic hexagons are of Type-III (i.e., there exists at least one right-angled hyperbolic hexagon with edge lengths given by (7)).
+In this case, the weight $\eta$ satisfies one of the conditions (i)–(iv) in Theorem 6.1 for those right-angled hyperbolic hexagons with edge lengths given by (48), and $\eta\in\mathbb{R}_{>0}^{E}$ for those with edge lengths given by (7).
+Then the admissible space $\mathcal{U}(\eta)=\bigcap_{\{ijk\}\in F}\mathcal{U}_{ijk}(\eta)$ is a convex polytope on $(\Sigma,\mathcal{T},\eta)$,
+Proof.
+The case (1) follows from Theorem 6.1.
+The case (2) follows from (47) and Theorem 6.1.
+Indeed, by (47), the admissible space $\mathcal{U}_{ijk}(\eta)$ of a right-angled hyperbolic hexagon $\{ijk\}$ with edge lengths given by (7) is a convex polytope.
+Q.E.D.
 Theorem 6.3.
-Under the same assumptions as in Theorem 5.1.
-The Jacobian $\Lambda_{ijk}=\frac{\partial(\theta^{jk}_{i},\theta^{ik}_{j},\theta^{ij}_{k})}{\partial(u_{i},u_{j},u_{k})}$ is non-degenerate, symmetric and negative definite on $\mathcal{U}_{ijk}(\zeta_{ij},\eta_{jk},\zeta_{ik})$.
-As a result, the Jacobian $\Lambda=\frac{\partial(K_{i},...,K_{N})}{\partial(u_{i},...,u_{N})}$ is symmetric and negative definite on $\mathcal{U}(\zeta,\eta)$.
+Under the same assumptions as those in Theorem 6.1, the Jacobian $\Lambda_{ijk}=\frac{\partial(\theta^{jk}_{i},\theta^{ik}_{j},\theta^{ij}_{k})}{\partial(u_{i},u_{j},u_{k})}$ is symmetric and negative definite on $\mathcal{U}_{ijk}(\eta)$.
 Proof.
-The formula (30) implies
-$\frac{\partial(f_{i},f_{j},f_{k})}{\partial(u_{i},u_{j},u_{k})}=\left(\begin{array}[]{ccc}-e^{f_{i}}&0&0\\ 0&e^{f_{j}}&0\\ 0&0&e^{f_{k}}\\ \end{array}\right):=Q_{5}.$
-Then
-$\Lambda_{ijk}=\frac{\partial(\theta_{i},\theta_{j},\theta_{k})}{\partial(u_{i},u_{j},u_{k})}=-\frac{1}{A}\left(\begin{array}[]{ccc}\sinh l_{i}&0&0\\ 0&\sinh l_{j}&0\\ 0&0&\sinh l_{k}\\ \end{array}\right)Q_{1}Q_{2}Q_{5},$
-where $Q_{1}$ is defined by (36) and $Q_{2}$ is defined by (38). Moreover, $\det Q_{1}>0$.
-Since $d_{ij}<0,d_{ik}<0$ and the others are positive for any right-angled hyperbolic hexagon $\{ijk\}\in F$ with edge lengths given by (43), then
+For simplicity, we adopt the notations introduced in Subsection 3.2. Formula (29) implies
+$\frac{\partial(f_{i},f_{j},f_{k})}{\partial(u_{i},u_{j},u_{k})}=\left(\begin{array}[]{ccc}-e^{f_{i}}&0&0\\ 0&e^{f_{j}}&0\\ 0&0&e^{f_{k}}\\ \end{array}\right):=Q_{6}.$
+Then $\Lambda_{ijk}$ can be expressed as
+$\Lambda_{ijk}=\frac{\partial(\theta_{i},\theta_{j},\theta_{k})}{\partial(u_{i},u_{j},u_{k})}=-\frac{1}{A}\left(\begin{array}[]{ccc}\sinh l_{i}&0&0\\ 0&\sinh l_{j}&0\\ 0&0&\sinh l_{k}\\ \end{array}\right)Q_{1}Q_{2}Q_{6},$
+where $Q_{1}$ and $Q_{2}$ are defined by (36) and (38), respectively. Furthermore, $\det Q_{1}>0$.
+For any right-angled hyperbolic hexagon $\{ijk\}\in F$ with edge lengths given by (48), since $d_{ij}<0$, $d_{ik}<0$, $d_{ik}<0$, $d_{ji}>0$, $d_{ki}>0$, $d_{jk}>0$, and $d_{kj}>0$, it follows that
 $\det Q_{2}$ $=\coth d_{jk}\coth d_{ki}\coth d_{ij}+\coth d_{kj}\coth d_{ik}\coth d_{ji}$
 $=\frac{\cosh d_{jk}\cosh d_{ki}\cosh d_{ij}+\cosh d_{kj}\cosh d_{ik}\cosh d_{ji}}{\sinh d_{ij}\sinh d_{jk}\sinh d_{ki}}<0,$
-where (1) is used in the second line. Thus
-(45) $\det\Lambda_{ijk}=-\frac{1}{A^{3}}\sinh l_{i}\sinh l_{j}\sinh l_{k}\cdot\det Q_{1}\cdot\det Q_{2}\cdot\det Q_{5}<0,$
+where (2) is used in the second line. Consequently,
+$\det\Lambda_{ijk}=-\frac{1}{A^{3}}\sinh l_{i}\sinh l_{j}\sinh l_{k}\cdot\det Q_{1}\cdot\det Q_{2}\cdot\det Q_{6}<0,$
 which implies the matrix $\Lambda_{ijk}$ is non-degenerate.
-The symmetry of the matrix $\Lambda_{ijk}$ follows from Lemma 2.22.
+The symmetry of $\Lambda_{ijk}$ follows from Lemma 2.21.
 Set $a_{rrs}=e^{f_{r}}\cdot\coth d_{rs}$.
-Then $a_{iij}<0,\ a_{iik}<0,\ a_{jjk}>0,\ a_{kkj}>0,\ a_{jji}>0,\ a_{kki}>0$.
+Then $a_{iij}<0$, $a_{iik}<0$, $a_{jjk}>0$, $a_{kkj}>0$, $a_{jji}>0$, and $a_{kki}>0$.
 Define
-$Q:=Q_{1}Q_{2}Q_{5}=\left(\begin{array}[]{ccc}-1&\cosh\theta_{k}&\cosh\theta_{j}\\ \cosh\theta_{k}&-1&\cosh\theta_{i}\\ \cosh\theta_{j}&\cosh\theta_{i}&-1\\ \end{array}\right)\times\left(\begin{array}[]{ccc}0&a_{jjk}&a_{kkj}\\ -a_{iik}&0&a_{kki}\\ -a_{iij}&a_{jji}&0\\ \end{array}\right).$
-By similar arguments in the proof of Theorem 3.2, we will show that the determinants of the $1\times 1$ and $2\times 2$ order principal minor of $Q$ are positive.
-Let $m_{rs}$ be the entry of $Q$ at $r$-th row and $s$-th column.
+$Q:=Q_{1}Q_{2}Q_{6}=\left(\begin{array}[]{ccc}-1&\cosh\theta_{k}&\cosh\theta_{j}\\ \cosh\theta_{k}&-1&\cosh\theta_{i}\\ \cosh\theta_{j}&\cosh\theta_{i}&-1\\ \end{array}\right)\times\left(\begin{array}[]{ccc}0&a_{jjk}&a_{kkj}\\ -a_{iik}&0&a_{kki}\\ -a_{iij}&a_{jji}&0\\ \end{array}\right).$
+By arguments analogous to those in the proof of Theorem 3.2, we show that the $1\times 1$ and $2\times 2$ leading principal minors of $Q$ are positive.
+Let $m_{rs}$ denote the entry of $Q$ at $r$-th row and $s$-th column.
 Then
@@
 $m_{21}=a_{iik}-\cosh\theta_{i}\cdot a_{iij}.$
-Then the determinants of the $1\times 1$ principal submatrices of $Q$ is $m_{11}>0$ by $a_{iik}<0,\ a_{iij}<0$. And the determinants of the $2\times 2$ principal submatrices of $Q$ is
+Thus the $1\times 1$ leading principal minor of $Q$ is $m_{11}>0$ due to $a_{iik}<0$ and $a_{iij}<0$. And the $2\times 2$ leading principal minor of $Q$ is
 $m_{11}m_{22}-m_{12}m_{21}=$ $-[\sinh^{2}\theta_{k}a_{iik}a_{jjk}+(\cosh\theta_{i}\cosh\theta_{k}+\cosh\theta_{j})a_{iik}a_{jji}$
@@
 This completes the proof. Q.E.D.
-The following theorem gives the rigidity of the mixed discrete conformal structure $\mathrm{I}$, which generalizes the case $\mathrm{(iii)}$ in Theorem 1.9.
-As the proof is similar to that of Theorem 3.4, we omit it here.
+Corollary 6.4.
+Under the same assumptions as those in Corollary 6.2, the Jacobian $\Lambda=\frac{\partial(K_{i},...,K_{N})}{\partial(u_{i},...,u_{N})}$ is symmetric and negative definite on $\mathcal{U}(\eta)$.
+Proof.
+The case (1) follows directly from Theorem 6.3.
+The proof of the case (2) is analogous to that of Corollary 3.3, and thus we omit the details here.
+Note that $\Lambda=\sum_{\{ijk\}\in F}\Lambda_{ijk}$, where each $\Lambda_{ijk}$ is derived from either Theorem 5.2 or Theorem 6.3.
+Q.E.D.
+The following theorem establishes the rigidity of the mixed discrete conformal structure III, which generalizes the rigidity part of Theorem 1.8 (vi).
+As its proof is analogous to that of Theorem 3.4, it is omitted here.
 Theorem 6.5.
-Under the same assumptions as in Theorem 5.1.
-The function $f$ is determined by its generalized combinatorial curvature $K\in\mathbb{R}_{>0}^{E}$.
-In particular, the map from $f$ to $K$ is a smooth embedding.
-5.2. Existence of the mixed discrete conformal structure $\mathrm{I}$
-This proof is similar to that of Theorem 3.5 for the discrete conformal structure (A1).
+Under the same assumptions as those in Corollary 6.2, the discrete conformal factor $f$ is uniquely determined by its generalized combinatorial curvature $K\in\mathbb{R}_{>0}^{N}$.
+6.2. Existence of the mixed discrete conformal structure III
+The following theorem characterizes the image of $K$ for the mixed discrete conformal structure III, which generalizes the existence part of Theorem 1.8 (vi).
 Theorem 6.6.
-Under the same assumptions as in Theorem 5.1. The image of $K$ is >$\mathbb{R}^{N}_{>0}$.
-To prove Theorem 5.4, we need the following two lemmas.
+Under the same assumptions as Corollary 6.2, the image of $K$ is >$\mathbb{R}^{N}_{>0}$.
+To prove Theorem 6.6, we need the following three lemmas.
 Lemma 6.7.
-Under the same assumptions as in Theorem 5.1. For any constants $c_{1},...,c_{6}\in\mathbb{R}$, if one of the following conditions is satisfied
+([5], Lemma 4.6) Let $\{ijk\}\in F$ be a right-angled hyperbolic hexagon with edge lengths given by (7).
+Then $\lim_{f_{j}\rightarrow+\infty}\theta^{ik}_{j}(f_{i},f_{j},f_{k})=0$, and the converge is uniform.
+Lemma 6.8.
+Under the same assumptions as those in Theorem 6.1, if one of the following conditions is satisfied
 (i):
@@
 (vi):
-$\lim f_{i}=-\infty,\ \lim f_{j}=c_{5},\ \lim f_{k}=c_{6}$.
-Then $\lim\theta^{jk}_{i}(f_{i},f_{j},f_{k})=0$.
+$\lim f_{i}=-\infty,\ \lim f_{j}=c_{5},\ \lim f_{k}=c_{6}$,
+where $c_{1},...,c_{6}\in\mathbb{R}$ are constants. Then $\lim\theta^{jk}_{i}(f_{i},f_{j},f_{k})=0$.
 Proof.
-By (43), we have
-$\cosh l_{ij}$ $=e^{f_{i}+f_{j}}(\zeta_{ij}+\frac{1}{2}e^{-2f_{i}}+\frac{1}{2}e^{-2f_{j}}),$
-$\cosh l_{ik}$ $=e^{f_{i}+f_{k}}(\zeta_{ik}+\frac{1}{2}e^{-2f_{i}}+\frac{1}{2}e^{-2f_{k}}),$
+For the case (i), by (48), we derive
+(50) $\cosh l_{ij}$ $=e^{f_{i}+f_{j}}(\eta_{ij}+\frac{1}{2}e^{-2f_{i}}+\frac{1}{2}e^{-2f_{j}}),$
+$\cosh l_{ik}$ $=e^{f_{i}+f_{k}}(\eta_{ik}+\frac{1}{2}e^{-2f_{i}}+\frac{1}{2}e^{-2f_{k}}),$
 $\cosh l_{jk}$ $=e^{f_{j}+f_{k}}(\eta_{jk}-\frac{1}{2}e^{-2f_{j}}-\frac{1}{2}e^{-2f_{k}}).$
-For the case $\mathrm{(i)}$, if $\lim f_{i}=+\infty,\ \lim f_{j}=+\infty,\ \lim f_{k}=+\infty$, then $\lim\cosh l_{ij}:=\lim c_{k}e^{f_{i}+f_{j}}$$\lim\cosh l_{jk}:=\lim c_{i}e^{f_{j}+f_{k}}$ and $\lim\cosh l_{ik}:=\lim c_{j}e^{f_{i}+f_{k}}$, where $c_{i},c_{j},c_{k}$ are positive constants.
-By the cosine law,
+Since $\lim f_{i}=+\infty$, $\lim f_{j}=+\infty$, and $\lim f_{k}=+\infty$, it follows that $\lim\cosh l_{ij}:=\lim c_{k}e^{f_{i}+f_{j}}$, $\lim\cosh l_{jk}:=\lim c_{i}e^{f_{j}+f_{k}}$, and $\lim\cosh l_{ik}:=\lim c_{j}e^{f_{i}+f_{k}}$, where $c_{i},c_{j},c_{k}$ are positive constants.
+By the cosine law, we have
 $\cosh\theta^{jk}_{i}=\frac{\cosh l_{jk}+\cosh l_{ij}\cosh l_{ik}}{\sinh l_{ij}\sinh l_{ik}}\rightarrow\frac{c_{i}e^{f_{j}+f_{k}}}{c_{k}e^{f_{i}+f_{j}}\cdot c_{j}e^{f_{i}+f_{k}}}+1\rightarrow 1,$
-which implies $\theta^{jk}_{i}\rightarrow 0$. The proofs of the cases $\mathrm{(ii)}$ and $\mathrm{(iii)}$ are similar, so we omit them here.
-For the case $\mathrm{(iv)}$, if $\lim f_{i}=-\infty,\ \lim f_{j}=+\infty,\ \lim f_{k}=+\infty$. By (43), we have
-$\cosh l_{ij}=e^{f_{j}-f_{i}}(\zeta_{ij}e^{2f_{i}}+\frac{1}{2}e^{2f_{i}-2f_{j}}+\frac{1}{2}),$
-$\cosh l_{ik}=e^{f_{k}-f_{i}}(\zeta_{ik}e^{2f_{i}}+\frac{1}{2}e^{2f_{i}-2f_{k}}+\frac{1}{2}),$
+which implies $\theta^{jk}_{i}\rightarrow 0$. The proofs for the cases (ii) and (iii) follow analogously and are omitted here.
+For the case (iv), by (48), we derive
+(51) $\cosh l_{ij}=e^{f_{j}-f_{i}}(\eta_{ij}e^{2f_{i}}+\frac{1}{2}e^{2f_{i}-2f_{j}}+\frac{1}{2}),$
+$\cosh l_{ik}=e^{f_{k}-f_{i}}(\eta_{ik}e^{2f_{i}}+\frac{1}{2}e^{2f_{i}-2f_{k}}+\frac{1}{2}),$
 $\cosh l_{jk}=e^{f_{j}+f_{k}}(\eta_{jk}-\frac{1}{2}e^{-2f_{j}}-\frac{1}{2}e^{-2f_{k}}).$
-Then $\lim\cosh l_{ij}:=\lim c_{k}e^{f_{j}-f_{i}},\ \lim\cosh l_{ik}:=\lim c_{j}e^{f_{k}-f_{i}}$ and $\lim\cosh l_{jk}:=\lim c_{i}e^{f_{j}+f_{k}}$, where $c_{i},\ c_{j},\ c_{k}$ are positive constants.
-By the cosine law,
+Since $\lim f_{i}=-\infty$, $\lim f_{j}=+\infty$, and $\lim f_{k}=+\infty$, it follows that $\lim\cosh l_{ij}:=\lim c_{k}e^{f_{j}-f_{i}}$, $\lim\cosh l_{ik}:=\lim c_{j}e^{f_{k}-f_{i}}$, and $\lim\cosh l_{jk}:=\lim c_{i}e^{f_{j}+f_{k}}$, where $c_{i},c_{j},c_{k}$ are positive constants.
+By the cosine law, we have
 $\cosh\theta^{jk}_{i}=\frac{\cosh l_{jk}+\cosh l_{ij}\cosh l_{ik}}{\sinh l_{ij}\sinh l_{ik}}\rightarrow\frac{c_{i}e^{f_{j}+f_{k}}}{c_{k}e^{f_{j}-f_{i}}\cdot c_{j}e^{f_{k}-f_{i}}}+1\rightarrow 1,$
-which implies $\theta^{jk}_{i}\rightarrow 0$. The same proofs are suitable for the cases $\mathrm{(v)}$ and $\mathrm{(vi)}$, and we omit them here. Q.E.D.
+which implies $\theta^{jk}_{i}\rightarrow 0$. The proofs for the cases (ii) and (iii) follow analogously and are omitted here. Q.E.D.
 Lemma 6.9.
-Under the same assumptions as in Theorem 5.1. For any constants $c_{1},...,c_{5}\in\mathbb{R}$, if one of the following conditions is satisfied
+Under the same assumptions as those in Theorem 6.1, if one of the following conditions is satisfied
 (i):
@@
 (vi):
-$\lim f_{j}=+\infty,\ \lim f_{i}=c_{4},\ \lim f_{k}=c_{5}$.
-Then $\lim\theta^{ik}_{j}(f_{i},f_{j},f_{k})=0$.
+$\lim f_{j}=+\infty,\ \lim f_{i}=c_{4},\ \lim f_{k}=c_{5}$,
+where $c_{1},...,c_{5}\in\mathbb{R}$ are constants. Then $\lim\theta^{ik}_{j}(f_{i},f_{j},f_{k})=0$.
 Proof.
-For the case $\mathrm{(i)}$, if $\lim f_{j}=+\infty,\ \lim f_{i}=+\infty,\ \lim f_{k}=+\infty$, then $\lim\cosh l_{ij}:=\lim c_{k}e^{f_{i}+f_{j}}$$\lim\cosh l_{jk}:=\lim c_{i}e^{f_{j}+f_{k}}$ and $\lim\cosh l_{ik}:=\lim c_{j}e^{f_{i}+f_{k}}$, where $c_{i},c_{j},c_{k}$ are positive constants.
-By the cosine law,
+For the case (i), it follows from (50) that $\lim\cosh l_{ij}:=\lim c_{k}e^{f_{i}+f_{j}}$$\lim\cosh l_{jk}:=\lim c_{i}e^{f_{j}+f_{k}}$, and $\lim\cosh l_{ik}:=\lim c_{j}e^{f_{i}+f_{k}}$, where $c_{i},c_{j},c_{k}$ are positive constants.
+By the cosine law, we have
 $\cosh\theta^{ik}_{j}=\frac{\cosh l_{ik}+\cosh l_{ij}\cosh l_{jk}}{\sinh l_{ij}\sinh l_{jk}}\rightarrow\frac{c_{j}e^{f_{i}+f_{k}}}{c_{k}e^{f_{i}+f_{j}}\cdot c_{i}e^{f_{i}+f_{j}}}+1\rightarrow 1,$
-which implies $\theta^{ik}_{j}\rightarrow 0$. The same proofs are suitable for the cases $\mathrm{(ii)}$, $\mathrm{(v)}$ and $\mathrm{(vi)}$, and we omit them.
-For the case $\mathrm{(iii)}$, if $\lim f_{j}=+\infty,\ \lim f_{i}=-\infty,\ \lim f_{k}=+\infty$, then $\lim\cosh l_{ij}:=\lim c_{k}e^{f_{j}-f_{i}}$$\lim\cosh l_{ik}:=\lim c_{j}e^{f_{k}-f_{i}}$ and $\lim\cosh l_{jk}:=\lim c_{i}e^{f_{j}+f_{k}}$, where $c_{i},\ c_{j},\ c_{k}$ are positive constants.
-By the cosine law,
+which implies $\theta^{ik}_{j}\rightarrow 0$. The proofs for the cases (ii), (v) and (vi) follow analogously and are omitted here.
+For the case (iii), it follows from (51) that $\lim\cosh l_{ij}:=\lim c_{k}e^{f_{j}-f_{i}}$$\lim\cosh l_{ik}:=\lim c_{j}e^{f_{k}-f_{i}}$, and $\lim\cosh l_{jk}:=\lim c_{i}e^{f_{j}+f_{k}}$, where $c_{i},c_{j},c_{k}$ are positive constants.
+By the cosine law, we have
 $\cosh\theta^{ik}_{j}=\frac{\cosh l_{ik}+\cosh l_{ij}\cosh l_{jk}}{\sinh l_{ij}\sinh l_{jk}}\rightarrow\frac{c_{j}e^{f_{k}-f_{i}}}{c_{k}e^{f_{j}-f_{i}}\cdot c_{i}e^{f_{j}+f_{k}}}+1\rightarrow 1,$
-which implies $\theta^{ik}_{j}\rightarrow 0$. The same proof is suitable for the case $\mathrm{(iv)}$, so we omit it. Q.E.D.
-Proof of Theorem 5.4: Let the image of $K$ is $X$.
-The openness of $X$ in >$\mathbb{R}^{N}_{>0}$ follows from the injectivity of the map $f$.
-To prove the closeness of $X$ in >$\mathbb{R}^{N}_{>0}$, we prove that if a sequence $f^{(m)}$ in $\Omega(\zeta,\eta)=\cap_{\{ijk\}\in F}\Omega_{ijk}(\zeta_{ij},\eta_{jk},\zeta_{ik})$ satisfies $\lim_{m\rightarrow+\infty}K^{(m)}\in\mathbb{R}^{N}_{>0}$, then there exists a subsequence, say $f^{(m)}$, such that $\lim_{m\rightarrow+\infty}f^{(m)}$ is in $\Omega(\zeta,\eta)$.
-Suppose otherwise, there exists a subsequence $f^{(m)}$, such that $\lim_{m\rightarrow+\infty}f^{(m)}$ is in the boundary of $\Omega(\zeta,\eta)$.
-The boundary of the admissible space $\Omega(\zeta,\eta)$ in $[-\infty,+\infty]^{N}$ consists of the following three parts:
+which implies $\theta^{ik}_{j}\rightarrow 0$. The proof for the case (iv) follow analogously and is omitted here.
+Q.E.D.
+Proof of Theorem 6.6: Let $X$ be the image of $K$.
+The openness of $X$ in >$\mathbb{R}^{N}_{>0}$ follows directly from the smooth injectivity of the map.
+To establish the closedness of $X$ in >$\mathbb{R}^{N}_{>0}$, we prove that every sequence $\{f^{(m)}\}\in\Omega(\eta)=\bigcap_{\{ijk\}\in F}\Omega_{ijk}(\eta)$ with $\lim_{m\rightarrow+\infty}K^{(m)}\in\mathbb{R}^{N}_{>0}$ admits a subsequence, still denoted by $\{f^{(m)}\}$, such that $\lim_{m\rightarrow+\infty}f^{(m)}\in\Omega(\eta)$.
+Assume, for the sake of contradiction, that there exists a subsequence $\{f^{(m)}\}$ such that $\lim_{m\rightarrow+\infty}f^{(m)}$ lies on the boundary of $\Omega(\eta)$.
+The boundary of the admissible space $\Omega(\eta)$ in $[-\infty,+\infty]^{N}$ consists of the following three parts:
 (i):
-There exists an ideal edge $\{rs\}\in E$ such that $l_{rs}=0$.
+There exists an ideal edge $\{rs\}\in E$ such that $l_{rs}=0$;
 (ii):
-There exists $i\in B$ in a right-angled hyperbolic hexagon $\{ijk\}\in F$ with edge lengths given by (43), such that $f_{i}^{(m)}\rightarrow\pm\infty$.
+There exists $i\in B$ in a right-angled hyperbolic hexagon $\{ijk\}\in F$ with edge lengths given by (48), such that $f_{i}^{(m)}\rightarrow\pm\infty$;
 (iii):
-There exists $j\in B$ in a right-angled hyperbolic hexagon $\{ijk\}\in F$ with edge lengths given by (43), such that $f_{j}^{(m)}\rightarrow+\infty$.
-For case $\mathrm{(i)}$, in a right-angled hyperbolic hexagon $\{ijk\}\in F$ with lengths $l_{ij},l_{jk},l_{ik}$ and opposite lengths of boundaries $\theta^{ij}_{k},\theta^{jk}_{i},\theta^{ik}_{j}$.
-Without loss of generality, we assume $l_{ij}\rightarrow 0$.
-By the cosine law,
-$\cosh\theta^{ik}_{j}=\frac{\cosh l_{ik}+\cosh l_{ij}\cosh l_{jk}}{\sinh l_{ij}\sinh l_{jk}}>\frac{\cosh l_{ij}\cosh l_{jk}}{\sinh l_{ij}\sinh l_{jk}}\geq\frac{\cosh l_{ij}}{\sinh l_{ij}}\rightarrow+\infty,$
-which contradicts the assumption that $\lim_{m\rightarrow+\infty}K^{(m)}\in\mathbb{R}^{N}_{>0}$.
-The case $\mathrm{(ii)}$ follows from Lemma 5.5 and the case $\mathrm{(iii)}$ follows from Lemma 5.6, which both contradict the assumption that $\lim_{m\rightarrow+\infty}K^{(m)}\in\mathbb{R}^{N}_{>0}$.
+There exists $j\in B$ in a right-angled hyperbolic hexagon $\{ijk\}\in F$ with edge lengths given by (48) or (7), such that $f_{j}^{(m)}\rightarrow+\infty$.
+The proof for the case (i) is identical to that of Theorem 3.5 and is thus omitted.
+For the case (ii), by Lemma 6.8, each generalized angle incident to the boundary component $i\in B$ converges to 0.
+Consequently, $\lim_{m\rightarrow+\infty}K_{i}^{(m)}=0$, which contradicts the assumption that $\lim_{m\rightarrow+\infty}K^{(m)}\in\mathbb{R}^{N}_{>0}$.
+For the case (iii), if all right-angled hyperbolic hexagons are Type-III, then Lemma 6.9 implies $\lim_{m\rightarrow+\infty}K_{j}^{(m)}=0$; if not all right-angled hyperbolic hexagons are Type-III, then by Lemma 6.7 and Lemma 6.9, we have $\lim_{m\rightarrow+\infty}K_{j}^{(m)}=0$.
+Both scenarios contradict the assumption that $\lim_{m\rightarrow+\infty}K^{(m)}\in\mathbb{R}^{N}_{>0}$.
 Q.E.D.
@@
 Note that in a right-angled hyperbolic hexagon $\{ijk\}\in F$ with edge lengths given by (48), there is no $j\in B$ such that $f_{j}^{(m)}\rightarrow-\infty$.
-Otherwise, $u_{j}=-e^{-f_{j}}\rightarrow-\infty$.
-According to the admissible space $\mathcal{U}_{ijk}(\zeta_{ij},\eta_{jk},\zeta_{ik})$ defined by (44), the equation $u_{j}+u_{k}>-2\sqrt{\eta_{jk}}$ implies $u_{k}\rightarrow+\infty$, which contradicts with the fact that $u_{k}\in\mathbb{R}_{<0}$.
-6. Rigidity and existence of the mixed discrete conformal structure II
-In this section, we consider the rigidity and the image of $K$ for the mixed discrete conformal structure $\mathrm{II}$, i.e., the discrete hyperbolic metrics $l_{ij},l_{jk},l_{ki}$ of a right-angled hyperbolic hexagon $\{ijk\}\in F$ are induced by (5) and (9).
-As mentioned in Remark 1.8, if $d_{rs}<0$, then $l_{rs},l_{rt}$ are given by (9) and $l_{st}$ is given by (5), where $\{r,s,t\}=\{i,j,k\}$.
-Note that $\eta\in\mathbb{R}_{>0}^{E}$ by (5) and $\zeta\in\mathbb{R}^{E}$ by (9).
-To simplify the notation, we use a new parameter $\xi$ to replace $\eta$ and $\zeta$.
-Without loss of generality, we can always assume $d_{ij}<0$ and $d_{ik}<0$.
-Then for any right-angled hyperbolic hexagon $\{ijk\}\in F$, the edge lengths $l_{ij}$ and $l_{ik}$ are given by (9) and the edge length $l_{jk}$ is given by (5), i.e.,
-(46) $\cosh l_{ij}$ $=\sqrt{(1+\alpha_{i}e^{2f_{i}})(1+\alpha_{j}e^{2f_{j}})}+\xi_{ij}e^{f_{i}+f_{j}},$
-$\cosh l_{ik}$ $=\sqrt{(1+\alpha_{i}e^{2f_{i}})(1+\alpha_{k}e^{2f_{k}})}+\xi_{ik}e^{f_{i}+f_{k}},$
-$\cosh l_{jk}$ $=-\sqrt{(1+\alpha_{j}e^{2f_{j}})(1+\alpha_{k}e^{2f_{k}})}+\xi_{jk}e^{f_{j}+f_{k}}.$
-6.1. Admissible space of mixed discrete conformal structure $\mathrm{II}$
-The admissible space of the discrete conformal factor $f$ for a right-angled hyperbolic hexagon $\{ijk\}\in F$ with edge lengths given by (46) is defined to be
-$\Omega_{ijk}(\xi_{ij},\xi_{jk},\xi_{ik})=\{(f_{i},f_{j},f_{k})\in\mathbb{R}^{3}|l_{ij}>0,\ l_{jk}>0,\ l_{ik}>0\}.$
-Since $\frac{\sinh d_{ij}}{\sinh d_{ji}}=-\sqrt{\frac{1+\alpha_{i}e^{2f_{i}}}{1+\alpha_{j}e^{2f_{j}}}}$, $\frac{\sinh d_{jk}}{\sinh d_{kj}}=\sqrt{\frac{1+\alpha_{j}e^{2f_{j}}}{1+\alpha_{k}e^{2f_{k}}}}$, $\frac{\sinh d_{ik}}{\sinh d_{ki}}=-\sqrt{\frac{1+\alpha_{i}e^{2f_{i}}}{1+\alpha_{k}e^{2f_{k}}}}$, then by by the proof of Lemma 2.22, we have
+Suppose, to the contrary, that such a $j$ exists, then $u_{j}=-e^{-f_{j}}\rightarrow-\infty$.
+By the definition of the admissible space $\mathcal{U}_{ijk}(\eta)$ in (49), the inequality $u_{j}+u_{k}>-2\sqrt{\eta_{jk}}$ implies $u_{k}\rightarrow+\infty$, which contradicts the fact that $u_{k}\in\mathbb{R}_{<0}$.
+An analogous argument applies to a right-angled hyperbolic hexagon $\{ijk\}\in F$ with edge lengths given by (7).
+7. Rigidity of the mixed discrete conformal structure II
+In this section, we investigate the rigidity of the mixed discrete conformal structure II, specifically, an ideally triangulated surface with boundary admitting both discrete conformal structures (A2) and (B2).
+All right-angled hyperbolic hexagons are either Type-II or a combination of Type-II right-angled hyperbolic hexagons and those with edge lengths given by (6).
+Here, a right-angled hyperbolic hexagon is Type-II if its edge lengths are given by both (6) and (9).
+By Remark 1.6, we have $\alpha\equiv-1$ for both (6) and (9).
+As noted in Remark 1.7, if $d_{rs}<0$, then $d_{rt}<0,\ d_{sr}>0,\ d_{tr}>0,\ d_{st}>0,\ d_{ts}>0$, and $l_{rs},l_{rt}$ are given by (9) and $l_{st}$ is given by (6), where $\{r,s,t\}=\{i,j,k\}$.
+Without loss of generality, we assume $d_{ij}<0$ in this section.
+This implies $d_{ik}<0$, $d_{ji}>0$, $d_{ki}>0$, $d_{jk}>0$, and $d_{kj}>0$.
+For the Type-II right-angled hyperbolic hexagon $\{ijk\}\in F$, the edge lengths $l_{ij},l_{ik}$ are given by (9), and $l_{jk}$ is given by (6), namely,
+(52) $\cosh l_{ij}$ $=-\sqrt{(e^{2f_{i}}-1)(e^{2f_{j}}-1)}+\eta_{ij}e^{f_{i}+f_{j}},$
+$\cosh l_{ik}$ $=-\sqrt{(e^{2f_{i}}-1)(e^{2f_{k}}-1)}+\eta_{ik}e^{f_{i}+f_{k}},$
+$\cosh l_{jk}$ $=\sqrt{(e^{2f_{j}}-1)(e^{2f_{k}}-1)}+\eta_{jk}e^{f_{j}+f_{k}}.$
+Here $\eta_{ij}\in\mathbb{R}_{>0}$, $\eta_{ik}\in\mathbb{R}_{>0}$, and $\eta_{jk}\in(-1,+\infty)$.
+For a right-angled hyperbolic hexagon $\{ijk\}\in F$ with edge lengths given by (52), the admissible space of the discrete conformal factor $f$ is defined as
+$\Omega_{ijk}(\eta)=\{(f_{i},f_{j},f_{k})\in\mathbb{R}_{>0}^{3}\mid l_{ij}>0,\ l_{jk}>0,\ l_{ik}>0\}.$
+By combining (30) and Remark 2.22 (iv), we consider $u\in(-\frac{\pi}{2},0)^{N}$ for simplicity.
+It follows that $\cosh l_{ij}=-(-\tan u_{i})\frac{-1}{\tan u_{j}}+\eta_{ij}\frac{1}{\cos u_{i}}\frac{-1}{\sin u_{j}}>1$, which implies $\sin(u_{i}+u_{j})>-\eta_{ij}$.
+Consequently, $u_{i}+u_{j}\in(-\pi,0)$ for $\eta_{ij}\in[1,+\infty)$, and $u_{i}+u_{j}>-\arcsin\eta_{ij}$ for $\eta_{ij}\in[0,1)$.
+Moreover, from (45), $\cosh l_{jk}>1$ implies $u_{j}+u_{k}\in(-\pi,0)$ for $\eta_{jk}\in[1,+\infty)$ and $u_{j}+u_{k}>-\arccos\eta_{jk}$ for $\eta_{jk}\in(-1,1)$.
+Theorem 7.1.
+Let $\{ijk\}\in F$ be a right-angled hyperbolic hexagon with edge lengths given by (52), and the weight $\eta_{ij}\in[0,+\infty)$, $\eta_{ik}\in[0,+\infty)$, $\eta_{jk}\in(-1,+\infty)$.
+Then the admissible space
+[…]
+$\sin(u_{i}+u_{k})>-\eta_{ik},\cos(u_{j}+u_{k})>-\eta_{jk}\}.$
+of $u$ is a convex polytope.
+Corollary 7.2.
+Suppose $(\Sigma,\mathcal{T},\eta)$ is a weighted triangulated surface with boundary. If one of the following conditions holds:
+(1):
+All right-angled hyperbolic hexagons are Type-II (i.e., the edge lengths of any right-angled hyperbolic hexagons are given by (52)), and the weight $\eta$ satisfies the condition in Theorem 7.1;
+(2):
+Not all right-angled hyperbolic hexagons are Type-II (i.e., there exist some right-angled hyperbolic hexagons with edge lengths given by (6)).
+In this case, the weight $\eta$ satisfies the condition in Theorem 7.1 for those right-angled hyperbolic hexagons with edge lengths given by (52), and $\eta\in(-1,+\infty)^{E}$ for those with edge lengths given by (6).
+Then the admissible space $\mathcal{U}(\eta)=\bigcap_{\{ijk\}\in F}\mathcal{U}_{ijk}(\eta)$ is a convex polytope on $(\Sigma,\mathcal{T},\eta)$.
+Proof.
+The case (1) follows from Theorem 7.1, and the case (2) follows from (46) and Theorem 7.1. Q.E.D.
+Theorem 7.3.
+Under the same assumptions as those in Theorem 7.1, the Jacobian $\Lambda_{ijk}=\frac{\partial(\theta^{jk}_{i},\theta^{ik}_{j},\theta^{ij}_{k})}{\partial(u_{i},u_{j},u_{k})}$ is symmetric and negative definite on $\mathcal{U}_{ijk}(\eta)$.
+Proof.
+For simplicity, we adopt the notations introduced in Subsection 3.2. Define
+$\frac{\partial(f_{i},f_{j},f_{k})}{\partial(u_{i},u_{j},u_{k})}=\left(\begin{array}[]{ccc}-\sqrt{e^{2f_{i}}-1}&0&0\\ 0&\sqrt{e^{2f_{j}}-1}&0\\ 0&0&\sqrt{e^{2f_{k}}-1}\\ \end{array}\right):=Q_{7}.$
+Then $\Lambda_{ijk}$ can be expressed as
+$\Lambda_{ijk}=\frac{\partial(\theta_{i},\theta_{j},\theta_{k})}{\partial(u_{i},u_{j},u_{k})}=-\frac{1}{A}\left(\begin{array}[]{ccc}\sinh l_{i}&0&0\\ 0&\sinh l_{j}&0\\ 0&0&\sinh l_{k}\\ \end{array}\right)Q_{1}Q_{2}Q_{7},$
+where $Q_{1}$ and $Q_{2}$ are defined by (36) and (38), respectively.
+Furthermore, $\det Q_{1}>0$.
+For any right-angled hyperbolic hexagon $\{ijk\}\in F$ with edge lengths given by (52), since $d_{ij}<0$, $d_{ik}<0$, $d_{ji}>0$, $d_{ki}>0$, $d_{jk}>0$, and $d_{kj}>0$, it follows from (39) that $\det Q_{2}<0$.
+The remaining part of the proof is analogously to that of Theorem 6.3, and is thus omitted here.
+Q.E.D.
+Combining Theorem 4.1 and Theorem 7.3, we obtain the following result.
+Corollary 7.4.
+Under the same assumptions as those in Corollary 7.2, the Jacobian $\Lambda=\frac{\partial(K_{i},...,K_{N})}{\partial(u_{i},...,u_{N})}$ is symmetric and negative definite on $\mathcal{U}(\eta)$.
+The following theorem establishes the rigidity of the mixed discrete conformal structure II, which generalizes Theorem 1.8 (v). As its proof is analogous to that of Theorem 3.4, it is omitted here.
+Theorem 7.5.
+Under the same assumptions as those in Corollary 7.2, the discrete conformal factor $f$ is uniquely determined by its generalized combinatorial curvature $K\in\mathbb{R}^{N}_{>0}$.
+Remark 7.6.
+The existence of the mixed discrete conformal structure II cannot be established.
+Specifically, the boundary of the admissible space $\Omega(\eta)=\bigcap_{\{ijk\}\in F}\Omega_{ijk}(\eta)$ in $[-\infty,+\infty]^{N}$ include the case that there exists $i\in B$ such that $f_{i}^{(m)}\rightarrow 0^{+}$.
+For this case, we are unable to prove the closedness of the curvature map.
+8. Rigidity and existence of the mixed discrete conformal structure I
+In this section, we investigate the rigidity and existence of the mixed discrete conformal structure I, specifically, an ideally triangulated surface with boundary admitting both discrete conformal structures (A1) and (B1).
+All right-angled hyperbolic hexagons are either Type-I or a combination of Type-I right-angled hyperbolic hexagons and those with edge lengths given by (5).
+Here, a right-angled hyperbolic hexagon is Type-I if its edge lengths are given by both (5) and (8).
+As noted in Remark 1.7, if $d_{rs}<0$, then $d_{rt}<0,\ d_{sr}>0,\ d_{tr}>0,\ d_{st}>0,\ d_{ts}>0$, and $l_{rs},l_{rt}$ are given by (8) and $l_{st}$ is given by (5), where $\{r,s,t\}=\{i,j,k\}$.
+Without loss of generality, we assume $d_{ij}<0$ in this section.
+This implies $d_{ik}<0$, $d_{ji}>0$, $d_{ki}>0$, $d_{jk}>0$, and $d_{kj}>0$.
+For the Type-I right-angled hyperbolic hexagon $\{ijk\}\in F$, the edge lengths $l_{ij},l_{ik}$ are given by (8), and $l_{jk}$ is given by (5), namely,
+(53) $\cosh l_{ij}$ $=\sqrt{(1+\alpha_{i}e^{2f_{i}})(1+\alpha_{j}e^{2f_{j}})}+\eta_{ij}e^{f_{i}+f_{j}},$
+$\cosh l_{ik}$ $=\sqrt{(1+\alpha_{i}e^{2f_{i}})(1+\alpha_{k}e^{2f_{k}})}+\eta_{ik}e^{f_{i}+f_{k}},$
+$\cosh l_{jk}$ $=-\sqrt{(1+\alpha_{j}e^{2f_{j}})(1+\alpha_{k}e^{2f_{k}})}+\eta_{jk}e^{f_{j}+f_{k}}.$
+8.1. Admissible space of the mixed discrete conformal structure I
+For a right-angled hyperbolic hexagon $\{ijk\}\in F$ with edge lengths given by (53), the admissible space of the discrete conformal factor $f$ is defined as
+$\Omega_{ijk}(\eta)=\{(f_{i},f_{j},f_{k})\in\mathbb{R}^{3}\mid l_{ij}>0,\ l_{jk}>0,\ l_{ik}>0\}.$
+Since $\frac{\sinh d_{ij}}{\sinh d_{ji}}=-\sqrt{\frac{1+\alpha_{i}e^{2f_{i}}}{1+\alpha_{j}e^{2f_{j}}}}$, $\frac{\sinh d_{jk}}{\sinh d_{kj}}=\sqrt{\frac{1+\alpha_{j}e^{2f_{j}}}{1+\alpha_{k}e^{2f_{k}}}}$, and $\frac{\sinh d_{ik}}{\sinh d_{ki}}=-\sqrt{\frac{1+\alpha_{i}e^{2f_{i}}}{1+\alpha_{k}e^{2f_{k}}}}$, it follows from the proof of Lemma 2.21 that
 (54) $\frac{\partial f_{i}}{\partial u_{i}}=-\sqrt{1+\alpha_{i}e^{2f_{i}}},\ \frac{\partial f_{j}}{\partial u_{j}}=\sqrt{1+\alpha_{j}e^{2f_{j}}},\ \frac{\partial f_{k}}{\partial u_{k}}=\sqrt{1+\alpha_{k}e^{2f_{k}}}.$
-To obtain the admissible space $\mathcal{U}_{ijk}(\xi_{ij},\xi_{jk},\xi_{ik})$ with respect to the discrete conformal factors $u$, we need to characterize $l_{ij}>0$ and $l_{jk}>0$.
-Since the case that $l_{jk}>0$ has been already discussed in Subsection 3.1, we just need to discuss the case $l_{ij}>0$.
-By Remark 2.23, we have
+To characterizes the admissible space $\mathcal{U}_{ijk}(\eta)$ of the discrete conformal factors $u$, it is necessary to derive the conditions under which the edge lengths $l_{ij}>0$ and $l_{jk}>0$.
+Since the case that $l_{jk}>0$ has already been addressed in Subsection 3.1, we focus on analyzing $l_{ij}>0$ here.
+By Remark 2.22, the following results are obtained:
 (i):
-If $\alpha_{i}=0$ and $\alpha_{j}=0$, then $\cosh l_{ij}=1+\xi_{ij}e^{u_{j}-u_{i}}>1$, which implies $\xi_{ij}>0$ and $(u_{j}-u_{i})\in\mathbb{R}$.
+If $\alpha_{i}=0$ and $\alpha_{j}=0$, then $\cosh l_{ij}=1+\eta_{ij}e^{u_{j}-u_{i}}>1$. This implies $\eta_{ij}>0$ and $(u_{j}-u_{i})\in\mathbb{R}$.
 (ii):
-If $\alpha_{i}=1$ and $\alpha_{j}=0$, then $\cosh l_{ij}=\sqrt{1+e^{2f_{i}}}+\xi_{ij}e^{f_{i}+f_{j}}=\coth u_{i}+\xi_{ij}e^{u_{j}}\frac{1}{\sinh u_{i}}>1$, which implies $\xi_{ij}e^{u_{i}+u_{j}}+1>0$ by $u_{i}>0$.
-Thus if $\xi_{ij}\geq 0$, then $(u_{i}+u_{j})\in\mathbb{R}$; if $\xi_{ij}<0$, then $u_{i}+u_{j}<\log(-\frac{1}{\xi_{ij}})$.
-For simplicity, we only consider the case $\xi_{ij}\geq 0$ in the following.
+If $\alpha_{i}=1$ and $\alpha_{j}=0$, then $\cosh l_{ij}=\sqrt{1+e^{2f_{i}}}+\eta_{ij}e^{f_{i}+f_{j}}=\coth u_{i}+\eta_{ij}e^{u_{j}}\frac{1}{\sinh u_{i}}>1$.
+Since $u_{i}>0$, it follows that $\eta_{ij}e^{u_{i}+u_{j}}+1>0$.
+Consequently, if $\eta_{ij}\geq 0$, then $(u_{i}+u_{j})\in\mathbb{R}$; if $\eta_{ij}<0$, then $u_{i}+u_{j}<\log(-\frac{1}{\eta_{ij}})$.
+For simplicity, we only consider the case $\eta_{ij}\geq 0$ in the following.
 (iii):
-If $\alpha_{i}=-1$ and $\alpha_{j}=0$, then $\cosh l_{ij}=\sqrt{1-e^{2f_{i}}}+\xi_{ij}e^{f_{i}+f_{j}}=\tanh u_{i}+\xi_{ij}e^{u_{j}}\frac{1}{\cosh u_{i}}>1$, which implies $\xi_{ij}e^{u_{i}+u_{j}}>1$.
-Thus $\xi_{ij}>0$ and $u_{i}+u_{j}>\log\frac{1}{\xi_{ij}}$.
+If $\alpha_{i}=-1$ and $\alpha_{j}=0$, then $\cosh l_{ij}=\sqrt{1-e^{2f_{i}}}+\eta_{ij}e^{f_{i}+f_{j}}=\tanh u_{i}+\eta_{ij}e^{u_{j}}\frac{1}{\cosh u_{i}}>1$.
+This implies $\eta_{ij}e^{u_{i}+u_{j}}>1$, so $\eta_{ij}>0$ and $u_{i}+u_{j}>\log\frac{1}{\eta_{ij}}$.
 (iv):
-If $\alpha_{i}=0$ and $\alpha_{j}=1$, then $\cosh l_{ij}=\sqrt{1+e^{2f_{j}}}+\xi_{ij}e^{f_{i}+f_{j}}=-\coth u_{j}+\xi_{ij}e^{-u_{i}}\frac{-1}{\sinh u_{j}}>1$, which implies $e^{u_{i}+u_{j}}+\xi_{ij}>0$ by $u_{j}<0$.
-Thus if $\xi_{ij}\geq 0$, then $(u_{i}+u_{j})\in\mathbb{R}$; if $\xi_{ij}<0$, then $u_{i}+u_{j}>\log(-\xi_{ij})$.
+If $\alpha_{i}=0$ and $\alpha_{j}=1$, then $\cosh l_{ij}=\sqrt{1+e^{2f_{j}}}+\eta_{ij}e^{f_{i}+f_{j}}=-\coth u_{j}+\eta_{ij}e^{-u_{i}}\frac{-1}{\sinh u_{j}}>1$.
+Since $u_{j}<0$, it follows that $e^{u_{i}+u_{j}}+\eta_{ij}>0$.
+Consequently, if $\eta_{ij}\geq 0$, then $(u_{i}+u_{j})\in\mathbb{R}$; if $\eta_{ij}<0$, then $u_{i}+u_{j}>\log(-\eta_{ij})$.
 (v):
-If $\alpha_{i}=1$ and $\alpha_{j}=1$, then $\cosh l_{ij}=\sqrt{(1+e^{2f_{i}})(1+e^{2f_{j}})}+\xi_{ij}e^{f_{i}+f_{j}}=-\coth u_{i}\coth u_{j}+\xi_{ij}\frac{1}{\sinh u_{i}}\frac{-1}{\sinh u_{j}}>1$, which implies $\cosh(u_{i}+u_{j})+\xi_{ij}>0$ by $u_{i}>0$ and $u_{j}<0$.
-Thus if $\xi_{ij}>-1$, then $(u_{i}+u_{j})\in\mathbb{R}$; if $\xi_{ij}<-1$, then $u_{i}+u_{j}>\mathrm{arccosh}\ (-\xi_{ij})$ or $u_{i}+u_{j}<-\mathrm{arccosh}\ (-\xi_{ij})$.
-For simplicity, we only consider the case $\xi_{ij}>-1$ in the following.
+If $\alpha_{i}=1$ and $\alpha_{j}=1$, then $\cosh l_{ij}=\sqrt{(1+e^{2f_{i}})(1+e^{2f_{j}})}+\eta_{ij}e^{f_{i}+f_{j}}=-\coth u_{i}\coth u_{j}+\eta_{ij}\frac{1}{\sinh u_{i}}\frac{-1}{\sinh u_{j}}>1$.
+Since $u_{i}>0$ and $u_{j}<0$, it follows that $\cosh(u_{i}+u_{j})+\eta_{ij}>0$.
+Consequently, if $\eta_{ij}>-1$, then $(u_{i}+u_{j})\in\mathbb{R}$; if $\eta_{ij}<-1$, then $u_{i}+u_{j}>\mathrm{arccosh}\,(-\eta_{ij})$ or $u_{i}+u_{j}<-\mathrm{arccosh}\,(-\eta_{ij})$.
+For simplicity, we only consider the case $\eta_{ij}>-1$ in the following.
 (vi):
-If $\alpha_{i}=-1$ and $\alpha_{j}=1$, then $\cosh l_{ij}=\sqrt{(1-e^{2f_{i}})(1+e^{2f_{j}})}+\xi_{ij}e^{f_{i}+f_{j}}=-\tanh u_{i}\coth u_{j}+\xi_{ij}\frac{1}{\cosh u_{i}}\frac{-1}{\sinh u_{j}}>1$, which implies $\sinh(u_{i}+u_{j})+\xi_{ij}>0$ by $u_{j}<0$.
-Thus $\xi_{ij}\in\mathbb{R}$ and $u_{i}+u_{j}>\mathrm{arcsinh}\ (-\xi_{ij})$.
+If $\alpha_{i}=-1$ and $\alpha_{j}=1$, then $\cosh l_{ij}=\sqrt{(1-e^{2f_{i}})(1+e^{2f_{j}})}+\eta_{ij}e^{f_{i}+f_{j}}=-\tanh u_{i}\coth u_{j}+\eta_{ij}\frac{1}{\cosh u_{i}}\frac{-1}{\sinh u_{j}}>1$.
+Since $u_{j}<0$, it follows that $\sinh(u_{i}+u_{j})+\eta_{ij}>0$.
+Thus $\eta_{ij}\in\mathbb{R}$ and $u_{i}+u_{j}>\mathrm{arcsinh}\ (-\eta_{ij})$.
 (vii):
-If $\alpha_{i}=0$ and $\alpha_{j}=-1$, then $\cosh l_{ij}=\sqrt{1-e^{2f_{j}}}+\xi_{ij}e^{f_{i}+f_{j}}=-\tanh u_{j}+\xi_{ij}e^{-u_{i}}\frac{1}{\cosh u_{j}}>1$, which implies $e^{u_{i}+u_{j}}<\xi_{ij}$.
-Thus $\xi_{ij}>0$ and $u_{i}+u_{j}<\log\xi_{ij}$.
+If $\alpha_{i}=0$ and $\alpha_{j}=-1$, then $\cosh l_{ij}=\sqrt{1-e^{2f_{j}}}+\eta_{ij}e^{f_{i}+f_{j}}=-\tanh u_{j}+\eta_{ij}e^{-u_{i}}\frac{1}{\cosh u_{j}}>1$.
+This implies $e^{u_{i}+u_{j}}<\eta_{ij}$, so $\eta_{ij}>0$ and $u_{i}+u_{j}<\log\eta_{ij}$.
 (viii):
-If $\alpha_{i}=1$ and $\alpha_{j}=-1$, then $\cosh l_{ij}=\sqrt{(1+e^{2f_{i}})(1-e^{2f_{j}})}+\xi_{ij}e^{f_{i}+f_{j}}=-\coth u_{i}\tanh u_{j}+\xi_{ij}\frac{1}{\sinh u_{i}}\frac{1}{\cosh u_{j}}>1$, which implies $\sinh(u_{i}+u_{j})<\xi_{ij}$.
-Thus $\xi_{ij}\in\mathbb{R}$ and $u_{i}+u_{j}<\mathrm{arcsinh}\ \xi_{ij}$.
+If $\alpha_{i}=1$ and $\alpha_{j}=-1$, then $\cosh l_{ij}=\sqrt{(1+e^{2f_{i}})(1-e^{2f_{j}})}+\eta_{ij}e^{f_{i}+f_{j}}=-\coth u_{i}\tanh u_{j}+\eta_{ij}\frac{1}{\sinh u_{i}}\frac{1}{\cosh u_{j}}>1$.
+This implies $\sinh(u_{i}+u_{j})<\eta_{ij}$, so $\eta_{ij}\in\mathbb{R}$ and $u_{i}+u_{j}<\mathrm{arcsinh}\,\eta_{ij}$.
 (ix):
-If $\alpha_{i}=-1$ and $\alpha_{j}=-1$, then $\cosh l_{ij}=\sqrt{(1-e^{2f_{i}})(1-e^{2f_{j}})}+\xi_{ij}e^{f_{i}+f_{j}}=-\tanh u_{i}\tanh u_{j}+\xi_{ij}\frac{1}{\cosh u_{i}}\frac{1}{\cosh u_{j}}>1$, which implies $\cosh(u_{i}+u_{j})<\xi_{ij}$.
-Thus $\xi_{ij}>1$ and $-\mathrm{arccosh}\ \xi_{ij}<u_{i}+u_{j}<\mathrm{arccosh}\ \xi_{ij}$.
-For simplicity, we use triples $(\alpha_{i},\alpha_{j},\alpha_{k})$ to represent different types of the admissible spaces of $u$ for a right-angled hyperbolic hexagon $\{ijk\}\in F$ with edge lengths given by (46).
-By permutations and combinations, the admissible space $\Omega_{ijk}(\xi_{ij},\xi_{jk},\xi_{ik})$ of $f$ can be transferred to 27 types of the admissible spaces of $u$.
-Note that the positions of $j$ and $k$ are inter-changeable in (46), we regard $(\alpha_{i},\alpha_{j},\alpha_{k})$ and $(\alpha_{i},\alpha_{k},\alpha_{j})$ as the same type.
-Therefore, there exist 18 types of the admissible spaces $\mathcal{U}^{(n)}_{ijk}(\xi_{ij},\xi_{jk},\xi_{ik})$ of $u$, as shown below.
-Here $n=1,2,...,18$.
-For simplicity, we write $\mathcal{U}^{(n)}_{ijk}(\xi_{ij},\xi_{jk},\xi_{ik})$ as $\mathcal{U}^{(n)}_{ijk}(\xi)$.
+If $\alpha_{i}=-1$ and $\alpha_{j}=-1$, then $\cosh l_{ij}=\sqrt{(1-e^{2f_{i}})(1-e^{2f_{j}})}+\eta_{ij}e^{f_{i}+f_{j}}=-\tanh u_{i}\tanh u_{j}+\eta_{ij}\frac{1}{\cosh u_{i}}\frac{1}{\cosh u_{j}}>1$.
+This implies $\cosh(u_{i}+u_{j})<\eta_{ij}$, so $\eta_{ij}>1$ and $-\mathrm{arccosh}\,\eta_{ij}<u_{i}+u_{j}<\mathrm{arccosh}\,\eta_{ij}$.
+For simplicity, we adopt the triple notation $(\alpha_{i},\alpha_{j},\alpha_{k})$ to classify the distinct types of admissible spaces of $u$ for a right-angled hyperbolic hexagon $\{ijk\}\in F$ with edge lengths given by (53).
+Through permutations and combinations, the admissible space $\Omega_{ijk}(\eta)$ of $f$ can be transformed into 27 distinct types of admissible spaces of $u$.
+However, due to the interchangeability of the positions of $j$ and $k$ in (53), the triples $(\alpha_{i},\alpha_{j},\alpha_{k})$ and $(\alpha_{i},\alpha_{k},\alpha_{j})$ are considered equivalent and thus belong to the same type.
+This reduction results in a total of 18 distinct types of admissible spaces of $u$, denoted by $\mathcal{U}^{(n)}_{ijk}(\eta)$, where $n=1,2,...,18$.
+These types are enumerated as follows.
 (I):
-Set $(\alpha_{i},\alpha_{j},\alpha_{k})=(0,0,0)$.
-Then $u_{j}+u_{k}>\log\frac{2}{\xi_{jk}}$ and $\xi_{jk}>0$ in Subsection 3.1.
-By the above discussion, we have $\xi_{ij}>0,\ (u_{j}-u_{i})\in\mathbb{R}$ and $\xi_{ik}>0,\ (u_{k}-u_{i})\in\mathbb{R}$.
-Therefore, the admissible space of $u$ is defined as
-(48) $\mathcal{U}^{(1)}_{ijk}(\xi)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}^{3}|u_{j}+u_{k}>\log\frac{2}{\xi_{jk}}\},$
-where $\xi_{jk}>0,\ \xi_{ij}>0,\ \xi_{ik}>0$.
+Assume $(\alpha_{i},\alpha_{j},\alpha_{k})=(0,0,0)$.
+In Subsection 3.1, it follows that $u_{j}+u_{k}>\log\frac{2}{\eta_{jk}}$ and $\eta_{jk}>0$.
+Building upon the preceding analysis, we deduce $\eta_{ij}>0$ with $(u_{j}-u_{i})\in\mathbb{R}$, and $\eta_{ik}>0$ with $(u_{k}-u_{i})\in\mathbb{R}$.
+Consequently, the admissible space of $u$ is defined as
+(55) $\mathcal{U}^{(1)}_{ijk}(\eta)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}^{3}\mid u_{j}+u_{k}>\log\frac{2}{\eta_{jk}}\},$
+where $\eta_{jk}>0$, $\eta_{ij}>0$, and $\eta_{ik}>0$.
 (II):
-Set $(\alpha_{i},\alpha_{j},\alpha_{k})=(1,0,0)$.
-Then $u_{j}+u_{k}>\log\frac{2}{\xi_{jk}}$ and $\xi_{jk}>0$.
-Moreover, $(u_{i}+u_{j})\in\mathbb{R}$ for $\xi_{ij}\geq 0$ and $(u_{i}+u_{k})\in\mathbb{R}$ for $\xi_{ik}\geq 0$.
-Then
+Assume $(\alpha_{i},\alpha_{j},\alpha_{k})=(1,0,0)$.
+It is known that $u_{j}+u_{k}>\log\frac{2}{\eta_{jk}}$ and $\eta_{jk}>0$.
+Moreover, $(u_{i}+u_{j})\in\mathbb{R}$ for $\eta_{ij}\geq 0$, and $(u_{i}+u_{k})\in\mathbb{R}$ for $\eta_{ik}\geq 0$.
+Consequently, the admissible space of $u$ is defined as
 […]
-where $\xi_{jk}>0,\ \xi_{ij}\geq 0,\ \xi_{ik}\geq 0$.
+where $\eta_{jk}>0$, $\eta_{ij}\geq 0$, and $\eta_{ik}\geq 0$.
 (III):
-Set $(\alpha_{i},\alpha_{j},\alpha_{k})=(-1,0,0)$.
-Then $u_{j}+u_{k}>\log\frac{2}{\xi_{jk}},\ \xi_{jk}>0$.
-Moreover, $\xi_{ij}>0$$u_{i}+u_{j}>\log\frac{1}{\xi_{ij}}$ and $\xi_{ik}>0,\ u_{i}+u_{k}>\log\frac{1}{\xi_{ik}}$.
-Then
-$\mathcal{U}^{(3)}_{ijk}(\xi)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}_{>0}\times\mathbb{R}^{2}|u_{j}+u_{k}>\log\frac{2}{\xi_{jk}},\ u_{i}+u_{j}>\log\frac{1}{\xi_{ij}},\ u_{i}+u_{k}>\log\frac{1}{\xi_{ik}}\},$
-where $\xi_{jk}>0,\ \xi_{ij}>0,\ \xi_{ik}>0$.
+Assume $(\alpha_{i},\alpha_{j},\alpha_{k})=(-1,0,0)$.
+It is known that $u_{j}+u_{k}>\log\frac{2}{\eta_{jk}}$ and $\eta_{jk}>0$.
+Moreover, $\eta_{ij}>0$ with $u_{i}+u_{j}>\log\frac{1}{\eta_{ij}}$, and $\eta_{ik}>0$ with $u_{i}+u_{k}>\log\frac{1}{\eta_{ik}}$.
+Consequently, the admissible space of $u$ is defined as
+$\mathcal{U}^{(3)}_{ijk}(\eta)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}_{>0}\times\mathbb{R}^{2}\mid u_{j}+u_{k}>\log\frac{2}{\eta_{jk}},\ u_{i}+u_{j}>\log\frac{1}{\eta_{ij}},\ u_{i}+u_{k}>\log\frac{1}{\eta_{ik}}\},$
+where $\eta_{jk}>0$, $\eta_{ij}>0$, and $\eta_{ik}>0$.
 (IV):
-Set $(\alpha_{i},\alpha_{j},\alpha_{k})=(0,0,1)$.
-Then $u_{j}+u_{k}>\log\frac{1}{\xi_{jk}}$ and $\xi_{jk}>0$ in Subsection 3.1.
-Moreover, $\xi_{ij}>0$ and $(u_{j}-u_{i})\in\mathbb{R}$.
-Similarly, $(u_{i}+u_{k})\in\mathbb{R}$ for $\xi_{ik}\geq 0$ and $u_{i}+u_{k}>\log(-\xi_{ik})$ for $\xi_{ik}<0$.
-Then
-(50) $\mathcal{U}^{(4)}_{ijk}(\xi)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}^{2}\times\mathbb{R}_{<0}|u_{j}+u_{k}>\log\frac{1}{\xi_{jk}},\ e^{u_{i}+u_{k}}>-\xi_{ik}\},$
-where $\xi_{jk}>0,\ \xi_{ij}>0,\ \xi_{ik}\in\mathbb{R}$.
-The admissible space (50) can be described explicitly as follows:
+Assume $(\alpha_{i},\alpha_{j},\alpha_{k})=(0,0,1)$.
+In Subsection 3.1, it follows that $u_{j}+u_{k}>\log\frac{1}{\eta_{jk}}$ and $\eta_{jk}>0$.
+Moreover, $\eta_{ij}>0$ and $(u_{j}-u_{i})\in\mathbb{R}$.
+Similarly, $(u_{i}+u_{k})\in\mathbb{R}$ for $\eta_{ik}\geq 0$, and $u_{i}+u_{k}>\log(-\eta_{ik})$ for $\eta_{ik}<0$.
+Consequently, the admissible space of $u$ is defined as
+(57) $\mathcal{U}^{(4)}_{ijk}(\eta)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}^{2}\times\mathbb{R}_{<0}\mid u_{j}+u_{k}>\log\frac{1}{\eta_{jk}},\ e^{u_{i}+u_{k}}>-\eta_{ik}\},$
+where $\eta_{jk}>0$, $\eta_{ij}>0$, and $\eta_{ik}\in\mathbb{R}$.
+The admissible space (57) can be explicitly characterized as follows:
 (i):
-For $\xi_{jk}>0,\ \xi_{ij}>0,\ \xi_{ik}\geq 0$, then
-(51) $\mathcal{U}^{(4)}_{ijk}(\xi)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}^{2}\times\mathbb{R}_{<0}|u_{j}+u_{k}>\log\frac{1}{\xi_{jk}}\}.$
+For $\eta_{jk}>0$, $\eta_{ij}>0$, and $\eta_{ik}\geq 0$, the admissible space is given by
+(58) $\mathcal{U}^{(4)}_{ijk}(\eta)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}^{2}\times\mathbb{R}_{<0}\mid u_{j}+u_{k}>\log\frac{1}{\eta_{jk}}\}.$
 (ii):
-For $\xi_{jk}>0,\ \xi_{ij}>0,\ \xi_{ik}<0$, then
-(52) $\mathcal{U}^{(4)}_{ijk}(\xi)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}^{2}\times\mathbb{R}_{<0}|u_{j}+u_{k}>\log\frac{1}{\xi_{jk}},\ u_{i}+u_{k}>\log(-\xi_{ik})\}.$
+For $\eta_{jk}>0$, $\eta_{ij}>0$, and $\eta_{ik}<0$, the admissible space is given by
+(59) $\mathcal{U}^{(4)}_{ijk}(\eta)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}^{2}\times\mathbb{R}_{<0}\mid u_{j}+u_{k}>\log\frac{1}{\eta_{jk}},\ u_{i}+u_{k}>\log(-\eta_{ik})\}.$
 (V):
-Set $(\alpha_{i},\alpha_{j},\alpha_{k})=(1,0,1)$.
-Then $u_{j}+u_{k}>\log\frac{1}{\xi_{jk}}$ and $\xi_{jk}>0$.
-Moreover, $(u_{i}+u_{j})\in\mathbb{R}$ for $\xi_{ij}\geq 0$ and $(u_{i}+u_{k})\in\mathbb{R}$ for $\xi_{ik}>-1$.
-Then
-(53) $\mathcal{U}^{(5)}_{ijk}(\xi)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}\times\mathbb{R}_{>0}\times\mathbb{R}_{<0}|u_{j}+u_{k}>\log\frac{1}{\xi_{jk}}\},$
-where $\xi_{jk}>0,\ \xi_{ij}\geq 0,\ \xi_{ik}>-1$.
+Assume $(\alpha_{i},\alpha_{j},\alpha_{k})=(1,0,1)$.
+It is known that $u_{j}+u_{k}>\log\frac{1}{\eta_{jk}}$ and $\eta_{jk}>0$.
+Moreover, $(u_{i}+u_{j})\in\mathbb{R}$ for $\eta_{ij}\geq 0$, and $(u_{i}+u_{k})\in\mathbb{R}$ for $\eta_{ik}>-1$.
+Consequently, the admissible space of $u$ is defined as
+(60) $\mathcal{U}^{(5)}_{ijk}(\eta)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}\times\mathbb{R}_{>0}\times\mathbb{R}_{<0}\mid u_{j}+u_{k}>\log\frac{1}{\eta_{jk}}\},$
+where $\eta_{jk}>0$, $\eta_{ij}\geq 0$, and $\eta_{ik}>-1$.
 (VI):
-Set $(\alpha_{i},\alpha_{j},\alpha_{k})=(-1,0,1)$.
-Then $u_{j}+u_{k}>\log\frac{1}{\xi_{jk}}$ and $\xi_{jk}>0$.
-Moreover, $\xi_{ij}>0$$u_{i}+u_{j}>\log\frac{1}{\xi_{ij}}$.
-Similarly, $\xi_{ik}\in\mathbb{R}$ and $u_{i}+u_{k}>\mathrm{arcsinh}\ (-\xi_{ik})$.
-Then
-$\mathcal{U}^{(6)}_{ijk}(\xi)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}_{>0}\times\mathbb{R}\times\mathbb{R}_{<0}|$ $u_{j}+u_{k}>\log\frac{1}{\xi_{jk}},\ u_{i}+u_{j}>\log\frac{1}{\xi_{ij}},$
-$u_{i}+u_{k}>\mathrm{arcsinh}\ (-\xi_{ik})\},$
-where $\xi_{jk}>0,\ \xi_{ij}>0,\ \xi_{ik}\in\mathbb{R}$.
+Assume $(\alpha_{i},\alpha_{j},\alpha_{k})=(-1,0,1)$.
+It is known that $u_{j}+u_{k}>\log\frac{1}{\eta_{jk}}$ and $\eta_{jk}>0$.
+Moreover, $\eta_{ij}>0$ and $u_{i}+u_{j}>\log\frac{1}{\eta_{ij}}$.
+Similarly, $\eta_{ik}\in\mathbb{R}$ and $u_{i}+u_{k}>\mathrm{arcsinh}\,(-\eta_{ik})$.
+Consequently, the admissible space of $u$ is defined as
+$\mathcal{U}^{(6)}_{ijk}(\eta)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}_{>0}\times\mathbb{R}\times\mathbb{R}_{<0}\mid$ $\ u_{j}+u_{k}>\log\frac{1}{\eta_{jk}},\ u_{i}+u_{j}>\log\frac{1}{\eta_{ij}},$
+$\ u_{i}+u_{k}>\mathrm{arcsinh}\,(-\eta_{ik})\},$
+where $\eta_{jk}>0$, $\eta_{ij}>0$, and $\eta_{ik}\in\mathbb{R}$.
 (VII):
-Set $(\alpha_{i},\alpha_{j},\alpha_{k})=(0,0,-1)$.
-Then $u_{j}+u_{k}>\log\frac{1}{\xi_{jk}}$ and $\xi_{jk}>0$ in Subsection 3.1.
-Moreover, $\xi_{ij}>0$$(u_{j}-u_{i})\in\mathbb{R}$.
-Similarly, $\xi_{ik}>0$ and $u_{i}+u_{k}<\log\xi_{ik}$.
-Then
-$\mathcal{U}^{(7)}_{ijk}(\xi)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}^{2}\times\mathbb{R}_{<0}|u_{j}+u_{k}>\log\frac{1}{\xi_{jk}},\ u_{i}+u_{k}<\log\xi_{ik}\},$
-where $\xi_{jk}>0,\ \xi_{ij}>0,\ \xi_{ik}>0$.
+Assume $(\alpha_{i},\alpha_{j},\alpha_{k})=(0,0,-1)$.
+In Subsection 3.1, it follows that $u_{j}+u_{k}>\log\frac{1}{\eta_{jk}}$ and $\eta_{jk}>0$.
+Moreover, $\eta_{ij}>0$ and $(u_{j}-u_{i})\in\mathbb{R}$.
+Similarly, $\eta_{ik}>0$ and $u_{i}+u_{k}<\log\eta_{ik}$.
+Consequently, the admissible space of $u$ is defined as
+$\mathcal{U}^{(7)}_{ijk}(\eta)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}^{2}\times\mathbb{R}_{<0}\mid u_{j}+u_{k}>\log\frac{1}{\eta_{jk}},\ u_{i}+u_{k}<\log\eta_{ik}\},$
+where $\eta_{jk}>0$, $\eta_{ij}>0$, and $\eta_{ik}>0$.
 (VIII):
-Set $(\alpha_{i},\alpha_{j},\alpha_{k})=(1,0,-1)$.
-Then $u_{j}+u_{k}>\log\frac{1}{\xi_{jk}}$ and $\xi_{jk}>0$.
-Moreover, $(u_{i}+u_{j})\in\mathbb{R}$ for $\xi_{ij}\geq 0$.
-Similarly, $\xi_{ik}\in\mathbb{R}$ and $u_{i}+u_{k}<\mathrm{arcsinh}\ \xi_{ik}$.
-Then
+Assume $(\alpha_{i},\alpha_{j},\alpha_{k})=(1,0,-1)$.
+It is known that $u_{j}+u_{k}>\log\frac{1}{\eta_{jk}}$ and $\eta_{jk}>0$.
+Moreover, $(u_{i}+u_{j})\in\mathbb{R}$ for $\eta_{ij}\geq 0$.
+Similarly, $\eta_{ik}\in\mathbb{R}$ and $u_{i}+u_{k}<\mathrm{arcsinh}\,\eta_{ik}$.
+Consequently, the admissible space of $u$ is defined as
 […]
-where $\xi_{jk}>0,\ \xi_{ij}\geq 0,\ \xi_{ik}\in\mathbb{R}$.
+where $\eta_{jk}>0$, $\eta_{ij}\geq 0$, and $\eta_{ik}\in\mathbb{R}$.
 (IX):
-Set $(\alpha_{i},\alpha_{j},\alpha_{k})=(-1,0,-1)$.
-Then $u_{j}+u_{k}>\log\frac{1}{\xi_{jk}}$ and $\xi_{jk}>0$.
-Moreover, $\xi_{ij}>0$ and $u_{i}+u_{j}>\log\frac{1}{\xi_{ij}}$.
-Similarly, $\xi_{ik}>1$ and $-\mathrm{arccosh}\ \xi_{ik}<u_{i}+u_{k}<\mathrm{arccosh}\ \xi_{ik}$.
-Then
-$\mathcal{U}^{(9)}_{ijk}(\xi)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}_{>0}\times\mathbb{R}\times\mathbb{R}_{<0}|$ $u_{j}+u_{k}>\log\frac{1}{\xi_{jk}},\ u_{i}+u_{j}>\log\frac{1}{\xi_{ij}},$
-$-\mathrm{arccosh}\ \xi_{ik}<u_{i}+u_{k}<\mathrm{arccosh}\ \xi_{ik}\},$
-where $\xi_{jk}>0,\ \xi_{ij}>0,\ \xi_{ik}>1$.
+Assume $(\alpha_{i},\alpha_{j},\alpha_{k})=(-1,0,-1)$.
+It is known that $u_{j}+u_{k}>\log\frac{1}{\eta_{jk}}$ and $\eta_{jk}>0$.
+Moreover, $\eta_{ij}>0$ and $u_{i}+u_{j}>\log\frac{1}{\eta_{ij}}$.
+Similarly, $\eta_{ik}>1$ and $-\mathrm{arccosh}\,\eta_{ik}<u_{i}+u_{k}<\mathrm{arccosh}\,\eta_{ik}$.
+Consequently, the admissible space of $u$ is defined as
+$\mathcal{U}^{(9)}_{ijk}(\eta)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}_{>0}\times\mathbb{R}\times\mathbb{R}_{<0}\mid$ $\ u_{j}+u_{k}>\log\frac{1}{\eta_{jk}},\ u_{i}+u_{j}>\log\frac{1}{\eta_{ij}},$
+$\ -\mathrm{arccosh}\,\eta_{ik}<u_{i}+u_{k}<\mathrm{arccosh}\,\eta_{ik}\},$
+where $\eta_{jk}>0$, $\eta_{ij}>0$, and $\eta_{ik}>1$.
 (X):
-Set $(\alpha_{i},\alpha_{j},\alpha_{k})=(0,1,1)$.
-Then $u_{j}+u_{k}>-\mathrm{arccosh}\ \xi_{jk}$ and $\xi_{jk}>1$ in Subsection 3.1.
-Moreover, $(u_{i}+u_{j})\in\mathbb{R}$ for $\xi_{ij}\geq 0$ and $u_{i}+u_{j}>\log(-\xi_{ij})$ for $\xi_{ij}<0$.
-Similarly, $(u_{i}+u_{k})\in\mathbb{R}$ for $\xi_{ik}\geq 0$ and $u_{i}+u_{k}>\log(-\xi_{ik})$ for $\xi_{ik}<0$.
-Then
-(54) $\mathcal{U}^{(10)}_{ijk}(\xi)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}\times\mathbb{R}^{2}_{<0}|$ $u_{j}+u_{k}>-\mathrm{arccosh}\ \xi_{jk},\ e^{u_{i}+u_{j}}>-\xi_{ij},$
-$e^{u_{i}+u_{k}}>-\xi_{ik}\},$
-where $\xi_{jk}>1,\ \xi_{ij}\in\mathbb{R},\ \xi_{ik}\in\mathbb{R}$.
-The admissible space (54) can also be described explicitly as follows:
+Assume $(\alpha_{i},\alpha_{j},\alpha_{k})=(0,1,1)$.
+In Subsection 3.1, it follows that $u_{j}+u_{k}>-\mathrm{arccosh}\,\eta_{jk}$ and $\eta_{jk}>1$.
+Moreover, $(u_{i}+u_{j})\in\mathbb{R}$ for $\eta_{ij}\geq 0$, and $u_{i}+u_{j}>\log(-\eta_{ij})$ for $\eta_{ij}<0$.
+Similarly, $(u_{i}+u_{k})\in\mathbb{R}$ for $\eta_{ik}\geq 0$, and $u_{i}+u_{k}>\log(-\eta_{ik})$ for $\eta_{ik}<0$.
+Consequently, the admissible space of $u$ is defined as
+(61) $\mathcal{U}^{(10)}_{ijk}(\eta)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}\times\mathbb{R}^{2}_{<0}\mid$ $\ u_{j}+u_{k}>-\mathrm{arccosh}\,\eta_{jk},\ e^{u_{i}+u_{j}}>-\eta_{ij},$
+$\ e^{u_{i}+u_{k}}>-\eta_{ik}\},$
+where $\eta_{jk}>1$, $\eta_{ij}\in\mathbb{R}$, and $\eta_{ik}\in\mathbb{R}$.
+The admissible space (61) can be explicitly characterized as follows:
 (i):
-For $\xi_{jk}>1,\ \xi_{ij}\geq 0,\ \xi_{ik}\geq 0$, then
-$\mathcal{U}^{(10)}_{ijk}(\xi)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}\times\mathbb{R}^{2}_{<0}|u_{j}+u_{k}>-\mathrm{arccosh}\ \xi_{jk}\}.$
+For $\eta_{jk}>1$, $\eta_{ij}\geq 0$, and $\eta_{ik}\geq 0$, the admissible space is given by
+$\mathcal{U}^{(10)}_{ijk}(\eta)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}\times\mathbb{R}^{2}_{<0}\mid u_{j}+u_{k}>-\mathrm{arccosh}\,\eta_{jk}\}.$
 (ii):
-For $\xi_{jk}>1,\ \xi_{ij}\geq 0,\ \xi_{ik}<0$, then
-$\mathcal{U}^{(10)}_{ijk}(\xi)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}\times\mathbb{R}^{2}_{<0}|u_{j}+u_{k}>-\mathrm{arccosh}\ \xi_{jk},\ u_{i}+u_{k}>\log(-\xi_{ik})\}.$
+For $\eta_{jk}>1$, $\eta_{ij}\geq 0$, and $\eta_{ik}<0$, the admissible space is given by
+$\mathcal{U}^{(10)}_{ijk}(\eta)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}\times\mathbb{R}^{2}_{<0}\mid u_{j}+u_{k}>-\mathrm{arccosh}\,\eta_{jk},\ u_{i}+u_{k}>\log(-\eta_{ik})\}.$
 (iii):
-For $\xi_{jk}>1,\ \xi_{ij}<0,\ \xi_{ik}\geq 0$, then
-$\mathcal{U}^{(10)}_{ijk}(\xi)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}\times\mathbb{R}^{2}_{<0}|u_{j}+u_{k}>-\mathrm{arccosh}\ \xi_{jk},\ u_{i}+u_{j}>\log(-\xi_{ij})\},$
+For $\eta_{jk}>1$, $\eta_{ij}<0$, and $\eta_{ik}\geq 0$, the admissible space is given by
+$\mathcal{U}^{(10)}_{ijk}(\eta)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}\times\mathbb{R}^{2}_{<0}\mid u_{j}+u_{k}>-\mathrm{arccosh}\,\eta_{jk},\ u_{i}+u_{j}>\log(-\eta_{ij})\},$
 (iv):
-For $\xi_{jk}>1,\ \xi_{ij}<0,\ \xi_{ik}<0$, then
+For $\eta_{jk}>1$, $\eta_{ij}<0$, and $\eta_{ik}<0$, the admissible space is given by
 […]
-$u_{i}+u_{j}>\log(-\xi_{ij}),\ u_{i}+u_{k}>\log(-\xi_{ik})\}.$
+$u_{i}+u_{j}>\log(-\eta_{ij}),\ u_{i}+u_{k}>\log(-\eta_{ik})\}.$
 (XI):
-Set $(\alpha_{i},\alpha_{j},\alpha_{k})=(1,1,1)$.
-Then $u_{j}+u_{k}>-\mathrm{arccosh}\ \xi_{jk}$ and $\xi_{jk}>1$.
-Moreover, $\xi_{ij}>-1$$(u_{i}+u_{j})\in\mathbb{R}$ and $\xi_{ik}>-1,\ (u_{i}+u_{k})\in\mathbb{R}$.
-Then
-(55) $\mathcal{U}^{(11)}_{ijk}(\xi)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}_{>0}\times\mathbb{R}^{2}_{<0}|u_{j}+u_{k}>-\mathrm{arccosh}\ \xi_{jk}\},$
-where $\xi_{jk}>1,\ \xi_{ij}>-1,\ \xi_{ik}>-1$.
+Assume $(\alpha_{i},\alpha_{j},\alpha_{k})=(1,1,1)$.
+It is known that $u_{j}+u_{k}>-\mathrm{arccosh}\,\eta_{jk}$ and $\eta_{jk}>1$.
+Moreover, $\eta_{ij}>-1$ with $(u_{i}+u_{j})\in\mathbb{R}$, and $\eta_{ik}>-1$ with $(u_{i}+u_{k})\in\mathbb{R}$.
+Consequently, the admissible space of $u$ is defined as
+(62) $\mathcal{U}^{(11)}_{ijk}(\eta)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}_{>0}\times\mathbb{R}^{2}_{<0}\mid u_{j}+u_{k}>-\mathrm{arccosh}\,\eta_{jk}\},$
+where $\eta_{jk}>1$, $\eta_{ij}>-1$, and $\eta_{ik}>-1$.
 (XII):
-Set $(\alpha_{i},\alpha_{j},\alpha_{k})=(-1,1,1)$.
-Then $u_{j}+u_{k}>-\mathrm{arccosh}\ \xi_{jk}$ and $\xi_{jk}>1$.
-Moreover, $\xi_{ij}\in\mathbb{R}$$u_{i}+u_{j}>\mathrm{arcsinh}\ (-\xi_{ij})$ and $\xi_{ik}\in\mathbb{R},\ u_{i}+u_{k}>\mathrm{arcsinh}\ (-\xi_{ik})$.
-Then
-$\mathcal{U}^{(12)}_{ijk}(\xi)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}_{>0}\times\mathbb{R}^{2}_{<0}|$ $u_{j}+u_{k}>-\mathrm{arccosh}\ \xi_{jk},u_{i}+u_{j}>\mathrm{arcsinh}\ (-\xi_{ij}),$
-$u_{i}+u_{k}>\mathrm{arcsinh}\ (-\xi_{ik})\},$
-where $\xi_{jk}>1,\ \xi_{ij}\in\mathbb{R},\ \xi_{ik}\in\mathbb{R}$.
+Assume $(\alpha_{i},\alpha_{j},\alpha_{k})=(-1,1,1)$.
+It is known that $u_{j}+u_{k}>-\mathrm{arccosh}\,\eta_{jk}$ and $\eta_{jk}>1$.
+Moreover, $\eta_{ij}\in\mathbb{R}$ with $u_{i}+u_{j}>\mathrm{arcsinh}\,(-\eta_{ij})$, and $\eta_{ik}\in\mathbb{R}$ with $u_{i}+u_{k}>\mathrm{arcsinh}\,(-\eta_{ik})$.
+Consequently, the admissible space of $u$ is defined as
+$\mathcal{U}^{(12)}_{ijk}(\eta)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}_{>0}\times\mathbb{R}^{2}_{<0}\mid$ $\ u_{j}+u_{k}>-\mathrm{arccosh}\,\eta_{jk},\ u_{i}+u_{j}>\mathrm{arcsinh}\,(-\eta_{ij}),$
+$\ u_{i}+u_{k}>\mathrm{arcsinh}\,(-\eta_{ik})\},$
+where $\eta_{jk}>1$, $\eta_{ij}\in\mathbb{R}$, and $\eta_{ik}\in\mathbb{R}$.
 (XIII):
-Set $(\alpha_{i},\alpha_{j},\alpha_{k})=(0,1,-1)$.
-Then $u_{j}+u_{k}>\mathrm{arcsinh}\ (-\xi_{jk})$ and $\xi_{jk}>0$ in Subsection 3.1.
-Moreover, $(u_{i}+u_{j})\in\mathbb{R}$ for $\xi_{ij}\geq 0$ and $u_{i}+u_{j}>\log(-\xi_{ij})$ for $\xi_{ij}<0$.
-Similarly, $\xi_{ik}>0$ and $u_{i}+u_{k}<\log\xi_{ik}$.
-If $\xi_{ij}<0$, then $\log(-\xi_{ij})+\mathrm{arcsinh}\ (-\xi_{jk})<u_{i}+2u_{j}+u_{k}<2u_{j}+\log\xi_{ik}<\log\xi_{ik}$, which implies $(\xi_{ik}-\xi_{ij})(\xi_{ik}+\xi_{ij})>2\xi_{jk}\xi_{ij}\xi_{ik}$.
-Thus we assume $\xi_{ik}+\xi_{ij}>0$ for simplicity.
-Then
-$\mathcal{U}^{(13)}_{ijk}(\xi)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}\times\mathbb{R}^{2}_{<0}|$ $u_{j}+u_{k}>\mathrm{arcsinh}\ (-\xi_{jk}),\ e^{u_{i}+u_{j}}>-\xi_{ij},$
-$u_{i}+u_{k}<\log\xi_{ik}\},$
-where $\xi_{jk}>0,\ \xi_{ik}>0,\ \xi_{ij}\geq 0$ or $\xi_{ij}<0$$\xi_{ij}+\xi_{ik}>0$.
-The admissible space $\mathcal{U}^{(13)}_{ijk}(\xi)$ can be described explicitly as follows:
+Assume $(\alpha_{i},\alpha_{j},\alpha_{k})=(0,1,-1)$.
+In Subsection 3.1, it follows that $u_{j}+u_{k}>\mathrm{arcsinh}\ (-\eta_{jk})$ and $\eta_{jk}>0$.
+Moreover, $(u_{i}+u_{j})\in\mathbb{R}$ for $\eta_{ij}\geq 0$, and $u_{i}+u_{j}>\log(-\eta_{ij})$ for $\eta_{ij}<0$.
+Similarly, $\eta_{ik}>0$ and $u_{i}+u_{k}<\log\eta_{ik}$.
+When $\eta_{ij}<0$, the inequalities $\log(-\eta_{ij})+\mathrm{arcsinh}\,(-\eta_{jk})<u_{i}+2u_{j}+u_{k}<2u_{j}+\log\eta_{ik}<\log\eta_{ik}$ implies $(\eta_{ik}-\eta_{ij})(\eta_{ik}+\eta_{ij})>2\eta_{jk}\eta_{ij}\eta_{ik}$.
+For simplicity, we assume $\eta_{ik}+\eta_{ij}>0$.
+Consequently, the admissible space of $u$ is defined as
+$\mathcal{U}^{(13)}_{ijk}(\eta)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}\times\mathbb{R}^{2}_{<0}\mid$ $\ u_{j}+u_{k}>\mathrm{arcsinh}\,(-\eta_{jk}),\ e^{u_{i}+u_{j}}>-\eta_{ij},$
+$\ u_{i}+u_{k}<\log\eta_{ik}\},$
+where $\eta_{jk}>0$, $\eta_{ik}>0$, and $\eta_{ij}$ satisfies either $\eta_{ij}\geq 0$ or $\eta_{ij}<0$ with $\eta_{ij}+\eta_{ik}>0$.
+The admissible space $\mathcal{U}^{(13)}_{ijk}(\eta)$ can be explicitly characterized as follows:
 (i):
-For $\xi_{jk}>0,\ \xi_{ik}>0,\ \xi_{ij}\geq 0$, then
-$\mathcal{U}_{ijk}^{(13)}(\xi)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}\times\mathbb{R}^{2}_{<0}|u_{j}+u_{k}>\mathrm{arcsinh}\ (-\xi_{jk}),u_{i}+u_{k}<\log\xi_{ik}\}.$
+For $\eta_{jk}>0$, $\eta_{ik}>0$, and $\eta_{ij}\geq 0$, the admissible space is given by
+$\mathcal{U}_{ijk}^{(13)}(\eta)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}\times\mathbb{R}^{2}_{<0}\mid u_{j}+u_{k}>\mathrm{arcsinh}\,(-\eta_{jk}),u_{i}+u_{k}<\log\eta_{ik}\}.$
 (ii):
-For $\xi_{jk}>0,\ \xi_{ik}>0,\ \xi_{ij}<0$ and $\xi_{ij}+\xi_{ik}>0$, then
-$\mathcal{U}_{ijk}^{(13)}(\xi)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}\times\mathbb{R}^{2}_{<0}|$ $u_{j}+u_{k}>\mathrm{arcsinh}\ (-\xi_{jk}),\ u_{i}+u_{k}<\log\xi_{ik},$
-$u_{i}+u_{j}>\log(-\xi_{ij})\}.$
+For $\eta_{jk}>0$, $\eta_{ik}>0$, $\eta_{ij}<0$, and $\eta_{ij}+\eta_{ik}>0$, the admissible space is given by
+$\mathcal{U}_{ijk}^{(13)}(\eta)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}\times\mathbb{R}^{2}_{<0}\mid$ $\ u_{j}+u_{k}>\mathrm{arcsinh}\,(-\eta_{jk}),\ u_{i}+u_{k}<\log\eta_{ik},$
+$\ u_{i}+u_{j}>\log(-\eta_{ij})\}.$
 (XIV):
-Set $(\alpha_{i},\alpha_{j},\alpha_{k})=(1,1,-1)$.
-Then $u_{j}+u_{k}>\mathrm{arcsinh}\ (-\xi_{jk})$ and $\xi_{jk}>0$.
-Moreover, $\xi_{ij}>-1$ and $(u_{i}+u_{j})\in\mathbb{R}$.
-Similarly, $\xi_{ik}\in\mathbb{R}$ and $u_{i}+u_{k}<\mathrm{arcsinh}\ \xi_{ik}$.
-Since $u_{i}>0,u_{j}<0$, then $\mathrm{arcsinh}\ (-\xi_{jk})<u_{j}+u_{k}<u_{i}+u_{k}<\mathrm{arcsinh}\ \xi_{ik}$ and hence $\xi_{ik}+\xi_{jk}>0$.
-Then
-$\mathcal{U}^{(14)}_{ijk}(\xi)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}_{>0}\times\mathbb{R}^{2}_{<0}|u_{j}+u_{k}>\mathrm{arcsinh}\ (-\xi_{jk}),u_{i}+u_{k}<\mathrm{arcsinh}\ \xi_{ik}\},$
-where $\xi_{jk}>0,\ \xi_{ij}>-1,\ \xi_{ik}\in\mathbb{R}$ and $\xi_{ik}+\xi_{jk}>0$.
+Assume $(\alpha_{i},\alpha_{j},\alpha_{k})=(1,1,-1)$.
+It is known that $u_{j}+u_{k}>\mathrm{arcsinh}\,(-\eta_{jk})$ and $\eta_{jk}>0$.
+Moreover, $\eta_{ij}>-1$ and $(u_{i}+u_{j})\in\mathbb{R}$.
+Similarly, $\eta_{ik}\in\mathbb{R}$ and $u_{i}+u_{k}<\mathrm{arcsinh}\,\eta_{ik}$.
+Since $u_{i}>0$ and $u_{j}<0$, it follows that $\mathrm{arcsinh}\,(-\eta_{jk})<u_{j}+u_{k}<u_{i}+u_{k}<\mathrm{arcsinh}\,\eta_{ik}$, and hence $\eta_{ik}+\eta_{jk}>0$.
+Consequently, the admissible space of $u$ is defined as
+$\mathcal{U}^{(14)}_{ijk}(\eta)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}_{>0}\times\mathbb{R}^{2}_{<0}\mid u_{j}+u_{k}>\mathrm{arcsinh}\,(-\eta_{jk}),\ u_{i}+u_{k}<\mathrm{arcsinh}\,\eta_{ik}\},$
+where $\eta_{jk}>0$, $\eta_{ij}>-1$, $\eta_{ik}\in\mathbb{R}$, and $\eta_{ik}+\eta_{jk}>0$.
 (XV):
-Set $(\alpha_{i},\alpha_{j},\alpha_{k})=(-1,1,-1)$.
-Then $u_{j}+u_{k}>\mathrm{arcsinh}\ (-\xi_{jk})$ and $\xi_{jk}>0$.
-Moreover, $\xi_{ij}\in\mathbb{R}$ and $u_{i}+u_{j}>\mathrm{arcsinh}\ (-\xi_{ij})$.
-Similarly, $\xi_{ik}>1$ and $-\mathrm{arccosh}\ \xi_{ik}<u_{i}+u_{k}<\mathrm{arccosh}\ \xi_{ik}$.
-Then $\mathrm{arcsinh}\ (-\xi_{jk})+\mathrm{arcsinh}(-\xi_{ij})<u_{i}+2u_{j}+u_{k}<2u_{j}+\mathrm{arccosh}\xi_{ik}<\mathrm{arccosh}\xi_{ik}$.
-For simplicity, we require $\xi_{ij}>0$.
-Then
-$\mathcal{U}^{(15)}_{ijk}(\xi)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}_{>0}\times\mathbb{R}^{2}_{<0}|$ $u_{j}+u_{k}>\mathrm{arcsinh}\ (-\xi_{jk}),\ u_{i}+u_{j}>\mathrm{arcsinh}\ (-\xi_{ij}),$
-$-\mathrm{arccosh}\ \xi_{ik}<u_{i}+u_{k}<\mathrm{arccosh}\ \xi_{ik}\},$
-where $\xi_{jk}>0,\ \xi_{ij}>0,\ \xi_{ik}>1$.
+Assume $(\alpha_{i},\alpha_{j},\alpha_{k})=(-1,1,-1)$.
+It is known that $u_{j}+u_{k}>\mathrm{arcsinh}\,(-\eta_{jk})$ and $\eta_{jk}>0$.
+Moreover, $\eta_{ij}\in\mathbb{R}$ and $u_{i}+u_{j}>\mathrm{arcsinh}\,(-\eta_{ij})$.
+Similarly, $\eta_{ik}>1$ and $-\mathrm{arccosh}\,\eta_{ik}<u_{i}+u_{k}<\mathrm{arccosh}\,\eta_{ik}$.
+Combining these results, we derive the inequality chain: $\mathrm{arcsinh}\,(-\eta_{jk})+\mathrm{arcsinh}\,(-\eta_{ij})<u_{i}+2u_{j}+u_{k}<2u_{j}+\mathrm{arccosh}\,\eta_{ik}<\mathrm{arccosh}\,\eta_{ik}$.
+For simplicity, we assume $\eta_{ij}>0$.
+Consequently, the admissible space of $u$ is defined as
+$\mathcal{U}^{(15)}_{ijk}(\eta)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}_{>0}\times\mathbb{R}^{2}_{<0}\mid$ $\ u_{j}+u_{k}>\mathrm{arcsinh}\,(-\eta_{jk}),\ u_{i}+u_{j}>\mathrm{arcsinh}\,(-\eta_{ij}),$
+$\ -\mathrm{arccosh}\,\eta_{ik}<u_{i}+u_{k}<\mathrm{arccosh}\,\eta_{ik}\},$
+where $\eta_{jk}>0$, $\eta_{ij}>0$, and $\eta_{ik}>1$.
 (XVI):
-Set $(\alpha_{i},\alpha_{j},\alpha_{k})=(0,-1,-1)$.
-Then $u_{j}+u_{k}>-\mathrm{arccosh}\ \xi_{jk}$ and $\xi_{jk}>1$ in Subsection 3.1.
-Moreover, $\xi_{ij}>0,\ u_{i}+u_{j}<\log\xi_{ij}$ and $\xi_{ik}>0,\ u_{i}+u_{k}<\log\xi_{ik}$.
-Then
-$\mathcal{U}^{(16)}_{ijk}(\xi)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}\times\mathbb{R}^{2}_{<0}|$ $u_{j}+u_{k}>-\mathrm{arccosh}\ \xi_{jk},\ u_{i}+u_{j}<\log\xi_{ij},$
-$u_{i}+u_{k}<\log\xi_{ik}\},$
-where $\xi_{jk}>1,\ \xi_{ij}>0,\ \xi_{ik}>0$.
+Assume $(\alpha_{i},\alpha_{j},\alpha_{k})=(0,-1,-1)$.
+In Subsection 3.1, it follows that $u_{j}+u_{k}>-\mathrm{arccosh}\,\eta_{jk}$ and $\eta_{jk}>1$.
+Moreover, $\eta_{ij}>0$ with $u_{i}+u_{j}<\log\eta_{ij}$, and $\eta_{ik}>0$ with $u_{i}+u_{k}<\log\eta_{ik}$.
+Consequently, the admissible space of $u$ is defined as
+$\mathcal{U}^{(16)}_{ijk}(\eta)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}\times\mathbb{R}^{2}_{<0}\mid$ $\ u_{j}+u_{k}>-\mathrm{arccosh}\,\eta_{jk},\ u_{i}+u_{j}<\log\eta_{ij},$
+$\ u_{i}+u_{k}<\log\eta_{ik}\},$
+where $\eta_{jk}>1$, $\eta_{ij}>0$, and $\eta_{ik}>0$.
 (XVII):
-Set $(\alpha_{i},\alpha_{j},\alpha_{k})=(1,-1,-1)$.
-Then $u_{j}+u_{k}>-\mathrm{arccosh}\ \xi_{jk}$ and $\xi_{jk}>1$.
-Moreover, $\xi_{ij}\in\mathbb{R}$ and $u_{i}+u_{j}<\mathrm{arcsinh}\ \xi_{ij}$.
-Since $u_{i}>0$ and $u_{k}<0$, then $-\mathrm{arccosh}\ \xi_{jk}<u_{j}+u_{k}<u_{i}+u_{j}<\mathrm{arcsinh}\ \xi_{ij}$.
-For simplicity, we assume $\xi_{ij}>0$.
-Similarly, $\xi_{ik}\in\mathbb{R}$ and $u_{i}+u_{k}<\mathrm{arcsinh}\ \xi_{ik}$, then we assume $\xi_{ik}>0$ for simplicity.
-Then
+Assume $(\alpha_{i},\alpha_{j},\alpha_{k})=(1,-1,-1)$.
+It is known that $u_{j}+u_{k}>-\mathrm{arccosh}\,\eta_{jk}$ and $\eta_{jk}>1$.
+Moreover, $\eta_{ij}\in\mathbb{R}$ and $u_{i}+u_{j}<\mathrm{arcsinh}\,\eta_{ij}$.
+Since $u_{i}>0$ and $u_{k}<0$, it follows that $-\mathrm{arccosh}\,\eta_{jk}<u_{j}+u_{k}<u_{i}+u_{j}<\mathrm{arcsinh}\,\eta_{ij}$.
+For simplicity, we assume $\eta_{ij}>0$.
+Similarly, $\eta_{ik}\in\mathbb{R}$ and $u_{i}+u_{k}<\mathrm{arcsinh}\,\eta_{ik}$, thus we assume $\eta_{ik}>0$.
+Consequently, the admissible space of $u$ is defined as
 […]
-$u_{i}+u_{j}<\mathrm{arcsinh}\ \xi_{ij},\ u_{i}+u_{k}<\mathrm{arcsinh}\ \xi_{ik}\},$
-where $\xi_{jk}>1,\ \xi_{ij}>0,\ \xi_{ik}>0$.
+$\ u_{i}+u_{j}<\mathrm{arcsinh}\,\eta_{ij},\ u_{i}+u_{k}<\mathrm{arcsinh}\,\eta_{ik}\},$
+where $\eta_{jk}>1$, $\eta_{ij}>0$, and $\eta_{ik}>0$.
 (XVIII):
-Set $(\alpha_{i},\alpha_{j},\alpha_{k})=(-1,-1,-1)$.
-Then $u_{j}+u_{k}>-\mathrm{arccosh}\ \xi_{jk}$ and $\xi_{jk}>1$.
-Moreover, $\xi_{ij}>1$ and $-\mathrm{arccosh}\ \xi_{ij}<u_{i}+u_{j}<\mathrm{arccosh}\ \xi_{ij}$.
-Similarly, $\xi_{ik}>1$ and $-\mathrm{arccosh}\ \xi_{ik}<u_{i}+u_{k}<\mathrm{arccosh}\ \xi_{ik}$.
-Then
+Assume $(\alpha_{i},\alpha_{j},\alpha_{k})=(-1,-1,-1)$.
+It is known that $u_{j}+u_{k}>-\mathrm{arccosh}\,\eta_{jk}$ and $\eta_{jk}>1$.
+Moreover, $\eta_{ij}>1$ and $-\mathrm{arccosh}\,\eta_{ij}<u_{i}+u_{j}<\mathrm{arccosh}\,\eta_{ij}$.
+Similarly, $\eta_{ik}>1$ and $-\mathrm{arccosh}\,\eta_{ik}<u_{i}+u_{k}<\mathrm{arccosh}\,\eta_{ik}$.
+Consequently, the admissible space of $u$ is defined as
 […]
-$-\mathrm{arccosh}\ \xi_{ij}<u_{i}+u_{j}<\mathrm{arccosh}\ \xi_{ij},$
-$-\mathrm{arccosh}\ \xi_{ik}<u_{i}+u_{k}<\mathrm{arccosh}\ \xi_{ik}\},$
-where $\xi_{jk}>1,\ \xi_{ij}>1,\ \xi_{ik}>1$.
-In the above different admissible spaces $\mathcal{U}^{(n)}_{ijk}(\xi)$, the weight $\xi$ has different range of values.
-We are unable to use a uniform formula to express them.
-For simplicity, we use a notation $\Xi$ to denote the set of the range of weight $\xi$ in the whole admissible spaces $\mathcal{U}^{(n)}_{ijk}(\xi)$.
+$-\mathrm{arccosh}\,\eta_{ij}<u_{i}+u_{j}<\mathrm{arccosh}\,\eta_{ij},$
+$-\mathrm{arccosh}\,\eta_{ik}<u_{i}+u_{k}<\mathrm{arccosh}\,\eta_{ik}\},$
+where $\eta_{jk}>1$, $\eta_{ij}>1$, and $\eta_{ik}>1$.
+Across the various admissible spaces $\mathcal{U}^{(n)}_{ijk}(\eta)$ discussed above, the weight $\eta$ takes on different value ranges.
+These ranges cannot be expressed through a single consistent formula.
+For simplicity, we introduce the notation $\Xi$ to denote the set of all such weight ranges associated with the entire family of admissible spaces $\mathcal{U}^{(n)}_{ijk}(\eta)$, where $n=1,2,\dots,18$.
 Theorem 8.1.
-Suppose $(\Sigma,\mathcal{T},\alpha,\xi)$ is a weighted triangulated surface with boundary, and the weights $\alpha:B\rightarrow\{-1,0,1\}$ and $\xi\in\Xi$.
 Let $\{ijk\}\in F$ be a right-angled hyperbolic hexagon with edge lengths given by (53).
-Then the admissible space $\mathcal{U}^{(n)}_{ijk}(\xi)$ of $u$ has 18 types, each of which is convex polytope.
-As a result, the admissible space
-$\mathcal{U}(\xi)=\bigcap_{\{ijk\}\in F}\mathcal{U}_{ijk}(\xi),\ \mathcal{U}_{ijk}(\xi)\in\{\mathcal{U}^{(n)}_{ijk}(\xi),\ n=1,2,...,18\}$
-is also a convex polytope.
-Remark 6.2.
-The admissible space $\mathcal{U}(\xi)$ is a finite intersection of 18 types of admissible spaces $\mathcal{U}^{(n)}_{ijk}(\xi)$ for $n=1,2,...,18$.
-But this finite intersection is not arbitrary because of the connectedness of the admissible space $\mathcal{U}(\xi)$.
-For example, if there exist the admissible space $\mathcal{U}^{(1)}_{ijk}(\xi)$ and $\mathcal{U}^{(18)}_{ijk}(\xi)$ for some right-angled hyperbolic hexagons $\{ijk\}\in F$ with edge lengths given by (46), then there must be the third type of the admissible space $\mathcal{U}^{(n)}_{ijk}(\xi)$ for $n\neq 1,18$.
-6.1.1. Negative definiteness of Jacobian
+Then the admissible space $\mathcal{U}^{(n)}_{ijk}(\eta)$ of $u$ admits 18 distinct types, each of which is a convex polytope.
+We rewrite the admissible space in (35) as $\mathcal{U}^{(0)}_{ijk}(\eta)$. Combining Theorem 3.1 and Theorem 8.1 gives the following result.
+Corollary 8.2.
+Suppose $(\Sigma,\mathcal{T},\alpha,\eta)$ is a weighted triangulated surface with boundary, where $\alpha:B\rightarrow\{-1,0,1\}$. If one of the following conditions holds:
+(1):
+All right-angled hyperbolic hexagons are Type-I (i.e., the edge lengths of any right-angled hyperbolic hexagons are given by (53)), and the weight $\eta$ belongs to $\Xi$;
+(2):
+Not all right-angled hyperbolic hexagons are Type-I (i.e., there exist some right-angled hyperbolic hexagons with edge lengths given by (5)).
+In this case, for those right-angled hyperbolic hexagons with edge lengths given by (53), the weight $\eta$ belongs to $\Xi$, while for those with edge lengths given by (5), $\eta$ satisfies the condition in Theorem 3.1.
+Then the admissible space $\mathcal{U}(\eta)=\bigcap_{\{ijk\}\in F}\mathcal{U}_{ijk}(\eta)$ is a convex polytope on $(\Sigma,\mathcal{T},\eta)$.
+Here $\mathcal{U}_{ijk}(\eta)\in\{\mathcal{U}^{(n)}_{ijk}(\eta),\ n=0,1,2,...,18\}$.
+8.2. Negative definiteness of the Jacobian
 Theorem 8.3.
-Under the same assumptions as in Theorem 6.1.
-The Jacobian $\Lambda_{ijk}=\frac{\partial(\theta^{jk}_{i},\theta^{ik}_{j},\theta^{ij}_{k})}{\partial(u_{i},u_{j},u_{k})}$ is non-degenerate, symmetric and negative definite.
-As a consequence, the Jacobian $\Lambda=\frac{\partial(K_{i},...,K_{N})}{\partial(u_{i},...,u_{N})}$ is symmetric and negative definite on the admissible space $\mathcal{U}(\xi)$.
+Under the same assumptions as those in Theorem 8.1, the Jacobian $\Lambda_{ijk}=\frac{\partial(\theta^{jk}_{i},\theta^{ik}_{j},\theta^{ij}_{k})}{\partial(u_{i},u_{j},u_{k})}$ is symmetric and negative definite on $\mathcal{U}_{ijk}(\eta)$.
 Proof.
-For simplicity, we set $l_{r}=l_{st},\ \theta_{r}=\theta^{st}_{r}$ and $A=\sinh l_{r}\sinh l_{s}\sinh\theta_{t}$, where $\{r,s,t\}=\{i,j,k\}$. By (47), we have
-$\frac{\partial(f_{i},f_{j},f_{k})}{\partial(u_{i},u_{j},u_{k})}=\left(\begin{array}[]{ccc}-C_{i}&0&0\\ 0&C_{j}&0\\ 0&0&C_{k}\\ \end{array}\right):=Q_{6},$
-where $C_{r}=\sqrt{1+\alpha_{r}e^{2f_{r}}}$. Then
-$\Lambda_{ijk}=\frac{\partial(\theta_{i},\theta_{j},\theta_{k})}{\partial(u_{i},u_{j},u_{k})}=-\frac{1}{A}\left(\begin{array}[]{ccc}\sinh l_{i}&0&0\\ 0&\sinh l_{j}&0\\ 0&0&\sinh l_{k}\\ \end{array}\right)Q_{1}Q_{2}Q_{6},$
-where $Q_{1}$ is defined by (36) and $Q_{2}$ is defined by (38).
-Note that $d_{ij}<0,d_{ik}<0$ and the others are positive for any right-angled hyperbolic hexagon $\{ijk\}\in F$ with edge lengths given by (46).
-The rest of the proof is similar to that of Theorem 5.2, so we omit it.
+For simplicity, we adopt the notations introduction in Subsection 3.2. Formula (54) implies
+$\frac{\partial(f_{i},f_{j},f_{k})}{\partial(u_{i},u_{j},u_{k})}=\left(\begin{array}[]{ccc}-C_{i}&0&0\\ 0&C_{j}&0\\ 0&0&C_{k}\\ \end{array}\right):=Q_{8},$
+where $C_{r}=\sqrt{1+\alpha_{r}e^{2f_{r}}}$. Then $\Lambda_{ijk}$ can be expressed as
+$\Lambda_{ijk}=\frac{\partial(\theta_{i},\theta_{j},\theta_{k})}{\partial(u_{i},u_{j},u_{k})}=-\frac{1}{A}\left(\begin{array}[]{ccc}\sinh l_{i}&0&0\\ 0&\sinh l_{j}&0\\ 0&0&\sinh l_{k}\\ \end{array}\right)Q_{1}Q_{2}Q_{8},$
+where $Q_{1}$ and $Q_{2}$ are defined by (36) and (38), respectively.
+Furthermore, $\det Q_{1}>0$.
+For any right-angled hyperbolic hexagon $\{ijk\}\in F$ with edge lengths given by (53), since $d_{ij}<0$, $d_{ik}<0$, $d_{ji}>0$, $d_{ki}>0$, $d_{jk}>0$, and $d_{kj}>0$, it follows from (39) that $\det Q_{2}<0$.
+The remaining part of the proof is analogously to that of Theorem 6.3, and is thus omitted here.
 Q.E.D.
-6.2. Rigidity of the mixed discrete conformal structure $\mathrm{II}$
-The following theorem gives the rigidity of mixed discrete conformal structure $\mathrm{II}$, which generalizes the case $\mathrm{(iv)}$ in Theorem 1.9.
-As the proof is similar to that of Theorem 3.4, we omit it.
+Corollary 8.4.
+Under the same assumptions as those in Corollary 8.2, the Jacobian $\Lambda=\frac{\partial(K_{i},...,K_{N})}{\partial(u_{i},...,u_{N})}$ is symmetric and negative definite on $\mathcal{U}(\eta)$.
+8.3. Rigidity of the mixed discrete conformal structure I
+The following theorem establishes the rigidity of the mixed discrete conformal structure I, which generalizes the rigidity part of Theorem 1.8 (iv).
+As its proof is similar to that of Theorem 3.4, it is omitted here.
 Theorem 8.5.
-Under the same assumptions as in Theorem 6.1.
-The function $f$ is determined by its generalized combinatorial curvature $K\in\mathbb{R}^{E}_{>0}$.
-In particular, the map from $f$ to $K$ is a smooth embedding.
-6.3. Existence of the mixed discrete conformal structure $\mathrm{II}$
-We consider the image of $K$ for four types of the admissible spaces $(0,0,0),(1,0,0),(0,0,1),(1,0,1)$.
-In other words, we do not consider the case that $\alpha_{j}=\alpha_{k}=1$ for $\alpha\in\{0,1\}$.
-For simplicity, we use a new notation $\widetilde{\Xi}$ to denote the set of the range of weights $\xi$ in the admissible spaces $\mathcal{U}^{(n)}_{ijk}$, i.e., $n=1,2,4,5.$
+Under the same assumptions as those in Corollary 8.2, the discrete conformal factor $f$ is uniquely determined by its generalized combinatorial curvature $K\in\mathbb{R}^{N}_{>0}$.
+8.4. Existence of the mixed discrete conformal structure I
+For Type-I right-angled hyperbolic hexagons, we restrict our consideration to four types of admissible spaces, specifically corresponding to $(0,0,0)$, $(1,0,0)$, $(0,0,1)$, and $(1,0,1)$.
+These restrictions are motivated by two considerations: first, all cases where $(\alpha_{i},\alpha_{j},\alpha_{k})$ includes $-1$ are excluded, as the inclusion of $-1$ introduces significant technical complexities in the proof.
+Second, the case $\alpha_{j}=\alpha_{k}=1$ is also excluded, with a detailed explanation provided in Remark 8.10.
+For simplicity, we introduce a new notation $\widetilde{\Xi}$ to denote the set consisting of the ranges of weights $\eta$ across all admissible spaces $\mathcal{U}^{(n)}_{ijk}$ with $n=1,2,4,5$.
+The following theorem characterizes the image of $K$ for the mixed discrete conformal structure I, which generalizes the existence part of Theorem 1.8 (iv).
 Theorem 8.6.
-Suppose $(\Sigma,\mathcal{T},\alpha,\xi)$ is a weighted triangulated surface with boundary, and the weights $\alpha:B\rightarrow\{0,1\}$ and $\xi\in\widetilde{\Xi}$.
-Let $\{ijk\}\in F$ be a right-angled hyperbolic hexagon with edge lengths given by (46).
-If $\alpha_{j}$ and $\alpha_{k}$ can not take 1 simultaneously, then the image of $K$ is >$\mathbb{R}^{N}_{>0}$.
+Suppose $(\Sigma,\mathcal{T},\alpha,\eta)$ is a weighted triangulated surface with boundary, where $\alpha:B\rightarrow\{0,1\}$. If one of the following conditions holds:
+(1):
+All right-angled hyperbolic hexagons are Type-I (i.e., the edge lengths of any right-angled hyperbolic hexagon are given by (53)), $\alpha_{j}$ and $\alpha_{k}$ cannot both be 1 simultaneously, and $\eta\in\widetilde{\Xi}$;
+(2):
+Not all right-angled hyperbolic hexagons are Type-I (i.e., there exist some right-angled hyperbolic hexagons with edge lengths given by (5)).
+In this case, for those right-angled hyperbolic hexagons with edge lengths given by (53), $\alpha_{j}$ and $\alpha_{k}$ cannot both be 1 simultaneously, and $\eta\in\widetilde{\Xi}$, while for those with edge lengths given by (5), $\eta$ satisfies the condition in Theorem 3.1.
+Then the image of $K$ is >$\mathbb{R}^{N}_{>0}$.
 To prove Theorem 8.6, we need the following four lemmas.
 Lemma 8.7.
-Under the same assumptions as in Theorem 6.5. For any constants $c_{1},c_{2},c_{3}\in\mathbb{R}$, if one of the following conditions is satisfied
+Suppose $(\Sigma,\mathcal{T},\alpha,\eta)$ is a weighted triangulated surface with boundary, where the weights are given by $\alpha:B\rightarrow\{0,1\}$ and $\eta\in\widetilde{\Xi}$.
+Let $\{ijk\}\in F$ be a right-angled hyperbolic hexagon with edge lengths given by (53), and assume $\alpha_{j}$ and $\alpha_{k}$ cannot both be 1 simultaneously.
+If one of the following conditions is satisfied
 (i):
@@
 (iii):
-$\lim f_{i}=+\infty,\ \lim f_{j}=c_{2},\ \lim f_{k}=c_{3}$.
-Then $\lim\theta^{jk}_{i}(f_{i},f_{j},f_{k})=0$.
+$\lim f_{i}=+\infty,\ \lim f_{j}=c_{2},\ \lim f_{k}=c_{3}$,
+[…]
 Proof.
-By (46), we have
-(56) $\cosh l_{ij}=\bigg{[}\xi_{ij}+\sqrt{(e^{-2f_{i}}+\alpha_{i})(e^{-2f_{j}}+\alpha_{j})}\bigg{]}e^{f_{i}+f_{j}},$
-$\cosh l_{ik}=\bigg{[}\xi_{ik}+\sqrt{(e^{-2f_{i}}+\alpha_{i})(e^{-2f_{k}}+\alpha_{k})}\bigg{]}e^{f_{i}+f_{k}},$
-$\cosh l_{jk}=\bigg{[}\xi_{jk}-\sqrt{(e^{-2f_{j}}+\alpha_{j})(e^{-2f_{k}}+\alpha_{k})}\bigg{]}e^{f_{j}+f_{k}}.$
-For the case $\mathrm{(i)}$, $\lim f_{i}=+\infty,\ \lim f_{j}=+\infty,\ \lim f_{k}=+\infty$.
-Then $\lim\cosh l_{ij}:=\lim(c_{k}e^{f_{i}+f_{j}})$$\lim\cosh l_{ik}:=\lim(c_{j}e^{f_{i}+f_{k}})$ and $\lim\cosh l_{jk}:=\lim(c_{i}e^{f_{j}+f_{k}})$, where $c_{i},\ c_{j},\ c_{k}$ are positive constants.
-By the cosine law,
+Formula (53) implies
+(63) $\cosh l_{ij}=\bigg{[}\eta_{ij}+\sqrt{(e^{-2f_{i}}+\alpha_{i})(e^{-2f_{j}}+\alpha_{j})}\bigg{]}e^{f_{i}+f_{j}},$
+$\cosh l_{ik}=\bigg{[}\eta_{ik}+\sqrt{(e^{-2f_{i}}+\alpha_{i})(e^{-2f_{k}}+\alpha_{k})}\bigg{]}e^{f_{i}+f_{k}},$
+$\cosh l_{jk}=\bigg{[}\eta_{jk}-\sqrt{(e^{-2f_{j}}+\alpha_{j})(e^{-2f_{k}}+\alpha_{k})}\bigg{]}e^{f_{j}+f_{k}}.$
+For the case (i), it follows that $\lim\cosh l_{ij}:=\lim(c_{k}e^{f_{i}+f_{j}})$, $\lim\cosh l_{ik}:=\lim(c_{j}e^{f_{i}+f_{k}})$, and $\lim\cosh l_{jk}:=\lim(c_{i}e^{f_{j}+f_{k}})$, where $c_{i},c_{j},c_{k}$ are positive constants.
+By the cosine law, we have
 $\cosh\theta^{jk}_{i}=\frac{\cosh l_{jk}+\cosh l_{ij}\cosh l_{ik}}{\sinh l_{ij}\sinh l_{ik}}\rightarrow\frac{c_{i}e^{f_{j}+f_{k}}}{c_{k}e^{f_{i}+f_{j}}\cdot c_{j}e^{f_{i}+f_{k}}}+1\rightarrow 1,$
-which implies $\theta^{jk}_{i}\rightarrow 0$. The proofs of the cases $\mathrm{(ii)}$ and $\mathrm{(iii)}$ are similar, so we omit them here.
-Q.E.D.
+which implies $\theta^{jk}_{i}\rightarrow 0$. The proofs for the cases (ii) and (iii) follow analogously and are omitted here. Q.E.D.
 Lemma 8.8.
-Under the same assumptions as in Theorem 6.5. If $\lim f_{k}=-\infty$, then $\lim f_{j}=+\infty$. Moreover, if one of the following conditions is satisfied
+Under the same assumptions as those in Lemma 8.7, if $\lim f_{k}=-\infty$, then $\lim f_{j}=+\infty$. Moreover, if one of the following conditions is satisfied
 (i):
-$\lim f_{i}=+\infty,\ \lim f_{j}=+\infty,\ \lim f_{k}=-\infty$;
+$\lim f_{i}=+\infty,\ \lim f_{j}=+\infty,\ \lim f_{k}=-\infty$,
 (ii):
@@
 Combining (54) and Remark 2.22, if $\lim f_{k}=-\infty$, then $\lim u_{k}=-\infty$.
-Note that $l_{jk}>0$ implies that $u_{j}+u_{k}>C(\xi_{jk})$, where $C(\xi_{jk})$ is a constant that depends on $\xi_{jk}$.
-Then $\lim u_{j}=+\infty$.
-This implies $\alpha_{j}=0$ and $\lim f_{j}=+\infty$ by Remark 2.23.
+As established in Subsection 8.1, the condition $l_{jk}>0$ implies $u_{j}+u_{k}>C(\eta_{jk})$, where $C(\eta_{jk})$ is a constant depending on $\eta_{jk}$.
+This, in turn, forces $\lim u_{j}=+\infty$.
+By combining (54) and Remark 2.22 again, we have $\alpha_{j}=0$ and $\lim f_{j}=+\infty$.
 Set $\widetilde{L}_{1}=\frac{\cosh l_{ij}}{\sinh l_{ik}\sinh l_{jk}}$ and $\widetilde{L}_{2}=\frac{\cosh l_{ik}\cosh l_{jk}}{\sinh l_{ik}\sinh l_{jk}}$, then $\cosh\theta_{k}^{ij}=\widetilde{L}_{1}+\widetilde{L}_{2}$.
-Moreover, (46) is reduced to
-$\cosh l_{ij}$ $=\sqrt{1+\alpha_{i}e^{2f_{i}}}+\xi_{ij}e^{f_{i}+f_{j}},$
-$\cosh l_{ik}$ $\rightarrow\sqrt{1+\alpha_{i}e^{2f_{i}}}+\xi_{ik}e^{f_{i}+f_{k}},$
-$\cosh l_{jk}$ $\rightarrow-1+\xi_{jk}e^{f_{j}+f_{k}}.$
-For the case $\mathrm{(i)}$, $\lim f_{i}=+\infty,\lim f_{j}=+\infty,\ \lim f_{k}=-\infty$. Then $\lim\cosh l_{ij}:=\lim c_{k}e^{f_{i}+f_{j}}$$\lim\cosh l_{jk}=\lim(-1+\xi_{jk}e^{f_{j}+f_{k}})$ and
-$\lim\cosh l_{ik}=\begin{cases}\lim(1+\xi_{ik}e^{f_{i}+f_{k}}),\ &\text{if}\ \alpha_{i}=0,\\ \lim(c_{j}e^{f_{i}}),\ &\text{if}\ \alpha_{i}=1,\end{cases}$
+Moreover, (53) reduces to
+$\cosh l_{ij}$ $=\sqrt{1+\alpha_{i}e^{2f_{i}}}+\eta_{ij}e^{f_{i}+f_{j}},$
+$\cosh l_{ik}$ $\rightarrow\sqrt{1+\alpha_{i}e^{2f_{i}}}+\eta_{ik}e^{f_{i}+f_{k}},$
+$\cosh l_{jk}$ $\rightarrow-1+\eta_{jk}e^{f_{j}+f_{k}}.$
+For the case (i), it follows that $\lim\cosh l_{ij}:=\lim c_{k}e^{f_{i}+f_{j}}$, $\lim\cosh l_{jk}=\lim(-1+\eta_{jk}e^{f_{j}+f_{k}})$, and
+$\lim\cosh l_{ik}=\begin{cases}\lim(1+\eta_{ik}e^{f_{i}+f_{k}}),\ &\text{if}\ \alpha_{i}=0,\\ \lim(c_{j}e^{f_{i}}),\ &\text{if}\ \alpha_{i}=1,\end{cases}$
 where $c_{j},c_{k}$ are positive constants. If $\alpha_{i}=0$, we have
-$\widetilde{L}_{1}\rightarrow$ $\frac{c_{k}e^{f_{i}+f_{j}}}{\sqrt{(-1+\xi_{jk}e^{f_{j}+f_{k}})^{2}-1}\cdot\sqrt{(1+\xi_{ik}e^{f_{i}+f_{k}})^{2}-1}}$
-$=$ $\frac{c_{k}e^{f_{i}+f_{j}}}{\sqrt{\xi_{jk}\xi_{ik}}e^{f_{i}+f_{j}}}\cdot\frac{1}{\sqrt{\xi_{jk}e^{2f_{k}}-2e^{f_{k}-f_{j}}}}\cdot\frac{1}{\sqrt{\xi_{ik}e^{2f_{k}}+2e^{f_{k}-f_{i}}}}\rightarrow+\infty.$
-And $\widetilde{L}_{2}>0$.
-Then $\cosh\theta^{ij}_{k}=\widetilde{L}_{1}+\widetilde{L}_{2}\rightarrow+\infty$, which implies $\theta^{ij}_{k}\rightarrow+\infty$.
+$\widetilde{L}_{1}\rightarrow$ $\frac{c_{k}e^{f_{i}+f_{j}}}{\sqrt{(-1+\eta_{jk}e^{f_{j}+f_{k}})^{2}-1}\cdot\sqrt{(1+\eta_{ik}e^{f_{i}+f_{k}})^{2}-1}}$
+$=$ $\frac{c_{k}e^{f_{i}+f_{j}}}{\sqrt{\eta_{jk}\eta_{ik}}e^{f_{i}+f_{j}}}\cdot\frac{1}{\sqrt{\eta_{jk}e^{2f_{k}}-2e^{f_{k}-f_{j}}}}\cdot\frac{1}{\sqrt{\eta_{ik}e^{2f_{k}}+2e^{f_{k}-f_{i}}}}\rightarrow+\infty.$
+Since $\widetilde{L}_{2}>0$, it follows that $\cosh\theta^{ij}_{k}=\widetilde{L}_{1}+\widetilde{L}_{2}\rightarrow+\infty$, which implies $\theta^{ij}_{k}\rightarrow+\infty$.
 The proof of the case that $\alpha_{i}=1$ is similar, so we omit it here.
-For the case $\mathrm{(ii)}$, $\lim f_{i}=c,\lim f_{j}=+\infty,\ \lim f_{k}=-\infty$. Then $\lim\cosh l_{ij}:=\lim c^{\prime}_{k}e^{f_{j}}$$\lim\cosh l_{jk}=\lim(-1+\xi_{jk}e^{f_{j}+f_{k}})$ and
+For the case (ii), it follows that $\lim\cosh l_{ij}:=\lim c^{\prime}_{k}e^{f_{j}}$, $\lim\cosh l_{jk}=\lim(-1+\eta_{jk}e^{f_{j}+f_{k}})$, and
 $\lim\cosh l_{ik}=\begin{cases}1,\ &\text{if}\ \alpha_{i}=0,\\ c^{\prime}_{j},\ &\text{if}\ \alpha_{i}=1,\end{cases}$
 where $c^{\prime}_{j},c^{\prime}_{k}$ are positive constants.
-If $\alpha_{i}=0$, then $\cosh l_{ik}\rightarrow 1$ and hence $l_{ik}\rightarrow 0$.
-Then $\cosh\theta^{ij}_{k}>\widetilde{L}_{2}>\frac{\cosh l_{ik}}{\sinh l_{ik}}\rightarrow+\infty$ and hence $\theta^{ij}_{k}\rightarrow+\infty$.
-If $\alpha_{i}=1$, similar to the proof of the case $\mathrm{(i)}$, we still have $\widetilde{L}_{1}\rightarrow+\infty$.
-This implies $\theta^{ij}_{k}\rightarrow+\infty$.
+If $\alpha_{i}=0$, then $\cosh l_{ik}\rightarrow 1$, and hence $l_{ik}\rightarrow 0$.
+This implies $\cosh\theta^{ij}_{k}>\widetilde{L}_{2}>\frac{\cosh l_{ik}}{\sinh l_{ik}}\rightarrow+\infty$, and hence $\theta^{ij}_{k}\rightarrow+\infty$.
+If $\alpha_{i}=1$, similar to the proof of the case (i), we derive $\widetilde{L}_{1}\rightarrow+\infty$.
+Consequently, $\theta^{ij}_{k}\rightarrow+\infty$.
 This completes the proof.
@@
 Lemma 8.9.
-Under the same assumptions as in Theorem 6.5. For any constants $c_{1},c_{2},c_{3}\in\mathbb{R}$, if one of the following conditions is satisfied
+Under the same assumptions as those in Lemma 8.7, if one of the following conditions is satisfied
 (i):
-$\lim f_{i}=-\infty,\ \lim f_{j}=+\infty,\ \lim f_{k}=+\infty$;
+$\lim f_{i}=-\infty,\ \lim f_{j}=+\infty,\ \lim f_{k}=+\infty$,
 (ii):
-$\lim f_{i}=-\infty,\ \lim f_{j}=+\infty,\ \lim f_{k}=c_{1}$;
+$\lim f_{i}=-\infty,\ \lim f_{j}=+\infty,\ \lim f_{k}=c_{1}$,
 (iii):
-$\lim f_{i}=-\infty,\ \lim f_{j}=c_{2},\ \lim f_{k}=c_{3}$;
+$\lim f_{i}=-\infty,\ \lim f_{j}=c_{2},\ \lim f_{k}=c_{3}$,
 (iv):
-$\lim f_{i}=-\infty,\ \lim f_{j}=+\infty,\ \lim f_{k}=-\infty$.
-Then $\lim\theta^{jk}_{i}(f_{i},f_{j},f_{k})=+\infty$.
+$\lim f_{i}=-\infty,\ \lim f_{j}=+\infty,\ \lim f_{k}=-\infty$,
+where $c_{1},c_{2},c_{3}\in\mathbb{R}$ are constants. Then $\lim\theta^{jk}_{i}(f_{i},f_{j},f_{k})=+\infty$.
 Proof.
 Set $L_{1}=\frac{\cosh l_{jk}}{\sinh l_{ij}\sinh l_{ik}}$ and $L_{2}=\frac{\cosh l_{ij}\cosh l_{ik}}{\sinh l_{ij}\sinh l_{ik}}$, then $\cosh\theta^{jk}_{i}=L_{1}+L_{2}$.
-Since $\lim f_{i}=-\infty$, then (46) is reduced to
-(57) $\cosh l_{ij}$ $\rightarrow\sqrt{1+\alpha_{j}e^{2f_{j}}}+\xi_{ij}e^{f_{i}+f_{j}},$
-$\cosh l_{ik}$ $\rightarrow\sqrt{1+\alpha_{k}e^{2f_{k}}}+\xi_{ik}e^{f_{i}+f_{k}},$
-$\cosh l_{jk}$ $=-\sqrt{(1+\alpha_{j}e^{2f_{j}})(1+\alpha_{k}e^{2f_{k}})}+\xi_{jk}e^{f_{j}+f_{k}}.$
-For the case $\mathrm{(i)}$, $\lim f_{j}=+\infty,\ \lim f_{k}=+\infty$. By (57), we have $\lim\cosh l_{jk}:=\lim(c_{i}e^{f_{j}+f_{k}})$,
-$\lim\cosh l_{ij}=\begin{cases}\lim(1+\xi_{ij}e^{f_{i}+f_{j}}),\ &\text{if}\ \alpha_{j}=0,\\ \lim(c_{k}e^{f_{j}}),\ &\text{if}\ \alpha_{j}=1,\end{cases}$
+Given that $\lim f_{i}=-\infty$, the formula (53) reduces to
+(64) $\cosh l_{ij}$ $\rightarrow\sqrt{1+\alpha_{j}e^{2f_{j}}}+\eta_{ij}e^{f_{i}+f_{j}},$
+$\cosh l_{ik}$ $\rightarrow\sqrt{1+\alpha_{k}e^{2f_{k}}}+\eta_{ik}e^{f_{i}+f_{k}},$
+$\cosh l_{jk}$ $=-\sqrt{(1+\alpha_{j}e^{2f_{j}})(1+\alpha_{k}e^{2f_{k}})}+\eta_{jk}e^{f_{j}+f_{k}}.$
+For the case (i), it follows from (64) that $\lim\cosh l_{jk}:=\lim(c_{i}e^{f_{j}+f_{k}})$,
+$\lim\cosh l_{ij}=\begin{cases}\lim(1+\eta_{ij}e^{f_{i}+f_{j}}),\ &\text{if}\ \alpha_{j}=0,\\ \lim(c_{k}e^{f_{j}}),\ &\text{if}\ \alpha_{j}=1,\end{cases}$
 and
-$\lim\cosh l_{ik}=\begin{cases}\lim(1+\xi_{ik}e^{f_{i}+f_{k}}),\ &\text{if}\ \alpha_{k}=0,\\ \lim(c_{j}e^{f_{k}}),\ &\text{if}\ \alpha_{k}=1,\end{cases}$
+$\lim\cosh l_{ik}=\begin{cases}\lim(1+\eta_{ik}e^{f_{i}+f_{k}}),\ &\text{if}\ \alpha_{k}=0,\\ \lim(c_{j}e^{f_{k}}),\ &\text{if}\ \alpha_{k}=1,\end{cases}$
 where $c_{i},c_{j},c_{k}$ are positive constants. If $\alpha_{j}=\alpha_{k}=0$, then
-$L_{1}\rightarrow$ $\frac{c_{i}e^{f_{j}+f_{k}}}{\sqrt{(1+\xi_{ij}e^{f_{i}+f_{j}})^{2}-1}\cdot\sqrt{(1+\xi_{ik}e^{f_{i}+f_{k}})^{2}-1}}$
-$=$ $\frac{c_{i}e^{f_{j}+f_{k}}}{\sqrt{\xi_{ij}\xi_{ik}}e^{f_{j}+f_{k}}}\cdot\frac{1}{\sqrt{\xi_{ij}e^{2f_{i}}+2e^{f_{i}-f_{j}}}}\cdot\frac{1}{\sqrt{\xi_{ik}e^{2f_{i}}+2e^{f_{i}-f_{k}}}}\rightarrow+\infty.$
-And $L_{2}>0$.
-Thus $\cosh\theta^{jk}_{i}=L_{1}+L_{2}\rightarrow+\infty$, so $\theta^{jk}_{i}\rightarrow+\infty$.
-For the cases that $\alpha_{j}=0,\ \alpha_{k}=1$ and $\alpha_{j}=1,\ \alpha_{k}=0$, we still have $\theta^{jk}_{i}\rightarrow+\infty$ by similar proofs.
-For the case $\mathrm{(ii)}$, $\lim f_{j}=+\infty,\ \lim f_{k}=c_{1}$. By (57), we have $\lim\cosh l_{jk}:=\lim(c^{\prime}_{i}e^{f_{j}})$ and
+$L_{1}\rightarrow$ $\frac{c_{i}e^{f_{j}+f_{k}}}{\sqrt{(1+\eta_{ij}e^{f_{i}+f_{j}})^{2}-1}\cdot\sqrt{(1+\eta_{ik}e^{f_{i}+f_{k}})^{2}-1}}$
+$=$ $\frac{c_{i}e^{f_{j}+f_{k}}}{\sqrt{\eta_{ij}\eta_{ik}}e^{f_{j}+f_{k}}}\cdot\frac{1}{\sqrt{\eta_{ij}e^{2f_{i}}+2e^{f_{i}-f_{j}}}}\cdot\frac{1}{\sqrt{\eta_{ik}e^{2f_{i}}+2e^{f_{i}-f_{k}}}}\rightarrow+\infty.$
+Since $L_{2}>0$, it follows that $\cosh\theta^{jk}_{i}=L_{1}+L_{2}\rightarrow+\infty$, which implies $\theta^{jk}_{i}\rightarrow+\infty$.
+For the cases that $\alpha_{j}=0,\ \alpha_{k}=1$ and $\alpha_{j}=1,\ \alpha_{k}=0$, analogous reasoning yields $\theta^{jk}_{i}\rightarrow+\infty$.
+For the case (ii), it follows from (64) that $\lim\cosh l_{jk}:=\lim(c^{\prime}_{i}e^{f_{j}})$, and
 $\lim\cosh l_{ik}=\begin{cases}1,\ &\text{if}\ \alpha_{k}=0,\\ c^{\prime}_{j},\ &\text{if}\ \alpha_{k}=1,\end{cases}$
 where $c^{\prime}_{i},c^{\prime}_{j}$ are positive constants.
-If $\alpha_{k}=0$, then $l_{ik}\rightarrow 0$ and hence $L_{2}>\frac{\cosh l_{ik}}{\sinh l_{ik}}\rightarrow+\infty$.
+If $\alpha_{k}=0$, then $l_{ik}\rightarrow 0$, and hence $L_{2}>\frac{\cosh l_{ik}}{\sinh l_{ik}}\rightarrow+\infty$.
 This implies $\cosh\theta^{jk}_{i}\rightarrow+\infty$.
-If $\alpha_{k}=1$, then by the assumption, $\alpha_{j}=0$.
-Hence, $\lim\cosh l_{ij}=\lim(1+\xi_{ij}e^{f_{i}+f_{j}})$, $\lim\cosh l_{jk}=\lim(c^{\prime}_{i}e^{f_{j}})$ and $\lim\cosh l_{ik}=c^{\prime}_{j}$.
-Then
-$L_{1}\rightarrow$ $\frac{c^{\prime}_{i}e^{f_{j}}}{\sqrt{(1+\xi_{ij}e^{f_{i}+f_{j}})^{2}-1}\cdot\sqrt{(c^{\prime}_{j})^{2}-1}}$
-$=$ $\frac{c^{\prime}_{i}e^{f_{j}}}{\sqrt{\xi_{ij}}\sqrt{(c^{\prime}_{j})^{2}-1}e^{f_{j}}}\cdot\frac{1}{\sqrt{\xi_{ij}e^{2f_{i}}+2e^{f_{i}-f_{j}}}}\rightarrow+\infty.$
-And $L_{2}>0$. Thus $\cosh\theta^{jk}_{i}=L_{1}+L_{2}\rightarrow+\infty$, so $\theta^{jk}_{i}\rightarrow+\infty$.
-For the case $\mathrm{(iii)}$, $\lim f_{j}=c_{2},\ \lim f_{k}=c_{3}$. By (57), we have $\lim\cosh l_{jk}:=c^{\prime\prime}_{i}$ and
+If $\alpha_{k}=1$, then by the assumption, we have $\alpha_{j}=0$.
+Hence, $\lim\cosh l_{ij}=\lim(1+\eta_{ij}e^{f_{i}+f_{j}})$, $\lim\cosh l_{jk}=\lim(c^{\prime}_{i}e^{f_{j}})$, and $\lim\cosh l_{ik}=c^{\prime}_{j}$.
+Consequently,
+$L_{1}\rightarrow$ $\frac{c^{\prime}_{i}e^{f_{j}}}{\sqrt{(1+\eta_{ij}e^{f_{i}+f_{j}})^{2}-1}\cdot\sqrt{(c^{\prime}_{j})^{2}-1}}$
+$=$ $\frac{c^{\prime}_{i}e^{f_{j}}}{\sqrt{\eta_{ij}}\sqrt{(c^{\prime}_{j})^{2}-1}e^{f_{j}}}\cdot\frac{1}{\sqrt{\eta_{ij}e^{2f_{i}}+2e^{f_{i}-f_{j}}}}\rightarrow+\infty.$
+Since $L_{2}>0$, it follows that $\cosh\theta^{jk}_{i}=L_{1}+L_{2}\rightarrow+\infty$, which implies $\theta^{jk}_{i}\rightarrow+\infty$.
+For the case (iii), it follows from (64) that $\lim\cosh l_{jk}:=c^{\prime\prime}_{i}$, and
 $\lim\cosh l_{ij}=\begin{cases}1,\ &\text{if}\ \alpha_{j}=0,\\ c^{\prime\prime}_{k},\ &\text{if}\ \alpha_{j}=1,\end{cases}$
 where $c^{\prime\prime}_{i},c^{\prime\prime}_{k}$ are positive constants.
-If $\alpha_{j}=0$, then $l_{ij}\rightarrow 0$ and hence $L_{2}\rightarrow+\infty$.
+If $\alpha_{j}=0$, then $l_{ij}\rightarrow 0$, and hence $L_{2}\rightarrow+\infty$.
 This implies $\theta^{jk}_{i}\rightarrow+\infty$.
 If $\alpha_{j}=1$, then by the assumption, $\alpha_{k}=0$.
-According to the arguments in the case $\mathrm{(ii)}$, we still have $\theta^{jk}_{i}\rightarrow+\infty$.
-For the case $\mathrm{(iv)}$, $\lim f_{j}=+\infty,\ \lim f_{k}=-\infty$.
-By (57), we have $\lim\cosh l_{ik}=1$ and hence $l_{ik}\rightarrow 0$.
-Then $\cosh\theta^{jk}_{i}>L_{2}\geq\frac{\cosh l_{ik}}{\sinh l_{ik}}\rightarrow+\infty$ and then $\theta^{jk}_{i}\rightarrow+\infty$.
+According to the arguments in the case (ii), we still have $\theta^{jk}_{i}\rightarrow+\infty$.
+For the case (iv), it follows from (64) that $\lim\cosh l_{ik}=1$, and hence $l_{ik}\rightarrow 0$.
+Consequently, $\cosh\theta^{jk}_{i}>L_{2}\geq\frac{\cosh l_{ik}}{\sinh l_{ik}}\rightarrow+\infty$, which implies $\theta^{jk}_{i}\rightarrow+\infty$.
 This completes the proof.
@@
 Remark 8.10.
-According to the proof of the case $\mathrm{(iii)}$ in Lemma 6.8, if $\alpha_{j}=\alpha_{k}=1$, $\lim\cosh l_{ij}=c^{\prime\prime}_{k}$, $\lim\cosh l_{jk}=c^{\prime\prime}_{i}$ and $\lim\cosh l_{ik}=c^{\prime}_{j}$, then $\cosh\theta^{jk}_{i}$, $\cosh\theta^{ik}_{j}$ and $\cosh\theta^{ij}_{k}$ all tend to finite constants, which are far away from 1 and +$\infty$+ ∞.
+As shown in the proof of case (iii) in Lemma 8.9, if $\alpha_{j}=\alpha_{k}=1$, then $\lim\cosh l_{ij}=c^{\prime\prime}_{k}$, $\lim\cosh l_{jk}=c^{\prime\prime}_{i}$, and $\lim\cosh l_{ik}=c^{\prime}_{j}$.
+In this case, $\cosh\theta^{jk}_{i}$, $\cosh\theta^{ik}_{j}$ and $\cosh\theta^{ij}_{k}$ all tend to finite constants, which neither approach 1 nor +$\infty$+ ∞.
+Consequently, no contradiction can be derived from this case in the subsequent analysis, thereby justifying the exclusion of the case that $\alpha_{j}=\alpha_{k}=1$.
 Lemma 8.11.
-Under the same assumptions as in Theorem 6.5. For any constants $c_{1},c_{2},c_{3}\in\mathbb{R}$, if one of the following conditions is satisfied
+Under the same assumptions as those in Lemma 8.7, if one of the following conditions is satisfied
 (i):
-$\lim f_{j}=+\infty,\ \lim f_{i}=c_{1},\ \lim f_{k}=+\infty$;
+$\lim f_{j}=+\infty,\ \lim f_{i}=c_{1},\ \lim f_{k}=+\infty$,
 (ii):
-$\lim f_{j}=+\infty,\ \lim f_{i}=c_{2},\ \lim f_{k}=c_{3}$.
-Then $\lim\theta^{ik}_{j}(f_{i},f_{j},f_{k})=0$.
+$\lim f_{j}=+\infty,\ \lim f_{i}=c_{2},\ \lim f_{k}=c_{3}$,
+where $c_{1},c_{2},c_{3}\in\mathbb{R}$ are constants. Then $\lim\theta^{ik}_{j}(f_{i},f_{j},f_{k})=0$.
 Proof.
-For the case $\mathrm{(i)}$, by (56), we have $\lim\cosh l_{ij}:=\lim(c_{k}e^{f_{j}})$, $\lim\cosh l_{ik}:=c_{j}e^{f_{k}}$ and $\lim\cosh l_{jk}:=\lim(c_{i}e^{f_{j}+f_{k}})$, where $c_{i},c_{j},c_{k}$ are positive constants.
+For the case (i), by (63), we have $\lim\cosh l_{ij}:=\lim(c_{k}e^{f_{j}})$, $\lim\cosh l_{ik}:=c_{j}e^{f_{k}}$, and $\lim\cosh l_{jk}:=\lim(c_{i}e^{f_{j}+f_{k}})$, where $c_{i},c_{j},c_{k}$ are positive constants.
 Then
 $\cosh\theta^{ik}_{j}=\frac{\cosh l_{ik}+\cosh l_{ij}\cosh l_{jk}}{\sinh l_{ij}\sinh l_{jk}}\rightarrow\frac{c_{j}e^{f_{k}}}{c_{k}e^{f_{j}}\cdot c_{i}e^{f_{j}+f_{k}}}+1\rightarrow 1,$
-which implies $\theta^{ik}_{j}\rightarrow 0$. The proof of the case $\mathrm{(ii)}$ is similar, so we omit it here. Q.E.D.
-Proof of Theorem 6.5: Let the image of $K$ is $X$.
-The openness of $X$ in >$\mathbb{R}^{N}_{>0}$ follows from the injectivity of the map $f$.
-To prove the closeness of $X$ in >$\mathbb{R}^{N}_{>0}$, we prove that if a sequence $f^{(m)}$ in $\Omega(\xi)=\bigcap_{\{ijk\}\in F}\Omega_{ijk}(\xi)$ satisfies $\lim_{m\rightarrow+\infty}K^{(m)}\in\mathbb{R}^{N}_{>0}$, then there exists a subsequence, say $f^{(m)}$, such that $\lim_{m\rightarrow+\infty}f^{(m)}$ is in $\Omega(\xi)$.
-Suppose otherwise, there exists a subsequence $f^{(m)}$, such that $\lim_{m\rightarrow+\infty}f^{(m)}$ is in the boundary of $\Omega(\xi)$.
-The boundary of the admissible space $\Omega(\xi)$ in $[-\infty,+\infty]^{N}$ consists of the following three parts:
+which implies $\theta^{ik}_{j}\rightarrow 0$. The proof of the case (ii) is similar, so we omit it here. Q.E.D.
+Proof of Theorem 8.6: Let $X$ be the image of $K$.
+The openness of $X$ in >$\mathbb{R}^{N}_{>0}$ follows from the smooth injectivity of the map.
+To prove the closedness of $X$ in >$\mathbb{R}^{N}_{>0}$, we prove that if a sequence $\{f^{(m)}\}\in\Omega(\eta)=\bigcap_{\{ijk\}\in F}\Omega_{ijk}(\eta)$ satisfies $\lim_{m\rightarrow+\infty}K^{(m)}\in\mathbb{R}^{N}_{>0}$, then there exists a subsequence, still denoted by $\{f^{(m)}\}$, such that $\lim_{m\rightarrow+\infty}f^{(m)}\in\Omega(\eta)$.
+Suppose otherwise, there exists a subsequence $\{f^{(m)}\}$, such that $\lim_{m\rightarrow+\infty}f^{(m)}$ lies on the boundary of $\Omega(\eta)$.
+The boundary of the admissible space $\Omega(\eta)$ in $[-\infty,+\infty]^{N}$ consists of the following five parts:
 (i):
@@
 There exists $k\in B$ in a right-angled hyperbolic hexagon with edge lengths given by (53), such that $f_{k}^{(m)}\rightarrow\pm\infty$.
-For the case $\mathrm{(i)}$, in a right-angled hyperbolic hexagon $\{ijk\}\in F$ with lengths $l_{ij},l_{jk},l_{ik}$ and opposite lengths of boundaries $\theta^{ij}_{k},\theta^{jk}_{i},\theta^{ik}_{j}$.
-Without loss of generality, we assume $l_{ij}\rightarrow 0$.
-By the cosine law,
-$\cosh\theta^{ik}_{j}=\frac{\cosh l_{ik}+\cosh l_{ij}\cosh l_{jk}}{\sinh l_{ij}\sinh l_{jk}}>\frac{\cosh l_{ij}\cosh l_{jk}}{\sinh l_{ij}\sinh l_{jk}}\geq\frac{\cosh l_{ij}}{\sinh l_{ij}}\rightarrow+\infty,$
-which contradicts the assumption that $\lim_{m\rightarrow+\infty}K^{(m)}\in\mathbb{R}^{N}_{>0}$.
-We will prove that the above cases $\mathrm{(ii)}$, $\mathrm{(iii)}$ and $\mathrm{(iv)}$ can not hold on $\mathcal{U}^{(n)}_{ijk}(\xi)$, $n=1,2,4,5.$
-(1): For $(0,0,0)$ type admissible space $\mathcal{U}^{(1)}_{ijk}(\xi)$ in (48), we have $u_{j}+u_{k}>\log\frac{2}{\xi_{jk}}$.
+(v):
+There exists $j\in B$ in a right-angled hyperbolic hexagon with edge lengths given by (5), such that $f_{j}^{(m)}\rightarrow\pm\infty$.
+The proof for case (i) is analogous to that of Theorem 3.5, and is thus omitted here.
+We first establish that the aforementioned cases (ii), (iii), and (iv) cannot hold on $\mathcal{U}^{(n)}(\eta)=\bigcap_{\{ijk\}\in F}\mathcal{U}^{(n)}_{ijk}(\eta)$, where $n=1,2,4,5$.
+The corresponding results are stated as follows.
+  • (1)
+    For $(0,0,0)$-type admissible space $\mathcal{U}^{(1)}_{ijk}(\eta)$ in (55), we have $u_{j}+u_{k}>\log\frac{2}{\eta_{jk}}$.
 By Remark 2.22, $u_{j}=f_{j}$ and $u_{k}=f_{k}$, then $f_{j}+f_{k}$ has a lower bound.
-Hence, $f_{j},\ f_{k}$ can not tend to $\infty$- ∞ simultaneously.
+Hence, both $f_{j}$ and $f_{k}$ cannot tend to $\infty$- ∞ simultaneously.
 […]
@@
 […]
-(2): For $(1,0,0)$ type admissible space $\mathcal{U}^{(2)}_{ijk}(\xi)$ in (49), we have $u_{j}+u_{k}>\log\frac{2}{\xi_{jk}}$.
-The proof is the same as that of $(0,0,0)$ type admissible space $\mathcal{U}^{(1)}_{ijk}(\xi)$, so we omit it here.
-(3): For $(0,0,1)$ type admissible space $\mathcal{U}^{(4)}_{ijk}(\xi)$ in (50), the positions of $j$ and $k$ are not interchangeable.
-(i):
-In the admissible space (51), we have $u_{j}+u_{k}>\log\frac{1}{\xi_{jk}}$.
-By (47) and Remark 2.23, $u_{j}=f_{j}\in\mathbb{R}$ and $u_{k}=-\mathrm{arcsinh}(e^{-f_{k}})\in\mathbb{R}_{0}$, then $f_{j}$ can not tend to $\infty$- ∞.
+  • (2)
+    For $(1,0,0)$-type admissible space $\mathcal{U}^{(2)}_{ijk}(\eta)$ in (56), we have $u_{j}+u_{k}>\log\frac{2}{\eta_{jk}}$.
+The proof follows identically to that of the $(0,0,0)$-type admissible space $\mathcal{U}^{(1)}_{ijk}(\eta)$, and thus is omitted here.
+  • (3)
+    For $(0,0,1)$-type admissible space $\mathcal{U}^{(4)}_{ijk}(\eta)$ in (57), the positions of $j$ and $k$ are not interchangeable.
+    • (a)
+      In the admissible space (58), we have $u_{j}+u_{k}>\log\frac{1}{\eta_{jk}}$.
+Combining (54) and Remark 2.22 yields $u_{j}=f_{j}\in\mathbb{R}$ and $u_{k}=-\mathrm{arcsinh}\,(e^{-f_{k}})\in\mathbb{R}_{<0}$.
+It follows that $f_{j}$ cannot tend to $\infty$- ∞.
 […]
@@
 If $f_{j}^{(m)}\rightarrow+\infty$ and $f_{k}^{(m)}\nrightarrow-\infty$, by Lemma 8.11, then $(\theta^{ik}_{j})^{(m)}\rightarrow 0$ and hence $K^{(m)}_{j}\rightarrow 0$.
-Hence, the subsequence $f^{(m)}_{j}$ can not tend to +$\infty$+ ∞ in the admissible space (51).
+Consequently, the subsequence $f^{(m)}_{j}$ cannot tend to +$\infty$+ ∞ in the admissible space (58).
 […]
-(ii):
-In the admissible space (52), we have $u_{j}+u_{k}>\log\frac{1}{\xi_{jk}},\ u_{i}+u_{k}>\log(-\xi_{ik})$.
-According to the case $\mathrm{(i)}$, $f_{j}$ can not tend to $\infty$- ∞.
+    • (b)
+      In the admissible space (59), we have $u_{j}+u_{k}>\log\frac{1}{\eta_{jk}},\ u_{i}+u_{k}>\log(-\eta_{ik})$.
+The case (a) shows that $f_{j}$ cannot tend to $\infty$- ∞.
 Similarly, combining (54) and Remark 2.22 gives $u_{i}=-f_{i}$.
-Then $f_{i}$ can not tend to +$\infty$+ ∞.
+Thus $f_{i}$ cannot tend to +$\infty$+ ∞.
       If $f_{i}^{(m)}\rightarrow-\infty$, by Lemma 8.9, then $(\theta^{jk}_{i})^{(m)}\rightarrow+\infty$ and hence $K^{(m)}_{i}\rightarrow+\infty$.
-Hence, the subsequence $f^{(m)}_{i}$ can not tend to $\infty$- ∞ in the admissible space (52).
-If $f_{j}^{(m)}\rightarrow+\infty$ and $f_{k}^{(m)}\rightarrow-\infty$, then $f_{i}^{(m)}\rightarrow-\infty$ to ensure $u_{i}+u_{k}>\log(-\xi_{ik})$.
+Consequently, the subsequence $f^{(m)}_{i}$ cannot tend to $\infty$- ∞ in the admissible space (59).
+      If $f_{j}^{(m)}\rightarrow+\infty$ and $f_{k}^{(m)}\rightarrow-\infty$, then $f_{i}^{(m)}\rightarrow-\infty$ must hold to satisfy the inequality $u_{i}+u_{k}>\log(-\eta_{ik})$.
 This case has been included in the above arguments.
 If $f_{j}^{(m)}\rightarrow+\infty$ and $f_{k}^{(m)}\nrightarrow-\infty$, by Lemma 8.11, then $(\theta^{ik}_{j})^{(m)}\rightarrow 0$ and hence $K^{(m)}_{j}\rightarrow 0$.
-Hence, the subsequence $f^{(m)}_{j}$ can not tend to +$\infty$+ ∞ in the admissible space (52).
+Consequently, the subsequence $f^{(m)}_{j}$ cannot tend to +$\infty$+ ∞ in the admissible space (59).
 […]
-(4): For $(1,0,1)$ type admissible space $\mathcal{U}^{(5)}_{ijk}(\xi)$ in (53), we have $u_{j}+u_{k}>\log\frac{1}{\xi_{jk}}$.
-The proof is the same as that of $(0,0,1)$ type admissible space (51), so we omit it here.
+  • (4)
+    For $(1,0,1)$-type admissible space $\mathcal{U}^{(5)}_{ijk}(\eta)$ in (60), we have $u_{j}+u_{k}>\log\frac{1}{\eta_{jk}}$.
+The proof follows identically to that of the $(0,0,1)$-type admissible space (58), and thus is omitted here.
+From the foregoing proofs, it follows that the generalized angle tends to either 00 or +$\infty$+ ∞ on the admissible space $\mathcal{U}^{(n)}(\eta)$, which is composed of the single admissible subspace $\mathcal{U}^{(n)}_{ijk}(\eta)$ for $n=1,2,4,5$.
+Furthermore, for the case (v), the proof of Theorem 3.5 demonstrates that the generalized angle $\theta_{j}^{ik}$ tends to either 00 or +$\infty$+ ∞ on $\mathcal{U}^{(0)}_{ijk}(\eta)$.
+Through a lengthy yet analogous analysis, one can establish that the generalized angle tends to either 00 or +$\infty$+ ∞ on $\widetilde{\mathcal{U}}(\eta)$, which is formed by mixed admissible subspaces $\mathcal{U}^{(n)}_{ijk}(\eta)$ with $n=0,1,2,4,5$.
+Accordingly, the aforementioned cases (ii), (iii), (iv), and (v) cannot hold on $\widetilde{\mathcal{U}}(\eta)$.
+This completes the proof.
 Q.E.D.
-Remark 6.11.
-For $(0,1,1)$ type admissible space $\mathcal{U}^{(10)}_{ijk}(\xi)$ in (54) and $(1,1,1)$ type admissible space $\mathcal{U}^{(11)}_{ijk}(\xi)$ in (55), we are unable to prove the image of $K$ is >$\mathbb{R}^{N}_{>0}$.
-The reasons follow from Remark 6.9.
 Appendix A Proofs of Lemmas
-For simplicity, we set $h_{t}=h_{rs},\ q_{t}=q_{rs}$ and assume $||v_{i}||=||v_{j}||=||v_{k}||=1$. Please refer to Figure 3.
+For simplicity, we set $h_{t}=h_{rs}$ and $q_{t}=q_{rs}$, and assume $\|v_{i}\|=\|v_{j}\|=\|v_{k}\|=1$. Please refer to Figure 3.
 Lemma A.1.
@@
 Without loss of generality, we assume $r=i,s=j,t=k$.
-In the generalized right-angled hyperbolic triangle $c_{ijk}c_{ik}v_{i}$, by Proposition 2.1 we have $-v_{i}\ast c_{ijk}=(c_{ik}\ast c_{ijk})(v_{i}\ast c_{ik})$.
-By Proposition 2.2 $\mathrm{(i)}$, $c_{ik}\ast c_{ijk}=-\cosh h_{j}$.
-Moreover, by Proposition 2.2 $\mathrm{(ii)}$, if $c_{ik}$ and $v_{i}$ are on same sides of the hyperplane $v_{i}^{\bot}$, then
+In the generalized right-angled hyperbolic triangle $c_{ijk}c_{ik}v_{i}$, by Proposition 2.1, we have $-v_{i}\ast c_{ijk}=(c_{ik}\ast c_{ijk})(v_{i}\ast c_{ik})$.
+By Proposition 2.2 (i), $c_{ik}\ast c_{ijk}=-\cosh h_{j}$.
+Moreover, by Proposition 2.2 (ii), if $c_{ik}$ and $v_{i}$ lie on the same side of the hyperplane $v_{i}^{\bot}$, then
 $v_{i}\ast c_{ik}=+\sinh(d_{\mathbb{H}}(c_{ik},v_{i}^{\bot}))=+\sinh(-d_{ik})=-\sinh d_{ik};$
-if $c_{ik}$ and $v_{i}$ are on opposite sides of the hyperplane $v_{i}^{\bot}$, then
+if $c_{ik}$ and $v_{i}$ lie on opposite sides of the hyperplane $v_{i}^{\bot}$, then
 $v_{i}\ast c_{ik}=-\sinh(d_{\mathbb{H}}(c_{ik},v_{i}^{\bot}))=-\sinh(+d_{ik})=-\sinh d_{ik}.$
-Thus $v_{i}\ast c_{ik}=-\sinh d_{ik}$.
-Similarly, if $c_{ijk}$ and $v_{i}$ are on same sides of the hyperplane $v_{i}^{\bot}$, then $v_{i}\ast c_{ijk}=+\sinh(d_{\mathbb{H}}(c_{ijk},v_{i}^{\bot}))=+\sinh(-q_{i})=-\sinh q_{i}$; if $c_{ijk}$ and $v_{i}$ are on opposite sides of the hyperplane $v_{i}^{\bot}$, then $v_{i}\ast c_{ijk}=-\sinh(d_{\mathbb{H}}(c_{ijk},v_{i}^{\bot}))=-\sinh(+q_{i})=-\sinh q_{i}$.
-Thus $v_{i}\ast c_{ijk}=-\sinh q_{i}$.
-Then $\sinh q_{i}=\cosh h_{j}\cdot\sinh d_{ik}$.
-In the generalized right-angled hyperbolic triangle $c_{ijk}c_{ij}v_{i}$, we have $-v_{i}\ast c_{ijk}=(c_{ij}\ast c_{ijk})(v_{i}\ast c_{ij})=-\cosh h_{k}\cdot(v_{i}\ast c_{jk})$.
+In either case, $v_{i}\ast c_{ik}=-\sinh d_{ik}$. Similarly, if $c_{ijk}$ and $v_{i}$ lie on the same side of the hyperplane $v_{i}^{\bot}$, then
+$v_{i}\ast c_{ijk}=+\sinh(d_{\mathbb{H}}(c_{ijk},v_{i}^{\bot}))=+\sinh(-q_{i})=-\sinh q_{i};$
+if $c_{ijk}$ and $v_{i}$ lie on opposite sides of the hyperplane $v_{i}^{\bot}$, then
+$v_{i}\ast c_{ijk}=-\sinh(d_{\mathbb{H}}(c_{ijk},v_{i}^{\bot}))=-\sinh(+q_{i})=-\sinh q_{i}.$
+In either case, $v_{i}\ast c_{ijk}=-\sinh q_{i}$. Consequently, $\sinh q_{i}=\cosh h_{j}\cdot\sinh d_{ik}$.
+In the generalized right-angled hyperbolic triangle $c_{ijk}c_{ij}v_{i}$, we have $-v_{i}\ast c_{ijk}=(c_{ij}\ast c_{ijk})(v_{i}\ast c_{ij})=-\cosh h_{k}\cdot(v_{i}\ast c_{ij})$.
 Similar arguments can imply $v_{i}\ast c_{ij}=-\sinh d_{ij}$.
-Then $\sinh q_{i}=\cosh h_{k}\cdot\sinh d_{ij}$.
+Thus $\sinh q_{i}=\cosh h_{k}\cdot\sinh d_{ij}$.
 In the generalized right-angled hyperbolic triangle $c_{ijk}c^{\prime}_{ik}v^{\prime}_{i}$, we have $-v^{\prime}_{i}\ast c_{ijk}=(c^{\prime}_{ik}\ast c_{ijk})(v^{\prime}_{i}\ast c^{\prime}_{ik})=-\cosh q_{j}\cdot(v^{\prime}_{i}\ast c^{\prime}_{ik})$.
 Similarly, $v^{\prime}_{i}\ast c_{ijk}=-\sinh h_{i}$ and $v^{\prime}_{i}\ast c^{\prime}_{ik}=-\sinh\theta_{ik}$.
-Then $\sinh h_{i}=\cosh q_{j}\cdot\sinh\theta_{ik}$.
+Thus $\sinh h_{i}=\cosh q_{j}\cdot\sinh\theta_{ik}$.
 In the generalized right-angled hyperbolic triangle $c_{ijk}c^{\prime}_{ij}v^{\prime}_{i}$, we have $-v^{\prime}_{i}\ast c_{ijk}=(c^{\prime}_{ij}\ast c_{ijk})(v^{\prime}_{i}\ast c^{\prime}_{ij})=-\cosh q_{k}\cdot(v^{\prime}_{i}\ast c^{\prime}_{ij})$.
 Similarly, $v^{\prime}_{i}\ast c_{ijk}=-\sinh h_{i}$ and $v^{\prime}_{i}\ast c^{\prime}_{ij}=-\sinh\theta_{ij}$.
-Then $\sinh h_{i}=\cosh q_{k}\cdot\sinh\theta_{ij}$.
+Thus $\sinh h_{i}=\cosh q_{k}\cdot\sinh\theta_{ij}$.
 Q.E.D.
@@
 $\cosh h_{k}=-\sinh q_{i}\cdot\sinh\theta_{kj}=\sinh q_{j}\cdot\sinh\theta_{ki}.$
-If $c_{ijk}\in\mathbb{R}^{3}$ lies in the open domain $D_{\mathrm{II}}$, then we swap $h$ and $q$, and $\theta$ and $d$.
+If $c_{ijk}\in\mathbb{R}^{3}$ lies in the open domain $D_{\mathrm{II}}$, then we swap $h$ with $q$, and $\theta$ with $d$.
 Proof.
 (1):
-Suppose $c_{ijk}$ lies in the domain $D_{\mathrm{V}}$.
+Assume $c_{ijk}\in D_{\mathrm{V}}$.
 In the generalized right-angled hyperbolic triangle $c_{ijk}c_{ik}v_{i}$, by Proposition 2.1, we have $-v_{i}\ast c_{ijk}=(c_{ik}\ast c_{ijk})(v_{i}\ast c_{ik})$.
-By Proposition 2.2 $\mathrm{(iii)}$, $v_{i}\ast c_{ijk}=\cosh(d_{\mathbb{H}}(v_{i}^{\bot},c_{ijk}^{\bot}))=\cosh(-q_{i})=\cosh q_{i}$.
-Moreover, by Proposition 2.2 $\mathrm{(ii)}$, $v_{i}\ast c_{ik}=+\sinh d_{\mathbb{H}}(v_{i}^{\bot},c_{ik})=+\sinh(-d_{ik})=-\sinh d_{ik}$.
-If $c_{ik}$ and $c_{ijk}$ are on same sides of the hyperplane $c_{ijk}^{\bot}$, then
+By Proposition 2.2 (iii), $v_{i}\ast c_{ijk}=\cosh(d_{\mathbb{H}}(v_{i}^{\bot},c_{ijk}^{\bot}))=\cosh(-q_{i})=\cosh q_{i}$.
+Moreover, by Proposition 2.2 (ii), $v_{i}\ast c_{ik}=+\sinh d_{\mathbb{H}}(v_{i}^{\bot},c_{ik})=+\sinh(-d_{ik})=-\sinh d_{ik}$.
+If $c_{ik}$ and $c_{ijk}$ lie on the same side of the hyperplane $c_{ijk}^{\bot}$, then
 $c_{ik}\ast c_{ijk}=+\sinh(d_{\mathbb{H}}(c_{ik},c_{ijk}^{\bot}))=+\sinh(-h_{j})=-\sinh h_{j}.$
-If $c_{ik}$ and $c_{ijk}$ are on opposite sides of the hyperplane $c_{ijk}^{\bot}$, then
+If $c_{ik}$ and $c_{ijk}$ lie on opposite sides of the hyperplane $c_{ijk}^{\bot}$, then
 $c_{ik}\ast c_{ijk}=-\sinh(d_{\mathbb{H}}(c_{ik},c_{ijk}^{\bot}))=-\sinh h_{j}.$
-Thus $c_{ik}\ast c_{ijk}=-\sinh h_{j}$.
-Then $\cosh q_{i}=-\sinh h_{j}\cdot\sinh d_{ik}$.
+In either case, $c_{ik}\ast c_{ijk}=-\sinh h_{j}$.
+Consequently, $\cosh q_{i}=-\sinh h_{j}\cdot\sinh d_{ik}$.
 By Lemma 2.5, it is easy to check that $d_{ik}<0$.
-Hence $h_{j}>0$.
-Then $d_{\mathbb{H}}(c_{ik},c_{ijk}^{\bot})=h_{j}$, which implies that $c_{ik}$ and $c_{ijk}$ are on opposite sides of the hyperplane $c_{ijk}^{\bot}$.
+It follows that $h_{j}>0$ and hence $d_{\mathbb{H}}(c_{ik},c_{ijk}^{\bot})=h_{j}$.
+This implies that $c_{ik}$ and $c_{ijk}$ lie on opposite sides of $c_{ijk}^{\bot}$.
 In the generalized right-angled hyperbolic triangle $c_{ijk}c_{ij}v_{i}$, we have $-v_{i}\ast c_{ijk}=(c_{ij}\ast c_{ijk})(v_{i}\ast c_{ij})=-\cosh q_{i}$.
 Similarly, $v_{i}\ast c_{ij}=-\sinh d_{ij}$.
-If $c_{ij}$ and $c_{ijk}$ are on same sides of the hyperplane $c_{ijk}^{\bot}$, then $c_{ij}\ast c_{ijk}=+\sinh(d_{\mathbb{H}}(c_{ij},c_{ijk}^{\bot}))=+\sinh(-h_{k})=-\sinh h_{k}$.
-If $c_{ij}$ and $c_{ijk}$ are on opposite sides of the hyperplane $c_{ijk}^{\bot}$, then $c_{ij}\ast c_{ijk}=-\sinh(d_{\mathbb{H}}(c_{ij},c_{ijk}^{\bot}))=-\sinh h_{k}$.
-Thus $c_{ij}\ast c_{ijk}=-\sinh h_{k}$.
-Then $\cosh q_{i}=-\sinh h_{k}\cdot\sinh d_{ij}$.
+If $c_{ij}$ and $c_{ijk}$ lie on the same side of the hyperplane $c_{ijk}^{\bot}$, then $c_{ij}\ast c_{ijk}=+\sinh(d_{\mathbb{H}}(c_{ij},c_{ijk}^{\bot}))=+\sinh(-h_{k})=-\sinh h_{k}$.
+If $c_{ij}$ and $c_{ijk}$ lie on opposite sides of the hyperplane $c_{ijk}^{\bot}$, then $c_{ij}\ast c_{ijk}=-\sinh(d_{\mathbb{H}}(c_{ij},c_{ijk}^{\bot}))=-\sinh h_{k}$.
+In either case, $c_{ij}\ast c_{ijk}=-\sinh h_{k}$.
+Consequently, $\cosh q_{i}=-\sinh h_{k}\cdot\sinh d_{ij}$.
 Similarly, $d_{ij}<0$ by Lemma 2.5 and hence $h_{k}>0$.
-So $d_{\mathbb{H}}(c_{ij},c_{ijk}^{\bot})=h_{k}$, which implies that $c_{ij}$ and $c_{ijk}$ are on opposite sides of the hyperplane $c_{ijk}^{\bot}$.
+So $d_{\mathbb{H}}(c_{ij},c_{ijk}^{\bot})=h_{k}$, which implies that $c_{ij}$ and $c_{ijk}$ lie on opposite sides of $c_{ijk}^{\bot}$.
 In the generalized right-angled hyperbolic triangle $c_{ijk}c_{jk}v_{j}$, we have $-v_{j}\ast c_{ijk}=(c_{jk}\ast c_{ijk})(v_{j}\ast c_{jk})$.
-By Proposition 2.2 $\mathrm{(iii)}$, $v_{j}\ast c_{ijk}=-\cosh(d_{\mathbb{H}}(v_{j}^{\bot},c_{ijk}^{\bot}))=-\cosh q_{j}$.
-Moreover, by Proposition 2.2 $\mathrm{(ii)}$, $c_{jk}\ast c_{ijk}=-\sinh(d_{\mathbb{H}}(c_{jk},c_{ijk}^{\bot}))=-\sinh h_{i}$.
+By Proposition 2.2 (iii), $v_{j}\ast c_{ijk}=-\cosh(d_{\mathbb{H}}(v_{j}^{\bot},c_{ijk}^{\bot}))=-\cosh q_{j}$.
+Moreover, by Proposition 2.2 (ii), $c_{jk}\ast c_{ijk}=-\sinh(d_{\mathbb{H}}(c_{jk},c_{ijk}^{\bot}))=-\sinh h_{i}$.
 Similarly, $v_{j}\ast c_{jk}=-\sinh d_{jk}$.
-Then $\cosh q_{j}=\sinh h_{i}\cdot\sinh d_{jk}$.
+Thus $\cosh q_{j}=\sinh h_{i}\cdot\sinh d_{jk}$.
 In the generalized right-angled hyperbolic triangle $c_{ijk}c_{ij}v_{j}$, we have $-v_{j}\ast c_{ijk}=(c_{ij}\ast c_{ijk})(v_{j}\ast c_{ij})$.
-Since $v_{j}\ast c_{ijk}=-\cosh q_{j}$, $v_{j}\ast c_{ij}=-\sinh d_{ji}$ and $c_{ij}\ast c_{ijk}=-\sinh h_{k}$, then $\cosh q_{j}=\sinh h_{k}\cdot\sinh d_{ji}$.
+Since $v_{j}\ast c_{ijk}=-\cosh q_{j}$, $v_{j}\ast c_{ij}=-\sinh d_{ji}$, and $c_{ij}\ast c_{ijk}=-\sinh h_{k}$, it follows that $\cosh q_{j}=\sinh h_{k}\cdot\sinh d_{ji}$.
 In the generalized right-angled hyperbolic triangle $c_{ijk}c_{jk}v_{k}$, we have $-v_{k}\ast c_{ijk}=(c_{jk}\ast c_{ijk})(v_{k}\ast c_{jk})$.
-Since $v_{k}\ast c_{ijk}=-\cosh q_{k}$, $v_{k}\ast c_{jk}=-\sinh d_{kj}$ and $c_{jk}\ast c_{ijk}=-\sinh h_{i}$, then $\cosh q_{k}=\sinh h_{i}\cdot\sinh d_{kj}$.
+Since $v_{k}\ast c_{ijk}=-\cosh q_{k}$, $v_{k}\ast c_{jk}=-\sinh d_{kj}$, and $c_{jk}\ast c_{ijk}=-\sinh h_{i}$, it follows that $\cosh q_{k}=\sinh h_{i}\cdot\sinh d_{kj}$.
 In the generalized right-angled hyperbolic triangle $c_{ijk}c_{ik}v_{k}$, we have $-v_{k}\ast c_{ijk}=(c_{ik}\ast c_{ijk})(v_{k}\ast c_{ik})$.
-Since $v_{k}\ast c_{ijk}=-\cosh q_{k}$, $v_{k}\ast c_{ik}=-\sinh d_{ki}$ and $c_{ik}\ast c_{ijk}=-\sinh h_{j}$, then $\cosh q_{k}=\sinh h_{j}\cdot\sinh d_{ki}$.
+Since $v_{k}\ast c_{ijk}=-\cosh q_{k}$, $v_{k}\ast c_{ik}=-\sinh d_{ki}$, and $c_{ik}\ast c_{ijk}=-\sinh h_{j}$, it follows that $\cosh q_{k}=\sinh h_{j}\cdot\sinh d_{ki}$.
 In the generalized right-angled hyperbolic triangle $c_{ijk}c^{\prime}_{ik}v^{\prime}_{i}$, we have $-v^{\prime}_{i}\ast c_{ijk}=(c^{\prime}_{ik}\ast c_{ijk})(v^{\prime}_{i}\ast c^{\prime}_{ik})$.
-Since $v^{\prime}_{i}\ast c_{ijk}=-\cosh h_{i}$, $v^{\prime}_{i}\ast c^{\prime}_{ik}=-\sinh\theta_{ik}$ and $c^{\prime}_{ik}\ast c_{ijk}=-\sinh q_{j}$, then $\cosh h_{i}=\sinh q_{j}\cdot\sinh\theta_{ik}$.
+Since $v^{\prime}_{i}\ast c_{ijk}=-\cosh h_{i}$, $v^{\prime}_{i}\ast c^{\prime}_{ik}=-\sinh\theta_{ik}$, and $c^{\prime}_{ik}\ast c_{ijk}=-\sinh q_{j}$, it follows that $\cosh h_{i}=\sinh q_{j}\cdot\sinh\theta_{ik}$.
 In the generalized right-angled hyperbolic triangle $c_{ijk}c^{\prime}_{ij}v^{\prime}_{i}$, we have $-v^{\prime}_{i}\ast c_{ijk}=(c^{\prime}_{ij}\ast c_{ijk})(v^{\prime}_{i}\ast c^{\prime}_{ij})$.
-Since $v^{\prime}_{i}\ast c_{ijk}=-\cosh h_{i}$, $v^{\prime}_{i}\ast c^{\prime}_{ij}=-\sinh\theta_{ij}$ and $c^{\prime}_{ij}\ast c_{ijk}=-\sinh q_{k}$, then $\cosh h_{i}=\sinh q_{k}\cdot\sinh\theta_{ij}$.
+Since $v^{\prime}_{i}\ast c_{ijk}=-\cosh h_{i}$, $v^{\prime}_{i}\ast c^{\prime}_{ij}=-\sinh\theta_{ij}$, and $c^{\prime}_{ij}\ast c_{ijk}=-\sinh q_{k}$, it follows that $\cosh h_{i}=\sinh q_{k}\cdot\sinh\theta_{ij}$.
 In the generalized right-angled hyperbolic triangle $c_{ijk}c^{\prime}_{jk}v^{\prime}_{j}$, we have $-v^{\prime}_{j}\ast c_{ijk}=(c^{\prime}_{jk}\ast c_{ijk})(v^{\prime}_{j}\ast c^{\prime}_{jk})$.
-By Proposition 2.2 $\mathrm{(iii)}$, $v^{\prime}_{j}\ast c_{ijk}=-\cosh(d_{\mathbb{H}}((v^{\prime}_{j})^{\bot},c_{ijk}^{\bot}))=-\cosh h_{j}$.
-Since $c^{\prime}_{jk}$ and $c_{ijk}$ are on opposite sides of the hyperplane $c_{ijk}^{\bot}$, then $c^{\prime}_{jk}\ast c_{ijk}=-\sinh(d_{\mathbb{H}}(c^{\prime}_{jk},c_{ijk}^{\bot}))=-\sinh(-q_{i})=+\sinh q_{i}$.
-Due to $v^{\prime}_{j}\ast c^{\prime}_{jk}=-\sinh\theta_{jk}$, then $\cosh h_{j}=-\sinh q_{i}\cdot\sinh\theta_{jk}$.
+By Proposition 2.2 (iii), $v^{\prime}_{j}\ast c_{ijk}=-\cosh(d_{\mathbb{H}}((v^{\prime}_{j})^{\bot},c_{ijk}^{\bot}))=-\cosh h_{j}$.
+Since $c^{\prime}_{jk}$ and $c_{ijk}$ lie on opposite sides of the hyperplane $c_{ijk}^{\bot}$, it follows that $c^{\prime}_{jk}\ast c_{ijk}=-\sinh(d_{\mathbb{H}}(c^{\prime}_{jk},c_{ijk}^{\bot}))=-\sinh(-q_{i})=+\sinh q_{i}$.
+Due to $v^{\prime}_{j}\ast c^{\prime}_{jk}=-\sinh\theta_{jk}$, we conclude that $\cosh h_{j}=-\sinh q_{i}\cdot\sinh\theta_{jk}$.
 In the generalized right-angled hyperbolic triangle $c_{ijk}c^{\prime}_{ij}v^{\prime}_{j}$, we have $-v^{\prime}_{j}\ast c_{ijk}=(c^{\prime}_{ij}\ast c_{ijk})(v^{\prime}_{j}\ast c^{\prime}_{ij})$.
-Note that $v^{\prime}_{j}\ast c_{ijk}=-\cosh h_{j}$, $v^{\prime}_{j}\ast c^{\prime}_{ij}=-\sinh\theta_{ji}$ and $c^{\prime}_{ij}\ast c_{ijk}=-\sinh q_{k}$, then $\cosh h_{j}=\sinh q_{k}\cdot\sinh\theta_{ji}$.
+Since $v^{\prime}_{j}\ast c_{ijk}=-\cosh h_{j}$, $v^{\prime}_{j}\ast c^{\prime}_{ij}=-\sinh\theta_{ji}$, and $c^{\prime}_{ij}\ast c_{ijk}=-\sinh q_{k}$, it follows that $\cosh h_{j}=\sinh q_{k}\cdot\sinh\theta_{ji}$.
 In the generalized right-angled hyperbolic triangle $c_{ijk}c^{\prime}_{jk}v^{\prime}_{k}$, we have $-v^{\prime}_{k}\ast c_{ijk}=(c^{\prime}_{jk}\ast c_{ijk})(v^{\prime}_{k}\ast c^{\prime}_{jk})$.
-Since $v^{\prime}_{k}\ast c_{ijk}=-\cosh h_{k}$, $v^{\prime}_{k}\ast c^{\prime}_{jk}=-\sinh\theta_{kj}$ and $c^{\prime}_{jk}\ast c_{ijk}=\sinh q_{i}$, then $\cosh h_{k}=-\sinh q_{i}\cdot\sinh\theta_{kj}$.
-In the generalized right-angled hyperbolic triangle $c_{ijk}c^{\prime}_{ik}v^{\prime}_{k}$, we have $-v^{\prime}_{k}\ast c_{ijk}=(c^{\prime}_{ik}\ast c_{ijk})(v^{\prime}_{k}\ast c^{\prime}_{ik})$.
-Since $v^{\prime}_{k}\ast c_{ijk}=-\cosh h_{k}$, $v^{\prime}_{k}\ast c^{\prime}_{ik}=-\sinh\theta_{ki}$ and $c^{\prime}_{ik}\ast c_{ijk}=-\sinh q_{j}$, then $\cosh h_{k}=\sinh q_{j}\cdot\sinh\theta_{ki}$.
+Since $v^{\prime}_{k}\ast c_{ijk}=-\cosh h_{k}$, $v^{\prime}_{k}\ast c^{\prime}_{jk}=-\sinh\theta_{kj}$, and $c^{\prime}_{jk}\ast c_{ijk}=\sinh q_{i}$, it follows that $\cosh h_{k}=-\sinh q_{i}\cdot\sinh\theta_{kj}$.
+[…]
 (2):
-Suppose $c_{ijk}$ lies in the domain $D_{\mathrm{II}}$.
+Assume $c_{ijk}\in D_{\mathrm{II}}$.
 In the generalized right-angled hyperbolic triangle $c_{ijk}c_{ik}v_{i}$, we have $-v_{i}\ast c_{ijk}=(c_{ik}\ast c_{ijk})(v_{i}\ast c_{ik})$.
-By Proposition 2.2 $\mathrm{(iii)}$, $v_{i}\ast c_{ijk}=-\cosh q_{i}$ and $c_{ik}\ast c_{ijk}=-\sinh h_{j}$.
-Since $v_{i}\ast c_{ik}=-\sinh d_{\mathbb{H}}(v_{i}^{\bot},c_{ik})=-\sinh(+d_{ik})=-\sinh d_{ik}$, then $\cosh q_{i}=\sinh h_{j}\cdot\sinh d_{ik}$.
+By Proposition 2.2 (iii), $v_{i}\ast c_{ijk}=-\cosh q_{i}$ and $c_{ik}\ast c_{ijk}=-\sinh h_{j}$.
+Since $v_{i}\ast c_{ik}=-\sinh d_{\mathbb{H}}(v_{i}^{\bot},c_{ik})=-\sinh(+d_{ik})=-\sinh d_{ik}$, it follows that $\cosh q_{i}=\sinh h_{j}\cdot\sinh d_{ik}$.
 In the generalized right-angled hyperbolic triangle $c_{ijk}c_{ij}v_{i}$, we have $-v_{i}\ast c_{ijk}=(c_{ij}\ast c_{ijk})(v_{i}\ast c_{ij})=\cosh q_{i}$.
-Similarly, $v_{i}\ast c_{ij}=-\sinh d_{ij}$ and $c_{ij}\ast c_{ijk}=-\sinh h_{k}$.
-Then $\cosh q_{i}=\sinh h_{k}\cdot\sinh d_{ij}$.
+Since $v_{i}\ast c_{ij}=-\sinh d_{ij}$ and $c_{ij}\ast c_{ijk}=-\sinh h_{k}$, it follows that $\cosh q_{i}=\sinh h_{k}\cdot\sinh d_{ij}$.
 In the generalized right-angled hyperbolic triangle $c_{ijk}c_{jk}v_{j}$, we have $-v_{j}\ast c_{ijk}=(c_{jk}\ast c_{ijk})(v_{j}\ast c_{jk})$.
-By Proposition 2.2 $\mathrm{(iii)}$, then $v_{j}\ast c_{ijk}=-\cosh(d_{\mathbb{H}}(v_{j}^{\bot},c_{ijk}^{\bot}))=-\cosh q_{j}$.
-Since $c_{jk}$ and $c_{ijk}$ are on opposite sides of the hyperplane $c_{ijk}^{\bot}$, then $c_{jk}\ast c_{ijk}=-\sinh(d_{\mathbb{H}}(c_{jk},c_{ijk}^{\bot}))=-\sinh(-h_{i})=+\sinh h_{i}$.
-Due to $v_{j}\ast c_{jk}=-\sinh d_{jk}$, then $\cosh q_{j}=-\sinh h_{i}\cdot\sinh d_{jk}$.
+By Proposition 2.2 (iii), $v_{j}\ast c_{ijk}=-\cosh(d_{\mathbb{H}}(v_{j}^{\bot},c_{ijk}^{\bot}))=-\cosh q_{j}$.
+Since $c_{jk}$ and $c_{ijk}$ lie on opposite sides of the hyperplane $c_{ijk}^{\bot}$, it follows that $c_{jk}\ast c_{ijk}=-\sinh(d_{\mathbb{H}}(c_{jk},c_{ijk}^{\bot}))=-\sinh(-h_{i})=+\sinh h_{i}$.
+Due to $v_{j}\ast c_{jk}=-\sinh d_{jk}$, we conclude that $\cosh q_{j}=-\sinh h_{i}\cdot\sinh d_{jk}$.
 In the generalized right-angled hyperbolic triangle $c_{ijk}c_{ij}v_{j}$, we have $-v_{j}\ast c_{ijk}=(c_{ij}\ast c_{ijk})(v_{j}\ast c_{ij})$.
-Since $v_{j}\ast c_{ijk}=-\cosh q_{j}$, $v_{j}\ast c_{ij}=-\sinh d_{ji}$ and $c_{ij}\ast c_{ijk}=-\sinh h_{k}$, then $\cosh q_{j}=\sinh h_{k}\cdot\sinh d_{ji}$.
+Since $v_{j}\ast c_{ijk}=-\cosh q_{j}$, $v_{j}\ast c_{ij}=-\sinh d_{ji}$, and $c_{ij}\ast c_{ijk}=-\sinh h_{k}$, it follows that $\cosh q_{j}=\sinh h_{k}\cdot\sinh d_{ji}$.
 In the generalized right-angled hyperbolic triangle $c_{ijk}c_{jk}v_{k}$, we have $-v_{k}\ast c_{ijk}=(c_{jk}\ast c_{ijk})(v_{k}\ast c_{jk})$.
-Since $v_{k}\ast c_{ijk}=-\cosh q_{k}$, $v_{k}\ast c_{jk}=-\sinh d_{kj}$ and $c_{jk}\ast c_{ijk}=\sinh h_{i}$, then $\cosh q_{k}=-\sinh h_{i}\cdot\sinh d_{kj}$.
+Since $v_{k}\ast c_{ijk}=-\cosh q_{k}$, $v_{k}\ast c_{jk}=-\sinh d_{kj}$, and $c_{jk}\ast c_{ijk}=\sinh h_{i}$, it follows that $\cosh q_{k}=-\sinh h_{i}\cdot\sinh d_{kj}$.
 In the generalized right-angled hyperbolic triangle $c_{ijk}c_{ik}v_{k}$, we have $-v_{k}\ast c_{ijk}=(c_{ik}\ast c_{ijk})(v_{k}\ast c_{ik})$.
-Since $v_{k}\ast c_{ijk}=-\cosh q_{k}$, $v_{k}\ast c_{ik}=-\sinh d_{ki}$ and $c_{ik}\ast c_{ijk}=-\sinh h_{j}$, then $\cosh q_{k}=\sinh h_{j}\cdot\sinh d_{ki}$.
+Since $v_{k}\ast c_{ijk}=-\cosh q_{k}$, $v_{k}\ast c_{ik}=-\sinh d_{ki}$, and $c_{ik}\ast c_{ijk}=-\sinh h_{j}$, it follows that $\cosh q_{k}=\sinh h_{j}\cdot\sinh d_{ki}$.
 In the generalized right-angled hyperbolic triangle $c_{ijk}c^{\prime}_{ik}v^{\prime}_{i}$, we have $-v^{\prime}_{i}\ast c_{ijk}=(c^{\prime}_{ik}\ast c_{ijk})(v^{\prime}_{i}\ast c^{\prime}_{ik})$.
-Since $v^{\prime}_{i}\ast c_{ijk}=\cosh h_{i}$ and $v^{\prime}_{i}\ast c^{\prime}_{ik}=-\sinh\theta_{ik}$.
-If $c^{\prime}_{ik}$ and $c_{ijk}$ are on same sides of the hyperplane $c_{ijk}^{\bot}$, then $c^{\prime}_{ik}\ast c_{ijk}=+\sinh(d_{\mathbb{H}}(c^{\prime}_{ik},c_{ijk}^{\bot}))=+\sinh(-q_{j})=-\sinh q_{j}$.
-If $c^{\prime}_{ik}$ and $c_{ijk}$ are on opposite sides of the hyperplane $c_{ijk}^{\bot}$, then $c^{\prime}_{ij}\ast c_{ijk}=-\sinh(d_{\mathbb{H}}(c^{\prime}_{ik},c_{ijk}^{\bot}))=-\sinh(+q_{j})=-\sinh q_{j}$.
-Thus $c^{\prime}_{ik}\ast c_{ijk}=-\sinh q_{j}$.
-Then $\cosh h_{i}=-\sinh q_{j}\cdot\sinh\theta_{ik}$.
+Similarly, $v^{\prime}_{i}\ast c_{ijk}=\cosh h_{i}$ and $v^{\prime}_{i}\ast c^{\prime}_{ik}=-\sinh\theta_{ik}$.
+If $c^{\prime}_{ik}$ and $c_{ijk}$ lie on the same side of the hyperplane $c_{ijk}^{\bot}$, then $c^{\prime}_{ik}\ast c_{ijk}=+\sinh(d_{\mathbb{H}}(c^{\prime}_{ik},c_{ijk}^{\bot}))=+\sinh(-q_{j})=-\sinh q_{j}$.
+If $c^{\prime}_{ik}$ and $c_{ijk}$ lie on opposite sides of the hyperplane $c_{ijk}^{\bot}$, then $c^{\prime}_{ij}\ast c_{ijk}=-\sinh(d_{\mathbb{H}}(c^{\prime}_{ik},c_{ijk}^{\bot}))=-\sinh(+q_{j})=-\sinh q_{j}$.
+In either case, $c^{\prime}_{ik}\ast c_{ijk}=-\sinh q_{j}$.
+Consequently, $\cosh h_{i}=-\sinh q_{j}\cdot\sinh\theta_{ik}$.
 By Lemma 2.6, it is easy to check that $\theta_{ik}<0$.
-Hence $q_{j}>0$.
-Then $d_{\mathbb{H}}(c^{\prime}_{ik},c_{ijk}^{\bot})=q_{j}$, which implies that $c^{\prime}_{ik}$ and $c_{ijk}$ are on opposite sides of the hyperplane $c_{ijk}^{\bot}$.
+It follows that $q_{j}>0$ and hence $d_{\mathbb{H}}(c^{\prime}_{ik},c_{ijk}^{\bot})=q_{j}$.
+This implies that $c^{\prime}_{ik}$ and $c_{ijk}$ lie on opposite sides of the hyperplane $c_{ijk}^{\bot}$.
 In the generalized right-angled hyperbolic triangle $c_{ijk}c^{\prime}_{ij}v^{\prime}_{i}$, we have $-v^{\prime}_{i}\ast c_{ijk}=(c^{\prime}_{ij}\ast c_{ijk})(v^{\prime}_{i}\ast c^{\prime}_{ij})$.
-Since $v^{\prime}_{i}\ast c_{ijk}=\cosh h_{i}$ and $v^{\prime}_{i}\ast c^{\prime}_{ij}=-\sinh\theta_{ij}$.
-If $c^{\prime}_{ij}$ and $c_{ijk}$ are on same sides of the hyperplane $c_{ijk}^{\bot}$, then $c^{\prime}_{ij}\ast c_{ijk}=+\sinh(d_{\mathbb{H}}(c^{\prime}_{ij},c_{ijk}^{\bot}))=+\sinh(-q_{k})=-\sinh q_{k}$.
-If $c^{\prime}_{ij}$ and $c_{ijk}$ are on opposite sides of the hyperplane $c_{ijk}^{\bot}$, then $c^{\prime}_{ij}\ast c_{ijk}=-\sinh(d_{\mathbb{H}}(c^{\prime}_{ij},c_{ijk}^{\bot}))=-\sinh(+q_{k})=-\sinh q_{k}$.
-Thus $c^{\prime}_{ij}\ast c_{ijk}=-\sinh q_{k}$.
-Then $\cosh h_{i}=-\sinh q_{k}\cdot\sinh\theta_{ij}$.
+Similarly, $v^{\prime}_{i}\ast c_{ijk}=\cosh h_{i}$ and $v^{\prime}_{i}\ast c^{\prime}_{ij}=-\sinh\theta_{ij}$.
+If $c^{\prime}_{ij}$ and $c_{ijk}$ lie on the same side of the hyperplane $c_{ijk}^{\bot}$, then $c^{\prime}_{ij}\ast c_{ijk}=+\sinh(d_{\mathbb{H}}(c^{\prime}_{ij},c_{ijk}^{\bot}))=+\sinh(-q_{k})=-\sinh q_{k}$.
+If $c^{\prime}_{ij}$ and $c_{ijk}$ lie on opposite sides of the hyperplane $c_{ijk}^{\bot}$, then $c^{\prime}_{ij}\ast c_{ijk}=-\sinh(d_{\mathbb{H}}(c^{\prime}_{ij},c_{ijk}^{\bot}))=-\sinh(+q_{k})=-\sinh q_{k}$.
+In either case, $c^{\prime}_{ij}\ast c_{ijk}=-\sinh q_{k}$.
+Consequently, $\cosh h_{i}=-\sinh q_{k}\cdot\sinh\theta_{ij}$.
 Similarly, $\theta_{ij}<0$ by Lemma 2.6 and hence $q_{k}>0$.
@@
 In the generalized right-angled hyperbolic triangle $c_{ijk}c^{\prime}_{jk}v^{\prime}_{j}$, we have $-v^{\prime}_{j}\ast c_{ijk}=(c^{\prime}_{jk}\ast c_{ijk})(v^{\prime}_{j}\ast c^{\prime}_{jk})$.
-Since $v^{\prime}_{j}\ast c_{ijk}=-\cosh h_{j}$, $v^{\prime}_{j}\ast c^{\prime}_{jk}=-\sinh\theta_{jk}$ and $c^{\prime}_{jk}\ast c_{ijk}=-\sinh q_{i}$.
-Then $\cosh h_{j}=\sinh q_{i}\cdot\sinh\theta_{jk}$.
+Since $v^{\prime}_{j}\ast c_{ijk}=-\cosh h_{j}$, $v^{\prime}_{j}\ast c^{\prime}_{jk}=-\sinh\theta_{jk}$, and $c^{\prime}_{jk}\ast c_{ijk}=-\sinh q_{i}$, it follows that $\cosh h_{j}=\sinh q_{i}\cdot\sinh\theta_{jk}$.
 In the generalized right-angled hyperbolic triangle $c_{ijk}c^{\prime}_{ij}v^{\prime}_{j}$, we have $-v^{\prime}_{j}\ast c_{ijk}=(c^{\prime}_{ij}\ast c_{ijk})(v^{\prime}_{j}\ast c^{\prime}_{ij})$.
-Since $v^{\prime}_{j}\ast c_{ijk}=-\cosh h_{j}$, $v^{\prime}_{j}\ast c^{\prime}_{ij}=-\sinh\theta_{ji}$ and $c^{\prime}_{ij}\ast c_{ijk}=-\sinh q_{k}$, then $\cosh h_{j}=\sinh q_{k}\cdot\sinh\theta_{ji}$.
+Since $v^{\prime}_{j}\ast c_{ijk}=-\cosh h_{j}$, $v^{\prime}_{j}\ast c^{\prime}_{ij}=-\sinh\theta_{ji}$, and $c^{\prime}_{ij}\ast c_{ijk}=-\sinh q_{k}$, it follows that $\cosh h_{j}=\sinh q_{k}\cdot\sinh\theta_{ji}$.
 In the generalized right-angled hyperbolic triangle $c_{ijk}c^{\prime}_{jk}v^{\prime}_{k}$, we have $-v^{\prime}_{k}\ast c_{ijk}=(c^{\prime}_{jk}\ast c_{ijk})(v^{\prime}_{k}\ast c^{\prime}_{jk})$.
-Since $v^{\prime}_{k}\ast c_{ijk}=-\cosh h_{k}$, $v^{\prime}_{k}\ast c^{\prime}_{jk}=-\sinh\theta_{kj}$ and $c^{\prime}_{jk}\ast c_{ijk}=-\sinh q_{i}$, then $\cosh h_{k}=\sinh q_{i}\cdot\sinh\theta_{kj}$.
+Since $v^{\prime}_{k}\ast c_{ijk}=-\cosh h_{k}$, $v^{\prime}_{k}\ast c^{\prime}_{jk}=-\sinh\theta_{kj}$, and $c^{\prime}_{jk}\ast c_{ijk}=-\sinh q_{i}$, it follows that $\cosh h_{k}=\sinh q_{i}\cdot\sinh\theta_{kj}$.
 In the generalized right-angled hyperbolic triangle $c_{ijk}c^{\prime}_{ik}v^{\prime}_{k}$, we have $-v^{\prime}_{k}\ast c_{ijk}=(c^{\prime}_{ik}\ast c_{ijk})(v^{\prime}_{k}\ast c^{\prime}_{ik})$.
-Since $v^{\prime}_{k}\ast c_{ijk}=-\cosh h_{k}$, $v^{\prime}_{k}\ast c^{\prime}_{ik}=-\sinh\theta_{ki}$ and $c^{\prime}_{ik}\ast c_{ijk}=-\sinh q_{j}$, then $\cosh h_{k}=\sinh q_{j}\cdot\sinh\theta_{ki}$.
+Since $v^{\prime}_{k}\ast c_{ijk}=-\cosh h_{k}$, $v^{\prime}_{k}\ast c^{\prime}_{ik}=-\sinh\theta_{ki}$, and $c^{\prime}_{ik}\ast c_{ijk}=-\sinh q_{j}$, it follows that $\cosh h_{k}=\sinh q_{j}\cdot\sinh\theta_{ki}$.
 Q.E.D.
@@
 In the generalized right-angled hyperbolic triangle $c_{ijk}c_{ik}v_{i}$, we have $-v_{i}\ast c_{ijk}=(c_{ik}\ast c_{ijk})(v_{i}\ast c_{ik})$.
-By Proposition 2.2 $\mathrm{(iii)}$, $v_{i}\ast c_{ijk}=\cosh d_{\mathbb{H}}(v_{i}^{\bot},c_{ijk}^{\bot})=\cosh q_{i}$.
-By Proposition 2.2 $\mathrm{(ii)}$, $v_{i}\ast c_{ik}=+\sinh d_{\mathbb{H}}(v_{i}^{\bot},c_{ik})=+\sinh(-d_{ik})=-\sinh d_{ik}$.
-If $c_{ik}$ and $c_{ijk}$ are on same sides of the hyperplane $c_{ijk}^{\bot}$, then $c_{ik}\ast c_{ijk}=+\sinh(d_{\mathbb{H}}(c_{ik},c_{ijk}^{\bot}))=+\sinh h_{j}$; If $c_{ik}$ and $c_{ijk}$ are on opposite sides of the hyperplane $c_{ijk}^{\bot}$, then $c_{ik}\ast c_{ijk}=-\sinh(d_{\mathbb{H}}(c_{ik},c_{ijk}^{\bot}))=-\sinh(-h_{j})=+\sinh h_{j}$.
-Thus $c_{ik}\ast c_{ijk}=\sinh h_{j}$.
-Then $\cosh q_{i}=\sinh h_{j}\cdot\sinh d_{ik}$.
+By Proposition 2.2 (iii), $v_{i}\ast c_{ijk}=\cosh d_{\mathbb{H}}(v_{i}^{\bot},c_{ijk}^{\bot})=\cosh q_{i}$.
+By Proposition 2.2 (ii), $v_{i}\ast c_{ik}=+\sinh d_{\mathbb{H}}(v_{i}^{\bot},c_{ik})=+\sinh(-d_{ik})=-\sinh d_{ik}$.
+If $c_{ik}$ and $c_{ijk}$ lie on the same side of the hyperplane $c_{ijk}^{\bot}$, then $c_{ik}\ast c_{ijk}=+\sinh(d_{\mathbb{H}}(c_{ik},c_{ijk}^{\bot}))=+\sinh h_{j}$; if $c_{ik}$ and $c_{ijk}$ lie on opposite sides of the hyperplane $c_{ijk}^{\bot}$, then $c_{ik}\ast c_{ijk}=-\sinh(d_{\mathbb{H}}(c_{ik},c_{ijk}^{\bot}))=-\sinh(-h_{j})=+\sinh h_{j}$.
+In either case, $c_{ik}\ast c_{ijk}=\sinh h_{j}$.
+Consequently, $\cosh q_{i}=\sinh h_{j}\cdot\sinh d_{ik}$.
 By Lemma 2.5, it is easy to check that $d_{ik}<0$.
-Then $h_{j}<0$ and hence $d_{\mathbb{H}}(c_{ik},c_{ijk}^{\bot})=-h_{j}$, which implies that $c_{ik}$ and $c_{ijk}$ are on opposite sides of the hyperplane $c_{ijk}^{\bot}$.
+It follows that $h_{j}<0$ and hence $d_{\mathbb{H}}(c_{ik},c_{ijk}^{\bot})=-h_{j}$.
+This implies that $c_{ik}$ and $c_{ijk}$ lie on opposite sides of $c_{ijk}^{\bot}$.
 In the generalized right-angled hyperbolic triangle $c_{ijk}c_{ij}v_{i}$, we have $-v_{i}\ast c_{ijk}=(c_{ij}\ast c_{ijk})(v_{i}\ast c_{ij})$.
-Similarly, $v_{i}\ast c_{ij}=-\sinh d_{ij}$ and $c_{ij}\ast c_{ijk}=-\sinh h_{k}$.
-Then $\cosh q_{i}=-\sinh h_{k}\cdot\sinh d_{ij}$.
+Since $v_{i}\ast c_{ij}=-\sinh d_{ij}$ and $c_{ij}\ast c_{ijk}=-\sinh h_{k}$, it follows that $\cosh q_{i}=-\sinh h_{k}\cdot\sinh d_{ij}$.
 In the generalized right-angled hyperbolic triangle $c_{ijk}c_{jk}v_{j}$, we have $-v_{j}\ast c_{ijk}=(c_{jk}\ast c_{ijk})(v_{j}\ast c_{jk})$.
-Since $v_{j}\ast c_{ijk}=-\cosh q_{j}$, $v_{j}\ast c_{jk}=-\sinh d_{jk}$ and $c_{jk}\ast c_{ijk}=-\sinh h_{i}$, then $\cosh q_{j}=\sinh h_{i}\cdot\sinh d_{jk}$.
+Since $v_{j}\ast c_{ijk}=-\cosh q_{j}$, $v_{j}\ast c_{jk}=-\sinh d_{jk}$, and $c_{jk}\ast c_{ijk}=-\sinh h_{i}$, it follows that $\cosh q_{j}=\sinh h_{i}\cdot\sinh d_{jk}$.
 In the generalized right-angled hyperbolic triangle $c_{ijk}c_{ij}v_{j}$, we have $-v_{j}\ast c_{ijk}=(c_{ij}\ast c_{ijk})(v_{j}\ast c_{ij})$.
-Since $v_{j}\ast c_{ijk}=-\cosh q_{j}$, $v_{j}\ast c_{ij}=-\sinh d_{ji}$ and $c_{ij}\ast c_{ijk}=-\sinh h_{k}$, then $\cosh q_{j}=\sinh h_{k}\cdot\sinh d_{ji}$.
+Since $v_{j}\ast c_{ijk}=-\cosh q_{j}$, $v_{j}\ast c_{ij}=-\sinh d_{ji}$, and $c_{ij}\ast c_{ijk}=-\sinh h_{k}$, it follows that $\cosh q_{j}=\sinh h_{k}\cdot\sinh d_{ji}$.
 In the generalized right-angled hyperbolic triangle $c_{ijk}c_{jk}v_{k}$, we have $-v_{k}\ast c_{ijk}=(c_{jk}\ast c_{ijk})(v_{k}\ast c_{jk})$.
-Since $v_{k}\ast c_{ijk}=-\cosh q_{k}$, $v_{k}\ast c_{jk}=-\sinh d_{kj}$ and $c_{jk}\ast c_{ijk}=-\sinh h_{i}$, then $\cosh q_{k}=\sinh h_{i}\cdot\sinh d_{kj}$.
+Since $v_{k}\ast c_{ijk}=-\cosh q_{k}$, $v_{k}\ast c_{jk}=-\sinh d_{kj}$, and $c_{jk}\ast c_{ijk}=-\sinh h_{i}$, it follows that $\cosh q_{k}=\sinh h_{i}\cdot\sinh d_{kj}$.
 In the generalized right-angled hyperbolic triangle $c_{ijk}c_{ik}v_{k}$, we have $-v_{k}\ast c_{ijk}=(c_{ik}\ast c_{ijk})(v_{k}\ast c_{ik})$.
-Since $v_{k}\ast c_{ijk}=-\cosh q_{k}$, $v_{k}\ast c_{ik}=-\sinh d_{ki}$ and $c_{ik}\ast c_{ijk}=\sinh h_{j}$, then $\cosh q_{k}=-\sinh h_{j}\cdot\sinh d_{ki}$.
+Since $v_{k}\ast c_{ijk}=-\cosh q_{k}$, $v_{k}\ast c_{ik}=-\sinh d_{ki}$, and $c_{ik}\ast c_{ijk}=\sinh h_{j}$, it follows that $\cosh q_{k}=-\sinh h_{j}\cdot\sinh d_{ki}$.
 In the generalized right-angled hyperbolic triangle $c_{ijk}c^{\prime}_{ik}v^{\prime}_{i}$, we have $-v^{\prime}_{i}\ast c_{ijk}=(c^{\prime}_{ik}\ast c_{ijk})(v^{\prime}_{i}\ast c^{\prime}_{ik})$.
-Since $v^{\prime}_{i}\ast c_{ijk}=-\cosh h_{i}$, $v^{\prime}_{i}\ast c^{\prime}_{ik}=-\sinh\theta_{ik}$ and $c^{\prime}_{ik}\ast c_{ijk}=-\sinh q_{j}$, then $\cosh h_{i}=\sinh q_{j}\cdot\sinh\theta_{ik}$.
+Since $v^{\prime}_{i}\ast c_{ijk}=-\cosh h_{i}$, $v^{\prime}_{i}\ast c^{\prime}_{ik}=-\sinh\theta_{ik}$, and $c^{\prime}_{ik}\ast c_{ijk}=-\sinh q_{j}$, it follows that $\cosh h_{i}=\sinh q_{j}\cdot\sinh\theta_{ik}$.
 In the generalized right-angled hyperbolic triangle $c_{ijk}c^{\prime}_{ij}v^{\prime}_{i}$, we have $-v^{\prime}_{i}\ast c_{ijk}=(c^{\prime}_{ij}\ast c_{ijk})(v^{\prime}_{i}\ast c^{\prime}_{ij})$.
-Since $v^{\prime}_{i}\ast c_{ijk}=-\cosh h_{i}$, $v^{\prime}_{i}\ast c^{\prime}_{ij}=-\sinh\theta_{ij}$ and $c^{\prime}_{ij}\ast c_{ijk}=-\sinh q_{k}$, then $\cosh h_{i}=\sinh q_{k}\cdot\sinh\theta_{ij}$.
+Since $v^{\prime}_{i}\ast c_{ijk}=-\cosh h_{i}$, $v^{\prime}_{i}\ast c^{\prime}_{ij}=-\sinh\theta_{ij}$, and $c^{\prime}_{ij}\ast c_{ijk}=-\sinh q_{k}$, it follows that $\cosh h_{i}=\sinh q_{k}\cdot\sinh\theta_{ij}$.
 In the generalized right-angled hyperbolic triangle $c_{ijk}c^{\prime}_{jk}v^{\prime}_{j}$, we have $-v^{\prime}_{j}\ast c_{ijk}=(c^{\prime}_{jk}\ast c_{ijk})(v^{\prime}_{j}\ast c^{\prime}_{jk})$.
-Since $v^{\prime}_{j}\ast c_{ijk}=\cosh h_{j}$ and $v^{\prime}_{j}\ast c^{\prime}_{jk}=-\sinh\theta_{jk}$.
-If $c^{\prime}_{jk}$ and $c_{ijk}$ are on same sides of the hyperplane $c_{ijk}^{\bot}$, then $c^{\prime}_{jk}\ast c_{ijk}=+\sinh(d_{\mathbb{H}}(c^{\prime}_{jk},c_{ijk}^{\bot}))=+\sinh q_{i}$; If $c^{\prime}_{jk}$ and $c_{ijk}$ are on opposite sides of the hyperplane $c_{ijk}^{\bot}$, then $c^{\prime}_{jk}\ast c_{ijk}=-\sinh(d_{\mathbb{H}}(c^{\prime}_{jk},c_{ijk}^{\bot}))=-\sinh(-q_{i})=+\sinh q_{i}$.
-Thus $c^{\prime}_{jk}\ast c_{ijk}=\sinh q_{i}$.
-Then $\cosh h_{j}=\sinh q_{i}\cdot\sinh\theta_{jk}$.
+Moreover, $v^{\prime}_{j}\ast c_{ijk}=\cosh h_{j}$ and $v^{\prime}_{j}\ast c^{\prime}_{jk}=-\sinh\theta_{jk}$.
+If $c^{\prime}_{jk}$ and $c_{ijk}$ lie on the same side of the hyperplane $c_{ijk}^{\bot}$, then $c^{\prime}_{jk}\ast c_{ijk}=+\sinh(d_{\mathbb{H}}(c^{\prime}_{jk},c_{ijk}^{\bot}))=+\sinh q_{i}$; if $c^{\prime}_{jk}$ and $c_{ijk}$ lie on opposite sides of the hyperplane $c_{ijk}^{\bot}$, then $c^{\prime}_{jk}\ast c_{ijk}=-\sinh(d_{\mathbb{H}}(c^{\prime}_{jk},c_{ijk}^{\bot}))=-\sinh(-q_{i})=+\sinh q_{i}$.
+In either case, $c^{\prime}_{jk}\ast c_{ijk}=\sinh q_{i}$.
+Consequently, $\cosh h_{j}=\sinh q_{i}\cdot\sinh\theta_{jk}$.
 By Lemma 2.6, it is easy to check that $\theta_{jk}<0$.
-Then $q_{i}<0$ and hence $d_{\mathbb{H}}(c^{\prime}_{jk},c_{ijk}^{\bot})=-q_{i}$, which implies that $c^{\prime}_{jk}$ and $c_{ijk}$ are on opposite sides of the hyperplane $c_{ijk}^{\bot}$.
+It follows that $q_{i}<0$ and hence $d_{\mathbb{H}}(c^{\prime}_{jk},c_{ijk}^{\bot})=-q_{i}$.
+This implies that $c^{\prime}_{jk}$ and $c_{ijk}$ lie on opposite sides of the hyperplane $c_{ijk}^{\bot}$.
 In the generalized right-angled hyperbolic triangle $c_{ijk}c^{\prime}_{ij}v^{\prime}_{j}$, we have $-v^{\prime}_{j}\ast c_{ijk}=(c^{\prime}_{ij}\ast c_{ijk})(v^{\prime}_{j}\ast c^{\prime}_{ij})$.
-Since $v^{\prime}_{j}\ast c_{ijk}=\cosh h_{j}$, $v^{\prime}_{j}\ast c^{\prime}_{ij}=-\sinh\theta_{ji}$ and $c^{\prime}_{ij}\ast c_{ijk}=-\sinh q_{k}$, then $\cosh h_{j}=-\sinh q_{k}\cdot\sinh\theta_{ji}$.
+Since $v^{\prime}_{j}\ast c_{ijk}=\cosh h_{j}$, $v^{\prime}_{j}\ast c^{\prime}_{ij}=-\sinh\theta_{ji}$, and $c^{\prime}_{ij}\ast c_{ijk}=-\sinh q_{k}$, it follows that $\cosh h_{j}=-\sinh q_{k}\cdot\sinh\theta_{ji}$.
 In the generalized right-angled hyperbolic triangle $c_{ijk}c^{\prime}_{jk}v^{\prime}_{k}$, we have $-v^{\prime}_{k}\ast c_{ijk}=(c^{\prime}_{jk}\ast c_{ijk})(v^{\prime}_{k}\ast c^{\prime}_{jk})$.
-Since $v^{\prime}_{k}\ast c_{ijk}=-\cosh h_{k}$, $v^{\prime}_{k}\ast c^{\prime}_{jk}=-\sinh\theta_{kj}$ and $c^{\prime}_{jk}\ast c_{ijk}=\sinh q_{i}$, then $\cosh h_{k}=-\sinh q_{i}\cdot\sinh\theta_{kj}$.
+Since $v^{\prime}_{k}\ast c_{ijk}=-\cosh h_{k}$, $v^{\prime}_{k}\ast c^{\prime}_{jk}=-\sinh\theta_{kj}$, and $c^{\prime}_{jk}\ast c_{ijk}=\sinh q_{i}$, it follows that $\cosh h_{k}=-\sinh q_{i}\cdot\sinh\theta_{kj}$.
 In the generalized right-angled hyperbolic triangle $c_{ijk}c^{\prime}_{ik}v^{\prime}_{k}$, we have $-v^{\prime}_{k}\ast c_{ijk}=(c^{\prime}_{ik}\ast c_{ijk})(v^{\prime}_{k}\ast c^{\prime}_{ik})$.
-Since $v^{\prime}_{k}\ast c_{ijk}=-\cosh h_{k}$, $v^{\prime}_{k}\ast c^{\prime}_{ik}=-\sinh\theta_{ki}$ and $c^{\prime}_{ik}\ast c_{ijk}=-\sinh q_{j}$, then $\cosh h_{k}=\sinh q_{j}\cdot\sinh\theta_{ki}$.
+Since $v^{\prime}_{k}\ast c_{ijk}=-\cosh h_{k}$, $v^{\prime}_{k}\ast c^{\prime}_{ik}=-\sinh\theta_{ki}$, and $c^{\prime}_{ik}\ast c_{ijk}=-\sinh q_{j}$, it follows that $\cosh h_{k}=\sinh q_{j}\cdot\sinh\theta_{ki}$.
 Q.E.D.